### v3

Abstract.
In this paper, we introduce the $F$-invariant in cluster algebras using tropicalization. This is an analog of the $E$-invariant introduced by Derksen, Weyman and Zelevinsky in the additive categorification of cluster algebras and the $\mathfrak{d}$-invariant introduced by Kang, Kashiwara, Kim and Oh in the monoidal categorification of (quantum) cluster algebras. We prove that the product of two cluster monomials is still a cluster monomial if and only if their $F$-invariant is zero. For cluster algebras with a compatible Poisson bracket, we prove that if two cluster variables are log-cannonical, then they are contained in the same cluster.
Inspired by $F$-invariant, we introduce the dominant sets for seeds of cluster algebras as an replacement of torsion classes for $\tau$-tilting pairs in $\tau$-tilting theory and as an replacement of inversion sets in the study of right weak order on Weyl groups. With the help of the dominant sets, we prove that the oriented exchange graphs of cluster algebras are acyclic. In particular, this implies that green mutations induce a partial order on the set of seeds (up to seed equivalence) of cluster algebras. We prove that the oriented exchange graphs of cluster algebras coincide with the Hasse quivers of the above posets of seeds.
2020 Mathematics Subject Classification:
13F60, 16G20, 20F55
1. Introduction
Cluster algebras. Cluster algebras were invented by Fomin and Zelevinsky [fz_2002] as a combinatorial approach to the upper global bases of quantum groups and to the theory of total positivity in algebraic groups. Such algebras often arise as the coordinate rings of spaces, such as double Bruhat cells [bfz_2005], unipotent cells [gls_2011], double Bott-Samelson cells [Shen-Weng-2021] and arise as the Grothendieck rings of certain monoidal subcategories of the representations of quantum affine algebras [HL_2010, HL_2013].
A cluster algebra is a $\mathbb{Z}$-subalgebra of a rational function field generated by a special set of generators called cluster variables, which are grouped into overlapping subsets, called clusters. A seed is a pair consisting of a cluster and a rectangular integer matrix with skew-symmetrizable principal part. One can obtain new seeds from a given one by a procedure called mutation. The sets of cluster variables and clusters of a cluster algebra are determined by an initial seed and the iterative mutations. A cluster momomial is a monomial in variables from the same cluster. In the main text, we actually use a broader definition of cluster monomials, which allows the presence of the inverse of frozen variables.
A fundamental problem in cluster algebras is to characterize when two cluster variables are contained in the same cluster. A more general problem is to characterize when the product of two cluster monomials is still a cluster monomial. Many progresses have been made by introducing different type of compatibility degrees or invariants.
Compatibility degrees in cluster algebras. In the study of Y-systems and generalized associahedra, Fomin and Zelevinsky [fz-2003y] introduce a function
$(-\mid\mid-):\Phi_{\geq-1}\times\Phi_{\geq-1}\rightarrow\mathbb{Z}_{\geq 0}$
defined on the set $\Phi_{\geq-1}$ of almost positive roots associated to a Cartan matrix $C$ of finite type. Let $\mathcal{A}$ be a cluster algebra of finite type with the same Cartan-Killing type as the Cartan matrix $C$. Fomin and Zelevinsky [fz_2003f] proved that there is a bijection from $\Phi_{\geq-1}$ to the set $\mathcal{X}$ of cluster variables of $\mathcal{A}$ and they proved that two cluster variables $x_{\alpha}$ and $x_{\beta}$ are contained in the same cluster if and only if their corresponding roots $\alpha,\beta\in\Phi_{\geq-1}$ are compatible, i.e., $(\alpha\mid\mid\beta)=0$.
Now let $\mathcal{A}$ be any skew-symmetrizable cluster algebra (not necessarily finite type). In [cao-li-2020], Li and the author of the present paper proved that there exists a well-defined function
$(-\mid\mid-)_{d}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{Z}_{\geq-1}$
defined on the set of cluster variables $\mathcal{X}$ of $\mathcal{A}$, which is called the $d$-compatibility degree. The values of the $d$-compatibility degree $(-\mid\mid-)_{d}$ are given by the components of the $d$-vectors [fomin_zelevinsky_2007] in cluster algebras. It is proved that two cluster variables $x,z\in\mathcal{X}$ are contained in the same cluster if and only if
$(x\mid\mid z)_{d}=0\;\;\;\text{or}\;\;\;(x\mid\mid z)_{d}=-1,$
moreover, $(x\mid\mid z)_{d}=-1$ if and only if $x=z$. This confirms a conjecture on $d$-vectors proposed by Fomin and Zelevinsky [fomin_zelevinsky_2007, Conjecture 7.4].
Recently, Fu and Gyoda [Fu-Gyoda] introduced the $f$-compatibility degree in cluster algebras using the components of $f$-vectors, which is a function
$(-\mid\mid-)_{f}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{Z}_{\geq 0}$
defined on the set $\mathcal{X}$ of cluster variables of $\mathcal{A}$. Based on the results in [cao-li-2020], Fu and Gyoda proved that two cluster variables $x,z\in\mathcal{X}$ are contained in the same cluster if and only if $(x\mid\mid z)_{f}=0$. We remark that when restricting to cluster algebras of finite type, the $f$-compatibility degree coincide with Fomin-Zelevinsky’s compatibility degree.
$E$-invariant. The $E$-invariant is introduced by Derksen, Weyman and Zelevinsky [DWZ10] in the additive categorification of cluster algebras using decorated representations of quivers with potentials. We refer to [DWZ08, Definition 10.1] for the definition of decorated representations. For any two decorated representations $\mathcal{M}$ and $\mathcal{N}$ of a quiver with potential $(Q,W)$, Derksen, Weyman and Zelevinsky [DWZ10, Section 7] defined an integer $E^{\rm inj}(\mathcal{M},\mathcal{N})$, which turns out to take non-negative values (see [DWZ10, Section 10]). The $E$-invariant $E^{\rm sym}(\mathcal{M},\mathcal{N})$ associated to $\mathcal{M}$ and $\mathcal{N}$ is defined to be the symmetrized sum
$E^{\rm sym}(\mathcal{M},\mathcal{N})=E^{\rm inj}(\mathcal{M},\mathcal{N})+E^{\rm inj}(\mathcal{N},\mathcal{M})\geq 0.$
Derksen, Weyman and Zelevinsky [DWZ10, Theorem 7.1] proved that the integer $E^{\rm sym}(\mathcal{M},\mathcal{N})$ is mutation-invariant in the sense that
$E^{\rm sym}(\mathcal{M},\mathcal{N})=E^{\rm sym}(\mu_{k}(\mathcal{M}),\mu_{k}(\mathcal{N})),$
where $\mu_{k}(\mathcal{M})$ and $\mu_{k}(\mathcal{N})$ are decorated representations of the new quiver with potential $\mu_{k}(Q,W)$ obtained by mutation in direction $k$. As a corollary, they obtained $E^{\rm sym}(\mathcal{M},\mathcal{M})=0$ for any negative-reachable decorated representations, i.e., those decorated representations corresponding to cluster monomials. This result plays a crucial role in their proofs of [DWZ10, Conjectures 1.3–1.6].
Let $(Q,W)$ be a quiver with potential. There is a dg algebra $\Gamma=\Gamma_{(Q,W)}$ and a triangulated category $\mathcal{C}_{(Q,W)}$ associated to $(Q,W)$, which are called (complete) Ginzburg dg algebra and cluster category, cf., [Amiot_2009] and [Keller-Yang_2011]. We remark that $\Gamma$ can be viewed as an object in $\mathcal{C}_{(Q,W)}$ satisfying $\operatorname{\mathsf{Hom}}_{\mathcal{C}_{(Q,W)}}(\Gamma,\Gamma[1])=0$. Plamondon [Plamondon_2011b, Section 4.1] constructed a bijection $\Psi_{(Q,W)}$ from the isomorphism classes of decorated representations of $(Q,W)$ to the isomorphism classes of objects in a full subcategory of $\mathcal{C}_{(Q,W)}$. For two decorated representations $\mathcal{M}$ and $\mathcal{N}$ of $(Q,W)$, we denote by $M=\Psi_{(Q,W)}(\mathcal{M})$ and $N=\Psi_{(Q,W)}(\mathcal{N})$ the corresponding objects in the cluster category $\mathcal{C}_{(Q,W)}$. In [Plamondon_2011b, Proposition 4.15], Plamondon proved
$E^{\rm inj}(\mathcal{M},\mathcal{N})=\dim(\Gamma[1])(M,N[1]),$
where $(\Gamma[1])(M,N[1])$ denotes the subspace of $\operatorname{\mathsf{Hom}}_{\mathcal{C}_{(Q,W)}}(M,N[1])$ consisting of morphisms that factor through an object of $\operatorname{\mathsf{add}}\nolimits(\Gamma[1])$. Thus we have
$E^{\rm sym}(\mathcal{M},\mathcal{N})=\dim(\Gamma[1])(M,N[1])+\dim(\Gamma[1])(N,M[1]).$
For the case of $\mathcal{M}=\mathcal{N}$, Plamondon also proved that the above equality becomes
$E^{\rm sym}(\mathcal{M},\mathcal{M})=\dim\operatorname{\mathsf{Hom}}_{\mathcal{C}_{(Q,W)}}(M,M[1]).$
One has that $E^{\rm sym}(\mathcal{M},\mathcal{M})=0$ if and only if the corresponding object $M$ in the cluster category $\mathcal{C}_{(Q,W)}$ is rigid, that is, $\operatorname{\mathsf{Hom}}_{\mathcal{C}_{(Q,W)}}(M,M[1])=0$. Notice that if $\mathcal{M}$ corresponds to a cluster monomial, then $M$ is rigid in $\mathcal{C}_{(Q,W)}$. Let us mention another very special case. If $N$ belongs to $\operatorname{\mathsf{add}}\nolimits(\Gamma)$, that is, $\mathcal{N}$ corresponds to a cluster monomial in initial cluster variables, then it is easy to see
(1.1) $E^{\rm sym}(\mathcal{M},\mathcal{N})=\dim\operatorname{\mathsf{Hom}}_{\mathcal{C}_{(Q,W)}}(M,N[1]).$
Inspired by Derksen-Weyman-Zelevinsky’s $E$-invariant for Jacobian algebras, the $E$-invariant for any finite dimensional algebra is introduced by Derksen and Fei [DK-2015] in the study of general presentations and by Adachi, Iyama and Reiten [air_2014] in the study of $\tau$-tilting theory.
Recently, tropical $F$-polynomials were studied by Fei [fei_2019a, fei_2019b] from the viewpoint of representation theory. It turns out that $E^{\rm inj}(\mathcal{M},\mathcal{N})$ is related to the evaluations of tropical $F$-polynomials, c.f., [fei_2019b, Theorem 3.22].
$\mathfrak{d}$-invariant. The $\mathfrak{d}$-invariant is introduced by Kang, Kashiwara, Kim and Oh [kkko-2018] in the monoidal categorification of (quantum) cluster algebras using representations of quiver Hecke algebras. Here we will mainly follow [kkko-2018, kk_2019]. Let
[…]
$\Lambda(M,N)=-c.$
The $\mathfrak{d}$-invariant $\mathfrak{d}(M,N)$ associated to $(M,N)$ is defined to be
$\mathfrak{d}(M,N)=\frac{1}{2}(\Lambda(M,N)+\Lambda(N,M)).$
[…]
Recently, Kashiwara, Kim, Oh and Park [kkop-2020] introduced the similar invariants $\Lambda(M,N)$ and $\mathfrak{d}(M,N)$ into the monoidal categorification of cluster algebras using representations of quantum affine algebras. In particular, in this setting, they proved that (i) $\mathfrak{d}(M,N)$ takes non-negative values; (ii) if $M$ is a real simple object, then $\mathfrak{d}(M,M)=0$, cf. [kkop-2020, Proposition 3.16, Corollary 3.17].
Tropical invariant and $F$-invariant. In this paper, we introduce the $F$-invariant in cluster algebras using tropicalization. This is an analog of the $E$-invariant introduced by Derksen, Weyman and Zelevinsky in the additive categorification of cluster algebras and the $\mathfrak{d}$-invariant introduced by Kang, Kashiwara, Kim and Oh in the monoidal categorification of (quantum) cluster algebras. The $F$-invariant is also a great improvement of Fu-Gyoda’s $f$-compatibility degree defined on the set of cluster variables.
[…]
For any choice of seed $({\bf x}_{w},\widetilde{B}_{w})$, we define an integer $\langle u,u^{\prime}\rangle_{w}$ for a pair of good elements $u$ and $u^{\prime}$ in Definition 4.7. Our first main result is as follows:
Theorem 1.1 (Theorem 4.8).
The integer $\langle u,u^{\prime}\rangle_{w}$ only depends on $u$ and $u^{\prime}$, not on the choice of seed $({\bf x}_{w},\widetilde{B}_{w})$.
From the proof of Theorem 4.8, we can see that the integer $\langle u,u^{\prime}\rangle_{w}$ can be interpreted as the evaluation of the positive function $u$ (on a cluster $\mathscr{A}$-variety in the sense of [FG-2009]) at a tropical point (associated to the same cluster $\mathscr{A}$-variety) determined by $u^{\prime}$. Different choices of seeds provide different charts to calculate this evaluation.
By Theorem 1.1, we get an invariant $\langle u,u^{\prime}\rangle:=\langle u,u^{\prime}\rangle_{w}$ for any pair $(u,u^{\prime})$ of good elements. We call $\langle u,u^{\prime}\rangle$ the tropical invariant between $u$ and $u^{\prime}$. We remark that the tropical invariant $\langle u,u^{\prime}\rangle=\langle u,u^{\prime}\rangle_{w}$ is an analog of the invariant $\Lambda(M,N)$ in the monoidal categorification of cluster algebras, because both $\langle u,u^{\prime}\rangle$ and $\Lambda(M,N)$ are related to the quantization matrices appearing in compatible pairs that are used to define quantum cluster algebras in the sense of [bz-2005], cf. Remark 4.14.
The $F$-invariant $(u\mid\mid u^{\prime})_{F}$ is defined to be the symmetrized sum of the tropical invariants.
(1.2) $(u\mid\mid u^{\prime})_{F}:=\langle u,u^{\prime}\rangle+\langle u^{\prime},u\rangle=\langle u,u^{\prime}\rangle_{w}+\langle u^{\prime},u\rangle_{w}.$
We remark that the $F$-invariant defined above is an analog of the $E$-invariant $E^{\rm sym}(\mathcal{M},\mathcal{N})$ in the additive categorification and $\mathfrak{d}$-invariant $\mathfrak{d}(M,N)$ in the monoidal categorification. Strictly speaking, the $F$-invariant $(u\mid\mid u^{\prime})_{F}$ is analogous to $2\mathfrak{d}(M,N)$, because in (1.2) we did not divide out ``$2"$ from the symmetrized sum $\langle u,u^{\prime}\rangle+\langle u^{\prime},u\rangle$.
We prove that the $F$-invariant admits the similar properties with the $E$-invarant and $\mathfrak{d}$-invariant.
Proposition 1.2 (Proposition 4.11).
The following statements hold.
  • (i)
    For any pair $(u,u^{\prime})$ of good elements, we have $(u\mid\mid u^{\prime})_{F}\in\mathbb{Z}_{\geq 0}$.
  • (ii)
    If $u$ is a cluster monomial, then $(u\mid\mid u)_{F}=0$.
We remark that the class of good elements satisfying $(u\mid\mid u)_{F}=0$ is analogous to the class of decorated representations satisfying $E^{\rm sym}(\mathcal{M},\mathcal{M})=0$ and to the class of finite-dimension graded modules satisfying $\mathfrak{d}(M,M)=0$.
As mentioned in [kkop-2020, Pages 1040–1041], the integer-valued invariants $\Lambda(M,N)$ and $\mathfrak{d}(M,N)$ provide important information in the representation theory of quiver Hecke algebra. However, in general, computing those values is quite difficult. We hope our tropical invariant $\langle u,u^{\prime}\rangle$ and $F$-invariant $(u\mid\mid u^{\prime})_{F}$ can shed some light on the calculation of the integer-valued invariants $\Lambda(M,N)$ and $\mathfrak{d}(M,N)$ in the monoidal categorification of cluster algebras.
Since the $F$-invariant $(u\mid\mid u^{\prime})_{F}$ always takes non-negative values, it is natural to ask when $(u\mid\mid u^{\prime})_{F}=0$ holds.
Theorem 1.3 (Theorem 4.22).
Let $u$ and $u^{\prime}$ be two cluster monomials in an upper cluster algebra. Then $(u\mid\mid u^{\prime})_{F}=0$ if and only if the product $uu^{\prime}$ is still a cluster monomial.
For arbitrary two good elements $u$ and $u^{\prime}$ (not necessarily cluster monomials), we give a necessary condition for $(u\mid\mid u^{\prime})_{F}=0$ in Theorem 4.27. Before stating the result, we need some preparations.
We use $m$ to denote the number of cluster variables (frozen and unfrozen) in a cluster and use $n$ to denote the number of unfrozen cluster variables in a cluster. By Definition 3.6, each good element $u$ corresponds to a tropical point $[{\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ (associated to a cluster $\mathscr{X}$-variety in the sense of [FG-2009]). Two tropical points $[{\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ and $[{\bf g}^{\prime}]=\{({\bf g}^{\prime})^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ are said to be sign-coherent if for any vertex $t\in\mathbb{T}_{n}$ and any $k\in[1,n]$, we have $g_{k}^{t}\cdot(g_{k}^{\prime})^{t}\geq 0$, where $g_{k}^{t}$ and $(g_{k}^{\prime})^{t}$ are the $k$th components of ${\bf g}^{t}$ and $({\bf g}^{\prime})^{t}$.
Theorem 1.4 (Theorem 4.27).
Let $u$ and $u^{\prime}$ be two good elements in an upper cluster algebra $\mathcal{U}$. If $(u\mid\mid u^{\prime})_{F}=0$, then the tropical points correspond to $u$ and $u^{\prime}$ are sign-coherent and the product $uu^{\prime}$ is still a good element in $\mathcal{U}$.
We also have a discussion on the link between the compatible Poisson structure of (upper) cluster algebras and the $F$-invariant.
Theorem 1.5 (Theorem 4.29).
The following statements hold.
  • (i)
    If a good element $u$ and an unfrozen cluster variable $x$ are log-canonical with respect to the given compatible Poisson bracket, then $(x\mid\mid u)_{F}=0$.
  • (ii)
    If two cluster variables are log-canonical, then they are contained in the same cluster.
Dominant sets and oriented exchange graphs. Two seeds of a cluster algebra are said to be equivalent, if they are the same up to a relabeling. The exchange graph $\mathcal{H}=\mathcal{H}(\mathcal{A})$ of a cluster algebra $\mathcal{A}$ is the graph whose vertices correspond to the seeds (up to seed equivalence) of $\mathcal{A}$ and whose edges correspond to seed mutations. Thanks to the sign-coherence of $C$-matrices [GHKK18], each mutation has two states: green mutation or red mutation. Fix an initial seed $({\bf x}_{t_{0}},\widetilde{B}_{t_{0}})$, the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}$ of $\mathcal{A}$ is the quiver whose underlying graph is the exchange graph $\mathcal{H}$ of $\mathcal{A}$ and whose orientation is induced by the green mutations.
$\tau$-tilting theory is introduced by Adachi, Iyama and Reiten [air_2014] as a far reaching generalization of cluster-tilting theory (one of the first major developments in representation theory discovered thanks to the relation with cluster algebras). Various of the fundamental concepts in cluster algebras, such as clusters/seeds, mutations, $C$-matrices, $G$-matrices, (oriented) exchange graphs, etc., were generalized to $\tau$-tilting theory.
The seeds in cluster algebras correspond to (basic) $\tau$-tilting pairs in $\tau$-tilting theory, which are in bijection with the functorially finite torsion classes in the module categories. The inclusion of subcategories induces a natural partial order on the set of functorially finite torsion classes and thus a partial order on the set of $\tau$-tilting pairs (up to isomorphisms).
Thanks to the existence of the partial order, the oriented exchange graphs in $\tau$-tilting theory are acyclic. Actually, Adachi, Iyama and Reiten [air_2014, Theorem 0.6] proved that the oriented exchange graphs in $\tau$-tilting theory coincide with the Hasse quivers of the posets of functorially finite torsion classes, equivalently, the posets of $\tau$-tilting pairs.
It is natural to ask whether the oriented exchange graphs of (skew-symmetrizable) cluster algebras are acyclic and whether these oriented exchange graphs coincide with Hasse quivers of some posets.
Inspired by $F$-invariant, we introduce the dominant sets for seeds of cluster algebras as an replacement of torsion classes for $\tau$-tilting pairs in $\tau$-tilting theory and as an replacement of inversion sets in the study of right weak order on Weyl groups. With the help of the dominant sets, we prove the following results on oriented exchange graphs.
Theorem 1.6 (Theorems 5.7, 5.12, Corollary 5.13).
The following statements hold.
  • (i)
    The oriented exchange graphs of cluster algebras are acyclic. In particular, green mutations induce a partial order on the set of seeds (up to seed equivalence) of cluster algebras.
  • (ii)
    The oriented exchange graphs of cluster algebras coincide with the Hasse quivers of the posets of seeds of cluster algebras.
Remark 1.7.
When I was a PhD student (2014-2019), I was interested in the above problems. But at that time, I was not able to prove them, because I did not know what is the right replacements of torsion classes in cluster algebras.
Finally, we would like to point out that by ignoring initial cluster variables, the dominant sets in cluster algebras correspond to the sets of indecomposable $\tau$-rigid modules contained in the functorially finite torsion classes (see Proposition 5.21). In this sense, we view the dominant sets as an replacement of torsion classes in $\tau$-titling theory and please see Section 5.4 for details. By ignoring initial cluster variables, the dominant sets also correspond to the inversion sets in the study of right weak order on Weyl groups (see Proposition 5.30).
Organization of the paper. In Section 2, we recall some well-known concepts in cluster algebras, including $Y$-patterns, cluster patterns, compatible pairs, and (upper) cluster algebras. In Section 3, we recall the definition of tropical points for $Y$-patterns (equivalently, cluster $\mathscr{X}$-varieties) and for cluster patterns (equivalently, cluster $\mathscr{A}$-varieties), and we also recall the (compatibly) pointed elements introduced by Qin [Qin_2017, qin_2019]. We introduce the notion of good elements in upper cluster algebras. The $F$-invariant will be defined for a pair of good elements.
In Section 4.1, we give two maps between the set of tropical points. Let $(\mathcal{S}_{X},\mathcal{S}_{Y})$ be a Langlands dual pair in the sense of Definition 2.6. The first map is from the set of tropical points associated to the cluster pattern $\mathcal{S}_{X}$ to the set of tropical points associated to the $Y$-pattern $\mathcal{S}_{Y}$ and the second map is from the set of tropical points associated to the $Y$-pattern $\mathcal{S}_{Y}$ to the set of tropical points associated to the cluster pattern $\mathcal{S}_{X}$. We need to emphasise that the second map will play an important role in the proof Theorem 4.8 which says that the pairing $\langle u,u^{\prime}\rangle_{w}$ is independent of the choice of seed $({\bf x}_{w},\widetilde{B}_{w})$.
In Section 4.2, we give the definition of the tropical invariant and $F$-invariant. The main results in this subsection are Theorem 4.8, Theorem 4.22 and Theorem 4.27. In Section 4.3, we discuss the link between the compatible Poisson structure of (upper) cluster algebras and the $F$-invariant and the main result in this subsection is Theorem 4.29.
In Section 5, we introduce the dominant sets for the seeds of cluster algebras as an replacement of torsion classes for $\tau$-tilting pairs in $\tau$-tilting theory and as an replacement of inversion sets in the study of right weak order on Weyl groups. Then we prove the results on oriented exchange graphs. The main results in this section are Theorems 5.7, 5.12 and Propositions 5.21, 5.30.
2. Preliminaries on cluster algebras
Throughout this paper, we fix a pair $(n,m)$ of integers with $m\geq n>0$. For any positive integer $r$, we denote $[1,r]:=\{1,\ldots,r\}$. The matrices in this paper are always integer matrices.
2.1. Mutation matrices and their mutations
Recall that an $n\times n$ integer matrix $B$ is said to be skew-symmetrizable, if there exists a diagonal integer matrix $S=diag(s_{1},\ldots,s_{n})$ with $s_{i}>0$ ($i\in[1,n]$) such that $SB$ is skew-symmetric. Such a diagonal matrix $S$ is called a skew-symmetrizer of $B$.
An $m\times n$ integer matrix $\widetilde{B}=\begin{bmatrix}B\\ P\end{bmatrix}=(b_{ij})$ is called a mutation matrix, if its top $n\times n$ submatrix $B$ is skew-symmetrizable. The submatrix $B$ is called the principal part of $\widetilde{B}$.
Let $\widetilde{B}=\begin{bmatrix}B\\ P\end{bmatrix}=(b_{ij})$ be an $m\times n$ mutation matrix. The mutation of $\widetilde{B}$ in direction $k\in[1,n]$ is defined to be the new integer matrix $\mu_{k}(\widetilde{B})=\widetilde{B}^{\prime}=\begin{bmatrix}B^{\prime}\\ P^{\prime}\end{bmatrix}=(b_{ij}^{\prime})$ given by
(2.1) $b_{ij}^{\prime}$ $=$ $\begin{cases}-b_{ij},&\text{if}\;i=k\;\text{or}\;j=k;\\ b_{ij}+[b_{ik}]_{+}[b_{kj}]_{+}-[-b_{ik}]_{+}[-b_{kj}]_{+},&\text{otherwise},\end{cases}$
where $[a]_{+}:=\max\{a,0\}$ for any $a\in\mathbb{R}$.
The following statements are well-known, c.f., [fz_2002, Proposition 4.5], [bfz_2005, Lemma 3.2].
  • $\mu_{k}(\widetilde{B})=\widetilde{B}^{\prime}$ is still a mutation matrix, namely, the submatrix $B^{\prime}$ of $\widetilde{B}^{\prime}$ is skew-symmetrizable;
  • $B$ and $B^{\prime}$ share the same skew-symmetrizer;
  • $\mu_{k}$ is an involution, i.e., $\mu_{k}^{2}=id$;
  • The rank of $\widetilde{B}^{\prime}=\mu_{k}(\widetilde{B})$ is equal to the rank of $\widetilde{B}$.
As shown in [bfz_2005, (3.2)], the matrix $\widetilde{B}^{\prime}=\mu_{k}(\widetilde{B})$ can be written as
(2.2) $\widetilde{B}^{\prime}=\mu_{k}(\widetilde{B})=E_{k,\varepsilon}^{\widetilde{B}}\widetilde{B}F_{k,\varepsilon}^{\widetilde{B}},$
where $\varepsilon\in\{\pm\}$ and
  • $E_{k,\varepsilon}^{\widetilde{B}}$ is the $m\times m$ matrix obtained by replacing the $k$th column of $I_{m}$ by the column vector $(e_{1},\ldots,e_{m})^{T}$, where
    $e_{i}=\begin{cases}-1,&\text{if }i=k;\\ [-\varepsilon b_{ik}]_{+},&\text{if }i\neq k.\end{cases}$
  • $F_{k,\varepsilon}^{\widetilde{B}}$ is the $n\times n$ matrix obtained by replacing the $k$th row of $I_{n}$ by the row vector $(f_{1},\ldots,f_{n})$, where
    $f_{i}=\begin{cases}-1,&\text{if }i=k;\\ [\varepsilon b_{ki}]_{+},&\text{if }i\neq k.\end{cases}$
It turns out that $\widetilde{B}^{\prime}$ is independent of the choice of a sign $\varepsilon\in\{\pm\}$ in (2.2). The matrices $E_{k,\varepsilon}^{\widetilde{B}}$ and $F_{k,\varepsilon}^{\widetilde{B}}$ defined above coincide with the ones used in [bz-2005, (3.2), (3.3)]. It is easy to check
(2.3) $(E_{k,\varepsilon}^{\widetilde{B}})^{2}=I_{m}\;\;\;\text{and}\;\;\;(F_{k,\varepsilon}^{\widetilde{B}})^{2}=I_{n}.$
2.2. Compatible pairs and their mutations
Let $\Lambda=(\lambda_{ij})$ be an $m\times m$ skew-symmetric integer matrix, and let $\widetilde{B}=\begin{bmatrix}B\\ P\end{bmatrix}=(b_{ij})$ be an $m\times n$ integer matrix, where $B$ is the top $n\times n$ submatrix of $\widetilde{B}$. Here we do not assume that $B$ is skew-symmetrizable.
Definition 2.1 (Compatible pair, [bz-2005]).
Keep $\Lambda$ and $\widetilde{B}$ as above. Let $S=diag(s_{1},\ldots,s_{n})$ be a diagonal integer matrix with $s_{j}>0,\;j\in[1,n]$. The pair $(\widetilde{B},\Lambda)$ is called a compatible pair of type $S$, if for any $j\in[1,n]$ and $i\in[1,m]$, we have
(2.4) $\sum_{k=1}^{m}b_{kj}\lambda_{ki}=\delta_{i,j}s_{j}.$
In other words, $\widetilde{B}^{T}\Lambda=(S\mid{\bf 0})$, where ${\bf 0}$ is the $n\times(m-n)$ zero matrix.
Notice that the diagonal matrix $S$ is uniquely determined by the compatible pair $(\widetilde{B},\Lambda)$. The matrix $\Lambda$ in a compatible pair $(\widetilde{B},\Lambda)$ is called a quantization matrix.
Proposition 2.2.
[bz-2005, Proposition 3.3] Let $(\widetilde{B},\Lambda)$ be a compatible pair of type $S$ and $B$ the top $n\times n$ submatrix of $\widetilde{B}$. Then the matrix $\widetilde{B}$ has the full rank $n$ and $SB=\widetilde{B}^{T}\Lambda\widetilde{B}$. In particular, $B$ is skew-symmetrizable.
Conversely, by [gsv-2003, Theorem 1.4], whenever $\widetilde{B}$ is a mutation matrix of full rank, we can find a skew-symmetric integer matrix $\Lambda$ such that $(\widetilde{B},\Lambda)$ is a compatible pair.
Let $(\widetilde{B},\Lambda)$ be a compatible pair of type $S$ and $k\in[1,n]$. We set $\widetilde{B}^{\prime}:=\mu_{k}(\widetilde{B})=E_{k,\varepsilon}^{\widetilde{B}}\widetilde{B}F_{k,\varepsilon}^{\widetilde{B}}$ and
$\Lambda^{\prime}:=(E_{k,\varepsilon}^{\widetilde{B}})^{T}\Lambda E_{k,\varepsilon}^{\widetilde{B}}.$
Proposition 2.3.
[bz-2005, Proposition 3.4] (i) The matrix $\Lambda^{\prime}$ is independent of the choice of $\varepsilon\in\{\pm\}$.
(ii) The pair $(\widetilde{B}^{\prime},\Lambda^{\prime})$ is still a compatible pair of type $S$.
Keep the above setting. We call $(\widetilde{B}^{\prime},\Lambda^{\prime})$ the mutation of the compatible pair $(\widetilde{B},\Lambda)$ in direction $k$ and denote $(\widetilde{B}^{\prime},\Lambda^{\prime})=\mu_{k}(\widetilde{B},\Lambda)$. It is known from [bz-2005, Proposition 3.6] that $\mu_{k}(\widetilde{B}^{\prime},\Lambda^{\prime})=(\widetilde{B},\Lambda)$, i.e., $\mu_{k}$ is an involution.
The following proposition will be not used in this paper, but it is very useful to construct many examples of compatible pairs from a given skew-symmetrizable matrix.
Proposition 2.4.
[zelevinsky-2005, Example 0.5] Let $B$ be an $n\times n$ skew-symmetrizable integer matrix and set $\widetilde{B}=\begin{bmatrix}B\\ I_{n}\end{bmatrix}$. Then the skew-symmetric matrices $\Lambda$ compatible with $\widetilde{B}$ in the sense of (2.4) are those of the form
$\Lambda=\begin{bmatrix}\Lambda_{0}&-S-\Lambda_{0}B\\ S-B^{T}\Lambda_{0}&-SB+B^{T}\Lambda_{0}B\end{bmatrix},$
where $\Lambda_{0}$ is an arbitrary skew-symmetric integer $n\times n$ matrix and $S$ is an arbitrary skew-symmetrizer of $B$.
2.3. $Y$-pattern and cluster pattern
Recall that we fixed a pair $(n,m)$ of integers with $m\geq n>0$. Let $\mathbb{F}$ be the field of rational functions over $\mathbb{Q}$ in $m$ variables.
A $Y$-seed of rank $n$ in $\mathbb{F}$ is a pair $({\bf y},\widehat{B})$, where
  • ${\bf y}=(y_{1},\ldots,y_{m})$ satisfies that $\{y_{1},\ldots,y_{m}\}$ is a free generating set of $\mathbb{F}$ over $\mathbb{Q}$;
  • $\widehat{B}=(B\mid Q)=(\hat{b}_{ij})$ is an $n\times m$ integer matrix such that its leftmost $n\times n$ submatrix $B$ is skew-symmetrizable.
The variables $y_{1},\ldots,y_{m}$ are called the $y$-variables of $({\bf y},\widehat{B})$.
Let $({\bf y},\widehat{B})$ be a $Y$-seed of rank $n$ in $\mathbb{F}$. The mutation of $({\bf y},\widehat{B})$ in direction $k\in[1,n]$ is the pair $({\bf y}^{\prime},\widehat{B}^{\prime}):=\mu_{k}({\bf y},\widehat{B})$ given as follows:
(2.5) $y_{i}^{\prime}$ $=$ $\begin{cases}y_{k}^{-1},&\text{if}\;i=k;\\ y_{i}y_{k}^{[\hat{b}_{ki}]_{+}}(1+y_{k})^{-\hat{b}_{ki}},&\text{otherwise}.\end{cases}$
$\hat{b}_{ij}^{\prime}$ $=$ $\begin{cases}-\hat{b}_{ij},&\text{if}\;i=k\;\text{or}\;j=k;\\ \hat{b}_{ij}+[\hat{b}_{ik}]_{+}[\hat{b}_{kj}]_{+}-[-\hat{b}_{ik}]_{+}[-\hat{b}_{kj}]_{+},&\text{otherwise}.\end{cases}$
One can check that the new pair $({\bf y}^{\prime},\widehat{B}^{\prime})=\mu_{k}({\bf y},\widehat{B})$ is still a $Y$-seed of rank $n$ and $({\bf y},\widehat{B})=\mu_{k}({\bf y}^{\prime},\widehat{B}^{\prime})$.
A cluster seed or simply a seed of rank $n$ in $\mathbb{F}$ is a pair $({\bf x},\widetilde{B})$, where
  • ${\bf x}=(x_{1},\ldots,x_{m})$ satisfies that $\{x_{1},\ldots,x_{m}\}$ is a free generating set of $\mathbb{F}$ over $\mathbb{Q}$;
  • $\widetilde{B}=\begin{bmatrix}B\\ P\end{bmatrix}=(b_{ij})$ is an $m\times n$ mutation matrix.
In this case, the tuple ${\bf x}$ is called the cluster of $({\bf x},\widetilde{B})$. Elements in ${\bf x}$ are called cluster variables. More precisely, we call $x_{1},\ldots,x_{n}$ unfrozen cluster variables and $x_{n+1},\ldots,x_{m}$ frozen cluster variables. The matrices $B$ and $P$ are respectively called the exchange matrix, coefficient matrix of $({\bf x},\widetilde{B})$. We denote by $\widehat{\bf y}:=(\hat{y}_{1},\ldots,\hat{y}_{n})$, where $\hat{y}_{k}:=\prod_{j=1}^{m}x_{j}^{b_{jk}}$. The variables $\hat{y}_{1},\ldots,\hat{y}_{n}$ are called $\hat{y}$-variables associated to $({\bf x},\widetilde{B})$.
Let $({\bf x},\widetilde{B})$ be a seed of rank $n$ in $\mathbb{F}$. The mutation of $({\bf x},\widetilde{B})$ in direction $k\in[1,n]$ is the pair $({\bf x}^{\prime},\widetilde{B}^{\prime})=\mu_{k}({\bf x},\widetilde{B})$ given by $\widetilde{B}^{\prime}=\mu_{k}(\widetilde{B})$ and
(2.6) $x_{i}^{\prime}=\begin{cases}x_{i},&\text{if}\;i\neq k;\\ x_{k}^{-1}\cdot(\prod_{j=1}^{m}x_{j}^{[b_{jk}]_{+}}+\prod_{j=1}^{m}x_{j}^{[-b_{jk}]_{+}}),&\text{if}\;i=k.\end{cases}$
One can check that the new pair $({\bf x}^{\prime},\widetilde{B}^{\prime})=\mu_{k}({\bf x},\widetilde{B})$ is still a seed of rank $n$ and $({\bf x},\widetilde{B})=\mu_{k}({\bf x}^{\prime},\widetilde{B}^{\prime})$.
Let $\mathbb{T}_{n}$ denote the $n$-regular tree. We label the edges of $\mathbb{T}_{n}$ by $1,\ldots,n$ such that the $n$ different edges adjacent to the same vertex of $\mathbb{T}_{n}$ receive different labels.
Definition 2.5.
(i) A $Y$-pattern $\mathcal{S}_{Y}=\{({\bf y}_{t},\widehat{B}_{t})\mid t\in\mathbb{T}_{n}\}$ of rank $n$ is an assignment of a $Y$-seed $({\bf y}_{t},\widehat{B}_{t})$ of rank $n$ to every vertex $t$ of $\mathbb{T}_{n}$ such that $({\bf y}_{t^{\prime}},\widehat{B}_{t^{\prime}})=\mu_{k}({\bf y}_{t},\widehat{B}_{t})$ whenever ${t}$${t^{\prime}}$$\scriptstyle{k}$ in $\mathbb{T}_{n}$.
(ii) A cluster pattern $\mathcal{S}_{X}=\{({\bf x}_{t},\widetilde{B}_{t})\mid t\in\mathbb{T}_{n}\}$ of rank $n$ is an assignment of a cluster seed $({\bf x}_{t},\widetilde{B}_{t})$ of rank $n$ to every vertex $t$ of $\mathbb{T}_{n}$ such that $({\bf x}_{t^{\prime}},\widetilde{B}_{t^{\prime}})=\mu_{k}({\bf x}_{t},\widetilde{B}_{t})$ whenever ${t}$${t^{\prime}}$$\scriptstyle{k}$ in $\mathbb{T}_{n}$.
We usually write ${\bf y}_{t}=(y_{1;t},\ldots,y_{m;t})$, $\widehat{B}_{t}=(B_{t}\mid Q_{t})=(\hat{b}_{ij;t})$, ${\bf x}_{t}=(x_{1;t},\ldots,x_{m;t})$ and
$\widetilde{B}_{t}=\begin{pmatrix}B_{t}\\ P_{t}\end{pmatrix}=(b_{ij;t}).$
We write $\widehat{\bf y}_{t}=(\hat{y}_{1;t},\ldots,\hat{y}_{n;t})$ for the collection of $\hat{y}$-variables associated to seed $({\bf x}_{t},\widetilde{B}_{t})$, where
$\hat{y}_{k;t}=\prod_{j=1}^{m}x_{j;t}^{b_{jk;t}}.$
In this paper, we usually fix a vertex $t_{0}$ as the rooted vertex of the $n$-regular tree $\mathbb{T}_{n}$. Clearly, both $Y$-pattern and cluster pattern are uniquely determined by the data at the rooted vertex $t_{0}$.
Definition 2.6.
Let $\mathcal{S}_{X}=\{({\bf x}_{t},\widetilde{B}_{t})\mid t\in\mathbb{T}_{n}\}$ be a cluster pattern, $\mathcal{S}_{Y}=\{({\bf y}_{t},\widehat{B}_{t})\mid t\in\mathbb{T}_{n}\}$ a $Y$-pattern, and let $\Lambda=\{\Lambda_{t}\mid t\in\mathbb{T}_{n}\}$ be a collection of $m\times m$ skew-symmetric integer matrices indexed by the vertices in $\mathbb{T}_{n}$.
  • (i)
    The pair $(\mathcal{S}_{X},\mathcal{S}_{Y})$ is called a Langlands dual pair, if $\widehat{B}_{t_{0}}=-\widetilde{B}_{t_{0}}^{T}$ holds for the rooted vertex $t_{0}$.
  • (ii)
    The triple $(\mathcal{S}_{X},\mathcal{S}_{Y},\Lambda)$ is called a Langlands-Poisson triple, if it satisfies that
    • $(\mathcal{S}_{X},\mathcal{S}_{Y})$ forms a Langlands dual pair;
    • $\{(\widetilde{B}_{t},\Lambda_{t})\mid t\in\mathbb{T}_{n}\}$ forms a collection of compatible pairs and $(\widetilde{B}_{t^{\prime}},\Lambda_{t^{\prime}})=\mu_{k}(\widetilde{B}_{t},\Lambda_{t})$ whenever ${t}$${t^{\prime}}$$\scriptstyle{k}$ in $\mathbb{T}_{n}$.
The above definition is purely for convenience of this paper. The following result can be checked easily.
Corollary 2.7.
Let $(\mathcal{S}_{X},\mathcal{S}_{Y})$ be a Langlands dual pair. Then $\widehat{B}_{t}=-\widetilde{B}_{t}^{T}$ holds for any vertex $t$ of $\mathbb{T}_{n}$.
[…]
$\{x_{i;t_{0}},x_{j;t_{0}}\}:=\lambda_{ij;t_{0}}\cdot x_{i;t_{0}}x_{j;t_{0}}.$
It turns out this Poisson bracket is compatible with the cluster pattern $\mathcal{S}_{X}$, that is, for any cluster ${\bf x}_{t}$ of $\mathcal{S}_{X}$, we have
$\{x_{i;t},x_{j;t}\}=\lambda_{ij;t}\cdot x_{i;t}x_{j;t},$
where $\lambda_{ij;t}$ is the $(i,j)$-entry of $\Lambda_{t}$, cf. [gsv-2003], [bz-2005, Remark 4.6].
Finally, let us give the definition of (upper) cluster algebras.
Definition 2.8.
Let $\mathcal{S}_{X}=\{({\bf x}_{t},\widetilde{B}_{t})\mid t\in\mathbb{T}_{n}\}$ be a cluster pattern of rank $n$ in $\mathbb{F}$.
  • (i)
    The cluster algebra $\mathcal{A}$ associated to $\mathcal{S}_{X}$ is the $\mathbb{Z}$-subalgebra of $\mathbb{F}$ given by
    $\mathcal{A}:=\mathbb{Z}[x_{1;t},\ldots,x_{n;t},x_{n+1;t}^{\pm 1},\ldots,x_{m;t}^{\pm 1}\mid t\in\mathbb{T}_{n}].$
  • (ii)
    The upper cluster algebra $\mathcal{U}$ associated to $\mathcal{S}_{X}$ is the $\mathbb{Z}$-subalgebra of $\mathbb{F}$ defined by
    $\mathcal{U}:=\bigcap\limits_{t\in\mathbb{T}_{n}}\mathcal{L}(t),$
    where $\mathcal{L}(t):=\mathbb{Z}[x_{1;t}^{\pm 1},\ldots,x_{m;t}^{\pm 1}]$.
Theorem 2.9.
[fz_2002]*Laurent phenomenon For any cluster variable $z$ and any seed $({\bf x}_{t},\widetilde{B}_{t})$ of $\mathcal{A}$, we have $z\in\mathbb{Z}[x_{1;t}^{\pm 1},\ldots,x_{n;t}^{\pm 1},x_{n+1;t},\ldots,x_{m;t}]$. In particular, $\mathcal{A}$ is contained in $\mathcal{U}$.
By the Laurent phenomenon, any cluster variable $z$ can be uniquely written as
$z=\frac{N(x_{1;t},\ldots,x_{n;t})}{x_{1;t}^{d_{1}^{t}}\ldots x_{n;t}^{d_{n}^{t}}},$
where $N$ is a polynomial in $x_{1;t},\ldots,x_{n;t}$ with coefficents in $\mathbb{Z}[x_{n+1;t}^{\pm 1},\ldots,x_{m;t}^{\pm 1}]$ with $x_{i;t}\nmid N$ for $i\in[1,n]$, and each $d_{i}^{t}$ is an integer. The vector ${\bf d}_{z}^{t}:=(d_{1}^{t},\ldots,d_{n}^{t})\in\mathbb{Z}^{n}$ is called the $d$-vector of $z$ with respect to seed at the vertex $t$.
Example 2.10.
Take $\widetilde{B}=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$ and ${\bf x}=(x_{1},x_{2})$. One can check that the cluster algebra $\mathcal{A}$ defined by the initial seed $({\bf x},\widetilde{B})$ has only five cluster variables:
$x_{1},\;x_{2},\;x_{3}:=\frac{x_{2}+1}{x_{1}},\;x_{4}:=\frac{x_{1}+x_{2}+1}{x_{1}x_{2}},\;x_{5}:=\frac{x_{1}+1}{x_{2}}.$
It can be seen that all the cluster variables are in $\mathbb{Z}[x_{1}^{\pm 1},x_{2}^{\pm 1}]$. The $d$-vectors with respect to the initial seed $({\bf x},\widetilde{B})$ are as follows:
${\bf d}_{x_{1}}=\begin{bmatrix}-1\\ 0\end{bmatrix},\;\;{\bf d}_{x_{2}}=\begin{bmatrix}0\\ -1\end{bmatrix},\;\;{\bf d}_{x_{3}}=\begin{bmatrix}1\\ 0\end{bmatrix},\;\;{\bf d}_{x_{4}}=\begin{bmatrix}1\\ 1\end{bmatrix},\;\;{\bf d}_{x_{5}}=\begin{bmatrix}0\\ 1\end{bmatrix}.$
3. Tropical points, pointed elements and good elements
3.1. Universal semi-field and tropical points
Recall that $(\mathbb{P},\cdot,\oplus)$ is called a semi-field if $(\mathbb{P},\cdot)$ is an abelian multiplicative group endowed with a binary operation of auxiliary addition $\oplus$ which is commutative, associative and satisfies that the multiplication distributes over the auxiliary addition. For example,
$\mathbb{Z}^{\rm min}:=(\mathbb{Z},+,\min\{-,-\})\;\;\;\text{and}\;\;\;\mathbb{Z}^{\rm max}:=(\mathbb{Z},+,\max\{-,-\})$
are semi-field and they are called tropical semi-field.
Let $\mathbb{Q}_{\rm sf}(u_{1},\ldots,u_{m})$ be the set of all non-zero rational functions in $u_{1},\ldots,u_{m}$ that have subtraction free expressions. The set $\mathbb{Q}_{\rm sf}(u_{1},\ldots,u_{m})$ is a semifield with respect to the usual operations of multiplication and addition. It is called an universal semi-field.
From now on, we fix $\mathbb{F}:=\mathbb{Q}(u_{1},\ldots,u_{m})$ and $\mathbb{F}_{>0}:=\mathbb{Q}_{\rm sf}(u_{1},\ldots,u_{m})$. For any semi-field $\mathbb{P}$ and ${\bf p}=(p_{1},\ldots,p_{m})^{T}\in\mathbb{P}^{m}$, there exists a unique semi-field homomorphism
${\pi}_{\bf p}:\mathbb{F}_{>0}\rightarrow\mathbb{P}$
induced by $u_{i}\mapsto p_{i}$ for any $i$. The map $\pi:{\bf p}\mapsto\pi_{\bf p}$ induces a bijection from $\mathbb{P}^{m}$ to the set $\operatorname{\mathsf{Hom}}_{\rm sf}(\mathbb{F}_{>0},\mathbb{P})$ of semifield homomorphisms from $\mathbb{F}_{>0}$ to $\mathbb{P}$.
In this paper, we often take $\mathbb{P}=\mathbb{Z}^{\rm max}=(\mathbb{Z},+,\max)$. The elements in $\operatorname{\mathsf{Hom}}_{\rm sf}(\mathbb{F}_{>0},\mathbb{Z}^{\rm max})$ are called tropical $\mathbb{Z}^{\rm max}$-points or simply tropical points, which are in bijection with the points in $\mathbb{Z}^{m}$.
Definition 3.1.
A chart in $\mathbb{F}_{>0}=\mathbb{Q}_{\rm sf}(u_{1},\ldots,u_{m})$ is an $m$-tuple ${\bf z}=(z_{1},\ldots,z_{m})$ of elements in $\mathbb{F}_{>0}$ that generate $\mathbb{F}_{>0}$ as a semifield.
Proposition 3.2.
Let $\mathcal{C}=\{{\bf u}_{t}\mid t\in\mathbb{T}\}$ be a collection of charts in $\mathbb{F}_{>0}=\mathbb{Q}_{\rm sf}(u_{1},\ldots,u_{m})$ indexed by any set $\mathbb{T}$. Then the following statements hold.
  • (i)
    For any chart ${\bf u}_{t}=(u_{1;t},\ldots,u_{m;t})$, denote by $u_{i}({\bf u}_{t})$ the expression of $u_{i}$ as an element in $\mathbb{F}(t)_{>0}:=\mathbb{Q}_{\rm sf}(u_{1;t},\ldots,u_{m;t})$. The map $\varphi_{t}:u_{i}\mapsto u_{i}({\bf u}_{t})$ induces an isomorphism of semifield from $\mathbb{F}_{>0}$ to $\mathbb{F}(t)_{>0}$.
  • (ii)
    For any chart ${\bf u}_{t}=(u_{1;t},\ldots,u_{m;t})$ and any $\beta\in\operatorname{\mathsf{Hom}}_{\rm sf}(\mathbb{F}_{>0},\mathbb{Z}^{\rm max})$, denote by
    ${\bf q}^{t}(\beta):=(\beta(u_{1;t}),\ldots,\beta(u_{m;t}))^{T}\in\mathbb{Z}^{m}.$
    Then the following diagram commutes.
    ${\mathbb{F}_{>0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi_{t}\;\;\;}$$\scriptstyle{\beta}$${\mathbb{F}(t)_{>0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{{\bf q}^{t}(\beta);t}}$${\mathbb{Z}^{\rm max}}$
    where $\pi_{{\bf q}^{t}(\beta);t}$ is the unique semifield homomorphism from $\mathbb{F}(t)_{>0}$ to $\mathbb{Z}^{\rm max}$ induced by $u_{i;t}\mapsto\beta(u_{i;t})$ for any $i$.
  • (iii)
    The map $\beta\mapsto{\bf q}^{t}(\beta)$ induces a bijection from the tropical points in $\operatorname{\mathsf{Hom}}_{\rm sf}(\mathbb{F}_{>0},\mathbb{Z}^{\rm max})$ and the points in $\mathbb{Z}^{m}$.
Proof.
Obviously. ∎
Keep the setting in the above proposition. We call the vector
${\bf q}^{t}(\beta)=(\beta(u_{1;t}),\ldots,\beta(u_{m;t}))^{T}\in\mathbb{Z}^{m}$
the coordinate vector of the tropical point $\beta$ under the chart ${\bf u}_{t}$. When a collection $\mathcal{C}=\{{\bf u}_{t}\mid t\in\mathbb{T}\}$ of charts in $\mathbb{F}_{>0}$ is given, we often identify a tropical point $\beta\in\operatorname{\mathsf{Hom}}_{\rm sf}(\mathbb{F}_{>0},\mathbb{Z}^{\rm max})$ with the collection $\{{\bf q}^{t}(\beta)\in\mathbb{Z}^{m}\mid t\in\mathbb{T}\}$ of coordinate vectors of $\beta$ under the charts.
Now let us go back to the situations that we are interested in.
  • If we work on a $Y$-pattern $\mathcal{S}_{Y}=\{({\bf y}_{t},\widehat{B}_{t})\mid t\in\mathbb{T}_{n}\}$, we take $\mathbb{F}_{>0}$ to be the universal semifield generated by the initial $y$-variables $y_{1;t_{0}},\ldots,y_{m;t_{0}}$. Then $\mathcal{C}:=\{{\bf y}_{t}\mid t\in\mathbb{T}_{n}\}$ is a collection of charts in $\mathbb{F}_{>0}$. We usually denote by
    $\{{\bf g}^{t}(\beta)\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$
    the collection of coordinate vectors of a tropical point $\beta\in\operatorname{\mathsf{Hom}}_{\rm sf}(\mathbb{F}_{>0},\mathbb{Z}^{\rm max})$ for the $Y$-pattern case.
  • If we work on a cluster pattern $\mathcal{S}_{X}=\{({\bf x}_{t},\widetilde{B}_{t})\mid t\in\mathbb{T}_{n}\}$, we take $\mathbb{F}_{>0}$ to be the universal semifield generated by the initial cluster variables $x_{1;t_{0}},\ldots,x_{m;t_{0}}$. Then $\mathcal{C}:=\{{\bf x}_{t}\mid t\in\mathbb{T}_{n}\}$ is a collection of charts in $\mathbb{F}_{>0}$. We usually denote by
    $\{{\bf a}^{t}(\beta)\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$
    the collection of coordinate vectors of a tropical point $\beta\in\operatorname{\mathsf{Hom}}_{\rm sf}(\mathbb{F}_{>0},\mathbb{Z}^{\rm max})$ for the cluster pattern case.
Since the transition maps between any two adjacent charts are clear in both $Y$-pattern and cluster pattern, the corresponding tropical points (after identifying with the corresponding coordinate vectors) can be defined using the tropical version of the transition maps in (2.5) and (2.6).
Definition 3.3 (Tropical points).
Let $\mathcal{S}_{Y}=\{({\bf y}_{t},\widehat{B}_{t})\mid t\in\mathbb{T}_{n}\}$ be a $Y$-pattern of rank $n$ in $\mathbb{F}$ and $\mathcal{S}_{X}=\{({\bf x}_{t},\widetilde{B}_{t})\mid t\in\mathbb{T}_{n}\}$ a cluster pattern of rank $n$ in $\mathbb{F}$.
  • (i)
    A tropical point $[{\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ associated to the $Y$-pattern $\mathcal{S}_{Y}$ is an assignment of a (column) vector ${\bf g}^{t}=(g_{1}^{t},\ldots,g_{m}^{t})^{T}$ in $\mathbb{Z}^{m}$ to each vertex $t$ of $\mathbb{T}_{n}$ such that
    (3.1) $g_{i}^{t^{\prime}}=\begin{cases}-g_{k}^{t},&\text{if}\;i=k;\\ g_{i}^{t}+[\hat{b}_{ki;t}]_{+}g_{k}^{t}+(-\hat{b}_{ki;t})[g_{k}^{t}]_{+},&\text{if}\;i\neq k.\end{cases}$
    whenever ${t}$${t^{\prime}}$$\scriptstyle{k}$ in $\mathbb{T}_{n}$. We denote by $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ the set of tropical points associated to the $Y$-pattern $\mathcal{S}_{Y}$.
  • (ii)
    A tropical point $[{\bf a}]=\{{\bf a}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ associated to the cluster pattern $\mathcal{S}_{X}$ is an assignment of a (column) vector ${\bf a}^{t}=(a_{1}^{t},\ldots,a_{m}^{t})^{T}$ in $\mathbb{Z}^{m}$ to each vertex $t$ of $\mathbb{T}_{n}$ such that
    (3.2) $a_{i}^{t^{\prime}}=\begin{cases}-a_{k}^{t}+\max\{\sum_{j=1}^{m}[b_{jk;t}]_{+}a_{j}^{t},\;\sum_{j=1}^{m}[-b_{jk;t}]_{+}a_{j}^{t}\},&\text{if}\;i=k;\\ a_{i}^{t},&\text{if}\;i\neq k.\end{cases}$
    whenever ${t}$${t^{\prime}}$$\scriptstyle{k}$ in $\mathbb{T}_{n}$. We denote by $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$ the set of tropical points associated to the cluster pattern $\mathcal{S}_{X}$.
Remark 3.4.
The relations in (3.1) and (3.2) are obtained from mutation relations (2.5) and (2.6) by tropicalization over the tropical semifield $\mathbb{Z}^{\rm max}=(\mathbb{Z},+,\max\{-,-\})$. Namely, we make the following replacements.
  • Replace the multiplication and addition in (2.5) and (2.6) by “+” and “max{–,–}” over $\mathbb{Z}$, respectively.
  • Replace “1” in (2.5) by “0”;
The following corollary is obvious.
Corollary 3.5.
The following statements hold.
  • (i)
    Each tropical point $[{\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ corresponds to a unique semifield homomorphism $\beta\in\operatorname{\mathsf{Hom}}_{\rm sf}(\mathbb{F}_{>0},\mathbb{Z}^{\rm max})$ such that
    $\beta({\bf y}_{t})=(\beta(y_{1;t}),\ldots,\beta(y_{m;t}))=({\bf g}^{t})^{T}$
    for any vertex $t\in\mathbb{T}$, where $\mathbb{F}_{>0}=\mathbb{Q}_{\rm sf}(y_{1;t_{0}},\ldots,y_{m;t_{0}})$.
  • (ii)
    Each tropical point $[{\bf a}]=\{{\bf a}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$ corresponds to a unique semifield homomorphism $\beta\in\operatorname{\mathsf{Hom}}_{\rm sf}(\mathbb{F}_{>0},\mathbb{Z}^{\rm max})$ such that
    $\beta({\bf x}_{t})=(\beta(x_{1;t}),\ldots,\beta(x_{m;t}))=({\bf a}^{t})^{T}$
    for any vertex $t\in\mathbb{T}$, where $\mathbb{F}_{>0}=\mathbb{Q}_{\rm sf}(x_{1;t_{0}},\ldots,x_{m;t_{0}})$.
3.2. Compatibly pointed elements and good elements
In this subsection, we fix a full rank upper cluster algebra $\mathcal{U}$, which means its initial mutation matrix $\widetilde{B}_{t_{0}}$ has the full rank $n$, equivalently, any mutation matrix $\widetilde{B}_{t}$ of $\mathcal{U}$ has the full rank $n$.
Since $\mathcal{U}$ is of full rank, each seed $({\bf x}_{t},\widetilde{B}_{t})$ of $\mathcal{U}$ defines a partial order $\preceq_{t}$ on $\mathbb{Z}^{m}$. For two vectors ${\bf g},{\bf g}^{\prime}\in\mathbb{Z}^{m}$, we write ${\bf g}^{\prime}\preceq_{t}{\bf g}$ if there exists some vector ${\bf v}=(v_{1},\ldots,v_{n})^{T}\in\mathbb{N}^{n}$ such that
${\bf g}^{\prime}={\bf g}+\widetilde{B}_{t}{\bf v},$
equivalently, ${\bf x}_{t}^{{\bf g}^{\prime}}={\bf x}_{t}^{\bf g}\cdot\widehat{\bf y}_{t}^{\bf v}$, where $\widehat{\bf y}_{t}^{\bf v}={\bf x}_{t}^{\widetilde{B}_{t}{\bf v}}$. We denote by ${\bf g}^{\prime}\prec_{t}{\bf g}$ for the case ${\bf g}^{\prime}\preceq_{t}{\bf g}$ and ${\bf g}^{\prime}\neq{\bf g}$.
The partial order $\preceq_{t}$ above is known as dominance order associated to seed $({\bf x}_{t},\widetilde{B}_{t})$ in [Qin_2017]. The dominance order plays an important role in study of the bases problem of (upper) cluster algebras.
For any seed $({\bf x}_{t},\widetilde{B}_{t})$ of $\mathcal{U}$, we denote by $R_{t}$ the collection of the formal Laurent series
$u=\sum_{{\bf h}\in\mathbb{Z}^{m}}b_{\bf h}{\bf x}_{t}^{\bf h},\;\;\;b_{\bf h}\in\mathbb{Z},$
such that the set $\{{\bf h}\in\mathbb{Z}^{m}\mid b_{\bf h}\neq 0\}$ has finitely many maximal elements with respect to the dominance order $\preceq_{t}$ on $\mathbb{Z}^{m}$. One can see that the collection $R_{t}$ has a ring structure.
Now we recall the definition of pointed elements introduced in [Qin_2017]. Let
$u=\sum_{{\bf h}\in\mathbb{Z}^{m}}b_{\bf h}{\bf x}_{t}^{\bf h}\in R_{t}.$
We say that $u$ is pointed for the seed $({\bf x}_{t},\widetilde{B}_{t})$ if the following two conditions are satisfied.
  • The set $\{{\bf h}\in\mathbb{Z}^{m}\mid b_{\bf h}\neq 0\}$ has a unique maximal element ${\bf g}$ under the dominance order $\preceq_{t}$ on $\mathbb{Z}^{m}$. We denote by $\operatorname{\mathsf{deg}}^{t}(u):={\bf g}$ and call it the degree of $u$ with respect to the seed $({\bf x}_{t},\widetilde{B}_{t})$.
  • The coefficient corresponds to the degree term is $1$, i.e., $b_{\bf g}=1$ for ${\bf g}=\operatorname{\mathsf{deg}}^{t}(u)$.
By definition, each pointed element $u$ for the seed $({\bf x}_{t},\widetilde{B}_{t})$ takes the following form.
$u={\bf x}_{t}^{\bf g}+\sum_{{\bf h}\prec_{t}{\bf g}}b_{\bf h}{\bf x}_{t}^{\bf h}={\bf x}_{t}^{\bf g}F(\hat{y}_{1;t},\ldots,\hat{y}_{n;t}),$
where $b_{\bf h}\in\mathbb{Z}$ and $F\in\mathbb{Z}[[y_{1},\ldots,y_{n}]]$ is a polynomial series with constant term $1$.
In cluster algebras, we are often interested in the pointed elements which can be controlled by tropical points associated to a $Y$-pattern. Such elements correspond to compatibly pointed elements introduced in [qin_2019].
Definition 3.6 (Compatibly pointed elements and good elements).
Let $(\mathcal{S}_{X},\mathcal{S}_{Y})$ be a Langlands dual pair and $\mathcal{U}$ the upper cluster algebra associated to the cluster pattern $\mathcal{S}_{X}$. Assume that $\mathcal{U}$ is of full rank.
  • (i)
    An element $u\in\mathcal{U}$ is said to be compatibly pointed, if $u$ is pointed for any seed $({\bf x}_{t},\widetilde{B}_{t})$ of $\mathcal{U}$ and the collection $[{\bf g}]:=\{\operatorname{\mathsf{deg}}^{t}(u)\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ forms a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$. In this case, we also say that $u$ is $[{\bf g}]$-pointed.
  • (ii)
    A $[{\bf g}]$-pointed element $u\in\mathcal{U}$ is said to be $[{\bf g}]$-good, if it is universally positive. In other words, the Laurent expansion of $u$ with respect to any seed $({\bf x}_{t},\widetilde{B}_{t})$ of $\mathcal{U}$ can be written as
    (3.3) $u={\bf x}_{t}^{{\bf g}_{u}^{t}}F_{u}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t}),$
    where ${\bf g}_{u}^{t}=\operatorname{\mathsf{deg}}^{t}(u)$ is the value of $[{\bf g}]$ at vertex $t$ and $F_{u}^{t}$ is a polynomial in $\mathbb{Z}_{\geq 0}[y_{1},\ldots,y_{n}]$ with constant term $1$.
Let $u$ be a $[{\bf g}]$-pointed element, then the Laurent expansion of $u$ with respect to any seed $({\bf x}_{t},\widetilde{B}_{t})$ of $\mathcal{U}$ has a canonical expression:
(3.4) $u={\bf x}_{t}^{{\bf g}_{u}^{t}}F_{u}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t}),$
where ${\bf g}_{u}^{t}=\operatorname{\mathsf{deg}}^{t}(u)$ is the value of $[{\bf g}]$ at vertex $t$ and $F_{u}^{t}$ is a polynomial in $\mathbb{Z}[y_{1},\ldots,y_{n}]$ with constant term $1$.
  • The degree vector ${\bf g}_{u}^{t}=\operatorname{\mathsf{deg}}^{t}(u)$ is often called the extended $g$-vector of $u$ with respect to vertex $t$;
  • The polynomial $F_{u}^{t}$ is called the $F$-polynomial of $u$ with respect to vertex $t$;
  • For $k\in[1,n]$, we denote by $f_{k;u}^{t}$ the highest degree of $y_{k}$ in $F_{u}^{t}$. The vector
    ${\bf f}_{u}^{t}:=(f_{1;u}^{t},\ldots,f_{n;u}^{t})^{T}\in\mathbb{N}^{n}$
    is called the $f$-vector of $u$ with respect to vertex $t$;
  • The $[{\bf g}]$-pointed element $u$ is said to be $[{\bf g}]$-bipointed, if for any vertex $t\in\mathbb{T}_{n}$ the monomial $y_{1}^{f_{1;u}^{t}}\cdots y_{n}^{f_{n;u}^{t}}$ appears in $F_{u}^{t}$ and it has coefficient $1$;
  • The $[{\bf g}]$-pointed element $u$ is said to be $[{\bf g}]$-bigood, if it is $[{\bf g}]$-bipointed and universally positive.
Recall that a cluster monomial in $\mathcal{U}$ is a Laurent monomial of the form
${\bf x}_{w}^{\bf v}:=\prod_{j=1}^{m}x_{j;w}^{v_{j}}$
for some cluster ${\bf x}_{w}$ of $\mathcal{U}$ and some vector ${\bf v}=(v_{1},\ldots,v_{m})^{T}\in\mathbb{Z}^{m}$ with $v_{i}\geq 0$ for any $i\in[1,n]$. Notice that we allow $v_{j}$ to be negative for $j\in[n+1,m]$.
Keep the setting as in Definition 3.6. Given a vertex $w\in\mathbb{T}_{n}$ and a vector ${\bf v}\in\mathbb{Z}^{m}$, we denote by $[({\bf v},w)]$ the unique tropical point $\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ determined by the condition ${\bf g}^{w}={\bf v}$.
Proposition 3.7.
Let ${\bf x}_{w}^{\bf v}$ be a cluster monomial in $\mathcal{U}$. Then ${\bf x}_{w}^{\bf v}$ is $[({\bf v},w)]$-bigood in $\mathcal{U}$.
Proof.
By the results in [GHKK18], we know that $u:={\bf x}_{w}^{\bf v}$ is $[({\bf v},w)]$-good in $\mathcal{U}$. Let ${\bf f}_{u}^{t}=(f_{1;u}^{t},\ldots,f_{n;u}^{t})^{T}$ be the $f$-vector of $u={\bf x}_{w}^{\bf v}$ with respect to vertex $t\in\mathbb{T}_{n}$. Since the $F$-polynomials of cluster variables have constant term $1$ and by [fomin_zelevinsky_2007, Proposition 5.3], we know that the monomial $y_{1}^{f_{1;u}^{t}}\cdots y_{n}^{f_{n;u}^{t}}$ appears in the $F$-polynomial $F_{u}^{t}$ with coefficient $1$. Hence, $u={\bf x}_{w}^{\bf v}$ is $[({\bf v},w)]$-bigood in $\mathcal{U}$. ∎
Example 3.8.
Let us continue with Example 2.10. We see that $\widetilde{B}=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$ has full rank and $\hat{y}_{1}=x_{2}^{-1},\;\hat{y}_{2}=x_{1}$. The canonical expressions of the five cluster variables
$x_{1},\;x_{2},\;x_{3}=\frac{x_{2}+1}{x_{1}},\;x_{4}=\frac{x_{1}+x_{2}+1}{x_{1}x_{2}},\;x_{5}=\frac{x_{1}+1}{x_{2}}$
with respect to the initial seed $({\bf x},\widetilde{B})$ are given as follows:
$x_{1}=x_{1}\cdot 1,\;\;\;x_{2}=x_{2}\cdot 1,\;\;\;x_{3}={x_{1}^{-1}x_{2}\cdot(1+\widehat{y}_{1}),}\;\;\;x_{4}={x_{1}^{-1}\cdot(1+\widehat{y}_{1}+\widehat{y}_{1}\widehat{y}_{2}),}\;\;\;x_{5}={x_{2}^{-1}\cdot(1+\widehat{y}_{2})}.$
3.3. Basic properties of pointed elements
As before, we fix a full rank upper cluster algebra $\mathcal{U}$. Let us recall some basic properties of pointed elements in $R_{t}$.
Lemma 3.9.
[qin_2019]*Lemmas 3.2.5, 3.2.6 Suppose that $u,u^{\prime}\in R_{t}$ are two pointed elements for seed $({\bf x}_{t},\widetilde{B}_{t})$. Then the following statements hold.
  • (i)
    The product $uu^{\prime}\in R_{t}$ is pointed for seed $({\bf x}_{t},\widetilde{B}_{t})$ and we have
    $\operatorname{\mathsf{deg}}^{t}(uu^{\prime})=\operatorname{\mathsf{deg}}^{t}(u)+\operatorname{\mathsf{deg}}^{t}(u^{\prime}).$
  • (ii)
    $u$ has a multiplicative inverse $u^{-1}$ in $R_{t}$. Moreover, $u^{-1}$ is pointed for $({\bf x}_{t},\widetilde{B}_{t})$ and
    $\operatorname{\mathsf{deg}}^{t}(u^{-1})=-\operatorname{\mathsf{deg}}^{t}(u).$
Let $({\bf x}_{t},\widetilde{B}_{t})$ and $({\bf x}_{w},\widetilde{B}_{w})$ be two seeds of $\mathcal{U}$ at vertices $t,w\in\mathbb{T}_{n}$. By Proposition 3.7, we know that any cluster variable $x_{j;t}$ ($j\in[1,m]$) is pointed for $({\bf x}_{w},\widetilde{B}_{w})$. We denote by $\widetilde{G}_{t}^{w}$ the $m\times m$ matrix whose $j$th column is given by ${\bf g}_{x_{j;t}}^{w}=\operatorname{\mathsf{deg}}^{w}(x_{j;t})\in\mathbb{Z}^{m}$. Notice that $\widetilde{G}_{t}^{w}$ takes the form
$\widetilde{G}_{t}^{w}=\begin{bmatrix}G_{t}^{w}&0\\ \ast&I_{m-n}\end{bmatrix},$
where $G_{t}^{w}$ is the $n\times n$ submatrix of $\widetilde{G}_{t}^{w}$. We call $\widetilde{G}_{t}^{w}$ the extended $G$-matrix and $G_{t}^{w}$ the $G$-matrix of vertex $t$ with respect to $w$. It is easy to see that if $t=w$, then $\widetilde{G}_{t}^{w}=I_{m}$.
Let $\overleftarrow{\mu}$ be the mutation sequence corresponding to the unique path from vertex $w$ to $t$ in $\mathbb{T}_{n}$. We apply the mutation sequence $\overleftarrow{\mu}$ to $\begin{pmatrix}B_{w}\\ I_{n}\end{pmatrix}$, where $B_{w}$ is the principal part of $\widetilde{B}_{w}$. The resulting matrix $\overleftarrow{\mu}\begin{pmatrix}B_{w}\\ I_{n}\end{pmatrix}$ takes the form $\begin{pmatrix}B_{t}\\ C_{t}^{w}\end{pmatrix}$. We call the $n\times n$ matrix $C_{t}^{w}$ the $C$-matrix of the vertex $t$ with respect to $w$.
Theorem 3.10.
Let $t,w$ be two vertices in $\mathbb{T}_{n}$. The following statements hold.
  • (i)
    [GHKK18] $\det(\widetilde{G}_{t}^{w})=\pm 1=\det(G_{t}^{w})$ and $\det(C_{t}^{w})=\pm 1$.
  • (ii)
    [GHKK18] Each column of $C_{t}^{w}$ is either a non-negative vector or a non-positive vector.
  • (iii)
    [NZ12, (3.11)] $SC_{t}^{w}S^{-1}(G_{t}^{w})^{\rm T}=I_{n}$, where $S$ is any skew-symmetrizer of $B_{w}$.
Lemma 3.11.
Let $\mathcal{U}$ be a full rank upper cluster algebra and $t,w$ two vertices of $\mathbb{T}_{n}$. The following statements hold.
  • (i)
    [qin_2019]*Lemma 3.3.6 Any Laurent monomial ${\bf x}_{t}^{\bf h}$ is pointed for the seed $({\bf x}_{w},\widetilde{B}_{w})$ and we have
    $\operatorname{\mathsf{deg}}^{w}({\bf x}_{t}^{\bf h})=\widetilde{G}_{t}^{w}{\bf h}=\operatorname{\mathsf{deg}}^{w}({\bf x}_{w}^{\widetilde{G}_{t}^{w}{\bf h}}).$
  • (ii)
    [qin_2019]*Proposition 3.3.10 Let $\widehat{\bf y}_{t}$ and $\widehat{\bf y}_{w}$ be the collections of $\hat{y}$-variables at vertices $t$ and $w$. Then for any ${\bf v}\in\mathbb{Z}^{n}$, the Laurent monomial $\widehat{\bf y}_{t}^{\bf v}={\bf x}_{t}^{\widetilde{B}_{t}{\bf v}}$ is pointed for the seed $({\bf x}_{w},\widetilde{B}_{w})$ and
    $\operatorname{\mathsf{deg}}^{w}(\widehat{\bf y}_{t}^{\bf v})=\widetilde{G}_{t}^{w}\widetilde{B}_{t}{\bf v}=\widetilde{B}_{w}C_{t}^{w}{\bf v}=\operatorname{\mathsf{deg}}^{w}(\widehat{\bf y}_{w}^{C_{t}^{w}{\bf v}}).$
Notice that any element in $R_{w}$ can be written as a finite $\mathbb{Z}$-linear combinations of pointed elements for seed $({\bf x}_{w},\widetilde{B}_{w})$. By the above lemma, any Laurent monomial ${\bf x}_{t}^{\bf h}$ in $\mathcal{L}(t):=\mathbb{Z}[x_{1;t}^{\pm 1},\ldots,x_{m;t}^{\pm 1}]$ is pointed for the seed $({\bf x}_{w},\widetilde{B}_{w})$. Hence, $\mathcal{L}(t)$ is contained in $R_{w}$.
Corollary 3.12.
Let $\mathcal{U}$ be a full rank upper cluster algebra and $t,w$ two vertices of $\mathbb{T}_{n}$. Let
$u=\sum\limits_{{\bf h}\in\mathbb{Z}^{m}}b_{\bf h}{\bf x}_{t}^{\bf h}$
be a Laurent polynomial in $\mathcal{L}(t)=\mathbb{Z}[x_{1;t}^{\pm 1},\ldots,x_{m;t}^{\pm 1}]$. Then the following statements hold.
  • (i)
    $u$ is pointed for the seed $({\bf x}_{w},\widetilde{B}_{w})$ if and only if there exists a (unique) ${\bf g}\in\mathbb{Z}^{m}$ with $b_{\bf g}=1$ such that $\operatorname{\mathsf{deg}}^{w}({\bf x}_{t}^{\bf h})\preceq_{w}\operatorname{\mathsf{deg}}^{w}({\bf x}_{t}^{\bf g})$ for any ${\bf h}\in\mathbb{Z}^{m}$ with $b_{\bf h}\neq 0$, equivalently,
    $\widetilde{G}_{t}^{w}{\bf h}\preceq_{w}\widetilde{G}_{t}^{w}{\bf g}$
    for any ${\bf h}\in\mathbb{Z}^{m}$ with $b_{\bf h}\neq 0$.
  • (ii)
    Suppose that $u$ is pointed for the seed $({\bf x}_{t},\widetilde{B}_{t})$, that is, $u$ can be written as
    $u={\bf x}_{t}^{\bf g}\cdot\sum\limits_{{\bf v}\in\mathbb{N}^{n}}c_{\bf v}\widehat{\bf y}_{t}^{\bf v},$
    where $c_{\bf v}\in\mathbb{Z}$ and $c_{0}=1$. Then $u$ is pointed for the new seed $({\bf x}_{w},\widetilde{B}_{w})$ if and only if there exists a (unique) ${\bf u}\in\mathbb{N}^{n}$ with $c_{\bf u}=1$ such that $\operatorname{\mathsf{deg}}^{w}(\widehat{\bf y}_{t}^{\bf v})\preceq_{w}\operatorname{\mathsf{deg}}^{w}(\widehat{\bf y}_{t}^{\bf u})$ for any ${\bf v}\in\mathbb{N}^{n}$ with $c_{\bf v}\neq 0$, equivalently,
    $C_{t}^{w}{\bf v}-C_{t}^{w}{\bf u}=C_{t}^{w}({\bf v}-{\bf u})\in\mathbb{N}^{n}$
    for any ${\bf v}\in\mathbb{N}^{n}$ with $c_{\bf v}\neq 0$.
Proof.
Thanks to Theorem 3.10 (i) and by Lemma 3.11, we know that $\operatorname{\mathsf{deg}}^{w}({\bf x}_{t}^{\bf h})=\operatorname{\mathsf{deg}}^{w}({\bf x}_{t}^{{\bf h}^{\prime}})$ if and only if ${\bf h}={\bf h}^{\prime}$. By the similar argument and since $\widetilde{B}_{w}$ has full rank, we know that $\operatorname{\mathsf{deg}}^{w}(\widehat{\bf y}_{t}^{\bf v})=\operatorname{\mathsf{deg}}^{w}(\widehat{\bf y}_{t}^{{\bf v}^{\prime}})$ if and only if ${\bf v}={\bf v}^{\prime}$.
Based on the above observations, the results in (i) and (ii) follow from the definition of pointed elements and Lemma 3.11. ∎
Let $(\mathcal{S}_{X},\mathcal{S}_{Y})$ be a Langlands dual pair and $\mathcal{U}$ the upper cluster algebra associated to the cluster pattern $\mathcal{S}_{X}$. As before, we assume that $\mathcal{U}$ is of full rank.
Corollary 3.13.
Keep the above setting. Let $[{\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ be a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$, and write ${\bf g}^{t}=(g_{1}^{t},\ldots,g_{m}^{t})^{T}$ for $t\in\mathbb{T}_{n}$. Let ${w}$${s}$$\scriptstyle{k}$ be an edge in $\mathbb{T}_{n}$. The following statements hold.
  • (i)
    $\operatorname{\mathsf{deg}}^{w}({\bf x}_{s}^{{\bf g}^{s}}\cdot\widehat{\bf y}_{s}^{[-g_{k}^{s}]_{+}{\bf e}_{k}})={\bf g}^{w}$, where ${\bf e}_{k}$ is the $k$th column of $I_{n}$.
  • (ii)
    Suppose that $u\in\mathcal{U}$ is a $[{\bf g}]$-pointed element. Let $u={\bf x}_{s}^{{\bf g}^{s}}\cdot\sum\limits_{{\bf v}\in\mathbb{N}^{n}}c_{\bf v}\widehat{\bf y}_{s}^{\bf v}$ be the Laurent expansion of $u$ with respect to seed $({\bf x}_{s},\widetilde{B}_{s})$. Then the coefficient of $\widehat{\bf y}_{s}^{[-g_{k}^{s}]_{+}{\bf e}_{k}}$ in $\sum\limits_{{\bf v}\in\mathbb{N}^{n}}c_{\bf v}\widehat{\bf y}_{s}^{\bf v}$ is $1$ and for any integer $d$ with $d>[-g_{k}^{s}]_{+}$, the coefficient of $\widehat{\bf y}_{s}^{d{\bf e}_{k}}$ in $\sum\limits_{{\bf v}\in\mathbb{N}^{n}}c_{\bf v}\widehat{\bf y}_{s}^{\bf v}$ is zero.
Proof.
(i) Let $\widetilde{G}_{s}^{w}$ and $C_{s}^{w}$ be the extended $G$-matrix and $C$-matrix of vertex $s$ with respect to $w$. By Lemma 3.9 and Lemma 3.11, we have
$\operatorname{\mathsf{deg}}^{w}({\bf x}_{s}^{{\bf g}^{s}}\cdot\widehat{\bf y}_{s}^{[-g_{k}^{s}]_{+}{\bf e}_{k}})=\widetilde{G}_{s}^{w}{\bf g}^{s}+\widetilde{B}_{w}C_{s}^{w}\cdot([-g_{k}^{s}]_{+}{\bf e}_{k}).$
Since the two vertices $w$ and $s$ are jointed in $\mathbb{T}_{n}$ by an edge labeled by $k$, we know that $\widetilde{G}_{s}^{w}=(g_{ij;s})$ and $C_{s}^{w}=(c_{ij;s})$ are given as follows:
$g_{ij;s}$ $=$ $\begin{cases}-1,&i=j=k;\\ [-b_{ik;w}]_{+}=[b_{ik;s}]_{+},&j=k,\;i\neq k;\\ \delta_{i,j},&\text{else},\end{cases}$
$c_{ij;s}$ $=$ $\begin{cases}-1,&i=j=k;\\ [b_{kj;w}]_{+}=[-b_{kj;s}]_{+},&i=k,\;j\neq k;\\ \delta_{i,j},&\text{else}.\end{cases}$
Because the $k$-th column of $C_{s}^{w}$ equals $-{\bf e}_{k}$, we have
$\widetilde{G}_{s}^{w}{\bf g}^{s}+\widetilde{B}_{w}C_{s}^{w}\cdot([-g_{k}^{s}]_{+}{\bf e}_{k})=\widetilde{G}_{s}^{w}{\bf g}^{s}-\widetilde{B}_{w}\cdot([-g_{k}^{s}]_{+}{\bf e}_{k}).$
Now we look at the $i$th component $g_{i}^{\prime}$ of the vector
${\bf g}^{\prime}:=\widetilde{G}_{s}^{w}{\bf g}^{s}-\widetilde{B}_{w}\cdot([-g_{k}^{s}]_{+}{\bf e}_{k}).$
We have
$g_{i}^{\prime}=\begin{cases}-g_{k}^{s}-0=-g_{k}^{s},&i=k;\\ (g_{i}^{s}+[b_{ik;s}]_{+}g_{k}^{s})-b_{ik;w}[-g_{k}^{s}]_{+}=g_{i}^{s}+[b_{ik;s}]_{+}g_{k}^{s}+b_{ik;s}[-g_{k}^{s}]_{+},&i\neq k.\end{cases}$
Since $\widehat{B}_{s}=-\widetilde{B}_{s}^{T}$, we can rewrite the above equality as follows:
$g_{i}^{\prime}=\begin{cases}-g_{k}^{s},&i=k;\\ g_{i}^{s}+[-\hat{b}_{ki;s}]_{+}g_{k}^{s}-\hat{b}_{ki;s}[-g_{k}^{s}]_{+},&i\neq k.\end{cases}$
Using the equality $[b]_{+}g-[-b]_{+}g=bg=b[g]_{+}-b[-g]_{+}$ for any $b,g\in\mathbb{R}$, one can see the above equality is the same as
$g_{i}^{\prime}=\begin{cases}-g_{k}^{s},&i=k;\\ g_{i}^{s}+[\hat{b}_{ki;s}]_{+}g_{k}^{s}-\hat{b}_{ki;s}[g_{k}^{s}]_{+},&i\neq k.\end{cases}$
Then by (3.1), we get ${\bf g}^{w}={\bf g}^{\prime}=\widetilde{G}_{s}^{w}{\bf g}^{s}-\widetilde{B}_{w}\cdot([-g_{k}^{s}]_{+}{\bf e}_{k})$. Hence, $\operatorname{\mathsf{deg}}^{w}({\bf x}_{s}^{{\bf g}^{s}}\cdot\widehat{\bf y}_{s}^{[-g_{k}^{s}]_{+}{\bf e}_{k}})={\bf g}^{w}$.
(ii) Since $u$ is $[{\bf g}]$-pointed, we know that it is pointed for the seed $({\bf x}_{w},\widetilde{B}_{w})$. By Corollary 3.12 (ii), there exists a unique ${\bf u}\in\mathbb{N}^{n}$ with $c_{\bf u}=1$ such that
$\operatorname{\mathsf{deg}}^{w}({\bf x}_{s}^{{\bf g}^{s}}\cdot\widehat{\bf y}_{s}^{\bf v})\preceq_{w}\operatorname{\mathsf{deg}}^{w}({\bf x}_{s}^{{\bf g}^{s}}\cdot\widehat{\bf y}_{s}^{\bf u})=\operatorname{\mathsf{deg}}^{w}(u)={\bf g}^{w}$
for any ${\bf v}\in\mathbb{N}^{n}$ with $c_{\bf v}\neq 0$. By the result in (i), we know that ${\bf u}=[-g_{k}^{s}]_{+}{\bf e}_{k}$.
Notice that $\operatorname{\mathsf{deg}}^{w}({\bf x}_{s}^{{\bf g}^{s}}\cdot\widehat{\bf y}_{s}^{\bf v})\preceq_{w}\operatorname{\mathsf{deg}}^{w}({\bf x}_{s}^{{\bf g}^{s}}\cdot\widehat{\bf y}_{s}^{\bf u})$ holds if and only if the following relation holds.
$C_{s}^{w}{\bf v}-C_{s}^{w}{\bf u}=C_{s}^{w}({\bf v}-{\bf u})\in\mathbb{N}^{n}.$
Since the $k$th column of $C_{s}^{w}$ is $-{\bf e}_{k}$, we obtain that $C_{s}^{w}(d{\bf e}_{k}-[-g_{k}^{s}]_{+}{\bf e}_{k})$ belongs to $\mathbb{N}^{n}$ if and only if $d\leq[-g_{k}^{s}]_{+}$. Hence, for any integer $d$ with $d>[-g_{k}^{s}]_{+}$, the coefficient of $\widehat{\bf y}_{s}^{d{\bf e}_{k}}$ in $\sum\limits_{{\bf v}\in\mathbb{N}^{n}}c_{\bf v}\widehat{\bf y}_{s}^{\bf v}$ is zero.
Corollary 3.14.
Let $u$ be a $[{\bf g}]$-pointed element in $\mathcal{U}$ and let ${\bf f}_{u}^{t}=(f_{1;u}^{t},\ldots,f_{n;u}^{t})^{T}$ be the $f$-vector of $u$ with respect to vertex $t\in\mathbb{T}_{n}$. The following statements hold.
  • (i)
    If $f_{k;u}^{t}=0$ for some $k\in[1,n]$, then the $k$th component $g_{k}^{t}$ of ${\bf g}_{u}^{t}=\operatorname{\mathsf{deg}}^{t}(u)$ is non-negative;
  • (ii)
    If $f_{i;u}^{t}=0$ for any $i\in[1,n]$, then $u$ is a cluster monomial in ${\bf x}_{t}$;
  • (iii)
    Let $k$ be an integer in $[1,n]$ and $({\bf x}_{t^{\prime}},\widetilde{B}_{t^{\prime}})=\mu_{k}({\bf x}_{t},\widetilde{B}_{t})$. Suppose that $f_{k;u}^{t}\neq 0$ but $f_{i;u}^{t}=0$ for any $i\in[1,n]\backslash\{k\}$. Then $u$ is a cluster monomial in ${\bf x}_{t^{\prime}}$.
Proof.
Since $u$ is $[{\bf g}]$-pointed in $\mathcal{U}$, the Laurent expansion of $u$ with respect to ${\bf x}_{t}$ has a cannonical expression:
$u={\bf x}_{t}^{{\bf g}_{u}^{t}}F_{u}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t}),$
where ${\bf g}_{u}^{t}=(g_{1}^{t},\ldots,g_{m}^{t})^{T}$ is the value of $[{\bf g}]$ at vertex $t$ and $F_{u}^{t}$ is a polynomial in $\mathbb{Z}[y_{1},\ldots,y_{n}]$ with constant term $1$.
(i) By Corollary 3.13 (ii), we know that $y_{k}^{[-g_{k}^{t}]_{+}}$ must be a monomial appearing in $F_{u}^{t}\in\mathbb{Z}[y_{1},\ldots,y_{n}]$. Recall that $f_{k;u}^{t}$ is the highest degree of $y_{k}$ in $F_{u}^{t}$. So we have $f_{k;u}^{t}\geq[-g_{k}^{t}]_{+}\geq 0$. Hence, if $f_{k;u}^{t}=0$, then $[-g_{k}^{t}]_{+}=0$. So $g_{k}^{t}$ is non-negative.
(ii) Since $f_{i;u}^{t}=0$ for any $i\in[1,n]$, we get $F_{u}^{t}=1$ and $u={\bf x}_{t}^{{\bf g}_{u}^{t}}$. On the other hand, by (i), we have $g_{i}^{t}\geq 0$ for any $i\in[1,n]$. Hence, $u={\bf x}_{t}^{{\bf g}_{u}^{t}}$ is a cluster monomial in ${\bf x}_{t}$.
(iii) Since $f_{i;u}^{t}=0$ holds for any $i\in[1,n]\setminus\{k\}$, we get that $F_{u}^{t}$ is a polynomial in the single variable $y_{k}$. In this case, Corollary 3.13 (ii) implies $f_{k;u}^{t}=[-g_{k}^{t}]_{+}$. Since $f_{k;u}^{t}\neq 0$, we have $g_{k}^{t}<0$ and thus $f_{k;u}^{t}=-g_{k}^{t}$.
Now $u$ has the following form:
$u=x_{1;t}^{g_{1}^{t}}\cdots x_{m;t}^{g_{m}^{t}}(1+\sum_{0<d<-g_{k}^{t}}c_{d}\hat{y}_{k;t}^{d}+\hat{y}_{k;t}^{-g_{k}^{t}}),$
where $c_{d}\in\mathbb{Z}$. By the exchange relations, we have $x_{j;t^{\prime}}=x_{j;t}$ for any $j\neq k$ and $x_{k;t}\cdot x_{k;t^{\prime}}=M\cdot(1+\hat{y}_{k;t})$, where $M=\prod_{j=1}^{m}x_{j;t}^{[-b_{jk;t}]_{+}}$. By substituting these relations into the above equality and using the fact $u\in\mathbb{Z}[x_{1;t^{\prime}}^{\pm 1},\ldots,x_{m;t^{\prime}}^{\pm 1}]$, one deduces that
$(1+\sum_{0<d<-g_{k}^{t}}c_{d}\hat{y}_{k;t}^{d}+\hat{y}_{k;t}^{-g_{k}^{t}})=(1+\hat{y}_{k;t})^{-g_{k}^{t}}$
and the Laurent expansion of $u$ with respect to seed $({\bf x}_{t^{\prime}},\widetilde{B}_{t^{\prime}})$ is a Laurent monomial in $\mathbb{Z}[x_{1;t^{\prime}}^{\pm 1},\ldots,x_{m;t^{\prime}}^{\pm 1}]$. Thus the $F$-polynomial $F_{u}^{t^{\prime}}$ of $u$ with respect to vertex $t^{\prime}$ is $1$. In particular, the $f$-vector ${\bf f}_{u}^{t^{\prime}}$ of $u$ with respect to vertex $t^{\prime}$ is zero. By (ii), we know that $u$ is a cluster monomial in ${\bf x}_{t^{\prime}}$.
4. $F$-invariant in cluster algebras
4.1. Maps between sets of tropical points
Let $(\mathcal{S}_{X},\mathcal{S}_{Y})$ be a Langlands dual pair. In this subsection, we give two maps. The first map is from the tropical points in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$ to those in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ and the second one is from the tropical points in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ to those in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$. The second map will play an important role in the proof Theorem 4.8.
Lemma 4.1.
Let $(\mathcal{S}_{X},\mathcal{S}_{Y})$ be a Langlands dual pair and $\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ a collection of vectors in $\mathbb{Z}^{m}$ indexed by the vertices in $\mathbb{T}_{n}$. Then the collection $\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ is a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ if and only if for any edge ${t}$${t^{\prime}}$$\scriptstyle{k}$ in $\mathbb{T}_{n}$, one has ${\bf g}^{t^{\prime}}=E_{k,\varepsilon(g_{k}^{t})}^{\widetilde{B}_{t}}{\bf g}^{t}$, where
$\varepsilon(g_{k}^{t})=\begin{cases}-,&\text{if }g_{k}^{t}\geq 0;\\ +,&\text{if }g_{k}^{t}<0,\end{cases}$
and $E_{k,\pm}^{\widetilde{B}_{t}}$ is as defined in (2.2).
Proof.
By definition, the collection $\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ is a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ if and only if the relations in (3.1) hold. Since $(\mathcal{S}_{X},\mathcal{S}_{Y})$ is a Langlands dual pair, we have $\widehat{B}_{t}=-\widetilde{B}_{t}^{T}$. Thus $\hat{b}_{ij;t}=-b_{ji;t}$ for any $i\in[1,n]$ and $j\in[1,m]$. Then the relations in (3.1) can be rewritten as
$g_{i}^{t^{\prime}}$ $=$ $\begin{cases}-g_{k}^{t},&\text{if}\;i=k;\\ g_{i}^{t}+[-b_{ik;t}]_{+}g_{k}^{t}+b_{ik;t}[g_{k}^{t}]_{+},&\text{if}\;i\neq k.\end{cases}$
$=$ $\begin{cases}-g_{k}^{t},&\text{if}\;i=k;\\ g_{i}^{t}+[b_{ik;t}]_{+}g_{k}^{t},&\text{if}\;i\neq k\;\;\text{and }\;g_{k}^{t}\geq 0;\\ g_{i}^{t}+[-b_{ik;t}]_{+}g_{k}^{t},&\text{if}\;i\neq k\;\;\text{and }\;g_{k}^{t}<0.\end{cases}$
This is exactly the same with ${\bf g}^{t^{\prime}}=E_{k,\varepsilon(g_{k}^{t})}^{\widetilde{B}_{t}}{\bf g}^{t}$. ∎
Lemma 4.2.
Let $\mathcal{S}_{X}$ be a cluster pattern of rank $n$ in $\mathbb{F}$ and $\{{\bf a}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ a collection of vectors in $\mathbb{Z}^{m}$ indexed by the vertices of $\mathbb{T}_{n}$. Then the collection $\{{\bf a}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ is a tropical point in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$ if and only if for any edge ${t}$${t^{\prime}}$$\scriptstyle{k}$ in $\mathbb{T}_{n}$, one has ${\bf a}^{t^{\prime}}=(E_{k,\varepsilon_{0}}^{\widetilde{B}_{t}})^{T}{\bf a}^{t}$, where
$\varepsilon_{0}=\begin{cases}-,&\text{if }\sum_{j=1}^{m}b_{jk;t}a_{j}^{t}\geq 0;\\ +,&\text{if }\sum_{j=1}^{m}b_{jk;t}a_{j}^{t}<0,\end{cases}$
and $E_{k,\pm}^{\widetilde{B}_{t}}$ is as defined in (2.2).
Proof.
By definition, the collection $\{{\bf a}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ is a tropical point in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$ if and only if the relations in (3.2) hold. The relations in (3.2) can be rewritten as follows:
$a_{i}^{t^{\prime}}=\begin{cases}-a_{k}^{t}+\sum_{j=1}^{m}[b_{jk;t}]_{+}a_{j}^{t},&\text{if}\;i=k\;\;\text{and }\;\sum_{j=1}^{m}b_{jk;t}a_{j}^{t}\geq 0;\\ -a_{k}^{t}+\sum_{j=1}^{m}[-b_{jk;t}]_{+}a_{j}^{t},&\text{if}\;i=k\;\;\text{and}\;\sum_{j=1}^{m}b_{jk;t}a_{j}^{t}<0;\\ a_{i}^{t},&\text{if}\;i\neq k.\end{cases}$
This is exactly the same with ${\bf a}^{t^{\prime}}=(E_{k,\varepsilon_{0}}^{\widetilde{B}_{t}})^{T}{\bf a}^{t}$. ∎
Let $(\mathcal{S}_{X},\mathcal{S}_{Y})$ be a Langlands dual pair. We extend $\widetilde{B}_{t_{0}}$ to an $m\times m$ skew-symmetrizable matrix $\widetilde{B}_{t_{0}}^{\operatorname{\mathsf{sq}}}=(\widetilde{B}_{t_{0}}\mid M)$, where $M$ is an $m\times(m-n)$ matrix. Denote by $\{\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}}\mid t\in\mathbb{T}_{n}\}$ the collection of matrices obtained from $\widetilde{B}_{t_{0}}^{\operatorname{\mathsf{sq}}}$ by any sequence of mutations in directions $1,\ldots,n$. Let $\widetilde{S}=diag(s_{1},\ldots,s_{m})$ be a skew-symmetrizer of the transposed matrix $(\widetilde{B}_{t_{0}}^{\operatorname{\mathsf{sq}}})^{T}$. In particular, $\widetilde{S}(\widetilde{B}_{t_{0}}^{\operatorname{\mathsf{sq}}})^{T}$ is a skew-symmetric matrix.
Proposition 4.3.
Keep the above setting. Then
$[{\bf a}]:=\{{\bf a}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}\mapsto\{\widetilde{S}(\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}})^{T}{\bf a}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$
defines a map from the tropical points in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$ to those in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$.
Proof.
Denote by ${\bf g}^{t}:=\widetilde{S}(\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}})^{T}{\bf a}^{t}=(g_{1}^{t},\ldots,g_{m}^{t})^{T}$ for any vertex $t\in\mathbb{T}_{n}$. We will use Lemma 4.1 to show the collection $\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ is a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$.
Let ${t}$${t^{\prime}}$$\scriptstyle{k}$ be an edge in $\mathbb{T}_{n}$. Since $[{\bf a}]=\{{\bf a}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ is a tropical point in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$ and by Lemma 4.2, we have ${\bf a}^{t^{\prime}}=(E_{k,\varepsilon_{0}}^{\widetilde{B}_{t}})^{T}{\bf a}^{t}$, where
$\varepsilon_{0}=\begin{cases}-,&\text{if }\sum_{j=1}^{m}b_{jk;t}a_{j}^{t}\geq 0;\\ +,&\text{if }\sum_{j=1}^{m}b_{jk;t}a_{j}^{t}<0.\end{cases}$
By (2.2), we have $\widetilde{B}_{t^{\prime}}^{\operatorname{\mathsf{sq}}}=E_{k,\varepsilon_{0}}^{\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}}}\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}}F_{k,\varepsilon_{0}}^{\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}}}$. Thus $(\widetilde{B}_{t^{\prime}}^{\operatorname{\mathsf{sq}}})^{T}=(F_{k,\varepsilon_{0}}^{\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}}})^{T}(\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}})^{T}(E_{k,\varepsilon_{0}}^{\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}}})^{T}$. Since $\widetilde{S}$ is a skew-symmetrizer of $(\widetilde{B}_{t_{0}}^{\operatorname{\mathsf{sq}}})^{T}$, we have
$\widetilde{S}(\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}})^{T}=-\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}}\widetilde{S}$
for any vertex $t\in\mathbb{T}_{n}$. In the following proofs, we will use the following obvious relations:
$\widetilde{S}(F_{k,\pm}^{\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}}})^{T}=(E_{k,\pm}^{\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}}})\widetilde{S},\;\;\;E_{k,\pm}^{\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}}}=E_{k,\pm}^{\widetilde{B}_{t}},\;\;\;\text{and}\;\;\;(E_{k,\pm}^{\widetilde{B}_{t}})^{2}=I_{m}.$
Now we have
${\bf g}^{t^{\prime}}$ $=$ $\widetilde{S}(\widetilde{B}_{t^{\prime}}^{\operatorname{\mathsf{sq}}})^{T}{\bf a}^{t^{\prime}}=\widetilde{S}\left((F_{k,\varepsilon_{0}}^{\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}}})^{T}(\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}})^{T}(E_{k,\varepsilon_{0}}^{\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}}})^{T}\right)\cdot(E_{k,\varepsilon_{0}}^{\widetilde{B}_{t}})^{T}{\bf a}^{t}$
$=$ $\widetilde{S}(F_{k,\varepsilon_{0}}^{\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}}})^{T}\cdot(\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}})^{T}\cdot(E_{k,\varepsilon_{0}}^{\widetilde{B}_{t}})^{T}(E_{k,\varepsilon_{0}}^{\widetilde{B}_{t}})^{T}\cdot{\bf a}^{t}$
$=$ $(E_{k,\varepsilon_{0}}^{\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}}})\widetilde{S}(\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}})^{T}{\bf a}^{t}$
$=$ $(E_{k,\varepsilon_{0}}^{\widetilde{B}_{t}}){\bf g}^{t}.$
Now we prove $\varepsilon_{0}=\varepsilon(g_{k}^{t})$, where $\varepsilon(g_{k}^{t})$ is as given in Lemma 4.1. By the definitions of $\varepsilon(g_{k}^{t})$ and $\varepsilon_{0}$, it suffices to show that $g_{k}^{t}\geq 0$ if and only if $\sum_{j=1}^{m}b_{jk;t}a_{j}^{t}\geq 0$. This is clear, because ${\bf g}^{t}=\widetilde{S}(\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}})^{T}{\bf a}^{t}$ implies that $g_{k}^{t}=s_{k}\cdot(\sum_{j=1}^{m}b_{jk;t}a_{j}^{t})$.
Hence, ${\bf g}^{t^{\prime}}=(E_{k,\varepsilon_{0}}^{\widetilde{B}_{t}}){\bf g}^{t}=(E_{k,\varepsilon(g_{k}^{t})}^{\widetilde{B}_{t}}){\bf g}^{t}$. By Lemma 4.1, we know the collection
$\{{\bf g}^{t}=\widetilde{S}(\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}})^{T}{\bf a}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$
is a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$. This completes the proof. ∎
Proposition 4.4.
Let $(\mathcal{S}_{X},\mathcal{S}_{Y},\Lambda)$ be a Langlands-Poisson triple. Then
$[{\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}\mapsto\{\Lambda_{t}{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$
defines a map from the tropical points in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ to those in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$.
Proof.
Denote by ${\bf a}^{t}:=\Lambda_{t}{\bf g}^{t}=(a_{1}^{t},\ldots,a_{m}^{t})^{T}$ for any vertex $t\in\mathbb{T}_{n}$. We will use Lemma 4.2 to show the collection $\{{\bf a}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ is a tropical point in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$.
Let ${t}$${t^{\prime}}$$\scriptstyle{k}$ be an edge in $\mathbb{T}_{n}$. Since $[{\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ is a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ and by Lemma 4.1, we have ${\bf g}^{t^{\prime}}=E_{k,\varepsilon(g_{k}^{t})}^{\widetilde{B}_{t}}{\bf g}^{t}$, where
$\varepsilon(g_{k}^{t})=\begin{cases}-,&\text{if }g_{k}^{t}\geq 0;\\ +,&\text{if }g_{k}^{t}<0.\end{cases}$
By Proposition 2.3 (i), we have $\Lambda_{t^{\prime}}=(E_{k,\varepsilon(g_{k}^{t})}^{\widetilde{B}_{t}})^{T}\Lambda_{t}E_{k,\varepsilon(g_{k}^{t})}^{\widetilde{B}_{t}}$. Hence,
[…]
Now we prove $\varepsilon(g_{k}^{t})=\varepsilon_{0}$, where $\varepsilon_{0}$ is as given in Lemma 4.2. By the definitions of $\varepsilon(g_{k}^{t})$ and $\varepsilon_{0}$, it suffices to show that $g_{k}^{t}\geq 0$ if and only if $\sum_{j=1}^{m}b_{jk;t}a_{j}^{t}\geq 0$. Since $(\widetilde{B}_{t},\Lambda_{t})$ is a compatible pair, we know that $\widetilde{B}_{t}^{T}\Lambda_{t}$ has the form
$\widetilde{B}_{t}^{T}\Lambda_{t}=(S\mid{\bf 0}),$
where $S=diag(s_{1},\ldots,s_{n})$ is a diagonal matrix with $s_{i}>0$, $i\in[1,n]$ and ${\bf 0}$ is the $n\times(m-n)$ zero matrix. Thus we have $\widetilde{B}_{t}^{T}{\bf a}^{t}=\widetilde{B}_{t}^{T}\Lambda_{t}{\bf g}^{t}=(S\mid{\bf 0}){\bf g}^{t}$. By comparing the $k$th component on both sides of the equality, we get
$\sum_{j=1}^{m}b_{jk;t}a_{j}^{t}=s_{k}g_{k}^{t}.$
Since $s_{k}$ is a positive integer, we get that $g_{k}^{t}\geq 0$ if and only if $\sum_{j=1}^{m}b_{jk;t}a_{j}^{t}\geq 0$. So we have $\varepsilon(g_{k}^{t})=\varepsilon_{0}$.
Hence, ${\bf a}^{t^{\prime}}=(E_{k,\varepsilon(g_{k}^{t})}^{\widetilde{B}_{t}})^{T}{\bf a}^{t}=(E_{k,\varepsilon_{0}}^{\widetilde{B}_{t}})^{T}{\bf a}^{t}$. By Lemma 4.2, we know the collection
$\{{\bf a}^{t}=\Lambda_{t}{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$
is a tropical point in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$. This completes the proof. ∎
4.2. Tropical invariant and $F$-invariant
Given a non-zero polynomial
$F=\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}{\bf y}^{{\bf v}}\in\mathbb{Z}[y_{1},\ldots,y_{n}]$
and a vector ${\bf r}\in\mathbb{Z}^{n}$, we denote by
$F[{\bf r}]:=\max\{{\bf v}^{T}{\bf r}\mid c_{{\bf v}}\neq 0\}\in\mathbb{Z}.$
We call the map $F[-]:\mathbb{Z}^{n}\rightarrow\mathbb{Z}$ a tropical polynomial.
Example 4.5.
Take $F=1+y_{1}+y_{1}y_{2}\in\mathbb{Z}[y_{1},y_{2}]$ and ${\bf r}=\begin{bmatrix}-2\\ 1\end{bmatrix}$. Then
$F[{\bf r}]=\max\{\begin{bmatrix}0,0\end{bmatrix}\begin{bmatrix}-2\\ 1\end{bmatrix},\begin{bmatrix}1,0\end{bmatrix}\begin{bmatrix}-2\\ 1\end{bmatrix},\begin{bmatrix}1,1\end{bmatrix}\begin{bmatrix}-2\\ 1\end{bmatrix}\}=\max\{0,-2,-1\}=0.$
The following proposition can be easily checked.
Proposition 4.6.
The following statements hold.
  • (i)
    If $0\neq F\in\mathbb{Z}[y_{1},\ldots,y_{n}]$ has constant term $1$, then $F[{\bf r}]\geq 0$ for any ${\bf r}\in\mathbb{Z}^{n}$.
  • (ii)
    If $0\neq F_{1},F_{2}\in\mathbb{Z}_{\geq 0}[y_{1},\ldots,y_{n}]$, then $(F_{1}F_{2})[{\bf r}]=F_{1}[{\bf r}]+F_{2}[{\bf r}]$ for any ${\bf r}\in\mathbb{Z}^{n}$.
Now we fix a Langlands-Poisson triple $(\mathcal{S}_{X},\mathcal{S}_{Y},\Lambda)$, and let $\mathcal{U}$ be the upper cluster algebra associated to the cluster pattern $\mathcal{S}_{X}$. Denote by $S=diag(s_{1},\ldots,s_{n})$ the type of the compatible pair $(\widetilde{B}_{t_{0}},\Lambda_{t_{0}})$, namely, $\widetilde{B}_{t_{0}}^{T}\Lambda_{t_{0}}=(S\mid{\bf 0})$, where ${\bf 0}$ is the $n\times(m-n)$ zero matrix. By Proposition 2.2, the mutation matrices of $\mathcal{U}$ always have full rank.
We denote by $\mathcal{U}^{[{\bf g}]}$ the set of $[{\bf g}]$-good elements in $\mathcal{U}$ and by
$\mathcal{U}^{\operatorname{\mathsf{good}}}:=\bigcup_{[{\bf g}]\in\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})}\mathcal{U}^{[\bf g]}$
the set of all good elements in $\mathcal{U}$. Notice that the union $\bigcup_{[{\bf g}]\in\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})}\mathcal{U}^{[\bf g]}$ is disjoint.
Definition 4.7 (Tropical pairing).
Let $u,u^{\prime}\in\mathcal{U}^{\operatorname{\mathsf{good}}}$ and consider their canonical expressions with respect to a vertex $w\in\mathbb{T}_{n}$:
$u={\bf x}_{w}^{{\bf g}_{u}^{w}}F_{u}^{w}(\hat{y}_{1;w},\ldots,\hat{y}_{n;w}),\;\;\;u^{\prime}={\bf x}_{w}^{{\bf g}_{u^{\prime}}^{w}}F_{u^{\prime}}^{w}(\hat{y}_{1;w},\ldots,\hat{y}_{n;w}).$
Define an integer $\langle u,u^{\prime}\rangle_{w}$ as follows:
$\langle u,u^{\prime}\rangle_{w}:=({\bf g}_{u}^{w})^{T}\Lambda_{w}{\bf g}_{u^{\prime}}^{w}+F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}].$
The pairing $\langle-,-\rangle_{w}:\mathcal{U}^{\operatorname{\mathsf{good}}}\times\mathcal{U}^{\operatorname{\mathsf{good}}}\rightarrow\mathbb{Z}$ is called the tropical pairing at vertex $w\in\mathbb{T}_{n}$.
Theorem 4.8.
For any two good elements $u$ and $u^{\prime}$, the integer $\langle u,u^{\prime}\rangle_{w}$ only depends on $u$ and $u^{\prime}$, not on the choice of $w\in\mathbb{T}_{n}$. We call $\langle u,u^{\prime}\rangle_{w}$ the tropical invariant between $u$ and $u^{\prime}$.
Proof.
For a vertex $t\in\mathbb{T}_{n}$, let $u={\bf x}_{t}^{{\bf g}_{u}^{t}}F_{u}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t})$ and $u^{\prime}={\bf x}_{t}^{{\bf g}_{u^{\prime}}^{t}}F_{u^{\prime}}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t})$ be the canonical expressions of $u$ and $u^{\prime}$ with respect to $t\in\mathbb{T}_{n}$. We know that
$[{\bf g}_{u^{\prime}}]:=\{{\bf g}_{u^{\prime}}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$
is a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$. By Proposition 4.4, the collection
$[{\bf a}]:=\{{\bf a}^{t}=\Lambda_{t}{\bf g}_{u^{\prime}}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$
is a tropical point in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$. By Corollary 3.5, the tropical point $[{\bf a}]\in\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$ corresponds to a unique semifield homomorphism
$\beta_{u^{\prime}}:\mathbb{F}_{>0}:=\mathbb{Q}_{\rm sf}(x_{1;t_{0}},\ldots,x_{m;t_{0}})\rightarrow\mathbb{Z}^{\rm max}$
such that $\beta_{u^{\prime}}({\bf x}_{t})=({\bf a}^{t})^{T}$ for any vertex $t\in\mathbb{T}_{n}$. Notice that the simifield homomorphism $\beta_{u^{\prime}}$ is uniquely determined by $u^{\prime}$.
By applying the semifield homomorphism $\beta_{u^{\prime}}$ to $u={\bf x}_{t}^{{\bf g}_{u}^{t}}F_{u}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t})\in\mathbb{F}_{>0}$, we see that
$\beta_{u^{\prime}}(u)=\beta_{u^{\prime}}({\bf x}_{t}^{{\bf g}_{u}^{t}})+\beta_{u^{\prime}}(F_{u}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t})).$
Since $\beta_{u^{\prime}}({\bf x}_{t})=({\bf a}^{t})^{T}=(\Lambda_{t}{\bf g}_{u^{\prime}}^{t})^{T}$, we have
$\beta_{u^{\prime}}({\bf x}_{t}^{{\bf g}_{u}^{t}})=(\Lambda_{t}{\bf g}_{u^{\prime}}^{t})^{T}{\bf g}_{u}^{t}=[(\Lambda_{t}{\bf g}_{u^{\prime}}^{t})^{T}{\bf g}_{u}^{t}]^{T}=({\bf g}_{u}^{t})^{T}\Lambda_{t}{\bf g}_{u^{\prime}}^{t}.$
We write $F_{u}^{t}=\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}{\bf y}^{{\bf v}}\in\mathbb{Z}_{\geq 0}[y_{1},\ldots,y_{n}]$, then we know that
$F_{u}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t})=\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}{\bf x}_{t}^{\widetilde{B}_{t}{\bf v}}.$
We see that
$\beta_{u^{\prime}}(\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}{\bf x}_{t}^{\widetilde{B}_{t}{\bf v}})$ $=$ $\max\{({\bf a}^{t})^{T}\widetilde{B}_{t}{\bf v}\mid c_{\bf v}\neq 0\}=\max\{[({\bf a}^{t})^{T}\widetilde{B}_{t}{\bf v}]^{T}\mid c_{\bf v}\neq 0\}$
$=$ $\max\{{\bf v}^{T}\widetilde{B}_{t}^{T}{\bf a}^{t}\mid c_{\bf v}\neq 0\}=\max\{{\bf v}^{T}\widetilde{B}_{t}^{T}\Lambda_{t}{\bf g}_{u^{\prime}}^{t}\mid c_{\bf v}\neq 0\}$
$=$ $\max\{{\bf v}^{T}(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{t}\mid c_{\bf v}\neq 0\}=F_{u}^{t}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{t}].$
Hence, $\beta_{u^{\prime}}(F_{u}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t}))=\beta_{u^{\prime}}(\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}{\bf x}_{t}^{\widetilde{B}_{t}{\bf v}})=F_{u}^{t}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{t}]$ and
$\beta_{u^{\prime}}(u)=\beta_{u^{\prime}}({\bf x}_{t}^{{\bf g}_{u}^{t}})+\beta_{u^{\prime}}(F_{u}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t}))=({\bf g}_{u}^{t})^{T}\Lambda_{t}{\bf g}_{u^{\prime}}^{t}+F_{u}^{t}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{t}]=\langle u,u^{\prime}\rangle_{t}.$
Since the semifield homomorphism $\beta_{u^{\prime}}$ is uniquely determined by $u^{\prime}$, we know that the integer $\beta_{u^{\prime}}(u)$ only depends on $u$ and $u^{\prime}$, not on the choice of an initial seed. The equality $\beta_{u^{\prime}}(u)=\langle u,u^{\prime}\rangle_{t}$ says that each choice of a vertex $t\in\mathbb{T}_{n}$ provides an explicit formula to calculate the integer $\beta_{u^{\prime}}(u)$. This completes the proof. ∎
Remark 4.9.
From the proof above, we see that the integer $\langle u,u^{\prime}\rangle_{w}$ can be interpreted as the evaluation of the positive function $u\in\mathbb{F}_{>0}$ (on cluster $\mathscr{A}$-variety in the sense of [FG-2009]) at the tropical point $[{\bf a}]:=\{{\bf a}^{t}=\Lambda_{t}{\bf g}_{u^{\prime}}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ (associated to the same cluster $\mathscr{A}$-variety) determined by $u^{\prime}$.
Since the tropical invariant $\langle u,u^{\prime}\rangle_{w}$ only depends on $u$ and $u^{\prime}$, not on the choice of $w\in\mathbb{T}_{n}$, we denote by $\langle u,u^{\prime}\rangle:=\langle u,u^{\prime}\rangle_{w}$ and call $\langle-,-\rangle:\mathcal{U}^{\operatorname{\mathsf{good}}}\times\mathcal{U}^{\operatorname{\mathsf{good}}}\rightarrow\mathbb{Z}$ the tropical pairing of $\mathcal{U}$.
Definition 4.10 ($F$-invariant and $F$-compatibility).
Let $u$ and $u^{\prime}$ be two good elements in $\mathcal{U}$.
  • (i)
    The $F$-invariant $(u\mid\mid u^{\prime})_{F}$ between $u$ and $u^{\prime}$ is defined to be the symmetrized sum of the tropical invariants:
    $(u\mid\mid u^{\prime})_{F}=\langle u,u^{\prime}\rangle+\langle u^{\prime},u\rangle=\langle u,u^{\prime}\rangle_{w}+\langle u^{\prime},u\rangle_{w},$
    where $w$ is any vertex in $\mathbb{T}_{n}$.
  • (ii)
    $u$ and $u^{\prime}$ are said to be $F$-compatible, if $(u\mid\mid u^{\prime})_{F}=0$.
Proposition 4.11.
Let $u$ and $u^{\prime}$ be two good elements in $\mathcal{U}$. The following statements hold.
  • (i)
    $(u\mid\mid u^{\prime})_{F}=F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}]+F_{u^{\prime}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]$, where $w$ is any vertex in $\mathbb{T}_{n}$.
  • (ii)
    $(u\mid\mid u^{\prime})_{F}\in\mathbb{Z}_{\geq 0}$.
  • (iii)
    If $u$ is a cluster monomial, then $(u\mid\mid u)_{F}=0$.
Proof.
(i) Let $w$ be a vertex in $\mathbb{T}_{n}$. By the definition of $(u\mid\mid u^{\prime})_{F}$, we see that
$(u\mid\mid u^{\prime})_{F}$ $=$ $\langle u,u^{\prime}\rangle_{w}+\langle u^{\prime},u\rangle_{w}$
$=$ $({\bf g}_{u}^{w})^{T}\Lambda_{w}{\bf g}_{u^{\prime}}^{w}+F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}]+({\bf g}_{u^{\prime}}^{w})^{T}\Lambda_{w}{\bf g}_{u}^{w}+F_{u^{\prime}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}].$
Since $\Lambda_{w}$ is skew-symmetric, we know that $({\bf g}_{u}^{w})^{T}\Lambda_{w}{\bf g}_{u^{\prime}}^{w}+({\bf g}_{u^{\prime}}^{w})^{T}\Lambda_{w}{\bf g}_{u}^{w}=0$. Thus
$(u\mid\mid u^{\prime})_{F}=F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}]+F_{u^{\prime}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}].$
(ii) Let $w$ be a vertex in $\mathbb{T}_{n}$. Since both $F_{u}^{w}$ and $F_{u^{\prime}}^{w}$ are polynomials with constant term $1$, we have
$F_{u}^{w}[{\bf h}],\;F_{u^{\prime}}^{w}[{\bf h}]\in\mathbb{Z}_{\geq 0}$
for any ${\bf h}\in\mathbb{Z}^{n}$. Then by (i), we have $(u\mid\mid u^{\prime})_{F}\in\mathbb{Z}_{\geq 0}$.
(iii) Say $u$ is a cluster monomial in seed $({{\bf x}_{w},\widetilde{B}_{w}})$. Then $F_{u}^{w}=1$, where $F_{u}^{w}$ is the $F$-polynomial of $u$ with respect to vertex $w\in\mathbb{T}_{n}$. Then by (i), we see that $(u\mid\mid u)_{F}=0$. ∎
Remark 4.12.
(i) The results in Proposition 4.11 (iii) are analogous to [DWZ10, Corollary 7.2] for the $E$-invariant in the additive categorification of cluster algebras and to [kkko-2018, Lemma 3.2.3] for the $\mathfrak{d}$-invariant in the monoidal categorification of cluster algebras.
(ii) The good elements satisfying $(u\mid\mid u)_{F}=0$ are analogous to the rigid objects in cluster categories and to the real simples in the monoidal categorification of cluster algebras.
Remark 4.13.
Recently, Fei introduced and studied the tropical $F$-polynomials in [fei_2019a, fei_2019b] from the viewpoint of representation theory. One can see that the $F$-invariant is the sum of the evaluations of two tropical $F$-polynomials. When restricting to cluster monomials, [fei_2019b, Theorem 3.22] implies that the $F$-invariant is essentially the same as the $E$-invariant in $\tau$-tilting theory. I will give more details about this in [cao_2025].
Remark 4.14.
(i) If we take $u=x_{i;t}$ and $u^{\prime}=x_{j;t}$, then ${\bf g}_{u}^{t}={\bf e}_{i},\;{\bf g}_{u^{\prime}}^{t}={\bf e}_{j}$ and $F_{u}^{t}=1$, where ${\bf e}_{k}$ is the $k$th column of $I_{m}$ for $k\in[1,m]$. By the definition of tropical invariant, we have
$\langle u,u^{\prime}\rangle=\langle u,u^{\prime}\rangle_{t}=\langle x_{i;t},x_{j;t}\rangle_{t}={\bf e}_{i}^{T}\Lambda_{t}{\bf e}_{j}+0=\lambda_{ij;t},$
where $\lambda_{ij;t}$ is the $(i,j)$-entry of $\Lambda_{t}$. So the tropical invariant $\langle x_{i;t},x_{j;t}\rangle$ is related to the quantization matrix $\Lambda_{t}$. This is analogous to [kkko-2018, Proposition 7.1.2 (a)] for the $\Lambda$-invariant $\Lambda(M,N)$ in the monoidal categorification of cluster algebras.
(ii) Notice that the $F$-invariant $(u\mid\mid u^{\prime})_{F}=\langle u,u^{\prime}\rangle+\langle u^{\prime},u\rangle$ in cluster algebras is defined to be the symmetrized sum of the tropical invariants while the $\mathfrak{d}$-invariant $\mathfrak{d}(M,N)=\frac{1}{2}(\Lambda(M,N)+\Lambda(N,M))$ in the monoidal categorification [kkko-2018] is defined to be half of the symmetrized sum of $\Lambda$-invariants. This suggests that the $F$-invariant in cluster algebras is related with the $\mathfrak{d}$-invariant in the monoidal categorification.
Proposition 4.15.
Let $x_{i;t}$ be a cluster variable of $\mathcal{U}$, where $i\in[1,m]$ and $t\in\mathbb{T}_{n}$. Then for any good element $u\in\mathcal{U}$, we have
$(x_{i;t}\mid\mid u)_{F}=\begin{cases}s_{i}f_{i;u}^{t},&i\in[1,n];\\ 0,&i\in[n+1,m],\end{cases}$
where $s_{i}$ is the $(i,i)$-entry of the diagonal matrix $S=diag(s_{1},\ldots,s_{n})$ and $f_{i;u}^{t}$ is the $i$th component of the $f$-vector of $u$ with respect vertex $t$.
Proof.
We know that $F_{x_{i;t}}^{t}=1$ and ${\bf g}_{x_{i;t}}^{t}={\bf e}_{i}$, where ${\bf e}_{i}$ is the $i$th column of $I_{m}$. Now we use the vertex $t$ to calculate $(x_{i;t}\mid\mid u)_{F}$ and by using the formula in Proposition 4.11 (i), we have
$(x_{i;t}\mid\mid u)_{F}=0+F_{u}^{t}[(S\mid{\bf 0}){\bf e}_{i}].$
Then the result follows. ∎
As we said that $F$-invariant can be viewed as an analog of the $E$-invariant in the additive categorification of cluster algebras and the $\mathfrak{d}$-invariant in the monoidal categorification of cluster algebras. We provide more reasons in the following remark.
Remark 4.16.
(i) By Proposition 4.15, we know that the $F$-invariant $(x_{i;t}\mid\mid u)_{F}$ is related to the components of $f$-vectors.
(ii) By the equality (1.1) in the introduction part, the $E$-invariant is related to the extension dimension in the cluster categories. Such dimensions are used to categorify the components of $f$-vectors.
(iii) The components of $f$-vectors measure the distance between the highest degree of some $y_{k}$ and the lowest degree of $y_{k}$ in $F$-polynomials, while the $\mathfrak{d}$-invariant has the similar meaning, cf., [kkko-2018, Corollary 4.2.4] and [qin-2020, Lemma 5.1].
Remark 4.17.
(i) The $F$-invariant $(-\mid\mid-)_{F}:\mathcal{U}^{\operatorname{\mathsf{good}}}\times\mathcal{U}^{\operatorname{\mathsf{good}}}\rightarrow\mathbb{Z}_{\geq 0}$ is a great improvement of the $f$-compatibility degree introduced by Fu and Gyoda in [Fu-Gyoda]. The $f$-compatibility degree, defined using the components of $f$-vectors of cluster variables, is a function
$(-\mid\mid-)_{f}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{Z}_{\geq 0},$
where $\mathcal{X}$ is the set of unfrozen cluster variables in $\mathcal{U}$. When restriced to $\mathcal{X}$, the $F$-invariant and the $f$-compatibility degree are related as follows:
$(x_{i;t}\mid\mid x_{j;w})_{F}=s_{i}(x_{i;t}\mid\mid x_{j;w})_{f}.$
(ii) By construction, $F$-invariant is symmetric, i.e., $(u\mid\mid u^{\prime})_{F}=(u^{\prime}\mid\mid u)_{F}$ for any two good elements in $\mathcal{U}$. Applying to the case $u=x_{i;t}$ and $u=x_{j;w}$ with $i,j\in[1,n]$, we get
$s_{i}(x_{i;t}\mid\mid x_{j;w})_{f}=(x_{i;t}\mid\mid x_{j;w})_{F}=(x_{j;w}\mid\mid x_{i;t})_{F}=s_{j}(x_{j;w}\mid\mid x_{i;t})_{f}.$
This recovers the result in [Fu-Gyoda, Proposition 4.14 (2)].
Corollary 4.18.
Let ${\bf x}_{t}$ be a cluster of $\mathcal{U}$ and $u\in\mathcal{U}^{\operatorname{\mathsf{good}}}$. The following statements hold.
  • (i)
    If $(x_{k;t}\mid\mid u)_{F}=0$ for some $k\in[1,n]$, then the $k$th component $g_{k}^{t}$ of ${\bf g}_{u}^{t}=\operatorname{\mathsf{deg}}^{t}(u)$ is non-negative.
  • (ii)
    If $(x_{i;t}\mid\mid u)_{F}=0$ holds for any $i\in[1,n]$, then $u$ is a cluster monomial in ${\bf x}_{t}$.
  • (iii)
    Let $k$ be an integer in $[1,n]$ and let $({\bf x}_{t^{\prime}},\widetilde{B}_{t^{\prime}})=\mu_{k}({\bf x}_{t},\widetilde{B}_{t})$. Suppose that $(x_{k;t}\mid\mid u)_{F}\neq 0$ but $(x_{i;t}\mid\mid u)_{F}=0$ for any $i\in[1,n]\backslash\{k\}$. Then $(x_{i;t^{\prime}}\mid\mid u)_{F}=0$ for any $i\in[1,n]$. In particular, $u$ is a cluster monomial in ${\bf x}_{t^{\prime}}$.
Proof.
The results follow from Proposition 4.15 and Corollary 3.14. ∎
Proposition 4.19.
Let $u$ be a $[{\bf g}]$-bigood element in $\mathcal{U}$, and let ${\bf x}_{t}^{\bf h}=\prod_{j=1}^{m}x_{j;t}^{h_{j}}$ be a cluster monomial of $\mathcal{U}$, i.e., ${\bf h}^{\circ}:=(h_{1},\ldots,h_{n})^{T}\in\mathbb{Z}_{\geq 0}^{n}$. Then
  • (i)
    $({\bf x}_{t}^{\bf h}\mid\mid u)_{F}=\sum_{j=1}^{n}h_{j}\cdot(x_{j;t}\mid\mid u)_{F}$.
  • (ii)
    $F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{{\bf x}_{t}^{\bf h}}^{w}]=\sum_{j=1}^{n}h_{j}\cdot F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{x_{j;t}}^{w}]$ for any vertex $w\in\mathbb{T}_{n}$.
Proof.
(i) Consider the canonical expression of $u$ with respect to the vertex $t$,
$u={\bf x}_{t}^{{\bf g}_{u}^{t}}F_{u}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t}),$
where ${\bf g}_{u}^{t}$ is the value of $[{\bf g}]$ at vertex $t$. Since $u$ is a $[{\bf g}]$-bigood element, $F_{u}^{t}$ has the form:
$F_{u}^{t}=1+\text{middle terms }+y_{1}^{f_{1}^{t}}\cdots y_{n}^{f_{n}^{t}},$
where ${\bf f}_{u}^{t}=(f_{1}^{t},\ldots,f_{n}^{t})^{T}$ is the $f$-vector of $u$ with respect to vertex $t$. Since ${\bf x}_{t}^{\bf h}$ is a cluster momomial, we know that $F_{{\bf x}_{t}^{\bf h}}^{t}=1$ and $(h_{1},\ldots,h_{n})\in\mathbb{Z}_{\geq 0}^{n}$. By using the chart ${\bf x}_{t}$ to calculate $({\bf x}_{t}^{\bf h}\mid\mid u)_{F}$, we have
$({\bf x}_{t}^{\bf h}\mid\mid u)_{F}$ $=$ $0+F_{u}^{t}[(S\mid{\bf 0}){\bf h}]=F_{u}^{t}[(s_{1}h_{1},\ldots,s_{n}h_{n})^{T}]$
$=$ $f_{1}^{t}s_{1}h_{1}+\ldots+f_{n}^{t}s_{n}h_{n}=\sum_{j=1}^{n}h_{j}\cdot(x_{j;t}\mid\mid u)_{F}.$
(ii) Let $w$ be a vertex in $\mathbb{T}_{n}$. We have
[…]
$=$ $\sum_{j=1}^{n}h_{j}\cdot\left(F_{x_{j;t}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]+F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{x_{j;t}}^{w}]\right)$
$=$ $\sum_{j=1}^{n}h_{j}\cdot F_{x_{j;t}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]+\sum_{j=1}^{n}h_{j}\cdot F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{x_{j;t}}^{w}]$
$=$ $F_{{\bf x}_{t}^{\bf h}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]+\sum_{j=1}^{n}h_{j}\cdot F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{x_{j;t}}^{w}].$
(The last equality comes from the facts that $F_{{\bf x}_{t}^{\bf h}}^{w}=\prod_{j=1}^{m}(F_{x_{j;t}}^{w})^{h_{j}}=\prod_{j=1}^{n}(F_{x_{j;t}}^{w})^{h_{j}}$ and that for any two polynomials $F_{1},F_{2}\in\mathbb{Z}_{\geq 0}[y_{1},\ldots,y_{n}]$ and any vector ${\bf r}\in\mathbb{Z}^{n}$, we have
$(F_{1}F_{2})[{\bf r}]=F_{1}[{\bf r}]+F_{2}[{\bf r}],$
by the definition of tropical polynomials.) On the other hand, we have
$({\bf x}_{t}^{\bf h}\mid\mid u)_{F}=F_{{\bf x}_{t}^{\bf h}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]+F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{{\bf x}_{t}^{\bf h}}^{w}].$
By comparing the two equalities, we obtain $F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{{\bf x}_{t}^{\bf h}}^{w}]=\sum_{j=1}^{n}h_{j}\cdot F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{x_{j;t}}^{w}]$. ∎
Lemma 4.20.
Let $x$ and $z$ be two unfrozen cluster variables of $\mathcal{U}$. If $(x\mid\mid z)_{F}=0$, then they are contained in the same cluster.
Proof.
Since $(x\mid\mid z)_{F}=0$ and by Remark 4.17, the $f$-compatibility degree $(x\mid\mid z)_{f}$ is zero. Then the result follows from [Fu-Gyoda, Theorem 3.3].
Let us give another proof here. Consider the tropical point
$[{\bf g}_{z}]:=\{{\bf g}_{z}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}\in\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$
corresponding to the cluster variable $z$. Since $(x\mid\mid z)_{F}=0$ and by Corollary 4.18 (i), we know that the $k$th component $g_{k;z}^{t}$ of ${\bf g}_{z}^{t}$ satisfies $g_{k;z}^{t}\geq 0$ whenever $x=x_{k;t}$.
On the other hand, by [cao-li-2020, Theorem 8], we know that there exists a seed $({\bf x}_{w},\widetilde{B}_{w})$ containing $x$ (say $x=x_{k;w}$) such that $g_{i;z}^{w}\geq 0$ for any $i\in[1,n]$ with $x_{i;t}\neq x$, i.e., for any $i\in[1,n]\setminus\{k\}$. Thus for the vertex $w$, we have $g_{j;z}^{w}\geq 0$ for $j=1,\ldots,n$, i.e., the principal part $({\bf g}_{z}^{t})^{\circ}$ of ${\bf g}_{z}^{t}$ is in $\mathbb{Z}_{\geq 0}^{n}$. Thus $u:={\bf x}_{w}^{{\bf g}_{z}^{w}}$ is a cluster monomial of $\mathcal{U}$. Hence, the two cluster monomials $u={\bf x}_{w}^{{\bf g}_{z}^{w}}$ and $z$ have the same extended $g$-vector ${\bf g}_{z}^{w}$ with respect to vertex $w$. It is well-known that different cluster monomials have different extended $g$-vectors, c.f., [GHKK18]. So we must have $z=u={\bf x}_{w}^{{\bf g}_{z}^{w}}$. This implies $z$ is a cluster variable in ${\bf x}_{w}$. In particular, $z$ and $x=x_{k;w}$ are contained in the same cluster ${\bf x}_{w}$. ∎
Lemma 4.21.
Let $\{z_{1},\ldots,z_{p}\}$ be a subset of unfrozen cluster variables of $\mathcal{U}$. If $(z_{i}\mid\mid z_{j})_{F}=0$ for any $i,j\in[1,p]$, then $\{z_{1},\ldots,z_{p}\}$ is a subset of some cluster of $\mathcal{U}$.
Proof.
By Lemma 4.20, we know that any two cluster variables in $\{z_{1},\ldots,z_{p}\}$ are contained in the same cluster. Then the result follows from [cao-li-2020, Theorem 13]. ∎
The following result says that $F$-invariant is a very concrete and easily computable numerical invariant to distinguish whether the product of two cluster monomials is still a cluster monomial.
Theorem 4.22.
Let $u$ and $u^{\prime}$ be two cluster monomials in $\mathcal{U}$. Then the product $uu^{\prime}$ is still a cluster monomial in $\mathcal{U}$ if and only if $(u\mid\mid u^{\prime})_{F}=0$, i.e.,
$F_{u}^{t}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{t}]+F_{u^{\prime}}^{t}[(S\mid{\bf 0}){\bf g}_{u}^{t}]=0,$
for some (equivalently, any) vertex $t\in\mathbb{T}_{n}$.
Proof.
$\Longrightarrow$”: Suppose that the product $uu^{\prime}$ is a cluster monomial in $\mathcal{U}$. Say, $uu^{\prime}$ is a cluster monomial in cluster ${\bf x}_{t}$. Then both $u$ and $u^{\prime}$ can be viewed as cluster monomials in ${\bf x}_{t}$. Thus $F_{u}^{t}=1=F_{u^{\prime}}^{t}$. So we have $(u\mid\mid u^{\prime})_{F}=F_{u}^{t}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{t}]+F_{u^{\prime}}^{t}[(S\mid{\bf 0}){\bf g}_{u}^{t}]=0$.
[…]
By Proposition 3.7 and Proposition 4.19, we have
$(u\mid\mid u^{\prime})_{F}=\sum_{i=1}^{n}\sum_{j=1}^{n}h_{i}r_{j}(x_{i;s}\mid\mid x_{j;s^{\prime}})_{F}.$
Since all the numbers $h_{i},r_{j}$ and $(x_{i;s}\mid\mid x_{j;s^{\prime}})_{F}$ are non-negative for $i,j\in[1,n]$, the assumption $(u\mid\mid u^{\prime})_{F}=0$ implies $h_{i}r_{j}(x_{i;s}\mid\mid x_{j;s^{\prime}})_{F}=0$ for any $i,j\in[1,n]$. Hence, $(x_{i;s}\mid\mid x_{j;s^{\prime}})_{F}=0$ for any $i,j\in[1,n]$ with $h_{i}r_{j}\neq 0$.
Let $V=\{x_{i;s}\mid i\in[1,n],h_{i}\neq 0\}$ and $V^{\prime}=\{x_{j;s^{\prime}}\mid j\in[1,n],r_{j}\neq 0\}$. We have $(z\mid\mid z^{\prime})_{F}=0$ for any $z\in V$ and $z^{\prime}\in V^{\prime}$. Since $V$ and $V^{\prime}$ are subsets of clusters of $\mathcal{U}$, we have
$(z_{1}\mid\mid z_{2})_{F}=0\;\;\;\text{and}\;\;\;(z_{1}^{\prime}\mid\mid z_{2}^{\prime})_{F}=0$
for any $z_{1},z_{2}\in V$ and $z_{1}^{\prime},z_{2}^{\prime}\in V^{\prime}$.
Hence, $(z\mid\mid z^{\prime})_{F}=0$ holds for any two cluster variables $z,z^{\prime}$ in $V\cup V^{\prime}$. Then by Lemma 4.21, we know that $V\cup V^{\prime}$ is a subset of some cluster ${\bf x}_{t}$ of $\mathcal{U}$. Hence, the product $uu^{\prime}$ is a cluster monomial in ${\bf x}_{t}$. ∎
Example 4.23.
Let continue with Example 3.8. Take $\Lambda=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}=\widetilde{B}$. We know that $\widetilde{B}^{T}\Lambda=(S\mid{\bf 0})=I_{2}$ and the canonical expressions of $x_{3},x_{4},x_{5}$ with respect to the initial seed $t_{0}=({\bf x},\widetilde{B})$ are given as follows:
$x_{3}={x_{1}^{-1}x_{2}\cdot(1+\widehat{y}_{1}),}\;\;\;x_{4}={x_{1}^{-1}\cdot(1+\widehat{y}_{1}+\widehat{y}_{1}\widehat{y}_{2}),}\;\;\;x_{5}={x_{2}^{-1}\cdot(1+\widehat{y}_{2})}.$
Thus we have
$(x_{3}\mid\mid x_{4})_{F}$ $=$ $F_{x_{3}}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{x_{4}}^{t_{0}}]+F_{x_{4}}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{x_{3}}^{t_{0}}]$
$=$ $(1+y_{1})\begin{bmatrix}-1\\ 0\end{bmatrix}+(1+y_{1}+y_{1}y_{2})\begin{bmatrix}-1\\ 1\end{bmatrix}$
[…]
$=$ $0$
$(x_{3}\mid\mid x_{5})_{F}$ $=$ $(1+y_{1})\begin{bmatrix}0\\ -1\end{bmatrix}+(1+y_{2})\begin{bmatrix}-1\\ 1\end{bmatrix}$
[…]
$=$ $1$
Then by Theorem 4.22, we know that $x_{3}x_{4}$ is a cluster monomial, while $x_{3}x_{5}$ is not.
Recall that the Newton polytope of a polynomial $F=\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}{\bf y}^{{\bf v}}\in\mathbb{Z}[y_{1},\ldots,y_{n}]$ is the convex hull of $\{{\bf v}\in\mathbb{N}^{n}\mid c_{\bf v}\neq 0\}$ in $\mathbb{R}^{n}$.
Corollary 4.24.
Let $u$ and $u^{\prime}$ be two cluster monomials in $\mathcal{U}$ and $t$ a vertex of $\mathbb{T}_{n}$. If the two $F$-polynomials $F_{u}^{t}$ and $F_{u^{\prime}}^{t}$ have the same Newton polytope, then the product $uu^{\prime}$ is a cluster monomial of $\mathcal{U}$.
Proof.
Since $u$ and $u^{\prime}$ are cluster monomials and by Proposition 4.11 (iii), we have
$(u\mid\mid u)_{F}=2F_{u}^{t}[(S\mid{\bf 0}){\bf g}_{u}^{t}]=0\;\;\;\text{and}\;\;\;(u^{\prime}\mid\mid u^{\prime})_{F}=2F_{u^{\prime}}^{t}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{t}]=0.$
We get $F_{u}^{t}[(S\mid{\bf 0}){\bf g}_{u}^{t}]=0=F_{u^{\prime}}^{t}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{t}]$. Since $F_{u}^{t}$ and $F_{u^{\prime}}^{t}$ have the same Newton polytope, we have $F_{u}^{t}[{\bf h}]=F_{u^{\prime}}^{t}[{\bf h}]$ for any ${\bf h}\in\mathbb{Z}^{n}$. Hence, we have
$(u\mid\mid u^{\prime})_{F}$ $=$ $F_{u}^{t}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{t}]+F_{u^{\prime}}^{t}[(S\mid{\bf 0}){\bf g}_{u}^{t}]$
$=$ $F_{u^{\prime}}^{t}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{t}]+F_{u}^{t}[(S\mid{\bf 0}){\bf g}_{u}^{t}]=0.$
Then by Theorem 4.22, we know that the product $uu^{\prime}$ is a cluster monomial. ∎
Definition 4.25.
Two tropical points $[{\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ and $[{\bf g}^{\prime}]=\{({\bf g}^{\prime})^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ are said to be sign-coherent, if for any vertex $t\in\mathbb{T}_{n}$ and any $k\in[1,n]$, we have $g_{k}^{t}\cdot(g_{k}^{\prime})^{t}\geq 0$, where $g_{k}^{t}$ and $(g_{k}^{\prime})^{t}$ are the $k$th components of ${\bf g}^{t}$ and $({\bf g}^{\prime})^{t}$.
Lemma 4.26.
Let $[{\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ and $[{\bf g}^{\prime}]=\{({\bf g}^{\prime})^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ be two tropical points in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$. If $[{\bf g}]$ and $[{\bf g}^{\prime}]$ are sign-coherent, then the collection $\{{\bf g}^{t}+({\bf g}^{\prime})^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ is still a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$. We denote it by $[{\bf g}\uplus{\bf g}^{\prime}]$.
Proof.
It suffices to show that the collection $\{{\bf g}^{t}+({\bf g}^{\prime})^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ satisfies the relations in (3.1). Thanks to the sign-coherence assumption, this is clear. ∎
Theorem 4.27.
Let $u$ be a $[{\bf g}]$-good element in $\mathcal{U}$ and $u^{\prime}$ a $[{\bf g}^{\prime}]$-good element in $\mathcal{U}$. Suppose that $(u\mid\mid u^{\prime})_{F}=0$. Then the following statements hold.
  • (i)
    The two tropical points $[{\bf g}]$ and $[{\bf g}^{\prime}]$ are sign-coherent. In particular, the tropical point $[{\bf g}\uplus{\bf g}^{\prime}]$ can be defined.
  • (ii)
    The product $uu^{\prime}$ is a $[{\bf g}\uplus{\bf g}^{\prime}]$-good element in $\mathcal{U}$.
Proof.
(i) We write $[{\bf g}]=\{{\bf g}_{u}^{t}=(g_{1;u}^{t},\ldots,g_{m;u}^{t})^{T}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ and $[{\bf g}^{\prime}]=\{{\bf g}_{u^{\prime}}^{t}=(g_{1;u^{\prime}}^{t},\ldots,g_{m;u^{\prime}}^{t})^{T}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$. Assume by contradiction that the two tropical points $[{\bf g}]$ and $[{\bf g}^{\prime}]$ are not sign-coherent. Then there exist some vertex $s\in\mathbb{T}_{n}$ and some $k\in[1,n]$ such that $g_{k;u}^{s}\cdot g_{k;u^{\prime}}^{s}<0$. Without loss of generality, we can assume $g_{k;u}^{s}<0$ and $g_{k;u^{\prime}}^{s}>0$.
By Corollary 3.13, we know that the monomial ${\bf y}^{[-g_{k;u}^{s}]_{+}{\bf e}_{k}}=y_{k}^{[-g_{k;u}^{s}]_{+}}$ appears in the $F$-polynomial $F_{u}^{s}\in\mathbb{Z}_{\geq 0}[y_{1},\ldots,y_{n}]$ with coefficient $1$, where ${\bf e}_{k}$ is the $k$th column of $I_{n}$. Thus we have
$F_{u}^{s}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{s}]\geq([-g_{k;u}^{s}]_{+}{\bf e}_{k})^{T}(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{s}=[-g_{k;u}^{s}]_{+}\cdot s_{k}g_{k;u^{\prime}}^{s}.$
Then by $g_{k;u}^{s}<0$, $g_{k;u^{\prime}}^{s}>0$ and $s_{k}>0$, we get $F_{u}^{s}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{s}]>0$.
Using the chart ${\bf x}_{s}$ to calculate the $F$-invariant $(u\mid\mid u^{\prime})_{F}$, we have
$(u\mid\mid u^{\prime})_{F}=F_{u}^{s}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{s}]+F_{u^{\prime}}^{s}[(S\mid{\bf 0}){\bf g}_{u}^{s}]\geq F_{u}^{s}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{s}]>0,$
which contradicts $(u\mid\mid u^{\prime})_{F}=0$. Hence, the two tropical points $[{\bf g}]$ and $[{\bf g}^{\prime}]$ must be sign-coherent.
(ii) For any vertex $t$, consider the Laurent expansions of $u$ and $u^{\prime}$ with respect to vertex $t$.
$u={\bf x}_{t}^{{\bf g}_{u}^{t}}F_{u}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t}),\;\;\;u^{\prime}={\bf x}_{t}^{{\bf g}_{u^{\prime}}^{t}}F_{u^{\prime}}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t}),$
where $F_{u}^{t}$ and $F_{u^{\prime}}^{t}$ are polynomials in $\mathbb{Z}_{\geq 0}[y_{1},\ldots,y_{n}]$ with constant term $1$. Hence, we have
$uu^{\prime}={\bf x}_{t}^{{\bf g}_{u}^{t}+{\bf g}_{u^{\prime}}^{t}}H_{uu^{\prime}}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t}),$
where $H_{uu^{\prime}}^{t}:=F_{u}^{t}F_{u^{\prime}}^{t}$ is a polynomial in $\mathbb{Z}_{\geq 0}[y_{1},\ldots,y_{n}]$ with constant term $1$.
By (i) and Lemma 4.26, we know that $[{\bf g}\uplus{\bf g}^{\prime}]:=\{{\bf g}_{u}^{t}+{\bf g}_{u^{\prime}}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ forms a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$. Then by the definition of good elements, we know that the product $uu^{\prime}$ is a $[{\bf g}\uplus{\bf g}^{\prime}]$-good element in $\mathcal{U}$. ∎
4.3. Log-canonicality and $F$-compatibility
In this subsection, we discuss the link between the compatible Poisson structure of (upper) cluster algebra and the $F$-compatibility.
As before, we fix a Langlands-Poisson triple $(\mathcal{S}_{X},\mathcal{S}_{Y},\Lambda)$, and let $\mathcal{U}$ be the upper cluster algebra associated to the cluster pattern $\mathcal{S}_{X}$. Denote by $S=diag(s_{1},\ldots,s_{n})$ the type of the compatible pair $(\widetilde{B}_{t_{0}},\Lambda_{t_{0}})$, namely, $\widetilde{B}_{t_{0}}^{T}\Lambda_{t_{0}}=(S\mid{\bf 0})$, where ${\bf 0}$ is the $n\times(m-n)$ zero matrix.
[…]
$\{x_{i;t_{0}},x_{j;t_{0}}\}:=\lambda_{ij;t_{0}}\cdot x_{i;t_{0}}x_{j;t_{0}}.$
This Poisson bracket is compatible with the cluster pattern $\mathcal{S}_{X}$, that is, for any cluster ${\bf x}_{t}$ of $\mathcal{S}_{X}$, we have
$\{x_{i;t},x_{j;t}\}=\lambda_{ij;t}\cdot x_{i;t}x_{j;t},$
where $\lambda_{ij;t}$ is the $(i,j)$-entry of $\Lambda_{t}$, cf. [gsv-2003], [bz-2005, Remark 4.6].
[…]
$\{f_{1},f_{2}\}=cf_{1}f_{2}$
for some $c\in\mathbb{Q}$.
Clearly, any two cluster variables in the same cluster are log-canonical. In Theorem 4.29 (ii) we will give the converse result.
Lemma 4.28.
Let $w$ be a vertex of $\mathbb{T}_{n}$ and $u$ an element in $\mathbb{F}$ with the form
$u={\bf x}_{w}^{\bf g}F(\hat{y}_{1;w},\ldots,\hat{y}_{n;w}),$
where $F=\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}{\bf y}^{{\bf v}}$ is a polynomial in $\mathbb{Z}[y_{1},\ldots,y_{n}]$ with constant term $1$. If $u$ and $x_{k;w}$ are log-canonical for some $k\in[1,n]$, then $f_{k}=0$, where $f_{k}$ is the highest degree of $y_{k}$ in $F$.
Proof.
[…]
$\{\hat{y}_{i;w},x_{j;w}\}$ $=$ $\{{\bf x}_{w}^{\widetilde{B}_{w}{\bf e}_{i}},{\bf x}_{w}^{\widetilde{\bf e}_{j}}\}=\left((\widetilde{B}_{w}{\bf e}_{i})^{T}\Lambda_{w}\widetilde{\bf e}_{j}\right)\hat{y}_{i;w}x_{j;w}$
$=$ $\left({\bf e}_{i}^{T}(\widetilde{B}_{w})^{T}\Lambda_{w}\widetilde{\bf e}_{j}\right)\hat{y}_{i;w}x_{j;w}=\left({\bf e}_{i}^{T}(S\mid{\bf 0})\widetilde{\bf e}_{j}\right)\hat{y}_{i;w}x_{j;w}$
$=$ $\begin{cases}s_{j}\cdot\hat{y}_{j;w}x_{j;w},&i=j;\\ 0,&i\neq j.\end{cases}$
Now we consider the integer $k\in[1,n]$ such that $u$ and $x_{k;w}$ are log-canonical. For ${\bf v}=(v_{1},\ldots,v_{n})^{T}\in\mathbb{N}^{n}$, we have
$\{\widehat{\bf y}_{w}^{\bf v},x_{k;w}\}=s_{k}(v_{k}\widehat{\bf y}_{w}^{\bf v})x_{k;w}.$
For simplicity, we write $F(\widehat{\bf y}_{w})=F(\hat{y}_{1;w},\ldots,\hat{y}_{n;w})$. We have
$\{F(\widehat{\bf y}_{w}),x_{k;w}\}=\{\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}\widehat{\bf y}_{w}^{{\bf v}},x_{k;w}\}=s_{k}(\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}v_{k}\widehat{\bf y}_{w}^{\bf v})x_{k;w}.$
Now let us calculate $\{u,x_{k;w}\}=\{{\bf x}_{w}^{\bf g}F(\widehat{\bf y}_{w}),x_{k;w}\}$.
$\{{\bf x}_{w}^{\bf g}F(\widehat{\bf y}_{w}),x_{k;w}\}$ $=$ $\{{\bf x}_{w}^{\bf g},x_{k;w}\}F(\widehat{\bf y}_{w})+{\bf x}_{w}^{\bf g}\{F(\widehat{\bf y}_{w}),x_{k;w}\}$
$=$ $({\bf g}^{T}\Lambda_{w}\widetilde{\bf e}_{k})\cdot{\bf x}_{w}^{\bf g}\cdot x_{k;w}\cdot F(\widehat{\bf y}_{w})+{\bf x}_{w}^{\bf g}\cdot s_{k}(\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}v_{k}\widehat{\bf y}_{w}^{\bf v})x_{k;w}.$
Since $u$ and $x_{k;w}$ are log-canonical, there exists some $c\in\mathbb{Q}$ such that
$\{u,x_{k;w}\}=c\cdot ux_{k;w}=c\cdot{\bf x}_{w}^{\bf g}\cdot F(\widehat{\bf y}_{w})\cdot x_{k;w}.$
Thus we have the following equality in $\mathbb{F}$.
$({\bf g}^{T}\Lambda_{w}\widetilde{\bf e}_{k})\cdot{\bf x}_{w}^{\bf g}\cdot x_{k;w}\cdot F(\widehat{\bf y}_{w})+{\bf x}_{w}^{\bf g}\cdot s_{k}(\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}v_{k}\widehat{\bf y}_{w}^{\bf v})x_{k;w}=c\cdot{\bf x}_{w}^{\bf g}\cdot F(\widehat{\bf y}_{w})\cdot x_{k;w}.$
We get
(4.1) $s_{k}(\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}v_{k}\widehat{\bf y}_{w}^{\bf v})=(c-{\bf g}^{T}\Lambda_{w}\widetilde{\bf e}_{k})\cdot F(\widehat{\bf y}_{w}).$
Assume by contradiction that the highest degree $f_{k}$ of $y_{k}$ in $F=\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}{\bf y}^{{\bf v}}$ is non-zero. Then the left side of (4.1) is non-zero, which implies $c-{\bf g}^{T}\Lambda_{w}\widetilde{\bf e}_{k}\neq 0$. Now we compare the constant term on both sides of (4.1). The constant term on the left side is 00 while the constant term on the right side is $(c-{\bf g}^{T}\Lambda_{w}\widetilde{\bf e}_{k})\cdot 1=c-{\bf g}^{T}\Lambda_{w}\widetilde{\bf e}_{k}\neq 0$. This is a contradiction. Hence, $f_{k}=0$. ∎
Theorem 4.29.
Let $u\in\mathcal{U}^{\operatorname{\mathsf{good}}}$ and $w\in\mathbb{T}_{n}$. The following statements hold.
  • (i)
    If $u$ and $x_{k;w}$ are log-canonical for some $k\in[1,n]$, then $(x_{k;w}\mid\mid u)_{F}=0$.
  • (ii)
    If $u$ is a cluster monomial and it is log-canonical with $x_{k;w}$, then the product $u\cdot x_{k;w}$ is a cluster monomial of $\mathcal{U}$. In particular, if two cluster variables are log-canonical, then they are contained in the same cluster.
  • (iii)
    If $u$ and $x_{k;w}$ are log-canonical for any $k\in[1,n]$, then $u$ is a cluster monomial in ${\bf x}_{w}$.
  • (iv)
    Let $k$ be an integer in $[1,n]$ and let $({\bf x}_{t},\widetilde{B}_{t})=\mu_{k}({\bf x}_{w},\widetilde{B}_{w})$. Suppose that $u$ and $x_{k;w}$ are not log-canonical but $u$ and $x_{i;w}$ are log-canonical for any $i\in[1,n]\backslash\{k\}$, then $u$ is a cluster monomial in ${\bf x}_{t}$.
Proof.
(i) This follows from Lemma 4.28 and Proposition 4.15.
(ii) By (i), we have $(x_{k;u}\mid\mid u)_{F}=0$. Since $u$ is a cluster monomial and by Theorem 4.22, we obtain that the product $u\cdot x_{k;w}$ is a cluster monomial of $\mathcal{U}$.
(iii) and (iv) follow from (i) and Corollary 4.18. ∎
5. Dominant sets and oriented exchange graphs
In this section, we fix a Langlands-Poisson triple $(\mathcal{S}_{X},\mathcal{S}_{Y},\Lambda)$, and let $\mathcal{A}$ be the cluster algebra associated to the cluster pattern $\mathcal{S}_{X}$. We fix a rooted vertex $t_{0}$ of $\mathbb{T}_{n}$, that is, we fix an initial seed $({\bf x}_{t_{0}},\widetilde{B}_{t_{0}})$ of $\mathcal{A}$. We will first introduce the dominant sets for seeds of cluster algebras as an replacement of torsion classes for $\tau$-tilting pairs in $\tau$-tilting theory and as an replacement of inversion sets in the study of right weak order on Weyl groups. Then we prove the main results on oriented exchange graphs.
5.1. Oriented exchange graphs are acyclic
Two seeds $({\bf x}_{t},\widetilde{B}_{t})$ and $({\bf x}_{w},\widetilde{B}_{w})$ of $\mathcal{A}$ are said to be equivalent if they are the same up to a relabeling. We denote by $[{\bf x}_{t},\widetilde{B}_{t}]$ the equivalent class of $({\bf x}_{t},\widetilde{B}_{t})$.
The exchange graph $\mathcal{H}=\mathcal{H}(\mathcal{A})$ of $\mathcal{A}$ is the graph whose vertices correspond to the seeds (up to seed equivalence) of $\mathcal{A}$ and whose edges correspond to seed mutations.
For a seed $({\bf x}_{t},\widetilde{B}_{t})$ of $\mathcal{A}$, we denote by $C_{t}=C_{t}^{t_{0}}$ the $C$-matrix of the vertex $t$ with respect to the rooted vertex $t_{0}$.
Definition 5.1 (Green mutation and exchange quiver).
(i) A seed mutation $\mu_{k}({\bf x}_{t},\widetilde{B}_{t})$ in $\mathcal{A}$ is called a green mutation, if the $k$th column of the $C$-matrix $C_{t}=C_{t}^{t_{0}}$ is a non-negative vector. Otherwise, it is called a red mutation.
(ii) The exchange quiver (or oriented exchange graph) of $\mathcal{A}$ with initial seed $({\bf x}_{t_{0}},\widetilde{B}_{t_{0}})$ is the quiver $\overrightarrow{\mathcal{H}}^{t_{0}}=\overrightarrow{\mathcal{H}}^{t_{0}}(\mathcal{A})$ whose vertices correspond to the seeds (up to seed equivalence) of $\mathcal{A}$ and whose arrows correspond to green mutations.
We remark that if two seeds $({\bf x}_{t},\widetilde{B}_{t})$ and $({\bf x}_{w},\widetilde{B}_{w})$ are related by once mutation, then either the mutation from $({\bf x}_{t},\widetilde{B}_{t})$ to $({\bf x}_{w},\widetilde{B}_{w})$ is a green mutation or the mutation from $({\bf x}_{w},\widetilde{B}_{w})$ to $({\bf x}_{t},\widetilde{B}_{t})$ is a green mutation.
For a vector ${\bf h}=(h_{1},\ldots,h_{m})^{T}\in\mathbb{Z}^{m}$, we denote by ${\bf h}^{\circ}:=(h_{1},\ldots,h_{n})^{T}\in\mathbb{Z}^{n}$ the principal part of ${\bf h}$.
Denote by $\widetilde{\mathcal{X}}$ the set of cluster variables (frozen and unfrozen) of $\mathcal{A}$ and by $\mathcal{X}$ the set of unfrozen cluster variables of $\mathcal{A}$. Given a cluster monomial $u={\bf x}_{t}^{\bf h}$ in ${\bf x}_{t}$, that is, ${\bf h}^{\circ}=(h_{1},\ldots,h_{n})\in\mathbb{Z}_{\geq 0}^{n}$, we define two subsets of $\mathcal{X}$.
  • The subset $\operatorname{\mathsf{supp}}(u)=\{x_{i;t}\mid i\in[1,n],\;h_{i}\neq 0\}\subseteq\mathcal{X}$ is called the support of $u={\bf x}_{t}^{\bf h}$. We remark that this set only depends on $u$, not on the choice of $t\in\mathbb{T}_{n}$ and ${\bf h}\in\mathbb{Z}^{m}$ such that $u={\bf x}_{t}^{\bf h}$.
  • For a vertex $w\in\mathbb{T}_{n}$ and a cluster variable $z$, recall that
    $(z\mid\mid u)_{F}=F_{z}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]+F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{z}^{w}].$
    The subset $\operatorname{\mathsf{dom}}^{w}(u)=\{z\in\mathcal{X}\mid F_{z}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]=0\}\subseteq\mathcal{X}$ is called the dominant set of $u={\bf x}_{t}^{\bf h}$ with respect to vertex $w\in\mathbb{T}_{n}$.
Definition 5.2 (Dominant set of a seed).
Let $({\bf x}_{t},\widetilde{B}_{t})$ be a seed of $\mathcal{A}$. The dominant set $\operatorname{\mathsf{dom}}^{w}[t]$ of $({\bf x}_{t},\widetilde{B}_{t})$ with respect to a vertex $w\in\mathbb{T}_{n}$ is defined to be the dominant set of the multiplicity free cluster monomial $u_{t}=\prod_{i=1}^{n}x_{i;t}$ supported on the unfrozen part of ${\bf x}_{t}$, i.e.,
$\operatorname{\mathsf{dom}}^{w}[t]:=\operatorname{\mathsf{dom}}^{w}(u_{t})=\{z\in\mathcal{X}\mid F_{z}^{w}[(S\mid{\bf 0}){\bf g}_{u_{t}}^{w}]=0\}.$
Clearly, if $[{\bf x}_{t^{\prime}},\widetilde{B}_{t^{\prime}}]=[{\bf x}_{t},\widetilde{B}_{t}]$, then $\operatorname{\mathsf{dom}}^{w}[t^{\prime}]=\operatorname{\mathsf{dom}}^{w}[t]$. Recall that for a cluster ${\bf x}_{t}$, we denote by ${\bf x}_{t}^{\circ}=(x_{1;t},\ldots,x_{n;t})$ the unfrozen part of ${\bf x}_{t}=(x_{1;t},\ldots,x_{m;t})$. If we want to view ${\bf x}_{t},{\bf x}_{t}^{\circ}$ as sets, we will write
$[{\bf x}_{t}]:=\{x_{i;t}\mid i\in[1,m]\}\;\;\;\text{and}\;\;\;[{\bf x}_{t}^{\circ}]:=\{x_{j;t}\mid j\in[1,n]\}.$
Proposition 5.3.
Let $u={\bf x}_{t}^{\bf h}$ be a cluster monomial of $\mathcal{A}$. Then the following statements hold for any vertex $w\in\mathbb{T}_{n}$.
  • (i)
    $[{\bf x}_{w}^{\circ}]\subseteq\operatorname{\mathsf{dom}}^{w}(u)$ and in particular, $[{\bf x}_{w}^{\circ}]\subseteq\operatorname{\mathsf{dom}}^{w}[t]$.
  • (ii)
    $[{\bf x}_{t}^{\circ}]\subseteq\operatorname{\mathsf{dom}}^{w}(u)$ and in particular, $[{\bf x}_{t}^{\circ}]\subseteq\operatorname{\mathsf{dom}}^{w}[t]$.
  • (iii)
    $\operatorname{\mathsf{dom}}^{w}(u)=\bigcap_{x\in\operatorname{\mathsf{supp}}(u)}\operatorname{\mathsf{dom}}^{w}(x)$ and in particular, $\operatorname{\mathsf{dom}}^{w}[t]=\bigcap_{i=1}^{n}\operatorname{\mathsf{dom}}^{w}(x_{i;t})$.
Proof.
(i) Since $F_{x_{i;w}}^{w}=1$ for $i\in[1,n]$, we have $F_{x_{i;w}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]=0$. Thus $x_{i;w}\in\operatorname{\mathsf{dom}}^{w}(u)$ for $i\in[1,n]$, that is, $[{\bf x}_{w}^{\circ}]\subseteq\operatorname{\mathsf{dom}}^{w}(u)$. In particular, if we take $u=u_{t}=\prod_{i=1}^{n}x_{i;t}$, we can get
$[{\bf x}_{w}^{\circ}]\subseteq\operatorname{\mathsf{dom}}^{w}(u_{t})=\operatorname{\mathsf{dom}}^{w}[t].$
(ii) Fix $i\in[1,n]$. By using the vertex $t$ to calculate $(x_{i;t}\mid\mid u)_{F}=(x_{i;t}\mid\mid{\bf x}_{t}^{\bf h})_{F}$, we see $(x_{i;t}\mid\mid u)_{F}=0$. On the other hand, we know that
$(x_{i;t}\mid\mid u)_{F}=F_{x_{i;t}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]+F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{x_{i;t}}^{w}]$
and each term in this sum is non-negative. Thus we obtain $F_{x_{i;t}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]=0$, i.e., $x_{i;t}\in\operatorname{\mathsf{dom}}^{w}(u)$. Since $i$ is an arbitrary integer in $[1,n]$, we get $[{\bf x}_{t}^{\circ}]\subseteq\operatorname{\mathsf{dom}}^{w}(u)$. In particular, if we take $u=u_{t}=\prod_{i=1}^{n}x_{i;t}$, we can get $[{\bf x}_{t}^{\circ}]\subseteq\operatorname{\mathsf{dom}}^{w}(u_{t})=\operatorname{\mathsf{dom}}^{w}[t]$.
(iii) By definition, an unfrozen cluster variable $z\in\mathcal{X}$ belongs to $\operatorname{\mathsf{dom}}^{w}(u)$ if and only if
$F_{z}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]=0.$
By Proposition 4.19 (ii), we have
$F_{z}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]=F_{z}^{w}[(S\mid{\bf 0}){\bf g}_{{\bf x}_{t}^{\bf h}}^{w}]=\sum_{j=1}^{n}h_{j}\cdot F_{z}^{w}[(S\mid{\bf 0}){\bf g}_{x_{j;t}}^{w}].$
Since $h_{j}\geq 0$ and $F_{z}^{w}[(S\mid{\bf 0}){\bf g}_{x_{j;t}}^{w}]\geq 0$ for any $j\in[1,n]$, we know that $F_{z}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]=0$ if and only if $F_{z}^{w}[(S\mid{\bf 0}){\bf g}_{x_{j;t}}^{w}]=0$ for any $j\in[1,n]$ with $h_{j}\neq 0$, i.e., for any $x_{j;t}\in\operatorname{\mathsf{supp}}(u)$. This is equivalent to say that $z$ belongs to $\bigcap_{x\in\operatorname{\mathsf{supp}}(u)}\operatorname{\mathsf{dom}}^{w}(x)$. In particular, if we take $u=u_{t}=\prod_{i=1}^{n}x_{i;t}$, we see that $\operatorname{\mathsf{dom}}^{w}[t]=\operatorname{\mathsf{dom}}^{w}(u_{t})=\bigcap_{i=1}^{n}\operatorname{\mathsf{dom}}^{w}(x_{i;t})$. ∎
Corollary 5.4.
Let $u$ and $u^{\prime}$ be two cluster monomials of $\mathcal{A}$ such that $\operatorname{\mathsf{supp}}(u)=\operatorname{\mathsf{supp}}(u^{\prime})$. Then $\operatorname{\mathsf{dom}}^{w}(u)=\operatorname{\mathsf{dom}}^{w}(u^{\prime})$ for any vertex $w\in\mathbb{T}_{n}$. In particular, $\operatorname{\mathsf{dom}}^{w}[t]=\operatorname{\mathsf{dom}}^{w}(u)$ for any cluster monomial $u$ with $\operatorname{\mathsf{supp}}(u)=[{\bf x}_{t}^{\circ}]$.
Proof.
This follows from Proposition 5.3 (iii). ∎
For a seed $({\bf x}_{t},\widetilde{B}_{t})$ of $\mathcal{A}$, we denote by $G_{t}=G_{t}^{t_{0}}$ the $G$-matrix of the vertex $t$ with respect to the rooted vertex $t_{0}$ and by
$\mathbb{R}_{\geq 0}G_{t}:=\mathbb{R}_{\geq 0}{\bf g}_{1;t}+\ldots+\mathbb{R}_{\geq 0}{\bf g}_{n;t}\subseteq\mathbb{R}^{n}$
the cone in $\mathbb{R}^{n}$ spanned by the columns of $G_{t}=({\bf g}_{1;t},\ldots,{\bf g}_{n;t})$. We remark that ${\bf g}_{i;t}\in\mathbb{Z}^{n}$ is the principal part of ${\bf g}_{x_{i;t}}^{t_{0}}=\deg^{t_{0}}(x_{i;t})\in\mathbb{Z}^{m}$, i.e., ${\bf g}_{i;t}=({\bf g}_{x_{i;t}}^{t_{0}})^{\circ}$.
We also denote by
$\mathbb{Q}_{>0}G_{t}$ ::=:= $\mathbb{Q}_{>0}{\bf g}_{1;t}+\ldots+\mathbb{Q}_{>0}{\bf g}_{n;t}\subseteq\mathbb{Q}^{n},$
$\mathbb{Z}_{>0}G_{t}$ ::=:= $\mathbb{Z}_{>0}{\bf g}_{1;t}+\ldots+\mathbb{Z}_{>0}{\bf g}_{n;t}\subseteq\mathbb{Z}^{n}.$
Lemma 5.5.
Let $({\bf x}_{t^{\prime}},\widetilde{B}_{t^{\prime}})=\mu_{k}({\bf x}_{t},\widetilde{B}_{t})$ be a green mutation in $\mathcal{A}$. Then there exists ${\bf v},{\bf v}^{\prime}\in\mathbb{Z}_{>0}^{n}$ such that $G_{t}{\bf v}-G_{t^{\prime}}{\bf v}^{\prime}\in\mathbb{Z}_{\geq 0}^{n}$.
Proof.
Without loss of generality, we can assume that $k=1$. We write $G_{t}=({\bf g}_{1},{\bf g}_{2},\ldots,{\bf g}_{n})$, where ${\bf g}_{i}\in\mathbb{Z}^{n}$. Since $({\bf x}_{t^{\prime}},\widetilde{B}_{t^{\prime}})=\mu_{k}({\bf x}_{t},\widetilde{B}_{t})=\mu_{1}({\bf x}_{t},\widetilde{B}_{t})$, the $G$-matrix $G_{t^{\prime}}$ has the form $G_{t^{\prime}}=({\bf g}_{1}^{\prime},{\bf g}_{2},\ldots,{\bf g}_{n})$.
By Theorem 3.10 (i), we know that $\mathbb{R}_{\geq 0}G_{t}$ and $\mathbb{R}_{\geq 0}G_{t^{\prime}}$ are two cones of dimension $n$ in $\mathbb{R}^{n}$ and they have a common facet $\mathcal{C}$ (cone of dimension $n-1$) given by $\mathcal{C}=\mathbb{R}_{\geq 0}{\bf g}_{2}+\ldots+\mathbb{R}_{\geq 0}{\bf g}_{n}$.
Let $S=diag(s_{1},\ldots,s_{n})$ be the type of the compatible pair $(\widetilde{B}_{t_{0}},\Lambda_{t_{0}})$, i.e., $\widetilde{B}_{t_{0}}^{T}\Lambda_{t_{0}}=(S\mid{\bf 0})$. We know that $S$ is a skew-symmetrizer of $B_{t_{0}}$. Let ${\bf c}_{1;t}$ be the first column of the $C$-matrix $C_{t}=C_{t}^{t_{0}}$.
By Theorem 3.10 (iii), we have $SC_{t}S^{-1}(G_{t})^{\rm T}=I_{n}$, equivalently, $(G_{t})^{\rm T}(SC_{t}S^{-1})=I_{n}$. So the first column vector $s_{1}^{-1}(S{\bf c}_{1;t})$ of $SC_{t}S^{-1}$ is an inner normal vector of the facet $\mathcal{C}$ of the cone $\mathbb{R}_{\geq 0}G_{t}$. So when we go along the direction $s_{1}^{-1}(S{\bf c}_{1;t})$, we can cross the common facet $\mathcal{C}$ from the cone $\mathbb{R}_{\geq 0}G_{t^{\prime}}$ to the cone $\mathbb{R}_{\geq 0}G_{t}$.
Now we take a vector ${\bf h}^{\prime}\in\mathbb{Z}_{>0}G_{t^{\prime}}\subseteq\mathbb{R}_{\geq 0}G_{t^{\prime}}$, there exists $\lambda\in\mathbb{Q}_{>0}$ such that
${\bf h}:={\bf h}^{\prime}+\lambda(s_{1}^{-1}(S{\bf c}_{1;t}))\in\mathbb{Q}_{>0}G_{t}\subseteq\mathbb{R}_{\geq 0}G_{t}.$
We can take a positive integer $p$ large enough such that ${\bf a}:=p{\bf h}\in\mathbb{Z}_{>0}G_{t}$ and set ${\bf a}^{\prime}=p{\bf h}^{\prime}\in\mathbb{Z}_{>0}G_{t^{\prime}}$. Then we have $p\lambda s_{1}^{-1}(S{\bf c}_{1;t})={\bf a}-{\bf a}^{\prime}\in\mathbb{Z}^{n}$. Since $({\bf x}_{t^{\prime}},\widetilde{B}_{t^{\prime}})=\mu_{k}({\bf x}_{t},\widetilde{B}_{t})$ is a green mutation, we have ${\bf c}_{1;t}\in\mathbb{Z}_{\geq 0}^{n}$. So we have ${\bf a}-{\bf a}^{\prime}=p\lambda s_{1}^{-1}(S{\bf c}_{1;t})\in\mathbb{Z}_{\geq 0}^{n}$.
Since ${\bf a}\in\mathbb{Z}_{>0}G_{t}$ and ${\bf a}^{\prime}\in\mathbb{Z}_{>0}G_{t^{\prime}}$, there exist ${\bf v},{\bf v}^{\prime}\in\mathbb{Z}_{>0}^{n}$ such that ${\bf a}=G_{t}{\bf v},{\bf a}^{\prime}=G_{t^{\prime}}{\bf v}^{\prime}$. Thus $G_{t}{\bf v}-G_{t^{\prime}}{\bf v}^{\prime}={\bf a}-{\bf a}^{\prime}\in\mathbb{Z}_{\geq 0}^{n}$. This completes the proof. ∎
Lemma 5.6.
Let $F$ be a polynomial in $\mathbb{Z}[y_{1},\ldots,y_{n}]$ with constant term $1$. If $({\bf r}-{\bf r}^{\prime})\in\mathbb{Z}_{\geq 0}^{n}$ for some ${\bf r},{\bf r}^{\prime}\in\mathbb{Z}^{n}$ and $F[{\bf r}]=0$, then $F[{\bf r}^{\prime}]=0$.
Proof.
Let ${\bf y}^{\bf u}$ be a monomial appearing in $F$. We know that ${\bf u}\in\mathbb{Z}_{\geq 0}^{n}$. Since $({\bf r}-{\bf r}^{\prime})\in\mathbb{Z}_{\geq 0}^{n}$, we have ${\bf u}^{T}({\bf r}^{\prime}-{\bf r})=-{\bf u}^{T}({\bf r}-{\bf r}^{\prime})\leq 0$. Since $F[{\bf r}]=0$, we have ${\bf u}^{T}{\bf r}\leq 0$. Thus
${\bf u}^{T}{\bf r}^{\prime}={\bf u}^{T}({\bf r}^{\prime}-{\bf r}+{\bf r})={\bf u}^{T}({\bf r}^{\prime}-{\bf r})+{\bf u}^{T}{\bf r}\leq 0.$
Since $F$ has constant term $1$ and by the definition of $F[{\bf r}^{\prime}]$, we obtain $F[{\bf r}^{\prime}]=0$. ∎
Theorem 5.7.
If $({\bf x}_{t^{\prime}},\widetilde{B}_{t^{\prime}})=\mu_{k}({\bf x}_{t},\widetilde{B}_{t})$ is a green mutation, then $\operatorname{\mathsf{dom}}^{t_{0}}[t]\subsetneq\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$. In particular, the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}=\overrightarrow{\mathcal{H}}^{t_{0}}(\mathcal{A})$ of $\mathcal{A}$ is acyclic.
Proof.
Since $({\bf x}_{t^{\prime}},\widetilde{B}_{t^{\prime}})=\mu_{k}({\bf x}_{t},\widetilde{B}_{t})$ is a green mutation and by Lemma 5.5, there exists
${\bf v}=(v_{1},\ldots,v_{n})^{T},\;{\bf v}^{\prime}=(v_{1}^{\prime},\ldots,v_{n}^{\prime})^{T}\in\mathbb{Z}_{>0}^{n}$
such that $G_{t}{\bf v}-G_{t^{\prime}}{\bf v}^{\prime}\in\mathbb{Z}_{\geq 0}^{n}$. Set $u=\prod_{i=1}^{n}x_{i;t}^{v_{i}}$ and $u^{\prime}=\prod_{i=1}^{n}x_{i;t^{\prime}}^{v_{i}^{\prime}}$. Notice that we have
$({\bf g}_{u}^{t_{0}})^{\circ}=G_{t}{\bf v},\;\;({\bf g}_{u^{\prime}}^{t_{0}})^{\circ}=G_{t^{\prime}}{\bf v}^{\prime}.$
Since $\operatorname{\mathsf{supp}}(u)=[{\bf x}_{t}^{\circ}],\;\operatorname{\mathsf{supp}}(u^{\prime})=[{\bf x}_{t^{\prime}}^{\circ}]$ and by Corollary 5.4, we have that
$\operatorname{\mathsf{dom}}^{t_{0}}[t]$ $=$ $\operatorname{\mathsf{dom}}^{t_{0}}(u)=\{z\in\mathcal{X}\mid F_{z}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{u}^{t_{0}}]=0\}=\{z\in\mathcal{X}\mid F_{z}^{t_{0}}[SG_{t}{\bf v}]=0\},$
$\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$ $=$ $\operatorname{\mathsf{dom}}^{t_{0}}(u^{\prime})=\{z\in\mathcal{X}\mid F_{z}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{t_{0}}]=0\}=\{z\in\mathcal{X}\mid F_{z}^{t_{0}}[SG_{t^{\prime}}{\bf v}^{\prime}]=0\}.$
Let $z\in\operatorname{\mathsf{dom}}^{t_{0}}[t]$, that is, we have $F_{z}^{t_{0}}[SG_{t}{\bf v}]=0$. Since
$SG_{t}{\bf v}-SG_{t^{\prime}}{\bf v}^{\prime}=S(G_{t}{\bf v}-G_{t^{\prime}}{\bf v}^{\prime})\in\mathbb{Z}_{\geq 0}^{n}$
and by Lemma 5.6, we get $F_{z}^{t_{0}}[SG_{t^{\prime}}{\bf v}^{\prime}]=0$. This implies that $z\in\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$. So we have $\operatorname{\mathsf{dom}}^{t_{0}}[t]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$.
In rest part of the proof, we will show $x_{k;t^{\prime}}\in\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$ but $x_{k;t^{\prime}}\notin\operatorname{\mathsf{dom}}^{t_{0}}[t]$. Thus the inclusion $\operatorname{\mathsf{dom}}^{t_{0}}[t]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$ is strict.
By Proposition 5.3 (ii), we have $[{\bf x}_{t}^{\circ}]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t]$ and $[{\bf x}_{t^{\prime}}^{\circ}]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$. In particular,
$x_{k;t}\in\operatorname{\mathsf{dom}}^{t_{0}}[t]\;\;\;\text{and}\;\;\;x_{k;t^{\prime}}\in\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}].$
Since $\operatorname{\mathsf{dom}}^{t_{0}}[t]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$, we get $x_{k;t}\in\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$. Then by Proposition 5.3 (iii), we have $x_{k;t}\in\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}(x_{k;t^{\prime}})$. This implies
$F_{x_{k;t}}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{x_{k;t^{\prime}}}^{t_{0}}]=0.$
By using the vertex $t_{0}$ to calculate the $F$-invariant $(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}$, we have
$(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}=F_{x_{k;t}}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{x_{k;t}^{\prime}}^{t_{0}}]+F_{x_{k;t^{\prime}}}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{x_{k;t}}^{t_{0}}]=F_{x_{k;t^{\prime}}}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{x_{k;t}}^{t_{0}}].$
On the other hand, if we use the vertex $t$ to calculate $(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}$, we see that
$(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}=0+F_{x_{k;t^{\prime}}}^{t}[(S\mid{\bf 0}){\bf g}_{x_{k;t}}^{t}]=(1+y_{k})[(S\mid{\bf 0}){\bf e}_{k}]=s_{k},$
where ${\bf e}_{k}$ is the $k$th column of $I_{m}$ and $s_{k}$ is the $(k,k)$-entry of $S=diag(s_{1},\ldots,s_{n})$. Thus
$F_{x_{k;t^{\prime}}}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{x_{k;t}}^{t_{0}}]=(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}=s_{k}>0,$
which means $x_{k;t^{\prime}}\notin\operatorname{\mathsf{dom}}^{t_{0}}(x_{k;t})$. By Proposition 5.3 (iii), we know that $\operatorname{\mathsf{dom}}^{t_{0}}[t]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}(x_{k;t})$. Thus $x_{k;t^{\prime}}\notin\operatorname{\mathsf{dom}}^{t_{0}}(x_{k;t})$ implies that $x_{k;t^{\prime}}\notin\operatorname{\mathsf{dom}}^{t_{0}}[t]$.
Since $x_{k;t^{\prime}}\in\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$ but $x_{k;t^{\prime}}\notin\operatorname{\mathsf{dom}}^{t_{0}}[t]$, we know that the inclusion $\operatorname{\mathsf{dom}}^{t_{0}}[t]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$ is strict, i.e., we have $\operatorname{\mathsf{dom}}^{t_{0}}[t]\subsetneq\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$. ∎
Example 5.8.
Take $\widetilde{B}=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}=\Lambda$ and ${\bf x}=(x_{1},x_{2})$. Clearly, $\widetilde{B}^{T}\Lambda=(S\mid{\bf 0})=I_{2}$. The oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}$ of $\mathcal{A}=\mathcal{A}({\bf x},\widetilde{B})$ is as follows:
${t_{0}=\{x_{1},x_{2}\}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{x_{1}}}$$\scriptstyle{\mu_{x_{2}}}$${t_{1}=\{{x_{3}},x_{2}\}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{x_{2}}}$${t_{4}=\{x_{5},x_{1}\}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{x_{1}}}$${t_{2}=\{x_{3},{x_{4}}\}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{x_{3}}}$${t_{3}=\{{x_{5}},x_{4}\}}$
where $x_{3}={x_{1}^{-1}x_{2}\cdot(1+\widehat{y}_{1}),}\;\;\;x_{4}={x_{1}^{-1}\cdot(1+\widehat{y}_{1}+\widehat{y}_{1}\widehat{y}_{2}),}\;\;\;x_{5}={x_{2}^{-1}\cdot(1+\widehat{y}_{2})}$. It is easy to check that
$\operatorname{\mathsf{dom}}^{t_{0}}(x_{1})$ $=$ $\{x_{1},x_{2},x_{5}\},\;\;\;\operatorname{\mathsf{dom}}^{t_{0}}(x_{2})=\{x_{1},x_{2},x_{3}\},\;\;\;\operatorname{\mathsf{dom}}^{t_{0}}(x_{3})=\{x_{1},x_{2},x_{3},x_{4}\},$
$\operatorname{\mathsf{dom}}^{t_{0}}(x_{4})$ $=$ $\{x_{1},x_{2},x_{3},x_{4},x_{5}\}=\operatorname{\mathsf{dom}}^{t_{0}}(x_{5}).$
Then by Proposition 5.3 (iii), we can get that
$\operatorname{\mathsf{dom}}^{t_{0}}[t_{0}]$ $=$ $\{x_{1},x_{2}\},\;\;\operatorname{\mathsf{dom}}^{t_{0}}[t_{1}]=\{x_{1},x_{2},x_{3}\},\;\;\operatorname{\mathsf{dom}}^{t_{0}}[t_{2}]=\{x_{1},x_{2},x_{3},x_{4}\},$
$\operatorname{\mathsf{dom}}^{t_{0}}[t_{3}]$ $=$ $\{x_{1},x_{2},x_{3},x_{4},x_{5}\},\;\;\operatorname{\mathsf{dom}}^{t_{0}}[t_{4}]=\{x_{1},x_{2},x_{5}\}.$
We can see that the initial cluster variables $x_{1},x_{2}$ are contained in any dominant set with respect to the vertex $t_{0}$, as shown in Proposition 5.3 (i).
Let $\mathcal{H}_{0}$ be the set of seeds (up to seed equivalence) of $\mathcal{A}$, equivalently, $\mathcal{H}_{0}$ is the vertex set of the exchange graph of $\mathcal{A}$. For the rooted vertex $t_{0}\in\mathbb{T}_{n}$, we define an relation $\leq_{t_{0}}$ on $\mathcal{H}_{0}$ as follows: $[{\bf x}_{w},\widetilde{B}_{w}]\leq_{t_{0}}[{\bf x}_{t},\widetilde{B}_{t}]$ if and only if there exists a sequence $\overleftarrow{\mu}$ of green mutations such that
$[\overleftarrow{\mu}({\bf x}_{w},\widetilde{B}_{w})]=[{\bf x}_{t},\widetilde{B}_{t}],$
i.e., there exists a path from the vertex $[{\bf x}_{w},\widetilde{B}_{w}]$ to the vertex $[{\bf x}_{t},\widetilde{B}_{t}]$ in the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}=\overrightarrow{\mathcal{H}}^{t_{0}}(\mathcal{A})$ of $\mathcal{A}$.
If $[{\bf x}_{w},\widetilde{B}_{w}]\leq_{t_{0}}[{\bf x}_{t},\widetilde{B}_{t}]$ and $[{\bf x}_{t},\widetilde{B}_{t}]\neq[{\bf x}_{w},\widetilde{B}_{w}]$, we write $[{\bf x}_{w},\widetilde{B}_{w}]<_{t_{0}}[{\bf x}_{t},\widetilde{B}_{t}]$.
Corollary 5.9.
The relation $\leq_{t_{0}}$ gives a partial order on $\mathcal{H}_{0}$, i.e., $(\mathcal{H}_{0},\leq_{t_{0}})$ is a poset.
Proof.
By Theorem 5.7, the oriented exchange graph of $\mathcal{A}$ is acyclic. The result follows. ∎
The Hasse quiver of the poset set $(\mathcal{H}_{0},\leq_{t_{0}})$ is a quiver $\overrightarrow{\mathcal{H}}^{\prime}$ defined as follows:
  • the vertex set of $\mathcal{H}^{\prime}$ is $\mathcal{H}_{0}$;
  • $[{\bf x}_{w},\widetilde{B}_{w}]\rightarrow[{\bf x}_{t},\widetilde{B}_{t}]$ is an arrow in $\overrightarrow{\mathcal{H}}^{\prime}$ if and only if $[{\bf x}_{w},\widetilde{B}_{w}]<_{t_{0}}[{\bf x}_{t},\widetilde{B}_{t}]$ and there exists no seed $({\bf x}_{s},\widetilde{B}_{s})$ such that $[{\bf x}_{w},\widetilde{B}_{w}]<_{t_{0}}[{\bf x}_{s},\widetilde{B}_{s}]<_{t_{0}}[{\bf x}_{t},\widetilde{B}_{t}]$.
Notice that the Hasse quiver $\overrightarrow{\mathcal{H}}^{\prime}$ and the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}$ of $\mathcal{A}$ have the same vertex set $\mathcal{H}_{0}$. It is easy to see that the arrow set of $\overrightarrow{\mathcal{H}}^{\prime}$ is a subset of the arrow set of $\overrightarrow{\mathcal{H}}^{t_{0}}$ of $\mathcal{A}$. It is natural to ask when the two quivers coincide. The next subsection aims to answer this question.
5.2. Oriented exchange graphs are Hasse quivers
In this subsection, we will show that the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}=\overrightarrow{\mathcal{H}}^{t_{0}}(\mathcal{A})$ of $\mathcal{A}$ coincides with the Hasse quiver $\overrightarrow{\mathcal{H}}^{\prime}$ of the poset $(\mathcal{H}_{0},\leq_{t_{0}})$.
Proposition 5.10.
Let $({\bf x}_{t},\widetilde{B}_{t})$ be a seed of $\mathcal{A}$. Denote by
$\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])$ $=$ $\{x\in\operatorname{\mathsf{dom}}^{t_{0}}[t]\mid F_{z}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{x}^{t_{0}}]=0,\;\forall z\in\operatorname{\mathsf{dom}}^{t_{0}}[t]\}$
$=$ $\{x\in\operatorname{\mathsf{dom}}^{t_{0}}[t]\mid\operatorname{\mathsf{dom}}^{t_{0}}[t]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}(x)\}.$
Then $\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])=[{\bf x}_{t}^{\circ}]$ and in particular, $[{\bf x}_{t}]$ is uniquely determined by the dominant set $\operatorname{\mathsf{dom}}^{t_{0}}[t]$.
Proof.
Fix $k\in[1,n]$. By Proposition 5.3 (ii), (iii), we know that
$x_{k;t}\in\operatorname{\mathsf{dom}}^{t_{0}}[t]\;\;\;\text{and}\;\;\;\operatorname{\mathsf{dom}}^{t_{0}}[t]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}(x_{k;t}).$
So we have $x_{k;t}\in\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])$. Since $k$ is an arbitrary integer $[1,n]$, we get $[{\bf x}_{t}^{\circ}]\subseteq\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])$.
Now let us take $x\in\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])$. We know that $F_{z}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{x}^{t_{0}}]=0$ for any $z\in\operatorname{\mathsf{dom}}^{t_{0}}[t]$. Since $x_{k;t}$ is in $\operatorname{\mathsf{dom}}^{t_{0}}[t]$ for $k\in[1,n]$, we have $F_{x_{k;t}}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{x}^{t_{0}}]=0$ for $k\in[1,n]$.
Since $x\in\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t]$ and $\operatorname{\mathsf{dom}}^{t_{0}}[t]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}(x_{k;t})$, we have $x\in\operatorname{\mathsf{dom}}^{t_{0}}(x_{k;t})$, i.e., $F_{x}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{x_{k;t}}^{t_{0}}]=0$. Thus $(x_{k;t}\mid\mid x)_{F}=F_{x_{k;t}}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{x}^{t_{0}}]+F_{x}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{x_{k;t}}^{t_{0}}]=0$. Thus for any two cluster variables $z_{1},z_{2}$ in $[{\bf x}_{t}^{\circ}]\cup\{x\}$, we have $(z_{1}\mid\mid z_{2})_{F}=0$. Then by Lemma 4.21, we know that $[{\bf x}_{t}^{\circ}]\cup\{x\}$ is a subset of some cluster of $\mathcal{A}$. Since $[{\bf x}_{t}^{\circ}]$ is the set of unfrozen cluster variables in ${\bf x}_{t}$ and $x$ is an unfrozen cluster variable, we must have $x=x_{i;t}$ for some $i\in[1,n]$. Thus we get
$\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])\subseteq[{\bf x}_{t}^{\circ}].$
Hence, $\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])=[{\bf x}_{t}^{\circ}]$ and in particular, $[{\bf x}_{t}]$ is given by the union of $\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])$ and the set of frozen variables. ∎
The following result is analogous to [air_2014, Proposition 2.26] in $\tau$-tilting theory.
Lemma 5.11.
Let $({\bf x}_{t},\widetilde{B}_{t})$, $({\bf x}_{s},\widetilde{B}_{s})$ and $({\bf x}_{w},\widetilde{B}_{w})$ be three seeds of $\mathcal{A}$. If
$[{\bf x}_{w},\widetilde{B}_{w}]\leq_{t_{0}}[{\bf x}_{s},\widetilde{B}_{s}]\leq_{t_{0}}[{\bf x}_{t},\widetilde{B}_{t}],$
then $[{\bf x}_{t}]\cap[{\bf x}_{w}]\subseteq[{\bf x}_{s}]$.
Proof.
It suffices to prove that $[{\bf x}_{t}^{\circ}]\cap[{\bf x}_{w}^{\circ}]\subseteq[{\bf x}_{s}^{\circ}]$, because frozen variables are contained in all the clusters of $\mathcal{A}$.
Since $[{\bf x}_{w},\widetilde{B}_{w}]\leq_{t_{0}}[{\bf x}_{s},\widetilde{B}_{s}]\leq_{t_{0}}[{\bf x}_{t},\widetilde{B}_{t}]$, we know that there is a sequence of green mutations from $[{\bf x}_{w},\widetilde{B}_{w}]$ to $[{\bf x}_{s},\widetilde{B}_{s}]$ and a sequence of green mutations from $[{\bf x}_{s},\widetilde{B}_{s}]$ to $[{\bf x}_{t},\widetilde{B}_{t}]$. Then by Theorem 5.7, we have $\operatorname{\mathsf{dom}}^{t_{0}}[w]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[s]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t]$.
By the definition of $\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[s])$, it is easy to see $\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])\cap\operatorname{\mathsf{dom}}^{t_{0}}[s]\subseteq\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[s])$. By Proposition 5.10, we know that $[{\bf x}_{t}^{\circ}]=\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])$ and $[{\bf x}_{s}^{\circ}]=\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[s])$. Thus we have $[{\bf x}_{t}^{\circ}]\cap\operatorname{\mathsf{dom}}^{t_{0}}[s]\subseteq[{\bf x}_{s}^{\circ}]$.
By Proposition 5.3 (ii), we know that $[{\bf x}_{w}^{\circ}]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[w]$. Then by $\operatorname{\mathsf{dom}}^{t_{0}}[w]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[s]$, we have $[{\bf x}_{w}^{\circ}]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[s]$. So we know that $[{\bf x}_{t}^{\circ}]\cap[{\bf x}_{w}^{\circ}]\subseteq[{\bf x}_{t}^{\circ}]\cap\operatorname{\mathsf{dom}}^{t_{0}}[s]\subseteq[{\bf x}_{s}^{\circ}]$. This implies $[{\bf x}_{t}]\cap[{\bf x}_{w}]\subseteq[{\bf x}_{s}]$. ∎
Theorem 5.12.
The following statements are equivalent:
  • (a)
    There exists an arrow $[{\bf x}_{w},\widetilde{B}_{w}]\rightarrow[{\bf x}_{t},\widetilde{B}_{t}]$ in the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}$ of $\mathcal{A}$.
  • (b)
    $[{\bf x}_{w},\widetilde{B}_{w}]<_{t_{0}}[{\bf x}_{t},\widetilde{B}_{t}]$ and there exists no seed $({\bf x}_{s},\widetilde{B}_{s})$ such that
    $[{\bf x}_{w},\widetilde{B}_{w}]<_{t_{0}}[{\bf x}_{s},\widetilde{B}_{s}]<_{t_{0}}[{\bf x}_{t},\widetilde{B}_{t}].$
Proof.
``$(a)\Rightarrow(b)"$: Suppose that $[{\bf x}_{w},\widetilde{B}_{w}]\rightarrow[{\bf x}_{t},\widetilde{B}_{t}]$ is an arrow in the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}$ of $\mathcal{A}$. Then we know that $[{\bf x}_{w},\widetilde{B}_{w}]<_{t_{0}}[{\bf x}_{t},\widetilde{B}_{t}]$.
Since the two seeds $({\bf x}_{t},\widetilde{B}_{t})$ and $({\bf x}_{w},\widetilde{B}_{w})$ differ by once mutation, for simplicity, we can assume $({\bf x}_{t},\widetilde{B}_{t})=\mu_{1}({\bf x}_{w},\widetilde{B}_{w})$. If we write ${\bf x}_{w}=(z_{1},z_{2},\ldots,z_{m})$, then ${\bf x}_{t}$ has the form ${\bf x}_{t}=(z_{1}^{\prime},z_{2},\ldots,z_{m})$.
If there exists a seed $({\bf x}_{s},\widetilde{B}_{s})$ such that $[{\bf x}_{w},\widetilde{B}_{w}]<_{t_{0}}[{\bf x}_{s},\widetilde{B}_{s}]\leq_{t_{0}}[{\bf x}_{t},\widetilde{B}_{t}]$. By Lemma 5.11, we have $\{z_{2},\ldots,z_{m}\}=[{\bf x}_{t}]\cap[{\bf x}_{w}]\subseteq[{\bf x}_{s}]$. Since $[{\bf x}_{w},\widetilde{B}_{w}]<_{t_{0}}[{\bf x}_{s},\widetilde{B}_{s}]$, we know that $[{\bf x}_{s}]\neq[{\bf x}_{w}]$. Thus the two clusters ${\bf x}_{s}$ and ${\bf x}_{w}$ have exactly $m-1$ common cluster variables $z_{2},\ldots,z_{m}$. Then by [cao-li-2020, Corollary 3], we get $[{\bf x}_{s},\widetilde{B}_{s}]=[\mu_{1}({\bf x}_{w},\widetilde{B}_{w})]=[{\bf x}_{t},\widetilde{B}_{t}]$. So there exists no seed $({\bf x}_{s},\widetilde{B}_{s})$ such that $[{\bf x}_{w},\widetilde{B}_{w}]<_{t_{0}}[{\bf x}_{s},\widetilde{B}_{s}]<_{t_{0}}[{\bf x}_{t},\widetilde{B}_{t}]$.
``$(b)\Rightarrow(a)"$: Since $[{\bf x}_{w},\widetilde{B}_{w}]<_{t_{0}}[{\bf x}_{t},\widetilde{B}_{t}]$, there exists a non-empty sequence $\overleftarrow{\mu}$ of green mutations from $[{\bf x}_{w},\widetilde{B}_{w}]$ to $[{\bf x}_{t},\widetilde{B}_{t}]$. Since there exists no seed $({\bf x}_{s},\widetilde{B}_{s})$ such that
$[{\bf x}_{w},\widetilde{B}_{w}]<_{t_{0}}[{\bf x}_{s},\widetilde{B}_{s}]<_{t_{0}}[{\bf x}_{t},\widetilde{B}_{t}],$
the length of $\overleftarrow{\mu}$ must be $1$, i.e., $[{\bf x}_{t},\widetilde{B}_{t}]$ is obtained from $[{\bf x}_{w},\widetilde{B}_{w}]$ by once green mutation. So there exists an arrow $[({\bf x}_{w},\widetilde{B}_{w})]\rightarrow[({\bf x}_{t},\widetilde{B}_{t})]$ in the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}$ of $\mathcal{A}$. ∎
The above theorem rules out the existence of the subquiver $\widetilde{A}_{1,d}$ ($d\geq 2$) of the form
[…]
in the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}$ of $\mathcal{A}$.
The following result is a direct consequence of the above theorem and it is analogous to the corresponding result [air_2014, Theorem 0.6] in the $\tau$-tilting theory.
Corollary 5.13.
The oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}$ of $\mathcal{A}$ coincides with the Hasse quiver $\overrightarrow{\mathcal{H}}^{\prime}$ of the poset $(\mathcal{H}_{0},\leq_{t_{0}})$.
5.3. Dominant sets and torsion classes
In this subsection we briefly explain why the dominant sets can be viewed as an replacement of torsion classes in $\tau$-tilting theory. More discussions will be given in [cao_2025].
We fix a finite dimensional basic algebra $A$ over a field $K$. Denote by $\operatorname{\mathsf{mod}}\nolimits A$ the category of finitely generated right $A$-modules, and by $\tau$ the Auslander-Reiten translation in $\operatorname{\mathsf{mod}}\nolimits A$. The isomorphism classes of indecomposable projective modules in $\operatorname{\mathsf{mod}}\nolimits A$ are denoted by $P_{1},\ldots,P_{n}$.
Given a module $M\in\operatorname{\mathsf{mod}}\nolimits A$, we denote by
  • $\operatorname{\mathsf{add}}\nolimits M$ the subcategory of $\operatorname{\mathsf{mod}}\nolimits A$ consisting of modules which are direct summands of $M^{\oplus p}$ for some $p>0$.
  • $\operatorname{\mathsf{Fac}}M$ the factor modules of the modules in $\operatorname{\mathsf{add}}\nolimits M$.
  • $\prescript{\bot}{}{M}:=\{X\in\operatorname{\mathsf{mod}}\nolimits A\mid\operatorname{\mathsf{Hom}}_{A}(X,M)=0\}$.
  • $M^{\bot}:=\{Y\in\operatorname{\mathsf{mod}}\nolimits A\mid\operatorname{\mathsf{Hom}}_{A}(M,Y)=0\}$.
  • $|M|$ the number of non-isomorphic indecomposable direct summands of $M$.
Definition 5.14.
Let $M$ be a module in $\operatorname{\mathsf{mod}}\nolimits A$ and $P$ a projective module in $\operatorname{\mathsf{mod}}\nolimits A$.
  • (i)
    $M$ is called $\tau$-rigid if $\operatorname{\mathsf{Hom}}_{A}(M,\tau M)=0$.
  • (ii)
    The pair $(M,P)$ is called $\tau$-rigid if $M$ is $\tau$-rigid and $\operatorname{\mathsf{Hom}}_{A}(P,M)=0$.
  • (iii)
    The pair $(M,P)$ is called $\tau$-tilting if $(M,P)$ is a $\tau$-rigid pair and $|M|+|P|=|A|=n$.
We will always consider modules and $\tau$-rigid pairs up to isomorphisms. A $\tau$-rigid pair $(M,P)$ is indecomposable if $M\oplus P$ is indecomposable in $\operatorname{\mathsf{mod}}\nolimits A$ and it is basic if both $M$ and $P$ are basic in $\operatorname{\mathsf{mod}}\nolimits A$. In a (basic) $\tau$-tilting pair $(M,P)$, it is known [air_2014, Proposition 2.3] that $P$ is uniquely determined by $M$.
A subcategory $\mathcal{T}$ of $\operatorname{\mathsf{mod}}\nolimits A$ is called a torsion class, if $\mathcal{T}$ is closed under quotients and extensions. A torsion class $\mathcal{T}$ is said to be functorially finite, if there exists a module $M\in\operatorname{\mathsf{mod}}\nolimits A$ such that $\mathcal{T}=\operatorname{\mathsf{Fac}}M$.
Let $\mathcal{C}$ be a subcategory of $\operatorname{\mathsf{mod}}\nolimits A$. A module $U\in\mathcal{C}$ is said to be Ext-projective in $\mathcal{C}$, if $\operatorname{\mathsf{Ext}}_{A}^{1}(U,\mathcal{C})=0$. We denote by $\mathcal{P}(\mathcal{C})$ the direct sum of one copy of each of the indecomposable Ext-projective objects in $\mathcal{C}$ up to isomorphisms.
Theorem 5.15.
[air_2014, Proposition 1.2 (b) and Theorem 2.7] The following statements hold.
  • (i)
    There is a well-defined map $\Psi$ from $\tau$-rigid pairs to functorially finite torsion classes in $\operatorname{\mathsf{mod}}\nolimits A$ given by $(M,P)\mapsto\operatorname{\mathsf{Fac}}M$.
  • (ii)
    The above map $\Psi$ is a bijection if we restrict it to basic $\tau$-tilting pairs, which induces a natural order on the set of basic $\tau$-tilting pairs
    $(M,P)\leq(M^{\prime},P^{\prime})\overset{{\rm def.}}{\Longleftrightarrow}\operatorname{\mathsf{Fac}}M\subseteq\operatorname{\mathsf{Fac}}M^{\prime}.$
  • (iii)
    Let $\mathcal{T}$ be a functorially finite torsion class and denote by $(M,P)$ the basic $\tau$-tilting pair such that $\operatorname{\mathsf{Fac}}M=\mathcal{T}$. Then $M=\mathcal{P}(\mathcal{T})$.
One can refer to [air_2014, Theorem 2.19] for the definition of mutations of $\tau$-tilting pairs. Adachi, Iyama and Reiten [air_2014, Theorem 0.6] proved that mutations of $\tau$-tilting pairs correspond to the cover relations in the poset of $\tau$-tilting pairs (or functorially finite torsion classes). Some correspondences between cluster algebras and $\tau$-tilting theory are given in Table 1.
$\text{Cluster algebras}&\hskip 5.69054pt\tau\text{-tilting theory}\\[5.69054pt] \hline\cr&\\[-8.53581pt] \text{Seeds}&\hskip 5.69054pt\tau\text{-tilting pairs}\\[5.69054pt] \hline\cr&\\[-5.69054pt] \text{Mutations of seeds}&\hskip 5.69054pt\text{Mutations of }\tau\text{-tilting pairs}\\[5.69054pt] \hline\cr&\\[-5.69054pt] \text{Initial unfrozen cluster variables }&\hskip 5.69054pt(0,P_{1}),\ldots,(0,P_{n})\\[5.69054pt] \hline\cr&\\[-8.53581pt] \text{Non-initial unfrozen cluster variables}&\hskip 5.69054pt\text{Indecomposable }\tau\text{-rigid modules}\\[5.69054pt] \hline\cr&\\[-8.53581pt] \text{Unfrozen cluster variables}&\hskip 5.69054pt\text{Indecomposable }\tau\text{-rigid pairs}\\[5.69054pt] \hline\cr\end{array}$
Table 1. Cluster algebras vs. $\tau$-tilting theory
Consider $M\in\operatorname{\mathsf{mod}}\nolimits A$ and let
$\bigoplus_{i=1}^{n}P_{i}^{b_{i}}\rightarrow\bigoplus_{i=1}^{n}P_{i}^{a_{i}}\rightarrow M\rightarrow 0$
be the minimal projective presentation of $M$ in $\operatorname{\mathsf{mod}}\nolimits A$. The vector
$\delta_{M}:=(a_{1}-b_{1},\ldots,a_{n}-b_{n})^{T}\in\mathbb{Z}^{n}$
is called the $\delta$-vector of $M$ and the vector ${\bf g}_{M}:=-\delta_{M}$ is called the $g$-vector of $M$.
For a $\tau$-rigid pair $(M,P)$, we define its $\delta$-vector and $g$-vector as follows:
$\delta_{(M,P)}:=\delta_{M}-\delta_{P},\;\;{\bf g}_{(M,P)}:=-\delta_{(M,P)}.$
With this definition, one can see that the $g$-vector ${\bf g}_{(0,P_{k})}$ of $(0,P_{k})$ is the $k$th column of $I_{n}$.
Remark 5.16.
Notice that the $\delta$-vectors defined here coincide with the $g$-vectors used in [air_2014]. For the considerations on the cluster algebra side, $g$-vectors defined here are the negative of the $\delta$-vectors.
Definition 5.17.
The $F$-polynomial $F_{M}$ of $M\in\operatorname{\mathsf{mod}}\nolimits A$ is defined to be
$F_{M}=\sum_{{\bf v}\in\mathbb{N}^{n}}\chi(\operatorname{\mathsf{Gr}}\nolimits_{\bf v}(M)){\bf y}^{\bf v}\in\mathbb{Z}[y_{1},\ldots,y_{n}],$
where $\operatorname{\mathsf{Gr}}\nolimits_{\bf v}(M)$ is the quotient module Grassmannian with dimensional vector ${\bf v}$ and $\chi$ is the Euler–Poincaré characteristic.
Since the zero module is a quotient module of $M$, $F_{M}$ is a polynomial with constant term $1$. Thus
$F_{M}[{\bf r}]=\max\{{\bf v}^{T}{\bf r}\mid\chi(\operatorname{\mathsf{Gr}}\nolimits_{\bf v}(M))\neq 0\}\geq 0$
for any ${\bf r}\in\mathbb{Z}^{n}$.
For a $\tau$-rigid pair $(M,P)$, we define its $F$-polynomial as follows:
$F_{(M,P)}:=F_{M}\in\mathbb{Z}[y_{1},\ldots,y_{n}].$
In particular, we have $F_{(0,P_{k})}=1$ for $k=1,\ldots,n$.
Denote by $\mathcal{X}_{+}$ the set of indecomposable $\tau$-rigid modules in $\operatorname{\mathsf{mod}}\nolimits A$ and $\mathcal{X}$ the set of indecomposable $\tau$-rigid pairs in $\operatorname{\mathsf{mod}}\nolimits A$. We have
$\mathcal{X}=\mathcal{X}_{+}\sqcup\{(0,P_{1}),\ldots,(0,P_{n})\}.$
Notice that when we write $U\in\mathcal{X}$, we mean that $U$ is an indecomposable $\tau$-rigid pair, that is, either $U\in\mathcal{X}_{+}$ or $U=(0,P_{k})$ for some $k\in[1,n]$.
Similarly to the dominant sets in cluster algebras, we can define dominant sets in $\tau$-tilting theory.
Definition 5.18.
Let $(M,P)$ be a $\tau$-tilting pair in $\operatorname{\mathsf{mod}}\nolimits A$. The dominant set $\operatorname{\mathsf{dom}}[M,P]$ and the reduced dominant set $\operatorname{\mathsf{dom}}_{+}[M,P]$ of $(M,P)$ are defined as follows:
$\operatorname{\mathsf{dom}}[M,P]$ ::=:= $\{U\in\mathcal{X}\mid F_{U}[{\bf g}_{(M,P)}]=0\},$
$\operatorname{\mathsf{dom}}_{+}[M,P]$ ::=:= $\{U\in\mathcal{X}_{+}\mid F_{U}[{\bf g}_{(M,P)}]=0\}.$
Since $F_{(0,P_{k})}=1$ for $k=1,\ldots,n$, we see that the set $\{(0,P_{1}),\ldots,(0,P_{n})\}$ is always contained in the dominant set $\operatorname{\mathsf{dom}}[M,P]$ for a $\tau$-tilting pair. Thus we have
(5.1) $\operatorname{\mathsf{dom}}[M,P]=\operatorname{\mathsf{dom}}_{+}[M,P]\sqcup\{(0,P_{1}),\ldots,(0,P_{n})\}.$
For a module $X\in\operatorname{\mathsf{mod}}\nolimits A$ and a vector $\delta\in\mathbb{R}^{n}$, we write $\delta(X)\in\mathbb{R}$ for the inner product of $\delta$ and the dimensional vector of $X$.
Lemma 5.19.
Denote by
$\overline{\mathcal{T}}_{\delta}=\{N\in\operatorname{\mathsf{mod}}\nolimits A\mid\text{for any quotient module X of }N,\;\;\delta(X)\geq 0\}$
for $\delta\in\mathbb{R}^{n}$. The following statements hold.
  • (i)
    [BKT-2014]*Proposition3.1 The subcategory $\overline{\mathcal{T}}_{\delta}$ is a torsion class in $\operatorname{\mathsf{mod}}\nolimits A$.
  • (ii)
    [asai-2021, Proposition 3.11] Let $(M,P)$ be a basic $\tau$-tilting pair in $\operatorname{\mathsf{mod}}\nolimits A$ and $\delta=\delta_{(M,P)}$ the $\delta$-vector of $(M,P)$. Then $\operatorname{\mathsf{Fac}}M=\overline{\mathcal{T}}_{\delta}$.
Torsion classes of the form $\overline{\mathcal{T}}_{\delta}$ are called semistable torsion classes. By Lemma 5.19 (ii), all functorially finite torsion classes are semistable torsion classes.
Lemma 5.20.
[fei_2019a, Theorem 1.4] Let $U\in\operatorname{\mathsf{mod}}\nolimits A$ and let
$F_{U}=\sum_{{\bf v}\in\mathbb{N}^{n}}\chi(\operatorname{\mathsf{Gr}}\nolimits_{\bf v}(U)){\bf y}^{\bf v}\in\mathbb{Z}[y_{1},\ldots,y_{n}]$
be the $F$-polynomial of $U$. Then the convex hull $\mathcal{N}(F_{U})$ of $\{{\bf v}\in\mathbb{N}^{n}\mid\chi(\operatorname{\mathsf{Gr}}\nolimits_{\bf v}(U))\neq 0\}$ and the convex hull $\mathcal{N}(U)$ of the dimensional vectors of the quotient modules of $U$ are the same.
Proof.
The corresponding result for submodules and $F$-polynomials defined using submodules is proved by Fei in [fei_2019a, Theorem 1.4]. The required result follows from Fei’s result by considering the modules over the opposite algebras. ∎
Proposition 5.21.
Let $(M,P)$ be a $\tau$-tilting pair in $\operatorname{\mathsf{mod}}\nolimits A$. Then we have
  • (i)
    $\operatorname{\mathsf{dom}}_{+}[M,P]=\mathcal{X}_{+}\cap\operatorname{\mathsf{Fac}}M=\{\text{indecomposable }\tau\text{-rigid modules in }\operatorname{\mathsf{Fac}}M\}.$
  • (ii)
    $\operatorname{\mathsf{dom}}[M,P]=\{\text{indecomposable }\tau\text{-rigid modules in }\operatorname{\mathsf{Fac}}M\}\sqcup\{(0,P_{1}),\ldots,(0,P_{n})\}$.
Proof.
(i) Let $\delta=\delta_{(M,P)}$ be the $\delta$-vector of $(M,P)$ and ${\bf g}={\bf g}_{(M,P)}=-\delta$ the $g$-vector of $(M,P)$. Let $U$ be an indecomposable $\tau$-rigid module and
$F_{U}=\sum_{{\bf v}\in\mathbb{N}^{n}}\chi(\operatorname{\mathsf{Gr}}\nolimits_{\bf v}(U)){\bf y}^{\bf v}\in\mathbb{Z}[y_{1},\ldots,y_{n}]$
the $F$-polynomial of $U$. We know that $U\in\operatorname{\mathsf{dom}}_{+}[M,P]$ if and only if
$F_{U}[{\bf g}]=\max\{{\bf v}^{T}{\bf g}\mid\chi(\operatorname{\mathsf{Gr}}\nolimits_{\bf v}(U))\neq 0\}=0.$
Thanks to Lemma 5.20, this is equivalent to
$\max\{{\bf g}(X)\mid X\text{ is a quotient module of }U\}=0,$
equivalently, ${\bf g}(X)\leq 0$ for any quotient module $X$ of $U$. This is also equivalent to $\delta(X)=-{\bf g}(X)\geq 0$ for any quotient module $X$ of $U$. Then by Lemma 5.19, this is equivalent to $U\in\operatorname{\mathsf{Fac}}M$. Hence, we have
$\operatorname{\mathsf{dom}}_{+}[M,P]=\mathcal{X}_{+}\cap\operatorname{\mathsf{Fac}}M=\{\text{indecomposable }\tau\text{-rigid modules in }\operatorname{\mathsf{Fac}}M\}.$
(ii) This follows from (i) and the equality (5.1). ∎
Since $\operatorname{\mathsf{Fac}}M$ is the smallest torsion classes containing all the indecomposable $\tau$-rigid modules in $\operatorname{\mathsf{Fac}}M$, we know that $\operatorname{\mathsf{Fac}}M$ is uniquely determined by the dominant set $\operatorname{\mathsf{dom}}[M,P]$. Thus we can view the dominant sets in cluster algebras as a kind of replacement of torsion classes in $\tau$-tilting theory.
Example 5.22.
Let $A$ be the path algebra of the quiver $1\rightarrow 2$. We have the exact sequence
$0\rightarrow P_{2}\rightarrow P_{1}\rightarrow S_{1}\rightarrow 0$
in $\operatorname{\mathsf{mod}}\nolimits A=\operatorname{\mathsf{add}}\nolimits(P_{1}\oplus P_{2}\oplus S_{1})$. The Hasse quiver of the poset of functorially torsion classes in $\operatorname{\mathsf{mod}}\nolimits A$ is as follows:
${\mathcal{T}_{0}=\{0\}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{T}_{1}=\operatorname{\mathsf{add}}\nolimits(S_{1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{T}_{4}=\operatorname{\mathsf{add}}\nolimits(P_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{T}_{2}=\operatorname{\mathsf{add}}\nolimits(S_{1}\oplus P_{1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{T}_{3}=\operatorname{\mathsf{mod}}\nolimits A}$
The oriented exchange graph of $\tau$-tilting pairs are given as follows:
${(0,P_{1}\oplus P_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{(0,P_{1})}}$$\scriptstyle{\mu_{(0,P_{2})}}$${(S_{1},P_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{(0,P_{2})}}$${(P_{2},P_{1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{(0,P_{1})}}$${(S_{1}\oplus P_{1},0)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{(S_{1},0)}}$${\mathcal{(}P_{1}\oplus P_{2},0)}$
The corresponding dominant sets are as follows:
$\operatorname{\mathsf{dom}}[0,P_{1}\oplus P_{2}]=\{(0,P_{1}),\;(0,P_{2})\},$
$\operatorname{\mathsf{dom}}[S_{1},P_{2}]=\{(0,P_{1}),\;(0,P_{2}),\;S_{1}\},$
$\operatorname{\mathsf{dom}}[S_{1}\oplus P_{1},0]=\{(0,P_{1}),\;(0,P_{2}),\;S_{1},\;P_{1}\},$
$\operatorname{\mathsf{dom}}[P_{1}\oplus P_{2},0]=\{(0,P_{1}),\;(0,P_{2}),\;S_{1},\;P_{1},\;P_{2}\},$
$\operatorname{\mathsf{dom}}[P_{2},P_{1}]=\{(0,P_{1}),\;(0,P_{2}),\;P_{2}\}.$
It is easy to check that
$F_{S_{1}}$ $=$ $1+y_{1},\;\;F_{P_{1}}=1+y_{1}+y_{1}y_{2},\;\;F_{P_{2}}=1+y_{2},$
${\bf g}_{S_{1}}$ $=$ $\begin{bmatrix}-1\\ 1\end{bmatrix},\;\;{\bf g}_{P_{1}}=\begin{bmatrix}-1\\ 0\end{bmatrix},\;\;{\bf g}_{P_{2}}=\begin{bmatrix}0\\ -1\end{bmatrix}.$
From the viewpoint of categorification of cluster algebras, the indecomposable $\tau$-rigid modules $S_{1},P_{1},P_{2}$ correspond to the non-initial cluster variables
$x_{3}={x_{1}^{-1}x_{2}\cdot(1+\widehat{y}_{1}),}\;\;\;x_{4}={x_{1}^{-1}\cdot(1+\widehat{y}_{1}+\widehat{y}_{1}\widehat{y}_{2}),}\;\;\;x_{5}={x_{2}^{-1}\cdot(1+\widehat{y}_{2})}$
in Example 5.8.
5.4. Dominant sets and inversion sets
The aim of this subsection is to point out there exists links between dominant sets and inversion sets so that the readers can have a better understanding on the dominant sets from the viewpoint of inversion sets. We will not go into depth and generality.
Let $C=(c_{ij})_{n\times n}$ be a Cartan matrix of finite type and $\Phi$ the corresponding root system. We fix a base $\Delta=\{\alpha_{1},\ldots,\alpha_{n}\}\subseteq$ of $\Phi$. Denote by $\Phi^{+}$ the set of positive roots of $(\Phi,\Delta)$ and $\Phi_{\geq-1}=\Phi^{+}\cup\{-\alpha_{1},\ldots,-\alpha_{n}\}$ the set of almost positive roots. Let $W$ be the Weyl group corresponding to $\Phi$ and denote by $s_{1},\ldots,s_{n}$ the simple reflections in $W$ corresponding to the simple roots $\alpha_{1},\ldots,\alpha_{n}$. There is a natural action of $W$ on the root lattice $Q:=\oplus_{i=1}^{n}\mathbb{Z}\alpha_{i}$ induced by
$s_{i}(\alpha_{j})=\alpha_{j}-c_{ij}\alpha_{i},$
where $c_{ij}$ is the $(i,j)$-entry of the Cartan matrix $C$. It is known that $w(\Phi)=\Phi$ for any $w\in W$.
We use ${\bf w}={\bf s}_{i_{1}}\ldots{\bf s}_{i_{r}}$ to denote a word in formal symbols ${\bf s}_{1},\ldots,{\bf s}_{n}$ and use $w=s_{i_{1}}\ldots s_{i_{r}}$ to denote the corresponding element in the Weyl group $W$. The length $\ell(w)$ of $w\in W$ is the smallest number $r\geq 0$ such that $w=s_{i_{1}}\cdots s_{i_{r}}$ for some $1\leq i_{j}\leq n$. In this case, ${\bf w}:={\bf s}_{i_{1}}\ldots{\bf s}_{i_{r}}$ is called an reduced word of $w$.
Definition 5.23 (Inversion set and decorated inversion set).
Let ${\bf w}={\bf s}_{i_{1}}\ldots{\bf s}_{i_{r}}$ be an reduced word of $w\in W$.
  • (i)
    The inversion set $\operatorname{\mathsf{inv}}(w)$ of $w$ is defined by
    $\operatorname{\mathsf{inv}}(w):=\{\beta_{1},\ldots,\beta_{r}\}\subseteq\Phi^{+},$
    where $\beta_{k}=s_{i_{1}}\ldots s_{i_{k-1}}(\alpha_{k})$ with $k\in[1,r]$.
  • (ii)
    The decorated inversion set $\widetilde{\operatorname{\mathsf{inv}}}(w)$ of $w$ is defined by $\widetilde{\operatorname{\mathsf{inv}}}(w):=\operatorname{\mathsf{inv}}(w)\cup\{-\alpha_{1},\ldots,-\alpha_{n}\}$, which is a subset of almost positive roots.
The inversion set $\operatorname{\mathsf{inv}}(w)$ is actually the set of positive roots which are sent to negative roots by $w^{-1}$, i.e.,
$\operatorname{\mathsf{inv}}(w)=\{\beta\in\Phi^{+}\mid w^{-1}(\beta)\in-\Phi^{+}\}=\Phi^{+}\cap w(-\Phi^{+}).$
Proposition 5.24.
[BB-2005, Corollaries 1.4.4, 1.4.5 and 3.1.4] All the reduced words of $w\in W$ give the same inversion set. Different elements of $W$ have different inversion sets. The size of the inversion set $\operatorname{\mathsf{inv}}(w)$ of $w$ is $\ell(w)$.
The inclusion relation on the inversion sets induces a partial order $\leq$ on $W$:
$v\leq w\;\;\;\text{if and only if}\;\;\;\operatorname{\mathsf{inv}}(v)\subseteq\operatorname{\mathsf{inv}}(w),$
equivalently, $v$ has a reduced word ${\bf v}$ such that ${\bf v}$ is a prefix of some reduced word ${\bf w}$ of $w$. This partial order is known as right weak order on $W$.
An element $c=s_{i_{1}}\ldots s_{i_{r}}\in W$ is called a Coxeter element, if $r=n$ and $\{i_{1},\ldots,i_{r}\}=\{1,\ldots,n\}$. For a word ${\bf w}={\bf s}_{i_{1}}\ldots{\bf s}_{i_{r}}$, define its support $\operatorname{\mathsf{supp}}(\bf w)$ by
$\operatorname{\mathsf{supp}}({\bf w}):=\{{\bf s}_{i_{1}},\ldots,{\bf s}_{i_{r}}\},$
which is a subset of $\{{\bf s}_{1},\ldots,{\bf s}_{n}\}$.
Definition 5.25 ($c$-sortable elements, [Reading-07]).
Let $c$ be a Coxeter element in $W$ and fix a reduced word ${\bf c}$ of $c\in W$. An element $w\in W$ is called $c$-sortable if there exists a reduced word ${\bf w}$ of $w$ with the form ${\bf w}={\bf c}^{(0)}{\bf c}^{(1)}\ldots{\bf c}^{(p)}$, where ${\bf c}^{(i)}$ are subwords of ${\bf c}$ satisfying
$\operatorname{\mathsf{supp}}({\bf c}^{(p)})\subseteq\ldots\subseteq\operatorname{\mathsf{supp}}({\bf c}^{(1)})\subseteq\operatorname{\mathsf{supp}}({\bf c}^{(0)})\subseteq\{{\bf s}_{1},\ldots,{\bf s}_{n}\}.$
Notice that $c$-sortability does not depend on the choice of reduced word of $c$. Sometimes, we prefer to write ${\bf w}={\bf c}^{(0)}{\bf c}^{(1)}\ldots{\bf c}^{(p)}$ as ${\bf w}={\bf c}^{(0)}|{\bf c}^{(1)}|\ldots|{\bf c}^{(p)}$.
With the weak order on $W$, the set $W_{c}$ of $c$-sortable elements in $W$ forms a poset, which is actually a lattice, known as Cambrian lattice.
Example 5.26.
Take $C$ to be the Cartan matrix of type $A_{2}$, i.e., $C=\begin{bmatrix}2&-1\\ -1&2\end{bmatrix}$. Then $\Phi^{+}=\{\alpha_{1},\alpha_{2},\alpha_{1}+\alpha_{2}\}$. It is easy to see that
$\operatorname{\mathsf{inv}}(e)=\emptyset,\;\;\operatorname{\mathsf{inv}}(s_{1})=\{\alpha_{1}\},\;\;\operatorname{\mathsf{inv}}(s_{2})=\{\alpha_{2}\},\;\;\operatorname{\mathsf{inv}}(s_{1}s_{2})=\{\alpha_{1},\;\alpha_{1}+\alpha_{2}\},$
$\operatorname{\mathsf{inv}}(s_{2}s_{1})=\{\alpha_{2},\;\alpha_{1}+\alpha_{2}\},\;\;\operatorname{\mathsf{inv}}(s_{1}s_{2}s_{1})=\{\alpha_{1},\;\alpha_{2},\;\alpha_{1}+\alpha_{2}\}.$
The Hasse quiver of the poset $(W,\;\leq)$ is as follows:
${e\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${s_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${s_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${s_{2}s_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${s_{1}s_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${s_{1}s_{2}s_{1}}$
Now we consider the Coxeter element $c=s_{1}s_{2}$, then the Hasse quiver of the poset $(W_{c},\;\leq)$ is as follows:
${e\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${s_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${s_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${s_{1}s_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${s_{1}s_{2}|s_{1}}$
The decorated inversion sets for elements in $W_{c}$ are given as follows:
$\widetilde{\operatorname{\mathsf{inv}}}(e)=\{-\alpha_{1},\;-\alpha_{2}\},\;\;\widetilde{\operatorname{\mathsf{inv}}}(s_{1})=\{-\alpha_{1},\;-\alpha_{2},\;\alpha_{1}\},\;\;\widetilde{\operatorname{\mathsf{inv}}}(s_{1}s_{2})=\{-\alpha_{1},\;-\alpha_{2},\;\alpha_{1},\;\alpha_{1}+\alpha_{2}\},$
$\widetilde{\operatorname{\mathsf{inv}}}(s_{1}s_{2}s_{1})=\{-\alpha_{1},\;-\alpha_{2},\;\alpha_{1},\;\alpha_{2},\;\alpha_{1}+\alpha_{2}\},\;\;\widetilde{\operatorname{\mathsf{inv}}}(s_{2})=\{-\alpha_{1},\;-\alpha_{2},\;\alpha_{2}\}.$
Let $Q$ be a Dynkin quiver with vertex set $\{1,\ldots,n\}$, namely, $Q$ is an acyclic quiver whose underlying graph is a Dynkin diagram. Let $C(Q)$ be the corresponding Cartan matrix, which is of finite type. Let $\Phi(Q)$ and $W(Q)$ be the root system and Weyl group associated to the Cartan matrix $C(Q)$.
An Coxeter element $c=s_{i_{1}}\ldots s_{i_{n}}\in W(Q)$ is said to be admissible with respect to the orientation of $Q$, if there is no arrow from the vertex $i_{q}$ to $i_{p}$ in $Q$ for any integers $p,q\in[1,n]$ with $p<q$. For example, $c=s_{1}s_{2}$ is an admissible Coxeter element for the quiver $Q=1\rightarrow 2$.
Let $A=\mathbb{C}Q$ be the path algebra of the Dynkin quiver $Q$ over $\mathbb{C}$. Denote by $\operatorname{\mathsf{tors}}(A)$ the poset of torsion classes in $\operatorname{\mathsf{mod}}\nolimits A$. Let $c$ be an admissible Coxeter element with respect to the orientation of $Q$. Denote by $W_{c}(Q)$ the poset of $c$-sortable elements in $W(Q)$.
A fundamental result is Gabriel’s Theorem, which can be found in [Ass-06, VII. 5]
Theorem 5.27.
Keep the above setting. The dimensional vector $\underline{\dim}(N)=(d_{1},\ldots,d_{n})^{T}$ of an indecomposable module $N$ in $\operatorname{\mathsf{mod}}\nolimits A$ gives a positive root $\beta_{N}:=d_{1}\alpha_{1}+\ldots+d_{n}\alpha_{n}$ in $\Phi(Q)$ and this correspondence gives a bijection from the indecomposable modules (up to isomorphisms) in $\operatorname{\mathsf{mod}}\nolimits A$ to the positive roots in $\Phi(Q)$.
For simplicity, sometimes we identify $\underline{\dim}(N)=(d_{1},\ldots,d_{n})^{T}$ with $\beta_{N}:=d_{1}\alpha_{1}+\ldots+d_{n}\alpha_{n}$ and thus view $\underline{\dim}(N)$ as a positive root in $\Phi(Q)$.
Theorem 5.28.
[Ingall-Thomas-09, Theorem 4.3] Keep the above setting. There is an isomorphism of poset $\varpi:\operatorname{\mathsf{tors}}(A)\rightarrow W_{c}(Q)$, $\mathcal{T}\mapsto w_{\mathcal{T}}$, where $w_{\mathcal{T}}\in W_{c}(Q)$ is defined by the property that
$\operatorname{\mathsf{inv}}(w_{\mathcal{T}})=\{\underline{\dim}(N)\mid N\text{is an indecomposable module in }\mathcal{T}\},$
where $\operatorname{\mathsf{inv}}(w_{\mathcal{T}})$ is the inversion set of $w_{\mathcal{T}}$.
Proposition 5.29.
Keep the above setting. Let $\mathcal{T}\in\operatorname{\mathsf{tors}}(A)$ and $w_{\mathcal{T}}$ the corresponding $c$-sortable element in $W_{Q}$. Then the following statements hold.
  • (i)
    There exists a $\tau$-tilting pair $(M,P)$ in $\operatorname{\mathsf{mod}}\nolimits A$ such that $\operatorname{\mathsf{Fac}}M=\mathcal{T}$.
  • (ii)
    Each indecomposable module in $\operatorname{\mathsf{mod}}\nolimits A$ is $\tau$-rigid.
  • (iii)
    The inversion set $\operatorname{\mathsf{inv}}(w_{\mathcal{T}})$ and the reduced dominant set
    $\operatorname{\mathsf{dom}}_{+}[(M,P)]=\{\text{indecomposable }\tau\text{-rigid modules in }\operatorname{\mathsf{Fac}}M\}$
    are related by $\operatorname{\mathsf{inv}}(w_{\mathcal{T}})=\{\underline{\dim}(N)\mid N\in\operatorname{\mathsf{dom}}_{+}[M,P]\}$.
Proof.
(i) Since $A=\mathbb{C}Q$ is representation-finite, the torsion class $\mathcal{T}\in\operatorname{\mathsf{tors}}(A)$ is functorially finite. Then by Theorem 5.15, there exists a $\tau$-tilting pair $(M,P)$ such that $\operatorname{\mathsf{Fac}}M=\mathcal{T}$.
(ii) Since $A=\mathbb{C}Q$ is the path algebra of a Dynkin quiver and by [Ass-06, Corollary 5.14], any indecomposable module $N\in\operatorname{\mathsf{mod}}\nolimits A$ is rigid, i.e., $\operatorname{\mathsf{Ext}}^{1}_{A}(N,N)=0$. Since $A$ is hereditary and by the Auslander-Reiten formula [Ass-06, Theorem 2.13], we have
$\operatorname{\mathsf{Hom}}_{A}(N,\tau N)\cong\operatorname{\mathsf{Ext}}^{1}_{A}(N,N)=0.$
Thus $N$ is $\tau$-rigid.
(iii) By (i) and (ii), we know that $\operatorname{\mathsf{dom}}_{+}[(M,P)]=\{\text{indecomposable modules in }\mathcal{T}\}$. Then the result follows from Theorem 5.28. ∎
For the Dynkin quiver $Q$, we define a skew-symmetrix matrix $B(Q)=(b_{ij})_{n\times n}$, where
$b_{ij}:=|i\rightarrow j|-|j\rightarrow i|,$
where $|i\rightarrow j|$ denotes the number of arrows from vertex $i$ to $j$ in $Q$. For example, if $Q=1\rightarrow 2$, then $B(Q)=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$.
Let $(\mathcal{S}_{X},\mathcal{S}_{Y},\Lambda)$ be a Langlands-Poisson triple with $B_{t_{0}}=B(Q)$ and let $\mathcal{A}_{Q}$ be the cluster algebra associated to the cluster pattern $\mathcal{S}_{X}$. It is known [bmrrt_2006, Theorem 4.5] that the unfrozen cluster variables (resp., seeds) of $\mathcal{A}_{Q}$ are in bijection with the indecomposable rigid objects (resp., basic cluster tilting objects) in the cluster category $\mathcal{C}_{Q}$ associated to $Q$. On the other hand, the indecomposable rigid objects (resp., basic cluster tilting objects) in $\mathcal{C}_{Q}$ are in bijection with the indecomposable $\tau$-rigid pairs (resp., basic $\tau$-tilting pairs) in $\operatorname{\mathsf{mod}}\nolimits A$ with $A=\mathbb{C}Q$, cf., [air_2014, Theorem 4.1]. Thus the cluster combinatorics of $\mathcal{A}_{Q}$ is categorified by the $\tau$-tilting combinatorics in $\operatorname{\mathsf{mod}}\nolimits A$.
Under the above bijections, the initial unfrozen cluster variables $x_{1;t_{0}},\ldots,x_{n;t_{0}}$ correspond to the indecomposable $\tau$-rigid pairs $(0,P_{1}),\ldots,(0,P_{n})$, and a non-initial unfrozen cluster variable $z$ corresponds to an indecomposable $\tau$-rigid module $M_{z}$ with the property
$\underline{\dim}(M_{z})={\bf d}_{z}^{t_{0}},$
where $\underline{\dim}(M_{z})$ is the dimension vector of $M_{z}\in\operatorname{\mathsf{mod}}\nolimits A$ and ${\bf d}_{z}^{t_{0}}$ is the $d$-vector of $z$ with respect to the initial seed of $\mathcal{A}_{Q}$.
Proposition 5.30.
Keep the above setting. Let $({\bf x}_{t},\widetilde{B}_{t})$ be a seed of $\mathcal{A}_{Q}$ and $(M_{t},P_{t})$ the corresponding $\tau$-tilting pair in $\operatorname{\mathsf{mod}}\nolimits A$. Let $w_{t}\in W_{c}(Q)$ be the $c$-sortable element in $W(Q)$ corresponding to the torsion class $\operatorname{\mathsf{Fac}}M_{t}\subseteq\operatorname{\mathsf{mod}}\nolimits A$. Then the correspondence
$\eta:z\mapsto d_{1;z}\alpha_{1}+\ldots+d_{n;z}\alpha_{n}$
is a bijection from the dominant set $\operatorname{\mathsf{dom}}^{t_{0}}[t]$ to the decorated inversion set
$\widetilde{\operatorname{\mathsf{inv}}}(w_{t})=\operatorname{\mathsf{inv}}(w_{t})\cup\{-\alpha_{1},\ldots,-\alpha_{n}\}$
of $w_{t}$, where ${\bf d}_{z}^{t_{0}}=(d_{1;z},\ldots,d_{n;z})^{T}$ is the $d$-vector of $z$ with respect to the initial seed.
Proof.
Under the categorification, the dominant set $\operatorname{\mathsf{dom}}^{t_{0}}[t]$ is identified with
$\operatorname{\mathsf{dom}}[M_{t},P_{t}]=\{\text{indecomposable }\tau\text{-rigid modules in }\operatorname{\mathsf{Fac}}M_{t}\}\cup\{(0,P_{1}),\ldots,(0,P_{n})\}.$
Let $z$ be a non-initial unfrozen cluster variable in $\operatorname{\mathsf{dom}}^{t_{0}}[t]$ and $M_{z}$ the corresponding indecomposable $\tau$-rigid module in $\operatorname{\mathsf{Fac}}M_{t}$ with $\underline{\dim}(M_{z})={\bf d}_{z}^{t_{0}}$. By Proposition 5.29 (iii), we know that $\eta(z)=d_{1;z}\alpha_{1}+\ldots+d_{n;z}\alpha_{n}$ belongs to $\operatorname{\mathsf{inv}}(w_{t})$.
Now let $z$ be an initial unfrozen cluster variable, say $z=x_{k;t}$, where $k\in[1,n]$. By Proposition 5.3 (i), we know that $z\in\operatorname{\mathsf{dom}}^{t_{0}}[t]$. Since ${\bf d}_{z}^{t_{0}}={\bf d}_{x_{k;t_{0}}}^{t_{0}}$ is the $k$th column of $-I_{n}$, we know that $\eta(z)=-\alpha_{k}$.
Therefore, the correspondence $\eta:z\mapsto d_{1;z}\alpha_{1}+\ldots+d_{n;z}\alpha_{n}$ gives a map from $\operatorname{\mathsf{dom}}^{t_{0}}[t]$ to the decorated inversion set $\widetilde{\operatorname{\mathsf{inv}}}(w_{t})$. By construction, this map is bijective. ∎
Example 5.31.
After taking the $d$-vectors (see Example 2.10), we can see that the dominant sets
$\operatorname{\mathsf{dom}}^{t_{0}}[t_{0}],\;\operatorname{\mathsf{dom}}^{t_{0}}[t_{1}],\;\operatorname{\mathsf{dom}}^{t_{0}}[t_{2}],\;\operatorname{\mathsf{dom}}^{t_{0}}[t_{3}],\;\operatorname{\mathsf{dom}}^{t_{0}}[t_{4}]$
in Example 5.8 get identified with the decorated inversion sets
$\widetilde{\operatorname{\mathsf{inv}}}(e),\;\widetilde{\operatorname{\mathsf{inv}}}(s_{1}),\;\widetilde{\operatorname{\mathsf{inv}}}(s_{1}s_{2}),\;\widetilde{\operatorname{\mathsf{inv}}}(s_{1}s_{2}\mid s_{1}),\;\widetilde{\operatorname{\mathsf{inv}}}(s_{2})$
in Example 5.26.
The more direct links between the dominant sets and inversion sets will be studied in a future work.
Acknowledgements
I would like to thank Antoine De Saint Germain, Professor Jiang-Hua Lu, and Jie Pan for their helpful discussions during my stay in Hong Kong and thank Professor Bernhard Keller for his email exchanges. I am grateful to Jiarui Fei for his valuable feedback on an earlier version of this paper.
This project have been partially supported by grants from the National Key R&D Program of China (2024YFA1013801), the Research Grants Council of the Hong Kong SAR, China (GRF 17307718, GRF 17306621, GRF 17303420), the National Natural Science Foundation of China (Grant No. 12071422), and the Guangdong Basic and Applied Basic Research Foundation (Grant No. 2021A1515012035).