### v4

# F-invariant in cluster algebras

Abstract.
We consider skew-symmetrizable (upper) cluster algebras with a compatible Poisson structure, called $\mathsf{\Lambda}$-(upper) cluster algebras. For any two good elements (e.g., cluster monomials) in a $\mathsf{\Lambda}$-upper cluster algebra, we introduce two invariants, called tropical invariant and $F$-invariant. We prove that (i) the product of two cluster monomials is still a cluster monomial if and only if their $F$-invariant is zero; (ii) if two cluster variables are log-canonical, then they are contained in the same cluster; and (iii) the notion of $F$-invariant for a pair of cluster monomials can be defined for any (upper) cluster algebra, regardless of whether it is a $\mathsf{\Lambda}$-(upper) cluster algebra.
When restricting to cluster monomials, we prove that the tropical invariant and $F$-invariant respectively coincide with the $\Lambda$-invariant and twice $\mathfrak{d}$-invariant in the monoidal cluster categorification using various monoidal subcategories of finite-dimensional modules over quantum affine algebras and quiver Hecke algebras; and we prove that the $F$-invariant coincides with the $E$-invariant in the additive cluster categorification using the theory of quivers with potentials.
Inspired by $F$-invariant, we introduce the dominant sets for seeds of cluster algebras as a replacement of torsion classes for $\tau$-tilting pairs in $\tau$-tilting theory. With the help of the dominant sets, we prove that the oriented exchange graphs of cluster algebras are acyclic. In particular, this implies that green mutations induce a partial order on the set of seeds (up to seed equivalence) of cluster algebras. We prove that the oriented exchange graphs of cluster algebras coincide with the Hasse quivers of the above posets of seeds.
2020 Mathematics Subject Classification:
13F60, 16G20, 18M05
1. Introduction
1.1. Backgrounds
Cluster algebras were introduced by Fomin and Zelevinsky [fz_2002] as a combinatorial approach to the dual canonical bases (or upper global bases) of quantum groups and to the theory of total positivity in algebraic groups. Such algebras often arise as the coordinate rings of spaces, such as double Bruhat cells [bfz_2005], unipotent cells [gls_2011], double Bott-Samelson cells [Shen-Weng-2021] and arise as the Grothendieck rings of certain monoidal subcategories of the representations of quantum affine algebras [HL_2010, HL_2013, kkop-2024] and quiver Hecke algebras [kkko-2018].
A cluster algebra is a $\mathbb{Z}$-subalgebra of a rational function field generated by a special set of generators called cluster variables, which are grouped into overlapping subsets, called clusters. A seed is a pair consisting of a cluster and a rectangular integer matrix with skew-symmetrizable principal part. One can obtain new seeds from a given one by a procedure called mutation. The sets of cluster variables and clusters of a cluster algebra are determined by an initial seed and the iterative mutations. A cluster momomial is a monomial in variables from the same cluster. In the main text, we actually use a broader definition of cluster monomials, which allows the presence of the inverse of frozen variables.
A fundamental problem in cluster algebras is to characterize when two cluster variables are contained in the same cluster. A more general problem is to characterize when the product of two cluster monomials is still a cluster monomial. Many progresses have been made by introducing different type of compatibility degrees or invariants.
Compatibility degrees in cluster algebras
In the study of Y-systems and generalized associahedra, Fomin and Zelevinsky [fz-2003y] introduce a function
$(-\mid\mid-)\colon\Phi_{\geq-1}\times\Phi_{\geq-1}\rightarrow\mathbb{Z}_{\geq 0}$
defined on the set $\Phi_{\geq-1}$ of almost positive roots associated to a Cartan matrix $C$ of finite type. Let $\mathcal{A}$ be a cluster algebra of finite type with the same Cartan-Killing type as the Cartan matrix $C$. Fomin and Zelevinsky [fz_2003f] proved that there is a bijection from $\Phi_{\geq-1}$ to the set $\mathcal{X}$ of (unfrozen) cluster variables of $\mathcal{A}$ and they proved that two cluster variables $x_{\alpha}$ and $x_{\beta}$ are contained in the same cluster if and only if their corresponding roots $\alpha,\beta\in\Phi_{\geq-1}$ are compatible, i.e., $(\alpha\mid\mid\beta)=0$.
Now let $\mathcal{A}$ be any skew-symmetrizable cluster algebra (not necessarily finite type). In [cao-li-2020], Li and the author of the present paper proved that there exists a function
$(-\mid\mid-)_{d}\colon\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{Z}_{\geq-1}$
defined on the set of (unfrozen) cluster variables $\mathcal{X}$ of $\mathcal{A}$, which is called the $d$-compatibility degree. The values of the $d$-compatibility degree $(-\mid\mid-)_{d}$ are given by the components of the $d$-vectors [fomin_zelevinsky_2007] in cluster algebras. It is proved that two cluster variables $x,z\in\mathcal{X}$ are contained in the same cluster if and only if
$(x\mid\mid z)_{d}=0\;\;\;\text{or}\;\;\;(x\mid\mid z)_{d}=-1,$
moreover, $(x\mid\mid z)_{d}=-1$ if and only if $x=z$. This confirms a conjecture on $d$-vectors proposed by Fomin and Zelevinsky [fomin_zelevinsky_2007, Conjecture 7.4].
Recently, Fu and Gyoda [Fu-Gyoda] introduced the $f$-compatibility degree in cluster algebras using the components of $f$-vectors, which is a function
$(-\mid\mid-)_{f}\colon\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{Z}_{\geq 0}$
defined on the set $\mathcal{X}$ of (unfrozen) cluster variables of $\mathcal{A}$. Based on the results in [cao-li-2020], Fu and Gyoda proved that two cluster variables $x,z\in\mathcal{X}$ are contained in the same cluster if and only if $(x\mid\mid z)_{f}=0$. When restricted to cluster algebras of finite type, the $f$-compatibility degree coincides with Fomin-Zelevinsky’s compatibility degree.
$\Lambda$-invariant and $\mathfrak{d}$-invariant
The $\Lambda$-invariant and $\mathfrak{d}$-invariant are introduced by Kang, Kashiwara, Kim and Oh [kkko-2018] in the representation theory of quiver Hecke algebras and by Kashiwara, Kim, Oh and Park [kkop-2020] in the representation theory of quantum affine algebras in their study of the monoidal categorification of (quantum) cluster algebras.
Here we briefly recall these notions in the quiver Hecke algebras case and we mainly follow [kkko-2018, kk_2019]. Let
$R\text{-gmod}=\bigoplus_{\beta\in Q^{+}}R(\beta)\text{-gmod}$
be the category of finite-dimensional graded modules with homomorphisms of degree 0 given in [kk_2019, Page 2270], which is a monoidal abelian category with the tensor product $\circ$} given there. Denote by $q$ the grading shift functor in $R\text{-gmod}$: for $M\in R\text{-gmod}$, we have $(qM)_{i}=M_{i-1}$ for $i\in\mathbb{Z}$. For each pair $(M,N)$ with $0\neq M\in R(\beta)\text{-gmod}$ and $0\neq N\in R(\gamma)\text{-gmod}$, Kang, Kashiwara, Kim and Oh [kkko-2018, Definition 2.2.1] constructed a renormalized $r$-matrix ${\bf r}_{M,N}$, which is a morphism from $M{\circ}N$ to $q^{c}N{\circ}M$ in $R\text{-gmod}$ for some integer $c$. The $\Lambda$-invariant $\Lambda(M,N)$ of the pair $(M,N)$ is defined as the homogeneous degree of ${\bf r}_{M,N}$, i.e.,
$\Lambda(M,N)=-c\;\in\mathbb{Z}.$
The $\mathfrak{d}$-invariant $\mathfrak{d}(M,N)$ of the pair $(M,N)$ is defined as half of the symmetrized sum:
$\mathfrak{d}(M,N)=\frac{1}{2}(\Lambda(M,N)+\Lambda(N,M)).$
It turns out that $\mathfrak{d}(M,N)\in\mathbb{Z}_{\geq 0}$, cf. [kkko-2018, Lemma 3.2.1].
Remark 1.1.
Denote by $R\text{-mod}^{0}$ the category obtained from $R\text{-gmod}$ by forgetting the gradings. By construction of the renormalized $r$-matrices, we have
$\Lambda(q^{a}M,q^{b}N)=\Lambda(M,N)\;\;\;\text{and}\;\;\;\mathfrak{d}(q^{a}M,q^{b}N)=\mathfrak{d}(M,N),\;\;\forall a,b\in\mathbb{Z}.$
Thus the $\Lambda$-invariant and $\mathfrak{d}$-invariant descend to the ungraded category $R\text{-mod}^{0}$.
A simple object $S$ in a monoidal abelian category is said to be real, if the tensor product of $S$ with itself is still simple. Kang, Kashiwara, Kim and Oh proved that if $M$ is a real simple object in $R\text{-gmod}$, then $\mathfrak{d}(M,M)=0$, cf. [kkko-2018, Lemma 3.2.3]. In monoidal categorification of cluster algebras, each cluster monomial corresponds to a real simple object. In particular, the $\mathfrak{d}$-invariant associated to a “cluster monomial” is zero.
Now let us turn to the $\mathfrak{d}$-invariant in the representation theory of quantum affine algebras. In this setting, Kashiwara, Kim, Oh and Park [kkop-2020] proved that (i) $\mathfrak{d}(M,N)$ takes non-negative values; (ii) if $M$ is a real simple object, then $\mathfrak{d}(M,M)=0$, cf. [kkop-2020, Proposition 3.16, Corollary 3.17].
$E$-invariant
The $E$-invariant is introduced by Derksen, Weyman and Zelevinsky [DWZ10] in the additive categorification of cluster algebras using decorated representations of quivers with potentials. For any two decorated representations $\mathcal{M}$ and $\mathcal{N}$ of a quiver with potential $(Q,W)$, Derksen, Weyman and Zelevinsky [DWZ10, Section 7] defined an integer $E^{\rm inj}(\mathcal{M},\mathcal{N})$, which is called the partial $E$-invariant of the pair $(\mathcal{M},\mathcal{N})$ in this paper (see Definition 6.2). The $E$-invariant $E^{\rm sym}(\mathcal{M},\mathcal{N})$ of the pair $(\mathcal{M},\mathcal{N})$ is defined to be the symmetrized sum:
$E^{\rm sym}(\mathcal{M},\mathcal{N})=E^{\rm inj}(\mathcal{M},\mathcal{N})+E^{\rm inj}(\mathcal{N},\mathcal{M})\;\;\in\mathbb{Z}.$
It is known from [DWZ10, Section 10] (or [CLS-2015, Section 3]) that
$E^{\rm inj}(\mathcal{M},\mathcal{N})\geq 0\;\;\;\text{and}\;\;\;E^{\rm sym}(\mathcal{M},\mathcal{N})\geq 0.$
Derksen, Weyman and Zelevinsky [DWZ10, Corollary 7.2] proved that $E^{\rm sym}(\mathcal{M},\mathcal{M})=0$ for any negative-reachable decorated representations, i.e., those decorated representations corresponding to cluster monomials.
Inspired by Derksen-Weyman-Zelevinsky’s $E$-invariant for quivers with potentials, i.e., for Jacobian algebras, the $E$-invariant for any finite-dimensional algebra is introduced by Derksen and Fei [DK-2015] in the study of general presentations and by Adachi, Iyama and Reiten [air_2014] in the study of $\tau$-tilting theory.
1.2. Results (I): Tropical invariant and $F$-invariant
We consider skew-symmetrizable (upper) cluster algebras with a compatible Poisson structure, called $\mathsf{\Lambda}$-(upper) cluster algebras. In a $\mathsf{\Lambda}$-upper cluster algebra, each seed $({\bf x}_{t},(\widetilde{B}_{t})_{m\times n})$ is endowed with a skew-symmetric integer matrix $(\Lambda_{t})_{m\times m}$, called Poisson coefficient matrix such that
(1.1) $\widetilde{B}_{t}^{T}\Lambda_{t}=(S\mid{\bf 0})$
for some fixed diagonal matrix $S=diag(s_{1},\ldots,s_{n})$ with $s_{i}>0$. We call the triple $({\bf x}_{t},\widetilde{B}_{t},\Lambda_{t})$ a $\mathsf{\Lambda}$-seed of $\mathcal{U}$.
For any two good elements (e.g., cluster monomials) in a $\mathsf{\Lambda}$-upper cluster algebra, we introduce two invariants, called tropical invariant and $F$-invariant, which generalize the $\Lambda$-invariant and $\mathfrak{d}$-invariant in the monoidal cluster categorification. The $F$-invariant also generalizes the $E$-invariant in the additive cluster categorification.
A good element is a compatibly pointed element in the sense of [qin_2019] satisfying the universally positive property (see Definition 3.6). We view such elements as a replacement of decorated representations of quivers with potentials and finite-dimensional graded modules of quiver Hecke algebras.
The Laurent expansion of a good element $u\in\mathcal{U}$ with respect to any seed $({\bf x}_{w},\widetilde{B}_{w})$ of $\mathcal{U}$ has a canonical expression:
$u={\bf x}_{w}^{{\bf g}_{u}^{w}}F_{u}^{w}(\hat{y}_{1;w},\ldots,\hat{y}_{n;w})$
where ${\bf g}_{u}^{w}\in\mathbb{Z}^{m}$ and $F_{u}^{w}\in\mathbb{Z}_{\geq 0}[y_{1},\ldots,y_{n}]$ is a polynomial with constant term $1$. They are respectively called the extended $g$-vector and $F$-polynomial of $u$ with respect to seed $({\bf x}_{w},\widetilde{B}_{w})$.
For each good element $u$ in a $\mathsf{\Lambda}$-upper cluster algebra $\mathcal{U}$, we introduce a semifield homomorphism
$\beta_{u}\colon\mathbb{F}_{>0}\coloneqq\mathbb{Q}_{\rm sf}(x_{1;t_{0}},\ldots,x_{m;t_{0}})\rightarrow\mathbb{Z}^{\rm max},$
where $\mathbb{F}_{>0}$ is the universal semifield generated by the initial cluster variables of $\mathcal{U}$ and $\mathbb{Z}^{\max}=(\mathbb{Z},+,\max)$ is the tropical semifield (see Proposition 4.4). Using such semifield homomorphisms, we introduce the tropical invariant and $F$-invariant.
Definition 1.2 (Tropical invariant and $F$-invariant).
Let $u$ and $u^{\prime}$ be two good elements in a $\mathsf{\Lambda}$-upper cluster algebra $\mathcal{U}$. The tropical invariant $\langle u,u^{\prime}\rangle_{\rm trop}$ and $F$-invariant $(u\mid\mid u^{\prime})_{F}$ of the pair $(u,u^{\prime})$ are defined as follows:
$\langle u,u^{\prime}\rangle_{\rm trop}\coloneqq\beta_{u^{\prime}}(u)\;\;\;\text{and}\;\;\;(u\mid\mid u^{\prime})_{F}=\beta_{u^{\prime}}(u)+\beta_{u}(u^{\prime})=\langle u,u^{\prime}\rangle_{\rm trop}+\langle u^{\prime},u\rangle_{\rm trop}.$
Given a non-zero polynomial
$F=\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}{\bf y}^{{\bf v}}\in\mathbb{Z}[y_{1},\ldots,y_{n}]$
and a vector ${\bf r}\in\mathbb{Z}^{n}$, we denote by
$F[{\bf r}]\coloneqq\max\{{\bf v}^{T}{\bf r}\mid c_{{\bf v}}\neq 0\}\;\in\mathbb{Z}.$
It is easy to see if $F$ has the constant term $1$, then $F[{\bf r}]\in\mathbb{Z}_{\geq 0}$ for any ${\bf r}\in\mathbb{Z}^{n}$.
The first main result in this paper is that for each $\mathsf{\Lambda}$-seed $({\bf x}_{w},\widetilde{B}_{w},\Lambda_{w})$ of $\mathcal{U}$, we have an explicit formula to calculate the tropical invariant and $F$-invariant.
Theorem 1.3 (Theorem 4.10).
Let $u$ and $u^{\prime}$ be two good elements in a $\mathsf{\Lambda}$-upper cluster algebra $\mathcal{U}$. Let
$u={\bf x}_{w}^{{\bf g}_{u}^{w}}F_{u}^{w}(\hat{y}_{1;w},\ldots,\hat{y}_{n;w}),\;\;\;u^{\prime}={\bf x}_{w}^{{\bf g}_{u^{\prime}}^{w}}F_{u^{\prime}}^{w}(\hat{y}_{1;w},\ldots,\hat{y}_{n;w})$
be the canonical expressions of $u$ and $u^{\prime}$ with respect to any vertex $w\in\mathbb{T}_{n}$. Then we have
  • (i)
    $\langle u,u^{\prime}\rangle_{\rm trop}=({\bf g}_{u}^{w})^{T}\Lambda_{w}{\bf g}_{u^{\prime}}^{w}+F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}]$, where $(S\mid{\bf 0})$ is as in (1.1).
  • (ii)
    $(u\mid\mid u^{\prime})_{F}=F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}]+F_{u^{\prime}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]$. In particular, $(u\mid\mid u^{\prime})_{F}\in\mathbb{Z}_{\geq 0}$.
  • (iii)
    $(u\mid\mid u)_{F}=0$ whenever $u$ is a cluster monomial.
Remark 1.4.
From the formula in (i), we see that the tropical invariant heavily depends on the Poisson coefficient matrix $\Lambda_{w}$. Since we define $F$-invariant by taking the symmetrized sum of tropical invariants, a prior, the $F$-invariant should also heavily depend on the Poisson coefficient matrix $\Lambda_{w}$. However, if we look at the formula in (ii), this formula eliminates the dependence of $(u\mid\mid u^{\prime})_{F}$ on the Poisson coefficient matrix and it only depends on the notion of $g$-vectors and $F$-polynomials. From this viewpoint, $F$-invariant can be defined for any skew-symmetrizable (upper) cluster algebra with trivial coefficients if we restrict to cluster monomials (see Proposition 4.30). The advantange to work on $\mathsf{\Lambda}$-(upper) cluster algebras is that we can define $F$-invariant for any pair of good elements and go beyond the class of cluster monomials.
Since the $F$-invariant $(u\mid\mid u^{\prime})_{F}$ can be written as
$(u\mid\mid u^{\prime})_{F}=F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}]+F_{u^{\prime}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}],$
we call the non-negative integer $F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}]$ the partial $F$-invariant of the pair $(u,u^{\prime})$ at vertex $w\in\mathbb{T}_{n}$.
As the $F$-invariant $(u\mid\mid u^{\prime})_{F}$ always takes non-negative values, it is natural to ask when $(u\mid\mid u^{\prime})_{F}=0$ holds.
Theorem 1.5 (Theorem 4.19).
Let $u$ and $u^{\prime}$ be two cluster monomials in an upper cluster algebra. Then $(u\mid\mid u^{\prime})_{F}=0$ if and only if the product $uu^{\prime}$ is still a cluster monomial.
For arbitrary two good elements $u$ and $u^{\prime}$ (not necessarily cluster monomials), we give a necessary condition for $(u\mid\mid u^{\prime})_{F}=0$ in Theorem 4.24. Before stating the result, we need some preparations.
Recall that $m$ is the number of cluster variables (frozen and unfrozen) in a cluster and $n$ is the number of unfrozen cluster variables in a cluster. By Definition 3.6, each good element $u$ corresponds to a tropical point ${\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ (associated to a $Y$-pattern). Two tropical points ${\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ and ${\bf g}^{\prime}]=\{({\bf g}^{\prime})^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ are said to be sign-coherent if for any vertex $t\in\mathbb{T}_{n}$ and any $k\in[1,n]$, we have $g_{k}^{t}\cdot(g_{k}^{\prime})^{t}\geq 0$, where $g_{k}^{t}$ and $(g_{k}^{\prime})^{t}$ are the $k$th components of ${\bf g}^{t}$ and $({\bf g}^{\prime})^{t}$.
Theorem 1.6 (Theorem 4.24).
Let $u$ and $u^{\prime}$ be two good elements in a $\mathsf{\Lambda}$-upper cluster algebra $\mathcal{U}$. If $(u\mid\mid u^{\prime})_{F}=0$, then the tropical points correspond to $u$ and $u^{\prime}$ are sign-coherent and the product $uu^{\prime}$ is still a good element in $\mathcal{U}$.
Since we work on a $\mathsf{\Lambda}$-upper cluster algebra $\mathcal{U}$, it has a natural compatible Poisson structure. By definition, if two cluster variables are in the same cluster, then they are log-canonical with respect to the compatible Poisson structure. We prove that the converse statement is also true.
Theorem 1.7 (Theorem 4.26).
Let $\mathcal{U}$ be a $\mathsf{\Lambda}$-upper cluster algebra. Then the following statements hold.
  • (i)
    Let $u\in\mathcal{U}$ be a good element and $x\in\mathcal{U}$ an unfrozen cluster variable. If $u$ and $x$ are log-canonical, then $(x\mid\mid u)_{F}=0$.
  • (ii)
    If two cluster variables are log-canonical, then they are contained in the same cluster.
1.3. Results (II): Comparison with some categorical invariants
As we have seen, the $F$-invariant shares the similar properties with the $E$-invariant in the additive cluster categorification and the $\mathfrak{d}$-invariant in the monoidal cluster categorification. It is natural to compare them.
We first focus on the monoidal cluster categorification [HL_2010, HL_2013]. A finite length, $\mathbf{k}$-linear, monoidal abelian category $\mathcal{C}$ is a monoidal categorification of a cluster algebra $\mathcal{A}^{+}$ (with non invertible frozen variables), if there exists a $\mathbb{Z}$-algebra isomorphism
$\varphi\colon\mathcal{K}_{0}(\mathcal{C})\rightarrow\mathcal{A}^{+}$
from the Grothendieck ring $\mathcal{K}_{0}(\mathcal{C})$ to the cluster algebra $\mathcal{A}^{+}$ such that the cluster monomials of $\mathcal{A}^{+}$ correspond to a subset of the simple objects (up to isomorphism) in $\mathcal{C}$. Such simple objects are said to be reachable. We denote by $(\mathcal{A}^{+},\mathcal{C},\varphi)$ for this monoidal categorification.
In order to unify the definition of $\Lambda$-invariant and $\mathfrak{d}$-invariant in different settings and unify the proofs for our results, we introduce the notion of $\mathsf{\Lambda}$-structure on a monoidal (abelian) category, which extracts the key properties of $\Lambda$-invariant appearing in the representation theory of quiver Hecke algebras [kkko-2018] and that of quantum affine algebras [kkop-2020].
Roughly speaking, a $\mathsf{\Lambda}$-structure on a monoidal category $\mathcal{C}$ is a pair $\mathsf{\Lambda}=(\Omega,\Lambda)$, where
  • $\Omega$ is a class of non-zero objects in $\mathcal{C}$ containing all the simple objects and closed under isomorphism, taking tensor products and non-zero subquotients;
  • ${\Lambda}:\Omega\times\Omega\rightarrow\mathbb{Z}$ is a map satisfying certain axiomatic conditions (see Definition 5.5).
If $\mathsf{\Lambda}=(\Omega,\Lambda)$ is a $\mathsf{\Lambda}$-structure on a monoidal category $\mathcal{C}$, we simply say that $\Lambda\colon\Omega\times\Omega\rightarrow\mathbb{Z}$ is a $\mathsf{\Lambda}$-structure on $\mathcal{C}$. For $M,N\in\Omega$, the numbers $\Lambda(M,N)$ and
$\mathfrak{d}(M,N)\coloneqq\frac{\Lambda(M,N)+\Lambda(N,M)}{2}$
are respectively called the $\Lambda$}-invariant and $\mathfrak{d}$-invariant of the pair $(M,N)$.
Definition 1.8 ($\mathsf{\Lambda}$-monoidal categorification [^1] We borrow this teminology from [kkop-2020, Definition 6.7].).
Let $(\mathcal{A}^{+},\mathcal{C},\varphi)$ be a monoidal categorification and $\Lambda\colon\Omega\times\Omega\rightarrow\mathbb{Z}$ a $\mathsf{\Lambda}$-structure on $\mathcal{C}$. We say that $(\mathcal{A}^{+},\mathcal{C},\varphi,\Omega,\Lambda)$ is a $\mathsf{\Lambda}$}-monoidal categorification if the following conditions hold.
  • (c1)
    Let $M,N\in\mathcal{C}$ be simple objects such that at least one of them is real. Then the head $M\nabla N$ of $M\otimes N$ is simple.
  • (c2)
    The $\mathsf{\Lambda}$-structure and cluster structure on $\mathcal{C}$ are compatible in the sense that there exists a monoidal cluster $\mathcal{M}_{w}=(M_{1;w},\ldots,M_{m;w})$ in $\mathcal{C}$ such that $(\widetilde{B}_{w},\Lambda_{w})$ forms a compatible pair, where $\widetilde{B}_{w}$ comes from the cluster structure and $\Lambda_{w}=(\lambda_{ij;w})$ with $\lambda_{ij;w}=\Lambda(M_{i;w},M_{j;w})$ comes from the $\mathsf{\Lambda}$-structure.
The examples of $\mathsf{\Lambda}$-monoidal categorifications contain those monoidal categorifications constructed by Kang, Kashiwara, Kim, and Oh [kkko-2018] using monoidal subcategories of representations of (symmetric) quiver Hecke algebras and by Kashiwara, Kim, Oh, and Park [kkop-2024] using monoidal subcategories of representations of quantum affine algebras.
Thanks to Proposition 5.13, the cluster algebra $\mathcal{A}^{+}$ above, together with the family of skew-symmetric matrices $\{\Lambda_{t}\mid t\in\mathbb{T}_{n}\}$ forms a $\mathsf{\Lambda}$-cluster algebra. Thus the tropical invariant $\langle u,u^{\prime}\rangle_{\rm trop}$ is defined for any pair of good elements in $\mathcal{A}^{+}$.
Theorem 1.9 (Theorem 5.16).
Let $(\mathcal{A}^{+},\mathcal{C},\varphi,\Omega,\Lambda)$ be a $\mathsf{\Lambda}$-monoidal categorification. Let $M,N$ be two simple objects in $\mathcal{C}$ such that $\varphi(M)$ and $\varphi(N)$ are good elements in $\mathcal{A}^{+}$. If $M$ or $N$ is reachable, then we have
$\Lambda(M,N)=\langle\varphi(M),\varphi(N)\rangle_{\rm trop}\;\;\;\text{and}\;\;\;2\mathfrak{d}(M,N)=(\varphi(M)\mid\mid\varphi(N))_{F}.$
Since the reachable simple objects correspond to cluster monomials, which are always good elements in $\mathcal{A}^{+}$, we have the following consequence.
Corollary 1.10.
Let $(\mathcal{A}^{+},\mathcal{C},\varphi,\Omega,\Lambda)$ be a $\mathsf{\Lambda}$-monoidal categorification. If $M$ and $N$ are reachable simple objects, then we have
$\Lambda(M,N)=\langle\varphi(M),\varphi(N)\rangle_{\rm trop}\;\;\;\text{and}\;\;\;2\mathfrak{d}(M,N)=(\varphi(M)\mid\mid\varphi(N))_{F}.$
As mentioned in [kkop-2020, Pages 1040–1041], the integer-valued invariants $\Lambda(M,N)$ and $\mathfrak{d}(M,N)$ provide important information in the representation theory of quiver Hecke algebra. However, in general, computing those values is quite difficult. Thanks to Theorem 1.9 and Theorem 1.3, we obtain explicit formulas expressing the $\Lambda$-invariant and $\mathfrak{d}$-invariant via $g$-vectors and $F$-polynomials, which can be easily implemented by a computer program.
Now, let us turn to the additive cluster categorification [DWZ08, DWZ10].
Theorem 1.11 (Theorem 6.11).
Let $(Q,W)$ be a quiver with non-degenerate potential and $\mathcal{A}_{Q}$ the corresponding cluster algebra with trivial coefficients. Let $u,u^{\prime}$ be two cluster monomials of $\mathcal{A}_{Q}$ and let $\mathcal{M}_{u},\mathcal{M}_{u^{\prime}}$ be the negative-reachable decorated representations of $(Q,W)$ corresponding to $u$ and $u^{\prime}$. Then
$(u\mid\mid u^{\prime})_{F}=E^{\rm sym}(\mathcal{M}_{u},\mathcal{M}_{u^{\prime}}).$
Notice that in the above theorem we take $S=I_{n}$ to be the fixed skew-symmetrizer for the exchange matrices of $\mathcal{A}_{Q}$, when we consider the $F$-invariant for cluster monomials in $\mathcal{A}_{Q}$.
1.4. Results (III): Dominant sets and oriented exchange graphs
Two seeds of a cluster algebra are equivalent, if they are the same up to relabeling. The exchange graph $\mathcal{H}=\mathcal{H}(\mathcal{A})$ of a cluster algebra $\mathcal{A}$ is the graph whose vertices correspond to the seeds (up to seed equivalence) of $\mathcal{A}$ and whose edges correspond to seed mutations. Thanks to the sign-coherence of $C$-matrices [GHKK18], each mutation has two states: green mutation or red mutation. Fix an initial seed at vertex $t_{0}\in\mathbb{T}_{n}$, the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}$ of $\mathcal{A}$ is the quiver whose underlying graph is the exchange graph $\mathcal{H}$ of $\mathcal{A}$ and whose orientation is induced by the green mutations.
$\tau$-tilting theory is introduced by Adachi, Iyama and Reiten [air_2014], which completes (classical) tilting theory from the viewpoint of mutation. Various of the fundamental concepts in cluster algebras, such as clusters/seeds, mutations, $C$-matrices, $G$-matrices, (oriented) exchange graphs, etc., were generalized to $\tau$-tilting theory.
The seeds in cluster algebras correspond to (basic) $\tau$-tilting pairs in $\tau$-tilting theory, which are in bijection with the functorially finite torsion classes in the module categories. The inclusion of subcategories induces a natural partial order on the set of functorially finite torsion classes and thus a partial order on the set of $\tau$-tilting pairs (up to isomorphism).
Thanks to the existence of the partial order, the oriented exchange graphs in $\tau$-tilting theory are acyclic. Actually, Adachi, Iyama and Reiten [air_2014, Theorem 0.6] proved that the oriented exchange graphs in $\tau$-tilting theory coincide with the Hasse quivers of the posets of functorially finite torsion classes, equivalently, the posets of $\tau$-tilting pairs.
It is natural to ask whether the oriented exchange graphs of (skew-symmetrizable) cluster algebras are acyclic and whether these oriented exchange graphs coincide with Hasse quivers of some posets.
Inspired by $F$-invariant, we introduce the dominant sets for seeds of cluster algebras as a replacement of torsion classes for $\tau$-tilting pairs in $\tau$-tilting theory. With the help of the dominant sets, we prove the following results on oriented exchange graphs.
Theorem 1.12 (Theorems 7.8, 7.13, Corollary 7.14).
The following statements hold.
  • (i)
    The oriented exchange graphs of cluster algebras are acyclic. In particular, green mutations induce a partial order on the set of seeds (up to seed equivalence) of cluster algebras.
  • (ii)
    The oriented exchange graphs of cluster algebras coincide with the Hasse quivers of the posets of seeds of cluster algebras.
We would like to point out that by ignoring initial cluster variables, the dominant sets in cluster algebras correspond to the sets of indecomposable $\tau$-rigid modules contained in the functorially finite torsion classes (see Proposition 7.23). In this sense, we view the dominant sets as a replacement of torsion classes in $\tau$-titling theory.
1.5. Conclusion and future work
In this paper, we introduce the tropical invariant and $F$-invariant for ($\mathsf{\Lambda}$)-cluster algebras with a focus on $F$-invariant. Up to a constant rescaling factor, we prove that the following three invariants coincide on cluster monomials: (i) the $\mathfrak{d}$-invariant in $\mathsf{\Lambda}$-monoidal cluster categorification; (ii) the $F$-invariant in cluster algebras; (iii) the $E$-invariant in additive cluster categorification. Notice that these three invariants are symmetric, e.g., $(u\mid\mid u^{\prime})_{F}=(u^{\prime}\mid\mid u)_{F}$. We prove the oriented exchange graphs of cluster algebras are acyclic by introducing the dominant sets for seeds of cluster algebras as a replacement of torsion classes for $\tau$-tilting pairs in $\tau$-tilting theory.
Now let us consider those $\mathsf{\Lambda}$-monoidal categorifications $(\mathcal{A}^{+},\mathcal{C},\varphi,\Omega,\Lambda)$ constructed by Kashiwara, Kim, Oh and Park [kkop-2024] using the representations of quantum affine algebras. For a pair $(M,N)$ of simple objects in $\mathcal{C}$, it is known [kkop-2020, Proposition 3.16] that the $\mathfrak{d}$-invariant can be written as
$\mathfrak{d}(M,N)=\mathfrak{o}(M,N)+\mathfrak{o}(N,M),$
where $\mathfrak{o}(M,N)\in\mathbb{Z}_{\geq 0}$ is the pole order of the normalized $R$-matrix $R^{\mathrm{norm}}_{M,N_{z}}$ at $z=1$. We may call $\mathfrak{o}(M,N)$ the partial $\mathfrak{d}$-invariant of the pair $(M,N)$. It is an important quantity in itself, particularly as a key ingredient in the construction of the generalized quantum affine Schur–Weyl duality functors due to [KKK-2018R].
In an ongoing work [^2] This work will be an extended version of the note [cao_2025]. with Ryo Fujita and Kota Murakami, we show that the following three (non-symmetric) partial invariants coincide on cluster monomials with a good choice of the reference seed: (i) the partial $\mathfrak{d}$-invariant in $\mathsf{\Lambda}$-monoidal cluster categorification using representations of quantum affine algebras; (ii) the partial $F$-invariant in cluster algebras; (iii) the partial $E$-invariant in additive cluster categorification. Theorem 1.9 will play an important role in the proof of the coincidence of the partial $F$-invariant and partial $\mathfrak{d}$-invariant. Such results will be used to prove the conjectural unified denominator formula [Fujita-Oh-2021, Conjecture 6.7] (equivalently, [Fujita-Murakami-2023, Conjecture 5.17]) for the normalized $R$-matrices between the KR-modules.
1.6. Organization
In Section 2, we recall some well-known concepts in cluster algebras, including $Y$-patterns, cluster patterns, compatible pairs, and $\mathsf{\Lambda}$-(upper) cluster algebras.
In Section 3, we recall the definition of tropical points for $Y$-patterns and cluster patterns, and compatibly pointed elements in upper cluster algebras. We give the definition of good elements for upper cluster algebras of full rank.
In Section 4, for each good element in a $\mathsf{\Lambda}$-upper cluster algebra, we introduce a semifield homomorphism (see Proposition 4.4) and we define the tropical invariant and $F$-invariant using such semifield homomorphisms. The main results in this section are Theorem 4.10, Theorem 4.19 , Theorem 4.24, Theorem 4.26. Notice that $F$-invariant can be defined for any skew-symmetrizable (upper) cluster algebra with trivial coefficients, when we only consider cluster monomials (see Proposition 4.30).
In Section 5 and Section 6, we compare the tropical invariant and $F$-invariant with some known invariants in additive/monoidal cluster categorification. When restricting to cluster monomials, we prove that the tropical invariant and $F$-invariant respectively coincide with the $\Lambda$-invariant and twice $\mathfrak{d}$-invariant in $\mathsf{\Lambda}$-monoidal categorification (see Theorem 5.16) and we prove that the $F$-invariant coincides with the $E$-invariant in additive cluster categorification (see Theorem 6.11).
In Section 7, we introduce the dominant sets for the seeds of cluster algebras as a replacement of torsion classes for $\tau$-tilting pairs in $\tau$-tilting theory. Then we prove the results on oriented exchange graphs (see Theorem 7.8 and Theorem 7.13).
2. Preliminaries on cluster algebras
Throughout this paper, we fix a pair $(n,m)$ of integers with $m\geq n>0$. For any positive integer $r$, we denote $1,r]\coloneqq\{1,\ldots,r\}$. The matrices in this paper are always integer matrices.
2.1. Mutation matrices and their mutations
Recall that an $n\times n$ integer matrix $B$ is said to be skew-symmetrizable, if there exists a diagonal integer matrix $S=diag(s_{1},\ldots,s_{n})$ with $s_{i}>0$ ($i\in[1,n]$) such that $SB$ is skew-symmetric. Such a diagonal matrix $S$ is called a skew-symmetrizer of $B$.
An $m\times n$ integer matrix $\widetilde{B}=\begin{bmatrix}B\\ P\end{bmatrix}=(b_{ij})$ is called a mutation matrix, if its top $n\times n$ submatrix $B$ is skew-symmetrizable. The submatrix $B$ is called the principal part of $\widetilde{B}$.
Let $\widetilde{B}=\begin{bmatrix}B\\ P\end{bmatrix}=(b_{ij})$ be an $m\times n$ mutation matrix. The mutation of $\widetilde{B}$ in direction $k\in[1,n]$ is defined to be the new integer matrix $\mu_{k}(\widetilde{B})=\widetilde{B}^{\prime}=\begin{bmatrix}B^{\prime}\\ P^{\prime}\end{bmatrix}=(b_{ij}^{\prime})$ given by
(2.1) $b_{ij}^{\prime}=\begin{cases}-b_{ij},&\text{if}\;i=k\;\text{or}\;j=k;\\ b_{ij}+[b_{ik}]_{+}[b_{kj}]_{+}-[-b_{ik}]_{+}[-b_{kj}]_{+},&\text{otherwise},\end{cases}$
where $a]_{+}\coloneqq\max\{a,0\}$ for any $a\in\mathbb{R}$.
The following statements are well-known, cf. [fz_2002, Proposition 4.5], [bfz_2005, Lemma 3.2].
  • $\mu_{k}(\widetilde{B})=\widetilde{B}^{\prime}$ is still a mutation matrix, namely, the submatrix $B^{\prime}$ of $\widetilde{B}^{\prime}$ is skew-symmetrizable;
  • $B$ and $B^{\prime}$ share the same skew-symmetrizer;
  • $\mu_{k}$ is an involution, i.e., $\mu_{k}^{2}=id$;
[…]
(2.2) $\widetilde{B}^{\prime}=\mu_{k}(\widetilde{B})=E_{k,\varepsilon}^{\widetilde{B}}\widetilde{B}F_{k,\varepsilon}^{\widetilde{B}}$,
where $\varepsilon\in\{\pm\}$ and
  • $E_{k,\varepsilon}^{\widetilde{B}}$ is the $m\times m$ matrix obtained by replacing the $k$th column of $I_{m}$ by the column vector $(e_{1},\ldots,e_{m})^{T}$, where
    $e_{i}=\begin{cases}-1,&\text{if }i=k;\\ [-\varepsilon b_{ik}]_{+},&\text{if }i\neq k.\end{cases}$
  • $F_{k,\varepsilon}^{\widetilde{B}}$ is the $n\times n$ matrix obtained by replacing the $k$th row of $I_{n}$ by the row vector $(f_{1},\ldots,f_{n})$, where
    $f_{i}=\begin{cases}-1,&\text{if }i=k;\\ [\varepsilon b_{ki}]_{+},&\text{if }i\neq k.\end{cases}$
It turns out that $\widetilde{B}^{\prime}$ is independent of the choice of a sign $\varepsilon\in\{\pm\}$ in (2.2). The matrices $E_{k,\varepsilon}^{\widetilde{B}}$ and $F_{k,\varepsilon}^{\widetilde{B}}$ defined above coincide with the ones used in [bz-2005, (3.2), (3.3)]. It is easy to check
(2.3) $(E_{k,\varepsilon}^{\widetilde{B}})^{2}=I_{m}\;\;\;\text{and}\;\;\;(F_{k,\varepsilon}^{\widetilde{B}})^{2}=I_{n}.$
2.2. Compatible pairs and their mutations
Let $\Lambda=(\lambda_{ij})$ be an $m\times m$ skew-symmetric integer matrix, and let $\widetilde{B}=\begin{bmatrix}B\\ P\end{bmatrix}=(b_{ij})$ be an $m\times n$ integer matrix, where $B$ is the top $n\times n$ submatrix of $\widetilde{B}$. Here we do not assume that $B$ is skew-symmetrizable.
Definition 2.1 (Compatible pair, [bz-2005]).
Keep $\Lambda$ and $\widetilde{B}$ as above. Let $S=diag(s_{1},\ldots,s_{n})$ be a diagonal integer matrix with $s_{j}>0,\;j\in[1,n]$. The pair $(\widetilde{B},\Lambda)$ is called a compatible pair of type $S$, if for any $j\in[1,n]$ and $i\in[1,m]$, we have
(2.4) $\sum_{k=1}^{m}b_{kj}\lambda_{ki}=\delta_{i,j}s_{j}.$
In other words, $\widetilde{B}^{T}\Lambda=(S\mid{\bf 0})$, where $\bf 0$} is the $n\times(m-n)$ zero matrix.
Notice that the diagonal matrix $S$ is uniquely determined by the compatible pair $(\widetilde{B},\Lambda)$. The matrix $\Lambda$ in a compatible pair $(\widetilde{B},\Lambda)$ is called a Poisson coefficient matrix.
Proposition 2.2 ([bz-2005, Proposition 3.3]).
Let $(\widetilde{B},\Lambda)$ be a compatible pair of type $S$ and $B$ the top $n\times n$ submatrix of $\widetilde{B}$. Then the matrix $\widetilde{B}$ has the full rank $n$ and $SB=\widetilde{B}^{T}\Lambda\widetilde{B}$. In particular, $B$ is skew-symmetrizable.
Conversely, by [gsv-2003, Theorem 1.4], whenever $\widetilde{B}$ is a mutation matrix of full rank, we can find a skew-symmetric integer matrix $\Lambda$ such that $(\widetilde{B},\Lambda)$ is a compatible pair.
Let $(\widetilde{B},\Lambda)$ be a compatible pair of type $S$ and $k\in[1,n]$. We set $\widetilde{B}^{\prime}\coloneqq\mu_{k}(\widetilde{B})=E_{k,\varepsilon}^{\widetilde{B}}\widetilde{B}F_{k,\varepsilon}^{\widetilde{B}}$ and
$\Lambda^{\prime}\coloneqq(E_{k,\varepsilon}^{\widetilde{B}})^{T}\Lambda E_{k,\varepsilon}^{\widetilde{B}}=(\lambda_{ij}^{\prime}).$
We have
(2.5) $\lambda_{ij}^{\prime}=\begin{cases}-\lambda_{ik}+\sum_{l=1}^{m}[-\varepsilon b_{lk}]_{+}\lambda_{il},&\text{if }j=k;\\ -\lambda_{kj}+\sum_{l=1}^{m}[-\varepsilon b_{lk}]_{+}\lambda_{lj},&\text{if }i=k;\\ \lambda_{ij},&\text{otherwise}.\end{cases}$
Proposition 2.3 ([bz-2005, Proposition 3.4]).
  • (i)
    The matrix $\Lambda^{\prime}$ is independent of the choice of $\varepsilon\in\{\pm\}$.
  • (ii)
    The pair $(\widetilde{B}^{\prime},\Lambda^{\prime})$ is still a compatible pair of type $S$.
Keep the above setting. We call $(\widetilde{B}^{\prime},\Lambda^{\prime})$ the mutation of the compatible pair $(\widetilde{B},\Lambda)$ in direction $k$ and denote $(\widetilde{B}^{\prime},\Lambda^{\prime})=\mu_{k}(\widetilde{B},\Lambda)$. It is known from [bz-2005, Proposition 3.6] that $\mu_{k}(\widetilde{B}^{\prime},\Lambda^{\prime})=(\widetilde{B},\Lambda)$, i.e., $\mu_{k}$ is an involution.
The following proposition is very useful to construct many examples of compatible pairs from a given skew-symmetrizable matrix.
Proposition 2.4 ([zelevinsky-2005, Example 0.5]).
Let $B$ be an $n\times n$ skew-symmetrizable integer matrix and set $\widetilde{B}=\begin{bmatrix}B\\ I_{n}\end{bmatrix}$. Then the skew-symmetric matrices $\Lambda$ compatible with $\widetilde{B}$ in the sense of (2.4) are those of the form
$\Lambda=\begin{bmatrix}\Lambda_{0}&-S-\Lambda_{0}B\\ S-B^{T}\Lambda_{0}&-SB+B^{T}\Lambda_{0}B\end{bmatrix}$,
where $\Lambda_{0}$ is an arbitrary skew-symmetric integer $n\times n$ matrix and $S$ is an arbitrary skew-symmetrizer of $B$.
2.3. $Y$-pattern and cluster pattern
Recall that we fixed a pair $(n,m)$ of integers with $m\geq n>0$. Let $\mathbb{F}$ be the field of rational functions over $\mathbb{Q}$ in $m$ variables.
[…]
  • ${\bf y}=(y_{1},\ldots,y_{m})$ satisfies that $\{y_{1},\ldots,y_{m}\}$ is a free generating set of $\mathbb{F}$ over $\mathbb{Q}$;
  • $\widehat{B}=(B\mid Q)=(\hat{b}_{ij})$ is an $n\times m$ integer matrix such that its leftmost $n\times n$ submatrix $B$ is skew-symmetrizable.
[…]
(2.6) $y_{i}^{\prime}$ $=\begin{cases}y_{k}^{-1},&\text{if}\;i=k;\\ y_{i}y_{k}^{[\hat{b}_{ki}]_{+}}(1+y_{k})^{-\hat{b}_{ki}},&\text{otherwise}.\end{cases}$
$\hat{b}_{ij}^{\prime}$ $=\begin{cases}-\hat{b}_{ij},&\text{if}\;i=k\;\text{or}\;j=k;\\ \hat{b}_{ij}+[\hat{b}_{ik}]_{+}[\hat{b}_{kj}]_{+}-[-\hat{b}_{ik}]_{+}[-\hat{b}_{kj}]_{+},&\text{otherwise}.\end{cases}$
[…]
  • ${\bf x}=(x_{1},\ldots,x_{m})$ satisfies that $\{x_{1},\ldots,x_{m}\}$ is a free generating set of $\mathbb{F}$ over $\mathbb{Q}$;
  • $\widetilde{B}=\begin{bmatrix}B\\ P\end{bmatrix}=(b_{ij})$ is an $m\times n$ mutation matrix.
[…]
(2.7) $x_{i}^{\prime}=\begin{cases}x_{i},&\text{if}\;i\neq k;\\ x_{k}^{-1}\cdot(\prod_{j=1}^{m}x_{j}^{[b_{jk}]_{+}}+\prod_{j=1}^{m}x_{j}^{[-b_{jk}]_{+}}),&\text{if}\;i=k.\end{cases}$
[…]
Let $\mathbb{T}_{n}$ denote the $n$-regular tree. We label the edges of $\mathbb{T}_{n}$ by $1,\ldots,n$ such that the $n$ different edges adjacent to the same vertex of $\mathbb{T}_{n}$ receive different labels.
Definition 2.5.
(i) A $Y$-pattern $\mathcal{S}_{Y}=\{({\bf y}_{t},\widehat{B}_{t})\mid t\in\mathbb{T}_{n}\}$ of rank $n$ is an assignment of a $Y$-seed $({\bf y}_{t},\widehat{B}_{t})$ of rank $n$ to every vertex $t$ of $\mathbb{T}_{n}$ such that $({\bf y}_{t^{\prime}},\widehat{B}_{t^{\prime}})=\mu_{k}({\bf y}_{t},\widehat{B}_{t})$ whenever ${t}$${t^{\prime}}$$\scriptstyle{k}$ in $\mathbb{T}_{n}$.
(ii) A cluster pattern $\mathcal{S}_{X}=\{({\bf x}_{t},\widetilde{B}_{t})\mid t\in\mathbb{T}_{n}\}$ of rank $n$ is an assignment of a cluster seed $({\bf x}_{t},\widetilde{B}_{t})$ of rank $n$ to every vertex $t$ of $\mathbb{T}_{n}$ such that $({\bf x}_{t^{\prime}},\widetilde{B}_{t^{\prime}})=\mu_{k}({\bf x}_{t},\widetilde{B}_{t})$ whenever ${t}$${t^{\prime}}$$\scriptstyle{k}$ in $\mathbb{T}_{n}$.
We usually write ${\bf y}_{t}=(y_{1;t},\ldots,y_{m;t})$, $\widehat{B}_{t}=(B_{t}\mid Q_{t})=(\hat{b}_{ij;t})$, ${\bf x}_{t}=(x_{1;t},\ldots,x_{m;t})$ and
$\widetilde{B}_{t}=\begin{pmatrix}B_{t}\\ P_{t}\end{pmatrix}=(b_{ij;t}).$
We write $\widehat{\bf y}_{t}=(\hat{y}_{1;t},\ldots,\hat{y}_{n;t})$ for the collection of $\hat{y}-$variables associated to seed $({\bf x}_{t},\widetilde{B}_{t})$, where
$\hat{y}_{k;t}=\prod_{j=1}^{m}x_{j;t}^{b_{jk;t}}.$
In this paper, we usually fix a vertex $t_{0}$ as the rooted vertex of the $n$-regular tree $\mathbb{T}_{n}$. Clearly, both $Y$-pattern and cluster pattern are uniquely determined by the data at the rooted vertex $t_{0}$.
Definition 2.6.
Let $\mathcal{S}_{X}=\{({\bf x}_{t},\widetilde{B}_{t})\mid t\in\mathbb{T}_{n}\}$ be a cluster pattern, $\mathcal{S}_{Y}=\{({\bf y}_{t},\widehat{B}_{t})\mid t\in\mathbb{T}_{n}\}$ a $Y$-pattern, and let $\mathsf{\Lambda}=\{\Lambda_{t}\mid t\in\mathbb{T}_{n}\}$ be a collection of $m\times m$ skew-symmetric integer matrices indexed by the vertices in $\mathbb{T}_{n}$.
  • (i)
    The pair $(\mathcal{S}_{X},\mathcal{S}_{Y})$ is called a Langlands dual pair, if $\widehat{B}_{t_{0}}=-\widetilde{B}_{t_{0}}^{T}$ holds for the rooted vertex $t_{0}$.
  • (ii)
    The pair $(\mathcal{S}_{X},\mathsf{\Lambda})$ is called a $\mathsf{\Lambda}$-cluster pattern, if $\{(\widetilde{B}_{t},\Lambda_{t})\mid t\in\mathbb{T}_{n}\}$ forms a collection of compatible pairs and $(\widetilde{B}_{t^{\prime}},\Lambda_{t^{\prime}})=\mu_{k}(\widetilde{B}_{t},\Lambda_{t})$ whenever ${t}$${t^{\prime}}$$\scriptstyle{k}$ in $\mathbb{T}_{n}$. In this case, we call the triple $({\bf x}_{t},\widetilde{B}_{t},\Lambda_{t})$ a $\mathsf{\Lambda}$-seed.
  • (iii)
    The triple $(\mathcal{S}_{X},\mathcal{S}_{Y},\mathsf{\Lambda})$ is called a Langlands-Poisson triple, if $(\mathcal{S}_{X},\mathsf{\Lambda})$ is a $\mathsf{\Lambda}$-cluster pattern and $(\mathcal{S}_{X},\mathcal{S}_{Y})$ is a Langlands dual pair.
The above definition is purely for convenience of this paper. Notice that each cluster pattern has a unique Langlands dual $Y$-pattern up to isomorphism. The following result can be checked easily.
Corollary 2.7.
Let $(\mathcal{S}_{X},\mathcal{S}_{Y})$ be a Langlands dual pair. Then $\widehat{B}_{t}=-\widetilde{B}_{t}^{T}$ holds for any vertex $t$ of $\mathbb{T}_{n}$.
Given a $\mathsf{\Lambda}$-cluster pattern $(\mathcal{S}_{X},\mathsf{\Lambda})$, we can define a Poisson bracket $\{-,-\}$ on the ambient field $\mathbb{F}$ using the $m\times m$ skew-symmetric matrix $\Lambda_{t_{0}}=(\lambda_{ij;t_{0}})$ at the rooted vertex $t_{0}$:
$\{x_{i;t_{0}},x_{j;t_{0}}\}\coloneqq\lambda_{ij;t_{0}}\cdot x_{i;t_{0}}x_{j;t_{0}}.$
It turns out this Poisson bracket is compatible with the cluster pattern $\mathcal{S}_{X}$, that is, for any cluster ${\bf x}_{t}$ of $\mathcal{S}_{X}$, we have
$\{x_{i;t},x_{j;t}\}=\lambda_{ij;t}\cdot x_{i;t}x_{j;t},$
where $\lambda_{ij;t}$ is the $(i,j)$-entry of $\Lambda_{t}$, cf. [gsv-2003], [bz-2005, Remark 4.6].
Finally, let us give the definition of (upper) cluster algebras.
Definition 2.8.
Let $\mathcal{S}_{X}=\{({\bf x}_{t},\widetilde{B}_{t})\mid t\in\mathbb{T}_{n}\}$ be a cluster pattern of rank $n$ in $\mathbb{F}$.
  • (i)
    The cluster algebra $\mathcal{A}$ associated to $\mathcal{S}_{X}$ is the $\mathbb{Z}$-subalgebra of $\mathbb{F}$ given by
    $\mathcal{A}\coloneqq\mathbb{Z}[x_{1;t},\ldots,x_{n;t},x_{n+1;t}^{\pm 1},\ldots,x_{m;t}^{\pm 1}\mid t\in\mathbb{T}_{n}].$
    Denote by $\mathcal{A}^{+}\coloneqq\mathbb{Z}[x_{1;t},\ldots,x_{m;t}\mid t\in\mathbb{T}_{n}]\subseteq\mathcal{A}$, the version of cluster algebra with frozen variables non inverted.
  • (ii)
    The upper cluster algebra $\mathcal{U}$ associated to $\mathcal{S}_{X}$ is the $\mathbb{Z}$-subalgebra of $\mathbb{F}$ defined by
    $\mathcal{U}\coloneqq\bigcap\limits_{t\in\mathbb{T}_{n}}\mathcal{L}(t),$
    where $\mathcal{L}(t)\coloneqq\mathbb{Z}[x_{1;t}^{\pm 1},\ldots,x_{m;t}^{\pm 1}]$.
  • (iii)
    If $(\mathcal{S}_{X},\mathsf{\Lambda})$ is a $\mathsf{\Lambda}$-cluster pattern, then the corresponding (upper) cluster algebra $\mathcal{A}$, $\mathcal{A}^{+}$ and $\mathcal{U}$ are endowed with an extra data $\mathsf{\Lambda}=\{\Lambda_{t}\mid t\in\mathbb{T}_{n}\}$. In this case, we call $\mathcal{A}$ or $\mathcal{A}^{+}$ a $\mathsf{\Lambda}$-cluster algebra and call $\mathcal{U}$ a $\mathsf{\Lambda}$-upper cluster algebra.
  • (iv)
    If $m=n$, i.e., there are no frozen variables, the corresponding (upper) cluster algebra is said to be with trivial coefficients.
Notice that the mutation matrices of $\mathsf{\Lambda}$-(upper) cluster algebras always have full rank, by Proposition 2.2.
Theorem 2.9 ([fz_2002]*Laurent phenomenon).
For any cluster variable $z$ and any seed $({\bf x}_{t},\widetilde{B}_{t})$ of a cluster algebra $\mathcal{A}$, we have
$z\in\mathbb{Z}[x_{1;t}^{\pm 1},\ldots,x_{n;t}^{\pm 1},x_{n+1;t},\ldots,x_{m;t}].$
In particular, $\mathcal{A}$ is contained in $\mathcal{U}$.
Example 2.10.
[…]
$x_{1},\;x_{2},\;x_{3}\coloneqq\frac{x_{2}+1}{x_{1}},\;x_{4}\coloneqq\frac{x_{1}+x_{2}+1}{x_{1}x_{2}},\;x_{5}\coloneqq\frac{x_{1}+1}{x_{2}}.$
It is clear that all the cluster variables are contained in $\mathbb{Z}[x_{1}^{\pm 1},x_{2}^{\pm 1}]$.
3. Tropical points, pointed elements and good elements
3.1. Universal semifield and tropical points
Recall that $(\mathbb{P},\cdot,\oplus)$ is called a semifield if $(\mathbb{P},\cdot)$ is an abelian multiplicative group endowed with a binary operation of auxiliary addition $\oplus$ which is commutative, associative and satisfies that the multiplication distributes over the auxiliary addition. For example,
$\mathbb{Z}^{\rm min}\coloneqq(\mathbb{Z},+,\min\{-,-\})\;\;\;\text{and}\;\;\;\mathbb{Z}^{\rm max}\coloneqq(\mathbb{Z},+,\max\{-,-\})$
are semifield and they are called tropical semifield.
Let $\mathbb{Q}_{\rm sf}(u_{1},\ldots,u_{m})$ be the set of all non-zero rational functions in $u_{1},\ldots,u_{m}$ that have subtraction free expressions. The set $\mathbb{Q}_{\rm sf}(u_{1},\ldots,u_{m})$ is a semifield with respect to the usual operations of multiplication and addition. It is called an universal semifield.
From now on, we fix $\mathbb{F}\coloneqq\mathbb{Q}(u_{1},\ldots,u_{m})$ and $\mathbb{F}_{>0}\coloneqq\mathbb{Q}_{\rm sf}(u_{1},\ldots,u_{m})$. For any semifield $\mathbb{P}$ and ${\bf p}=(p_{1},\ldots,p_{m})^{T}\in\mathbb{P}^{m}$, there exists a unique semifield homomorphism
${\pi}_{\bf p}:\mathbb{F}_{>0}\rightarrow\mathbb{P}$
induced by $u_{i}\mapsto p_{i}$ for any $i$. The map $\pi\colon{\bf p}\mapsto\pi_{\bf p}$ induces a bijection from $\mathbb{P}^{m}$ to the set $\operatorname{\mathsf{Hom}}_{\rm sf}(\mathbb{F}_{>0},\mathbb{P})$ of semifield homomorphisms from $\mathbb{F}_{>0}$ to $\mathbb{P}$.
In this paper, we often take $\mathbb{P}=\mathbb{Z}^{\rm max}=(\mathbb{Z},+,\max)$. The elements in $\operatorname{\mathsf{Hom}}_{\rm sf}(\mathbb{F}_{>0},\mathbb{Z}^{\rm max})$ are called tropical $\mathbb{Z}^{\rm max}$-points or simply tropical points, which are in bijection with the points in $\mathbb{Z}^{m}$.
Definition 3.1.
A chart in $\mathbb{F}_{>0}=\mathbb{Q}_{\rm sf}(u_{1},\ldots,u_{m})$ is an $m$-tuple ${\bf z}=(z_{1},\ldots,z_{m})$ of elements in $\mathbb{F}_{>0}$ that generate $\mathbb{F}_{>0}$ as a semifield.
The statements in the following proposition are clear.
Proposition 3.2.
Let $\mathcal{C}=\{{\bf u}_{t}\mid t\in\mathbb{T}\}$ be a collection of charts in $\mathbb{F}_{>0}=\mathbb{Q}_{\rm sf}(u_{1},\ldots,u_{m})$ indexed by a set $\mathbb{T}$. Then the following statements hold.
  • (i)
    For any chart ${\bf u}_{t}=(u_{1;t},\ldots,u_{m;t})$, denote by $u_{i}({\bf u}_{t})$ the expression of $u_{i}$ as an element in $\mathbb{F}(t)_{>0}\coloneqq\mathbb{Q}_{\rm sf}(u_{1;t},\ldots,u_{m;t})$. The map $\varphi_{t}\colon u_{i}\mapsto u_{i}({\bf u}_{t})$ induces an isomorphism of semifields from $\mathbb{F}_{>0}$ to $\mathbb{F}(t)_{>0}$.
  • (ii)
    For any chart ${\bf u}_{t}=(u_{1;t},\ldots,u_{m;t})$ and any $\beta\in\operatorname{\mathsf{Hom}}_{\rm sf}(\mathbb{F}_{>0},\mathbb{Z}^{\rm max})$, denote by
    ${\bf q}^{t}(\beta)\coloneqq(\beta(u_{1;t}),\ldots,\beta(u_{m;t}))^{T}\in\mathbb{Z}^{m}.$
    Then the following diagram commutes.
    ${\mathbb{F}_{>0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi_{t}\;\;\;}$$\scriptstyle{\beta}$${\mathbb{F}(t)_{>0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{{\bf q}^{t}(\beta);t}}$${\mathbb{Z}^{\rm max}}$
    where $\pi_{{\bf q}^{t}(\beta);t}$ is the unique semifield homomorphism from $\mathbb{F}(t)_{>0}$ to $\mathbb{Z}^{\rm max}$ induced by $u_{i;t}\mapsto\beta(u_{i;t})$ for any $i$.
  • (iii)
    The map $\beta\mapsto{\bf q}^{t}(\beta)$ induces a bijection from the tropical points in $\operatorname{\mathsf{Hom}}_{\rm sf}(\mathbb{F}_{>0},\mathbb{Z}^{\rm max})$ and the points in $\mathbb{Z}^{m}$.
Keep the setting in the above proposition. We call the column vector
${\bf q}^{t}(\beta)=(\beta(u_{1;t}),\ldots,\beta(u_{m;t}))^{T}\in\mathbb{Z}^{m}$
the coordinate vector of the tropical point $\beta$ under the chart ${\bf u}_{t}$. When a collection $\mathcal{C}=\{{\bf u}_{t}\mid t\in\mathbb{T}\}$ of charts in $\mathbb{F}_{>0}$ is given, we often identify a tropical point $\beta\in\operatorname{\mathsf{Hom}}_{\rm sf}(\mathbb{F}_{>0},\mathbb{Z}^{\rm max})$ with the collection $\{{\bf q}^{t}(\beta)\in\mathbb{Z}^{m}\mid t\in\mathbb{T}\}$ of coordinate vectors of $\beta$ under the charts.
Now let us go back to the situations that we are interested in.
  • If we work on a $Y$-pattern $\mathcal{S}_{Y}=\{({\bf y}_{t},\widehat{B}_{t})\mid t\in\mathbb{T}_{n}\}$, we take $\mathbb{F}_{>0}$ to be the universal semifield generated by the initial $y$-variables $y_{1;t_{0}},\ldots,y_{m;t_{0}}$. Then $\mathcal{C}\coloneqq\{{\bf y}_{t}\mid t\in\mathbb{T}_{n}\}$ is a collection of charts in $\mathbb{F}_{>0}$. We usually denote by
    $\{{\bf g}^{t}(\beta)\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$
    the collection of coordinate vectors of a tropical point $\beta\in\operatorname{\mathsf{Hom}}_{\rm sf}(\mathbb{F}_{>0},\mathbb{Z}^{\rm max})$ for the $Y$-pattern case.
  • If we work on a cluster pattern $\mathcal{S}_{X}=\{({\bf x}_{t},\widetilde{B}_{t})\mid t\in\mathbb{T}_{n}\}$, we take $\mathbb{F}_{>0}$ to be the universal semifield generated by the initial cluster variables $x_{1;t_{0}},\ldots,x_{m;t_{0}}$. Then $\mathcal{C}\coloneqq\{{\bf x}_{t}\mid t\in\mathbb{T}_{n}\}$ is a collection of charts in $\mathbb{F}_{>0}$. We usually denote by
    $\{{\bf a}^{t}(\beta)\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$
    the collection of coordinate vectors of a tropical point $\beta\in\operatorname{\mathsf{Hom}}_{\rm sf}(\mathbb{F}_{>0},\mathbb{Z}^{\rm max})$ for the cluster pattern case.
Since the transition maps between any two adjacent charts are clear in both $Y$-pattern and cluster pattern, the corresponding tropical points (after identifying with the corresponding coordinate vectors) can be defined using the tropical version of the transition maps in (2.6) and (2.7).
Definition 3.3 (Tropical points).
Let $\mathcal{S}_{Y}=\{({\bf y}_{t},\widehat{B}_{t})\mid t\in\mathbb{T}_{n}\}$ be a $Y$-pattern of rank $n$ in $\mathbb{F}$ and $\mathcal{S}_{X}=\{({\bf x}_{t},\widetilde{B}_{t})\mid t\in\mathbb{T}_{n}\}$ a cluster pattern of rank $n$ in $\mathbb{F}$.
  • (i)
    A tropical point ${\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ associated to the $Y$-pattern $\mathcal{S}_{Y}$ is an assignment of a (column) vector ${\bf g}^{t}=(g_{1}^{t},\ldots,g_{m}^{t})^{T}$ in $\mathbb{Z}^{m}$ to each vertex $t$ of $\mathbb{T}_{n}$ such that
    (3.1) $g_{i}^{t^{\prime}}=\begin{cases}-g_{k}^{t},&\text{if}\;i=k;\\ g_{i}^{t}+[\hat{b}_{ki;t}]_{+}g_{k}^{t}+(-\hat{b}_{ki;t})[g_{k}^{t}]_{+},&\text{if}\;i\neq k.\end{cases}$
    whenever ${t}$${t^{\prime}}$$\scriptstyle{k}$ in $\mathbb{T}_{n}$. We denote by $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ the set of tropical points associated to the $Y$-pattern $\mathcal{S}_{Y}$.
  • (ii)
    A tropical point ${\bf a}]=\{{\bf a}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ associated to the cluster pattern $\mathcal{S}_{X}$ is an assignment of a (column) vector ${\bf a}^{t}=(a_{1}^{t},\ldots,a_{m}^{t})^{T}$ in $\mathbb{Z}^{m}$ to each vertex $t$ of $\mathbb{T}_{n}$ such that
    (3.2) $a_{i}^{t^{\prime}}=\begin{cases}-a_{k}^{t}+\max\{\sum_{j=1}^{m}[b_{jk;t}]_{+}a_{j}^{t},\;\sum_{j=1}^{m}[-b_{jk;t}]_{+}a_{j}^{t}\},&\text{if}\;i=k;\\ a_{i}^{t},&\text{if}\;i\neq k.\end{cases}$
    whenever ${t}$${t^{\prime}}$$\scriptstyle{k}$ in $\mathbb{T}_{n}$. We denote by $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$ the set of tropical points associated to the cluster pattern $\mathcal{S}_{X}$.
Remark 3.4.
The relations in (3.1) and (3.2) are obtained from mutation relations (2.6) and (2.7) by tropicalization over the tropical semifield $\mathbb{Z}^{\rm max}=(\mathbb{Z},+,\max\{-,-\})$. Namely, we make the following replacements.
  • Replace the multiplication and addition in (2.6) and (2.7) by “+” and “max{–,–}” over $\mathbb{Z}$, respectively.
  • Replace “1” in (2.6) by “0”;
The following corollary is obvious.
Corollary 3.5.
The following statements hold.
  • (i)
    Each tropical point ${\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ corresponds to a unique semifield homomorphism $\beta\in\operatorname{\mathsf{Hom}}_{\rm sf}(\mathbb{F}_{>0},\mathbb{Z}^{\rm max})$ such that
    $\beta({\bf y}_{t})=(\beta(y_{1;t}),\ldots,\beta(y_{m;t}))=({\bf g}^{t})^{T}$
    for any vertex $t\in\mathbb{T}$, where $\mathbb{F}_{>0}=\mathbb{Q}_{\rm sf}(y_{1;t_{0}},\ldots,y_{m;t_{0}})$.
  • (ii)
    Each tropical point ${\bf a}]=\{{\bf a}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$ corresponds to a unique semifield homomorphism $\beta\in\operatorname{\mathsf{Hom}}_{\rm sf}(\mathbb{F}_{>0},\mathbb{Z}^{\rm max})$ such that
    $\beta({\bf x}_{t})=(\beta(x_{1;t}),\ldots,\beta(x_{m;t}))=({\bf a}^{t})^{T}$
    for any vertex $t\in\mathbb{T}$, where $\mathbb{F}_{>0}=\mathbb{Q}_{\rm sf}(x_{1;t_{0}},\ldots,x_{m;t_{0}})$.
3.2. Compatibly pointed elements and good elements
In this subsection, we fix a full rank upper cluster algebra $\mathcal{U}$, which means its initial mutation matrix $\widetilde{B}_{t_{0}}$ has the full rank $n$, equivalently, any mutation matrix $\widetilde{B}_{t}$ of $\mathcal{U}$ has the full rank $n$.
Since $\mathcal{U}$ is of full rank, each seed $({\bf x}_{t},\widetilde{B}_{t})$ of $\mathcal{U}$ defines a partial order $\preceq_{t}$ on $\mathbb{Z}^{m}$. For two vectors ${\bf g},{\bf g}^{\prime}\in\mathbb{Z}^{m}$, we write ${\bf g}^{\prime}\preceq_{t}{\bf g}$ if there exists some vector ${\bf v}=(v_{1},\ldots,v_{n})^{T}\in\mathbb{N}^{n}$ such that
${\bf g}^{\prime}={\bf g}+\widetilde{B}_{t}{\bf v},$
equivalently, ${\bf x}_{t}^{{\bf g}^{\prime}}={\bf x}_{t}^{\bf g}\cdot\widehat{\bf y}_{t}^{\bf v}$, where $\widehat{\bf y}_{t}^{\bf v}={\bf x}_{t}^{\widetilde{B}_{t}{\bf v}}$. We denote by ${\bf g}^{\prime}\prec_{t}{\bf g}$ for the case ${\bf g}^{\prime}\preceq_{t}{\bf g}$ and ${\bf g}^{\prime}\neq{\bf g}$.
The partial order $\preceq_{t}$ above is known as dominance order associated to seed $({\bf x}_{t},\widetilde{B}_{t})$ in [Qin_2017]. The dominance order plays an important role in study of the bases problem of (upper) cluster algebras.
For any seed $({\bf x}_{t},\widetilde{B}_{t})$ of $\mathcal{U}$, we denote by $R_{t}$ the collection of the formal Laurent series
$u=\sum_{{\bf h}\in\mathbb{Z}^{m}}b_{\bf h}{\bf x}_{t}^{\bf h},\;\;\;b_{\bf h}\in\mathbb{Z},$
such that the set $\{{\bf h}\in\mathbb{Z}^{m}\mid b_{\bf h}\neq 0\}$ has finitely many maximal elements with respect to the dominance order $\preceq_{t}$ on $\mathbb{Z}^{m}$. It is easy to see that the collection $R_{t}$ has a ring structure.
Now we recall the definition of pointed elements introduced in [Qin_2017]. Let
$u=\sum_{{\bf h}\in\mathbb{Z}^{m}}b_{\bf h}{\bf x}_{t}^{\bf h}\in R_{t}.$
We say that $u$ is pointed for the seed $({\bf x}_{t},\widetilde{B}_{t})$ if the following two conditions are satisfied.
  • The set $\{{\bf h}\in\mathbb{Z}^{m}\mid b_{\bf h}\neq 0\}$ has a unique maximal element $\bf g$} under the dominance order $\preceq_{t}$ on $\mathbb{Z}^{m}$. We denote by $\operatorname{\mathsf{deg}}^{t}(u)\coloneqq{\bf g}$ and call it the degree of $u$ with respect to the seed $({\bf x}_{t},\widetilde{B}_{t})$.
  • The coefficient corresponds to the degree term is $1$, i.e., $b_{\bf g}=1$ for ${\bf g}=\operatorname{\mathsf{deg}}^{t}(u)$.
By definition, each pointed element $u$ for the seed $({\bf x}_{t},\widetilde{B}_{t})$ takes the following form:
$u={\bf x}_{t}^{\bf g}+\sum_{{\bf h}\prec_{t}{\bf g}}b_{\bf h}{\bf x}_{t}^{\bf h}={\bf x}_{t}^{\bf g}F(\hat{y}_{1;t},\ldots,\hat{y}_{n;t}),$
where $b_{\bf h}\in\mathbb{Z}$ and $F\in\mathbb{Z}[[y_{1},\ldots,y_{n}]]$ is a polynomial series with constant term $1$.
In cluster algebras, we are often interested in the pointed elements which can be controlled by tropical points associated to a $Y$-pattern. Such elements correspond to compatibly pointed elements introduced in [qin_2019].
Definition 3.6 (Compatibly pointed elements and good elements).
Let $(\mathcal{S}_{X},\mathcal{S}_{Y})$ be a Langlands dual pair and $\mathcal{U}$ the upper cluster algebra associated to the cluster pattern $\mathcal{S}_{X}$. Assume that $\mathcal{U}$ is of full rank.
  • (i)
    An element $u\in\mathcal{U}$ is said to be compatibly pointed, if $u$ is pointed for any seed $({\bf x}_{t},\widetilde{B}_{t})$ of $\mathcal{U}$ and the collection ${\bf g}]\coloneqq\{\operatorname{\mathsf{deg}}^{t}(u)\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ forms a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$. In this case, we also say that $u$ is $\bf g$}]-pointed.
  • (ii)
    A $\bf g$}]-pointed element $u\in\mathcal{U}$ is said to be $\bf g$}]-good, if it is universally positive, i.e., $u\in\mathbb{Z}_{\geq 0}[x_{1;t}^{\pm 1},\ldots,x_{m;t}^{\pm 1}]$ for any vertex $t\in\mathbb{T}_{n}$.
Let $u$ be a $\bf g$}]-pointed element, then the Laurent expansion of $u$ with respect to any seed $({\bf x}_{t},\widetilde{B}_{t})$ of $\mathcal{U}$ has a canonical expression:
(3.3) $u={\bf x}_{t}^{{\bf g}_{u}^{t}}F_{u}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t}),$
where ${\bf g}_{u}^{t}=\operatorname{\mathsf{deg}}^{t}(u)$ is the value of $\bf g$}] at vertex $t$ and $F_{u}^{t}$ is a polynomial in $\mathbb{Z}[y_{1},\ldots,y_{n}]$ with constant term $1$.
  • The degree vector ${\bf g}_{u}^{t}=\operatorname{\mathsf{deg}}^{t}(u)$ is often called the extended $g$-vector of $u$ with respect to vertex $t$;
  • The polynomial $F_{u}^{t}$ is called the $F$-polynomial of $u$ with respect to vertex $t$;
  • For $k\in[1,n]$, we denote by $f_{k;u}^{t}$ the maximal degree of $y_{k}$ in $F_{u}^{t}$. The vector
    ${\bf f}_{u}^{t}\coloneqq(f_{1;u}^{t},\ldots,f_{n;u}^{t})^{T}\in\mathbb{N}^{n}$
    is called the $f$-vector of $u$ with respect to vertex $t$;
  • The $\bf g$}]-pointed element $u$ is said to be $\bf g$}]-bipointed, if for any vertex $t\in\mathbb{T}_{n}$ the monomial $y_{1}^{f_{1;u}^{t}}\cdots y_{n}^{f_{n;u}^{t}}$ appears in $F_{u}^{t}$ and it has coefficient $1$;
  • The $\bf g$}]-pointed element $u$ is said to be $\bf g$}]-bigood, if it is $\bf g$}]-bipointed and universally positive.
Recall that a cluster monomial in $\mathcal{U}$ is a Laurent monomial of the form
${\bf x}_{w}^{\bf v}\coloneqq\prod_{j=1}^{m}x_{j;w}^{v_{j}}$
for some cluster ${\bf x}_{w}$ of $\mathcal{U}$ and some vector ${\bf v}=(v_{1},\ldots,v_{m})^{T}\in\mathbb{Z}^{m}$ with $v_{i}\geq 0$ for any $i\in[1,n]$. Notice that we allow $v_{j}$ to be negative for $j\in[n+1,m]$.
Keep the setting as in Definition 3.6. Given a vertex $w\in\mathbb{T}_{n}$ and a vector ${\bf v}\in\mathbb{Z}^{m}$, we denote by $({\bf v},w)]$ the unique tropical point $\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ determined by the condition ${\bf g}^{w}={\bf v}$.
Proposition 3.7.
Let ${\bf x}_{w}^{\bf v}$ be a cluster monomial in $\mathcal{U}$. Then ${\bf x}_{w}^{\bf v}$ is $({\bf v},w)]-$bigood in $\mathcal{U}$.
Proof.
By the results in [GHKK18], we know that $u\coloneqq{\bf x}_{w}^{\bf v}$ is $({\bf v},w)]-$good in $\mathcal{U}$. Let ${\bf f}_{u}^{t}=(f_{1;u}^{t},\ldots,f_{n;u}^{t})^{T}$ be the $f$-vector of $u={\bf x}_{w}^{\bf v}$ with respect to vertex $t\in\mathbb{T}_{n}$. Since the $F$-polynomials of cluster variables have constant term $1$ and by [fomin_zelevinsky_2007, Proposition 5.3], we know that the monomial $y_{1}^{f_{1;u}^{t}}\cdots y_{n}^{f_{n;u}^{t}}$ appears in the $F$-polynomial $F_{u}^{t}$ with coefficient $1$. Hence, $u={\bf x}_{w}^{\bf v}$ is $({\bf v},w)]-$bigood in $\mathcal{U}$. ∎
Example 3.8.
Let us continue with Example 2.10. We see that $\widetilde{B}=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$ has full rank and $\hat{y}_{1}=x_{2}^{-1},\;\hat{y}_{2}=x_{1}$. The canonical expressions of the five cluster variables
$x_{1},\;x_{2},\;x_{3}=\frac{x_{2}+1}{x_{1}},\;x_{4}=\frac{x_{1}+x_{2}+1}{x_{1}x_{2}},\;x_{5}=\frac{x_{1}+1}{x_{2}}$
[…]
$x_{1}=x_{1}\cdot 1,\;\;\;x_{2}=x_{2}\cdot 1,\;\;\;x_{3}={x_{1}^{-1}x_{2}\cdot(1+\widehat{y}_{1}),}\;\;\;x_{4}={x_{1}^{-1}\cdot(1+\widehat{y}_{1}+\widehat{y}_{1}\widehat{y}_{2}),}\;\;\;x_{5}={x_{2}^{-1}\cdot(1+\widehat{y}_{2})}.$
3.3. Basic properties of pointed elements
As before, we fix a full rank upper cluster algebra $\mathcal{U}$. Let us recall some basic properties of pointed elements in $R_{t}$.
Lemma 3.9 ([qin_2019]*Lemmas 3.2.5, 3.2.6).
Suppose that $u,u^{\prime}\in R_{t}$ are two pointed elements for seed $({\bf x}_{t},\widetilde{B}_{t})$. Then the following statements hold.
  • (i)
    The product $uu^{\prime}\in R_{t}$ is pointed for seed $({\bf x}_{t},\widetilde{B}_{t})$ and we have
    $\operatorname{\mathsf{deg}}^{t}(uu^{\prime})=\operatorname{\mathsf{deg}}^{t}(u)+\operatorname{\mathsf{deg}}^{t}(u^{\prime}).$
  • (ii)
    $u$ has a multiplicative inverse $u^{-1}$ in $R_{t}$. Moreover, $u^{-1}$ is pointed for $({\bf x}_{t},\widetilde{B}_{t})$ and
    $\operatorname{\mathsf{deg}}^{t}(u^{-1})=-\operatorname{\mathsf{deg}}^{t}(u).$
Let $({\bf x}_{t},\widetilde{B}_{t})$ and $({\bf x}_{w},\widetilde{B}_{w})$ be two seeds of $\mathcal{U}$ at vertices $t,w\in\mathbb{T}_{n}$. By Proposition 3.7, we know that any cluster variable $x_{j;t}$ ($j\in[1,m]$) is pointed for $({\bf x}_{w},\widetilde{B}_{w})$. We denote by $\widetilde{G}_{t}^{w}$ the $m\times m$ matrix whose $j$th column is given by ${\bf g}_{x_{j;t}}^{w}=\operatorname{\mathsf{deg}}^{w}(x_{j;t})\in\mathbb{Z}^{m}$. Notice that $\widetilde{G}_{t}^{w}$ takes the form
$\widetilde{G}_{t}^{w}=\begin{bmatrix}G_{t}^{w}&0\\ \ast&I_{m-n}\end{bmatrix}$,
where $G_{t}^{w}$ is the $n\times n$ submatrix of $\widetilde{G}_{t}^{w}$. We call $\widetilde{G}_{t}^{w}$ the extended $G$-matrix and $G_{t}^{w}$ the $G$-matrix of vertex $t$ with respect to $w$. It is easy to see that if $t=w$, then $\widetilde{G}_{t}^{w}=I_{m}$.
Let $\overleftarrow{\mu}$ be the mutation sequence corresponding to the unique path from vertex $w$ to $t$ in $\mathbb{T}_{n}$. We apply the mutation sequence $\overleftarrow{\mu}$ to $\begin{pmatrix}B_{w}\\ I_{n}\end{pmatrix}$, where $B_{w}$ is the principal part of $\widetilde{B}_{w}$. The resulting matrix $\overleftarrow{\mu}\begin{pmatrix}B_{w}\\ I_{n}\end{pmatrix}$ takes the form $\begin{pmatrix}B_{t}\\ C_{t}^{w}\end{pmatrix}$. We call the $n\times n$ matrix $C_{t}^{w}$ the $C$-matrix of the vertex $t$ with respect to $w$.
Theorem 3.10 ([GHKK18], [NZ12, (3.11)]).
Let $t,w$ be two vertices in $\mathbb{T}_{n}$. The following statements hold.
  • (i)
    $\det(\widetilde{G}_{t}^{w})=\pm 1=\det(G_{t}^{w})$ and $\det(C_{t}^{w})=\pm 1$.
  • (ii)
    Each column of $C_{t}^{w}$ is either a non-negative vector or a non-positive vector.
  • (iii)
    $SC_{t}^{w}S^{-1}(G_{t}^{w})^{\rm T}=I_{n}$, where $S$ is any skew-symmetrizer of $B_{w}$.
Lemma 3.11 ([qin_2019]*Lem. 3.3.6, Prop. 3.3.10).
Let $\mathcal{U}$ be a full rank upper cluster algebra and $t,w$ two vertices of $\mathbb{T}_{n}$. The following statements hold.
  • (i)
    Any Laurent monomial ${\bf x}_{t}^{\bf h}$ is pointed for the seed $({\bf x}_{w},\widetilde{B}_{w})$ and we have
    $\operatorname{\mathsf{deg}}^{w}({\bf x}_{t}^{\bf h})=\widetilde{G}_{t}^{w}{\bf h}=\operatorname{\mathsf{deg}}^{w}({\bf x}_{w}^{\widetilde{G}_{t}^{w}{\bf h}}).$
  • (ii)
    Let $\widehat{\bf y}_{t}$ and $\widehat{\bf y}_{w}$ be the collections of $\hat{y}-$variables at vertices $t$ and $w$. Then for any ${\bf v}\in\mathbb{Z}^{n}$, the Laurent monomial $\widehat{\bf y}_{t}^{\bf v}={\bf x}_{t}^{\widetilde{B}_{t}{\bf v}}$ is pointed for the seed $({\bf x}_{w},\widetilde{B}_{w})$ and
    $\operatorname{\mathsf{deg}}^{w}(\widehat{\bf y}_{t}^{\bf v})=\widetilde{G}_{t}^{w}\widetilde{B}_{t}{\bf v}=\widetilde{B}_{w}C_{t}^{w}{\bf v}=\operatorname{\mathsf{deg}}^{w}(\widehat{\bf y}_{w}^{C_{t}^{w}{\bf v}}).$
Notice that any element in $R_{w}$ can be written as a finite $\mathbb{Z}$-linear combinations of pointed elements for seed $({\bf x}_{w},\widetilde{B}_{w})$. By the above lemma, any Laurent monomial ${\bf x}_{t}^{\bf h}$ in $\mathcal{L}(t)\coloneqq\mathbb{Z}[x_{1;t}^{\pm 1},\ldots,x_{m;t}^{\pm 1}]$ is pointed for the seed $({\bf x}_{w},\widetilde{B}_{w})$. Hence, $\mathcal{L}(t)$ is contained in $R_{w}$.
Corollary 3.12.
Let $\mathcal{U}$ be a full rank upper cluster algebra and $t,w$ two vertices of $\mathbb{T}_{n}$. Let
$u=\sum\limits_{{\bf h}\in\mathbb{Z}^{m}}b_{\bf h}{\bf x}_{t}^{\bf h}$
be a Laurent polynomial in $\mathcal{L}(t)=\mathbb{Z}[x_{1;t}^{\pm 1},\ldots,x_{m;t}^{\pm 1}]$. Then the following statements hold.
  • (i)
    $u$ is pointed for the seed $({\bf x}_{w},\widetilde{B}_{w})$ if and only if there exists a (unique) ${\bf g}\in\mathbb{Z}^{m}$ with $b_{\bf g}=1$ such that $\operatorname{\mathsf{deg}}^{w}({\bf x}_{t}^{\bf h})\preceq_{w}\operatorname{\mathsf{deg}}^{w}({\bf x}_{t}^{\bf g})$ for any ${\bf h}\in\mathbb{Z}^{m}$ with $b_{\bf h}\neq 0$, equivalently,
    $\widetilde{G}_{t}^{w}{\bf h}\preceq_{w}\widetilde{G}_{t}^{w}{\bf g}$
    for any ${\bf h}\in\mathbb{Z}^{m}$ with $b_{\bf h}\neq 0$.
  • (ii)
    Suppose that $u$ is pointed for the seed $({\bf x}_{t},\widetilde{B}_{t})$, that is, $u$ can be written as
    $u={\bf x}_{t}^{\bf g}\cdot\sum\limits_{{\bf v}\in\mathbb{N}^{n}}c_{\bf v}\widehat{\bf y}_{t}^{\bf v},$
    where $c_{\bf v}\in\mathbb{Z}$ and $c_{0}=1$. Then $u$ is pointed for the new seed $({\bf x}_{w},\widetilde{B}_{w})$ if and only if there exists a (unique) ${\bf u}\in\mathbb{N}^{n}$ with $c_{\bf u}=1$ such that $\operatorname{\mathsf{deg}}^{w}(\widehat{\bf y}_{t}^{\bf v})\preceq_{w}\operatorname{\mathsf{deg}}^{w}(\widehat{\bf y}_{t}^{\bf u})$ for any ${\bf v}\in\mathbb{N}^{n}$ with $c_{\bf v}\neq 0$, equivalently,
    $C_{t}^{w}{\bf v}-C_{t}^{w}{\bf u}=C_{t}^{w}({\bf v}-{\bf u})\in\mathbb{N}^{n}$
    for any ${\bf v}\in\mathbb{N}^{n}$ with $c_{\bf v}\neq 0$.
Proof.
Thanks to Theorem 3.10 (i) and by Lemma 3.11, we know that $\operatorname{\mathsf{deg}}^{w}({\bf x}_{t}^{\bf h})=\operatorname{\mathsf{deg}}^{w}({\bf x}_{t}^{{\bf h}^{\prime}})$ if and only if ${\bf h}={\bf h}^{\prime}$. By the similar argument and since $\widetilde{B}_{w}$ has full rank, we know that $\operatorname{\mathsf{deg}}^{w}(\widehat{\bf y}_{t}^{\bf v})=\operatorname{\mathsf{deg}}^{w}(\widehat{\bf y}_{t}^{{\bf v}^{\prime}})$ if and only if ${\bf v}={\bf v}^{\prime}$.
Based on the above observations, the results in (i) and (ii) follow from the definition of pointed elements and Lemma 3.11. ∎
Let $(\mathcal{S}_{X},\mathcal{S}_{Y})$ be a Langlands dual pair and $\mathcal{U}$ the upper cluster algebra associated to the cluster pattern $\mathcal{S}_{X}$. As before, we assume that $\mathcal{U}$ is of full rank.
Corollary 3.13.
Keep the above setting. Let ${\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ be a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$, and write ${\bf g}^{t}=(g_{1}^{t},\ldots,g_{m}^{t})^{T}$ for $t\in\mathbb{T}_{n}$. Let ${w}$${s}$$\scriptstyle{k}$ be an edge in $\mathbb{T}_{n}$. The following statements hold.
  • (i)
    $\operatorname{\mathsf{deg}}^{w}({\bf x}_{s}^{{\bf g}^{s}}\cdot\widehat{\bf y}_{s}^{[-g_{k}^{s}]_{+}{\bf e}_{k}})={\bf g}^{w}$, where ${\bf e}_{k}$ is the $k$th column of $I_{n}$.
  • (ii)
    Suppose that $u\in\mathcal{U}$ is a $\bf g$}]-pointed element. Let $u={\bf x}_{s}^{{\bf g}^{s}}\cdot\sum\limits_{{\bf v}\in\mathbb{N}^{n}}c_{\bf v}\widehat{\bf y}_{s}^{\bf v}$ be the Laurent expansion of $u$ with respect to seed $({\bf x}_{s},\widetilde{B}_{s})$. Then the coefficient of $\widehat{\bf y}_{s}^{[-g_{k}^{s}]_{+}{\bf e}_{k}}$ in $\sum\limits_{{\bf v}\in\mathbb{N}^{n}}c_{\bf v}\widehat{\bf y}_{s}^{\bf v}$ is $1$ and for any integer $d$ with $d>[-g_{k}^{s}]_{+}$, the coefficient of $\widehat{\bf y}_{s}^{d{\bf e}_{k}}$ in $\sum\limits_{{\bf v}\in\mathbb{N}^{n}}c_{\bf v}\widehat{\bf y}_{s}^{\bf v}$ is zero.
Proof.
(i) Let $\widetilde{G}_{s}^{w}$ and $C_{s}^{w}$ be the extended $G$-matrix and $C$-matrix of vertex $s$ with respect to $w$. By Lemma 3.9 and Lemma 3.11, we have
$\operatorname{\mathsf{deg}}^{w}({\bf x}_{s}^{{\bf g}^{s}}\cdot\widehat{\bf y}_{s}^{[-g_{k}^{s}]_{+}{\bf e}_{k}})=\widetilde{G}_{s}^{w}{\bf g}^{s}+\widetilde{B}_{w}C_{s}^{w}\cdot([-g_{k}^{s}]_{+}{\bf e}_{k}).$
Since the two vertices $w$ and $s$ are jointed in $\mathbb{T}_{n}$ by an edge labeled by $k$, we know that $\widetilde{G}_{s}^{w}=(g_{ij;s})$ and $C_{s}^{w}=(c_{ij;s})$ are given as follows:
$g_{ij;s}$ $=\begin{cases}-1,&i=j=k;\\ [-b_{ik;w}]_{+}=[b_{ik;s}]_{+},&j=k,\;i\neq k;\\ \delta_{i,j},&\text{else},\end{cases}$
$c_{ij;s}$ $=\begin{cases}-1,&i=j=k;\\ [b_{kj;w}]_{+}=[-b_{kj;s}]_{+},&i=k,\;j\neq k;\\ \delta_{i,j},&\text{else}.\end{cases}$
Because the $k$th column of $C_{s}^{w}$ equals ${\bf e}_{k}$, we have
$\widetilde{G}_{s}^{w}{\bf g}^{s}+\widetilde{B}_{w}C_{s}^{w}\cdot([-g_{k}^{s}]_{+}{\bf e}_{k})=\widetilde{G}_{s}^{w}{\bf g}^{s}-\widetilde{B}_{w}\cdot([-g_{k}^{s}]_{+}{\bf e}_{k}).$
Now we look at the $i$th component $g_{i}^{\prime}$ of the vector
${\bf g}^{\prime}\coloneqq\widetilde{G}_{s}^{w}{\bf g}^{s}-\widetilde{B}_{w}\cdot([-g_{k}^{s}]_{+}{\bf e}_{k}).$
We have
$g_{i}^{\prime}=\begin{cases}-g_{k}^{s}-0=-g_{k}^{s},&i=k;\\ (g_{i}^{s}+[b_{ik;s}]_{+}g_{k}^{s})-b_{ik;w}[-g_{k}^{s}]_{+}=g_{i}^{s}+[b_{ik;s}]_{+}g_{k}^{s}+b_{ik;s}[-g_{k}^{s}]_{+},&i\neq k.\end{cases}$
Since $\widehat{B}_{s}=-\widetilde{B}_{s}^{T}$, we can rewrite the above equality as follows:
$g_{i}^{\prime}=\begin{cases}-g_{k}^{s},&i=k;\\ g_{i}^{s}+[-\hat{b}_{ki;s}]_{+}g_{k}^{s}-\hat{b}_{ki;s}[-g_{k}^{s}]_{+},&i\neq k.\end{cases}$
Using the equality $b]_{+}g-[-b]_{+}g=bg=b[g]_{+}-b[-g]_{+}$ for any $b,g\in\mathbb{R}$, one can see the above equality is the same as
$g_{i}^{\prime}=\begin{cases}-g_{k}^{s},&i=k;\\ g_{i}^{s}+[\hat{b}_{ki;s}]_{+}g_{k}^{s}-\hat{b}_{ki;s}[g_{k}^{s}]_{+},&i\neq k.\end{cases}$
Then by (3.1), we get ${\bf g}^{w}={\bf g}^{\prime}=\widetilde{G}_{s}^{w}{\bf g}^{s}-\widetilde{B}_{w}\cdot([-g_{k}^{s}]_{+}{\bf e}_{k})$. Hence, $\operatorname{\mathsf{deg}}^{w}({\bf x}_{s}^{{\bf g}^{s}}\cdot\widehat{\bf y}_{s}^{[-g_{k}^{s}]_{+}{\bf e}_{k}})={\bf g}^{w}$.
(ii) Since $u$ is $\bf g$}]-pointed, we know that it is pointed for the seed $({\bf x}_{w},\widetilde{B}_{w})$. By Corollary 3.12 (ii), there exists a unique ${\bf u}\in\mathbb{N}^{n}$ with $c_{\bf u}=1$ such that
$\operatorname{\mathsf{deg}}^{w}({\bf x}_{s}^{{\bf g}^{s}}\cdot\widehat{\bf y}_{s}^{\bf v})\preceq_{w}\operatorname{\mathsf{deg}}^{w}({\bf x}_{s}^{{\bf g}^{s}}\cdot\widehat{\bf y}_{s}^{\bf u})=\operatorname{\mathsf{deg}}^{w}(u)={\bf g}^{w}$
for any ${\bf v}\in\mathbb{N}^{n}$ with $c_{\bf v}\neq 0$. By the result in (i), we know that ${\bf u}=[-g_{k}^{s}]_{+}{\bf e}_{k}$.
Notice that $\operatorname{\mathsf{deg}}^{w}({\bf x}_{s}^{{\bf g}^{s}}\cdot\widehat{\bf y}_{s}^{\bf v})\preceq_{w}\operatorname{\mathsf{deg}}^{w}({\bf x}_{s}^{{\bf g}^{s}}\cdot\widehat{\bf y}_{s}^{\bf u})$ holds if and only if the following relation holds.
$C_{s}^{w}{\bf v}-C_{s}^{w}{\bf u}=C_{s}^{w}({\bf v}-{\bf u})\in\mathbb{N}^{n}.$
Since the $k$th column of $C_{s}^{w}$ is ${\bf e}_{k}$, we obtain that $C_{s}^{w}(d{\bf e}_{k}-[-g_{k}^{s}]_{+}{\bf e}_{k})$ belongs to $\mathbb{N}^{n}$ if and only if $d\leq[-g_{k}^{s}]_{+}$. Hence, for any integer $d$ with $d>[-g_{k}^{s}]_{+}$, the coefficient of $\widehat{\bf y}_{s}^{d{\bf e}_{k}}$ in $\sum\limits_{{\bf v}\in\mathbb{N}^{n}}c_{\bf v}\widehat{\bf y}_{s}^{\bf v}$ is zero. ∎
Corollary 3.14.
Let $u$ be a $\bf g$}]-pointed element in $\mathcal{U}$ and let ${\bf f}_{u}^{t}=(f_{1;u}^{t},\ldots,f_{n;u}^{t})^{T}$ be the $f$-vector of $u$ with respect to vertex $t\in\mathbb{T}_{n}$. The following statements hold.
  • (i)
    If $f_{k;u}^{t}=0$ for some $k\in[1,n]$, then the $k$th component $g_{k}^{t}$ of ${\bf g}_{u}^{t}=\operatorname{\mathsf{deg}}^{t}(u)$ is non-negative;
  • (ii)
    If $f_{i;u}^{t}=0$ for any $i\in[1,n]$, then $u$ is a cluster monomial in ${\bf x}_{t}$;
  • (iii)
[…]
Proof.
Since $u$ is $\bf g$}]-pointed in $\mathcal{U}$, the Laurent expansion of $u$ with respect to ${\bf x}_{t}$ has a canonical expression:
$u={\bf x}_{t}^{{\bf g}_{u}^{t}}F_{u}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t}),$
where ${\bf g}_{u}^{t}=(g_{1}^{t},\ldots,g_{m}^{t})^{T}$ is the value of $\bf g$}] at vertex $t$ and $F_{u}^{t}$ is a polynomial in $\mathbb{Z}[y_{1},\ldots,y_{n}]$ with constant term $1$.
(i) By Corollary 3.13 (ii), we know that $y_{k}^{[-g_{k}^{t}]_{+}}$ must be a monomial appearing in $F_{u}^{t}\in\mathbb{Z}[y_{1},\ldots,y_{n}]$. Recall that $f_{k;u}^{t}$ is the maximal degree of $y_{k}$ in $F_{u}^{t}$. So we have $f_{k;u}^{t}\geq[-g_{k}^{t}]_{+}\geq 0$. Hence, if $f_{k;u}^{t}=0$, then $g_{k}^{t}]_{+}=0$. So $g_{k}^{t}$ is non-negative.
(ii) Since $f_{i;u}^{t}=0$ for any $i\in[1,n]$, we get $F_{u}^{t}=1$ and $u={\bf x}_{t}^{{\bf g}_{u}^{t}}$. On the other hand, by (i), we have $g_{i}^{t}\geq 0$ for any $i\in[1,n]$. Hence, $u={\bf x}_{t}^{{\bf g}_{u}^{t}}$ is a cluster monomial in ${\bf x}_{t}$.
(iii) Since $f_{i;u}^{t}=0$ holds for any $i\in[1,n]\setminus\{k\}$, we get that $F_{u}^{t}$ is a polynomial in the single variable $y_{k}$. In this case, Corollary 3.13 (ii) implies $f_{k;u}^{t}=[-g_{k}^{t}]_{+}$. Since $f_{k;u}^{t}\neq 0$, we have $g_{k}^{t}<0$ and thus $f_{k;u}^{t}=-g_{k}^{t}$. Thus $u$ has the following form:
$u=x_{1;t}^{g_{1}^{t}}\cdots x_{m;t}^{g_{m}^{t}}(1+\sum_{0<d<-g_{k}^{t}}c_{d}\hat{y}_{k;t}^{d}+\hat{y}_{k;t}^{-g_{k}^{t}}),$
where $c_{d}\in\mathbb{Z}$. By the exchange relations, we have $x_{j;t^{\prime}}=x_{j;t}$ for any $j\neq k$ and $x_{k;t}\cdot x_{k;t^{\prime}}=M\cdot(1+\hat{y}_{k;t})$, where $M=\prod_{j=1}^{m}x_{j;t}^{[-b_{jk;t}]_{+}}$. By substituting these relations into the above equality and using the fact $u\in\mathbb{Z}[x_{1;t^{\prime}}^{\pm 1},\ldots,x_{m;t^{\prime}}^{\pm 1}]$, we obtain that
$(1+\sum_{0<d<-g_{k}^{t}}c_{d}\hat{y}_{k;t}^{d}+\hat{y}_{k;t}^{-g_{k}^{t}})=(1+\hat{y}_{k;t})^{-g_{k}^{t}}$
and the Laurent expansion of $u$ with respect to seed $({\bf x}_{t^{\prime}},\widetilde{B}_{t^{\prime}})$ is a Laurent monomial in $\mathbb{Z}[x_{1;t^{\prime}}^{\pm 1},\ldots,x_{m;t^{\prime}}^{\pm 1}]$. Thus the $F$-polynomial $F_{u}^{t^{\prime}}$ of $u$ with respect to vertex $t^{\prime}$ is $1$. In particular, the $f$-vector ${\bf f}_{u}^{t^{\prime}}$ of $u$ with respect to vertex $t^{\prime}$ is zero. By (ii), we know that $u$ is a cluster monomial in ${\bf x}_{t^{\prime}}$. ∎
4. Tropical invariant and $F$-invariant in cluster algebras
4.1. Semifield homomorphisms arising from good elements
Let $(\mathcal{S}_{X},\mathcal{S}_{Y},\mathsf{\Lambda})$ be a Langlands-Poisson triple. In this subsection, we give a map from the tropical points in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ to those in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$. Thanks to this map, for each good element in a $\mathsf{\Lambda}$-upper cluster algebra, we introduce a semifield homomorphism, which will be used to define tropical invariant and $F$-invariant.
Lemma 4.1.
Let $(\mathcal{S}_{X},\mathcal{S}_{Y})$ be a Langlands dual pair and $\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ a collection of vectors in $\mathbb{Z}^{m}$ indexed by the vertices in $\mathbb{T}_{n}$. Then the collection $\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ is a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ if and only if for any edge ${t}$${t^{\prime}}$$\scriptstyle{k}$ in $\mathbb{T}_{n}$, one has ${\bf g}^{t^{\prime}}=E_{k,\varepsilon(g_{k}^{t})}^{\widetilde{B}_{t}}{\bf g}^{t}$, where
$\varepsilon(g_{k}^{t})=\begin{cases}-,&\text{if }g_{k}^{t}\geq 0;\\ +,&\text{if }g_{k}^{t}<0,\end{cases}$
and $E_{k,\pm}^{\widetilde{B}_{t}}$ is as defined in (2.2).
Proof.
By definition, the collection $\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ is a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ if and only if the relations in (3.1) hold. Since $(\mathcal{S}_{X},\mathcal{S}_{Y})$ is a Langlands dual pair, we have $\widehat{B}_{t}=-\widetilde{B}_{t}^{T}$. Thus $\hat{b}_{ij;t}=-b_{ji;t}$ for any $i\in[1,n]$ and $j\in[1,m]$. Then the relations in (3.1) can be rewritten as
$g_{i}^{t^{\prime}}$ $=\begin{cases}-g_{k}^{t},&\text{if}\;i=k;\\ g_{i}^{t}+[-b_{ik;t}]_{+}g_{k}^{t}+b_{ik;t}[g_{k}^{t}]_{+},&\text{if}\;i\neq k.\end{cases}$
$=\begin{cases}-g_{k}^{t},&\text{if}\;i=k;\\ g_{i}^{t}+[b_{ik;t}]_{+}g_{k}^{t},&\text{if}\;i\neq k\;\;\text{and }\;g_{k}^{t}\geq 0;\\ g_{i}^{t}+[-b_{ik;t}]_{+}g_{k}^{t},&\text{if}\;i\neq k\;\;\text{and }\;g_{k}^{t}<0.\end{cases}$
This is exactly the same with ${\bf g}^{t^{\prime}}=E_{k,\varepsilon(g_{k}^{t})}^{\widetilde{B}_{t}}{\bf g}^{t}$. ∎
Lemma 4.2.
Let $\mathcal{S}_{X}$ be a cluster pattern of rank $n$ in $\mathbb{F}$ and $\{{\bf a}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ a collection of vectors in $\mathbb{Z}^{m}$ indexed by the vertices of $\mathbb{T}_{n}$. Then the collection $\{{\bf a}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ is a tropical point in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$ if and only if for any edge ${t}$${t^{\prime}}$$\scriptstyle{k}$ in $\mathbb{T}_{n}$, one has ${\bf a}^{t^{\prime}}=(E_{k,\varepsilon_{0}}^{\widetilde{B}_{t}})^{T}{\bf a}^{t}$, where
$\varepsilon_{0}=\begin{cases}-,&\text{if }\sum_{j=1}^{m}b_{jk;t}a_{j}^{t}\geq 0;\\ +,&\text{if }\sum_{j=1}^{m}b_{jk;t}a_{j}^{t}<0,\end{cases}$
and $E_{k,\pm}^{\widetilde{B}_{t}}$ is as defined in (2.2).
Proof.
By definition, the collection $\{{\bf a}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ is a tropical point in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$ if and only if the relations in (3.2) hold. The relations in (3.2) can be rewritten as follows:
$a_{i}^{t^{\prime}}=\begin{cases}-a_{k}^{t}+\sum_{j=1}^{m}[b_{jk;t}]_{+}a_{j}^{t},&\text{if}\;i=k\;\;\text{and }\;\sum_{j=1}^{m}b_{jk;t}a_{j}^{t}\geq 0;\\ -a_{k}^{t}+\sum_{j=1}^{m}[-b_{jk;t}]_{+}a_{j}^{t},&\text{if}\;i=k\;\;\text{and }\;\sum_{j=1}^{m}b_{jk;t}a_{j}^{t}<0;\\ a_{i}^{t},&\text{if}\;i\neq k.\end{cases}$
This is exactly the same with ${\bf a}^{t^{\prime}}=(E_{k,\varepsilon_{0}}^{\widetilde{B}_{t}})^{T}{\bf a}^{t}$. ∎
Proposition 4.3.
Let $(\mathcal{S}_{X},\mathcal{S}_{Y},\mathsf{\Lambda})$ be a Langlands-Poisson triple. Then
${\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}\mapsto\{\Lambda_{t}{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$
defines a map from the tropical points in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ to those in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$.
Proof.
Denote by ${\bf a}^{t}\coloneqq\Lambda_{t}{\bf g}^{t}=(a_{1}^{t},\ldots,a_{m}^{t})^{T}$ for any vertex $t\in\mathbb{T}_{n}$. We will use Lemma 4.2 to show the collection $\{{\bf a}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ is a tropical point in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$.
Let ${t}$${t^{\prime}}$$\scriptstyle{k}$ be an edge in $\mathbb{T}_{n}$. Since ${\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ is a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ and by Lemma 4.1, we have ${\bf g}^{t^{\prime}}=E_{k,\varepsilon(g_{k}^{t})}^{\widetilde{B}_{t}}{\bf g}^{t}$, where
$\varepsilon(g_{k}^{t})=\begin{cases}-,&\text{if }g_{k}^{t}\geq 0;\\ +,&\text{if }g_{k}^{t}<0.\end{cases}$
By Proposition 2.3 (i), we have $\Lambda_{t^{\prime}}=(E_{k,\varepsilon(g_{k}^{t})}^{\widetilde{B}_{t}})^{T}\Lambda_{t}E_{k,\varepsilon(g_{k}^{t})}^{\widetilde{B}_{t}}$. Hence,
[…]
Now we prove $\varepsilon(g_{k}^{t})=\varepsilon_{0}$, where $\varepsilon_{0}$ is as given in Lemma 4.2. By the definitions of $\varepsilon(g_{k}^{t})$ and $\varepsilon_{0}$, it suffices to show that $g_{k}^{t}\geq 0$ if and only if $\sum_{j=1}^{m}b_{jk;t}a_{j}^{t}\geq 0$. Since $(\widetilde{B}_{t},\Lambda_{t})$ is a compatible pair, we know that $\widetilde{B}_{t}^{T}\Lambda_{t}$ has the form
$\widetilde{B}_{t}^{T}\Lambda_{t}=(S\mid{\bf 0}),$
where $S=diag(s_{1},\ldots,s_{n})$ is a diagonal matrix with $s_{i}>0$, $i\in[1,n]$ and $\bf 0$} is the $n\times(m-n)$ zero matrix. Thus we have $\widetilde{B}_{t}^{T}{\bf a}^{t}=\widetilde{B}_{t}^{T}\Lambda_{t}{\bf g}^{t}=(S\mid{\bf 0}){\bf g}^{t}$. By comparing the $k$th component on both sides of the equality, we get
$\sum_{j=1}^{m}b_{jk;t}a_{j}^{t}=s_{k}g_{k}^{t}.$
Since $s_{k}$ is a positive integer, we get that $g_{k}^{t}\geq 0$ if and only if $\sum_{j=1}^{m}b_{jk;t}a_{j}^{t}\geq 0$. So we have $\varepsilon(g_{k}^{t})=\varepsilon_{0}$.
Hence, ${\bf a}^{t^{\prime}}=(E_{k,\varepsilon(g_{k}^{t})}^{\widetilde{B}_{t}})^{T}{\bf a}^{t}=(E_{k,\varepsilon_{0}}^{\widetilde{B}_{t}})^{T}{\bf a}^{t}$. By Lemma 4.2, we know the collection
$\{{\bf a}^{t}=\Lambda_{t}{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$
is a tropical point in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$. This completes the proof. ∎
Proposition 4.4.
Let $(\mathcal{S}_{X},\mathcal{S}_{Y},\mathsf{\Lambda})$ be a Langlands-Poisson triple and let $\mathcal{U}$ be the corresponding $\mathsf{\Lambda}$-upper cluster algebra. Then any good element $u\in\mathcal{U}$ defines a unique semifield homomorphism
$\beta_{u}:\;\mathbb{F}_{>0}:=\mathbb{Q}_{\rm sf}(x_{1;t_{0}},\ldots,x_{m;t_{0}})\rightarrow\mathbb{Z}^{\rm max}$
such that $\beta_{u}({\bf x}_{t})=(\Lambda_{t}{\bf g}_{u}^{t})^{T}$ for any vertex $t\in\mathbb{T}_{n}$, where ${\bf g}_{u}^{t}\in\mathbb{Z}^{m}$ is the extended $g$-vector of $u$ with respect to vertex $t$.
Proof.
Since $u$ is a good element, we know that the collection
${\bf g}_{u}]\coloneqq\{{\bf g}_{u}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$
is a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$. Then by Proposition 4.3, the collection
${\bf a}]\coloneqq\{{\bf a}^{t}=\Lambda_{t}{\bf g}_{u}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$
is a tropical point in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$. By Corollary 3.5, the tropical point ${\bf a}]\in\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$ corresponds to a unique semifield homomorphism
$\beta_{u}\colon\mathbb{F}_{>0}\coloneqq\mathbb{Q}_{\rm sf}(x_{1;t_{0}},\ldots,x_{m;t_{0}})\rightarrow\mathbb{Z}^{\rm max}$
such that $\beta_{u}({\bf x}_{t})=({\bf a}^{t})^{T}=(\Lambda_{t}{\bf g}_{u}^{t})^{T}$ for any vertex $t\in\mathbb{T}_{n}$. ∎
Remark 4.5.
We remark that the above proof still works if we replace the good elements by compatibly pointed elements. But in this paper, we only use good elements.
4.2. Tropical invariant and $F$-invariant
In this subsection, we fix a Langlands-Poisson triple $(\mathcal{S}_{X},\mathcal{S}_{Y},\mathsf{\Lambda})$, and let $\mathcal{U}$ be the corresponding $\mathsf{\Lambda}$-upper cluster algebra. Denote by $S=diag(s_{1},\ldots,s_{n})$ the type of the compatible pair $(\widetilde{B}_{t_{0}},\Lambda_{t_{0}})$, namely, $\widetilde{B}_{t_{0}}^{T}\Lambda_{t_{0}}=(S\mid{\bf 0})$, where $\bf 0$} is the $n\times(m-n)$ zero matrix. By Proposition 2.2, the mutation matrices of $\mathcal{U}$ always have full rank.
We denote by $\mathcal{U}^{[{\bf g}]}$ the set of $\bf g$}]-good elements in $\mathcal{U}$ and by
$\mathcal{U}^{\operatorname{\mathsf{good}}}\coloneqq\bigcup_{[{\bf g}]\in\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})}\mathcal{U}^{[\bf g]}$
the set of all good elements in $\mathcal{U}$. Notice that the union $\bigcup_{[{\bf g}]\in\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})}\mathcal{U}^{[\bf g]}$ is disjoint.
Definition 4.6 (Tropical invariant and $F$-invariant).
Let $\mathcal{U}$ be a $\mathsf{\Lambda}$-upper cluster algebra.
(i) For $u,u^{\prime}\in\mathcal{U}^{\operatorname{\mathsf{good}}}$, the tropical invariant $\langle u,u^{\prime}\rangle_{\rm trop}$ of the pair $(u,u^{\prime})$ is defined by
$\langle u,u^{\prime}\rangle_{\rm trop}\coloneqq\beta_{u^{\prime}}(u)\;\in\mathbb{Z},$
where $\beta_{u^{\prime}}\colon\mathbb{F}_{>0}\coloneqq\mathbb{Q}_{\rm sf}(x_{1;t_{0}},\ldots,x_{m;t_{0}})\rightarrow\mathbb{Z}^{\rm max}$ is the semifield homomorphism defined in Proposition 4.4. The pairing $\langle-,-\rangle_{\rm trop}\colon\mathcal{U}^{\operatorname{\mathsf{good}}}\times\mathcal{U}^{\operatorname{\mathsf{good}}}\rightarrow\mathbb{Z}$ is called the tropical pairing on the $\mathsf{\Lambda}$-upper cluster algebra $\mathcal{U}$.
(ii) For $u,u^{\prime}\in\mathcal{U}^{\operatorname{\mathsf{good}}}$, the $F$-invariant $(u\mid\mid u^{\prime})_{F}$ of the pair $(u,u^{\prime})$ is defined by
$(u\mid\mid u^{\prime})_{F}\coloneqq\beta_{u^{\prime}}(u)+\beta_{u}(u^{\prime})=\langle u,u^{\prime}\rangle_{\rm trop}+\langle u^{\prime},u\rangle_{\rm trop}.$
In the case $(u\mid\mid u^{\prime})_{F}=0$, we say that $u$ and $u^{\prime}$ are $F$-compatible.
Remark 4.7.
Notice that in general $\langle u,u^{\prime}\rangle_{\rm trop}\neq\langle u^{\prime},u\rangle_{\rm trop}$, but we always have $(u\mid\mid u^{\prime})_{F}=(u^{\prime}\mid\mid u)_{F}$, by definition.
Given a non-zero polynomial
$F=\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}{\bf y}^{{\bf v}}\in\mathbb{Z}[y_{1},\ldots,y_{n}]$
and a vector ${\bf r}\in\mathbb{Z}^{n}$, we denote by
$F[{\bf r}]\coloneqq\max\{{\bf v}^{T}{\bf r}\mid c_{{\bf v}}\neq 0\}\in\mathbb{Z}.$
We call the map $F[-]\colon\mathbb{Z}^{n}\rightarrow\mathbb{Z}$ a tropical polynomial.
Example 4.8.
Take $F=1+y_{1}+y_{1}y_{2}\in\mathbb{Z}[y_{1},y_{2}]$ and ${\bf r}=\begin{bmatrix}-2\\ 1\end{bmatrix}$. Then
$F[{\bf r}]=\max\{\begin{bmatrix}0,0\end{bmatrix}\begin{bmatrix}-2\\ 1\end{bmatrix},\begin{bmatrix}1,0\end{bmatrix}\begin{bmatrix}-2\\ 1\end{bmatrix},\begin{bmatrix}1,1\end{bmatrix}\begin{bmatrix}-2\\ 1\end{bmatrix}\}=\max\{0,-2,-1\}=0.$
Proposition 4.9.
The following statements hold.
  • (i)
    If $F\in\mathbb{Z}[y_{1},\ldots,y_{n}]$ has constant term $1$, then $F[{\bf r}]\geq 0$ for any ${\bf r}\in\mathbb{Z}^{n}$.
  • (ii)
    If $0\neq F_{1},F_{2}\in\mathbb{Z}_{\geq 0}[y_{1},\ldots,y_{n}]$, then $(F_{1}F_{2})[{\bf r}]=F_{1}[{\bf r}]+F_{2}[{\bf r}]$ for any ${\bf r}\in\mathbb{Z}^{n}$.
Proof.
(i) This is clear. (ii) Fix ${\bf r}=(r_{1},\ldots,r_{n})^{T}\in\mathbb{Z}^{n}$ and let
$\pi_{\bf r}\colon\mathbb{Q}_{\rm sf}(y_{1},\ldots,y_{n})\rightarrow\mathbb{Z}^{\max}$
be the semifield homomorphism induced by $y_{i}\mapsto r_{i}$. For any $0\neq F\in\mathbb{Z}_{\geq 0}[y_{1},\ldots,y_{n}]$, we have $F\in\mathbb{Q}_{\rm sf}(y_{1},\ldots,y_{n})$ and $\pi_{\bf r}(F)=F[{\bf r}]$. Since $0\neq F_{1},F_{2}\in\mathbb{Z}_{\geq 0}[y_{1},\ldots,y_{n}]$, we have $\pi_{\bf r}(F_{1}F_{2})=\pi_{\bf r}(F_{1})+\pi_{\bf r}(F_{2})$, i.e., $(F_{1}F_{2})[{\bf r}]=F_{1}[{\bf r}]+F_{2}[{\bf r}]$. ∎
The following result shows that for each $\mathsf{\Lambda}$-seed $({\bf x}_{w},\widetilde{B}_{w},\Lambda_{w})$ of $\mathcal{U}$, we have an explicit formula to calculate the tropical invariant and $F$-invariant.
Theorem 4.10.
Let $u$ and $u^{\prime}$ be two good elements in a $\mathsf{\Lambda}$-upper cluster algebra $\mathcal{U}$. Let
$u={\bf x}_{w}^{{\bf g}_{u}^{w}}F_{u}^{w}(\hat{y}_{1;w},\ldots,\hat{y}_{n;w}),\;\;\;u^{\prime}={\bf x}_{w}^{{\bf g}_{u^{\prime}}^{w}}F_{u^{\prime}}^{w}(\hat{y}_{1;w},\ldots,\hat{y}_{n;w})$
be the canonical expressions of $u$ and $u^{\prime}$ with respect to any vertex $w\in\mathbb{T}_{n}$. Then we have
  • (i)
    $\langle u,u^{\prime}\rangle_{\rm trop}=({\bf g}_{u}^{w})^{T}\Lambda_{w}{\bf g}_{u^{\prime}}^{w}+F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}]$.
  • (ii)
    $(u\mid\mid u^{\prime})_{F}=F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}]+F_{u^{\prime}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]$. In particular, $(u\mid\mid u^{\prime})_{F}\in\mathbb{Z}_{\geq 0}$.
  • (iii)
    $(u\mid\mid u)_{F}=0$ whenever $u$ is a cluster monomial.
Proof.
(i) By definition, $\langle u,u^{\prime}\rangle_{\rm trop}=\beta_{u^{\prime}}(u)$, where $\beta_{u^{\prime}}$ is the unique semifield homomorphism
$\beta_{u^{\prime}}\colon\mathbb{F}_{>0}\coloneqq\mathbb{Q}_{\rm sf}(x_{1;t_{0}},\ldots,x_{m;t_{0}})\rightarrow\mathbb{Z}^{\rm max}$
such that $\beta_{u^{\prime}}({\bf x}_{t})=(\Lambda_{t}{\bf g}_{u^{\prime}}^{t})^{T}$ for any vertex $t\in\mathbb{T}_{n}$. By applying the semifield homomorphism $\beta_{u^{\prime}}$ to $u={\bf x}_{w}^{{\bf g}_{u}^{w}}F_{u}^{w}(\widehat{\bf y}_{w})\in\mathbb{F}_{>0}$, we have
$\beta_{u^{\prime}}(u)=\beta_{u^{\prime}}({\bf x}_{w}^{{\bf g}_{u}^{w}})+\beta_{u^{\prime}}(F_{u}^{w}(\widehat{\bf y}_{w})).$
Since $\beta_{u^{\prime}}({\bf x}_{w})=(\Lambda_{w}{\bf g}_{u^{\prime}}^{w})^{T}$, we have
$\beta_{u^{\prime}}({\bf x}_{w}^{\bf h})=(\Lambda_{w}{\bf g}_{u^{\prime}}^{w})^{T}{\bf h}=[(\Lambda_{w}{\bf g}_{u^{\prime}}^{w})^{T}{\bf h}]^{T}={\bf h}^{T}\Lambda_{w}{\bf g}_{u^{\prime}}^{w},\quad\forall{\bf h}\in\mathbb{Z}^{m}.$
We write $F_{u}^{w}({\bf y})=\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}{\bf y}^{{\bf v}}\in\mathbb{N}[y_{1},\ldots,y_{n}]$. Then $F_{u}^{w}(\widehat{\bf y}_{w})=\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}{\bf x}_{w}^{\widetilde{B}_{w}{\bf v}}$. Thus
$\beta_{u^{\prime}}(F_{u}^{w}(\widehat{\bf y}_{w}))$ $=\beta_{u^{\prime}}(\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}{\bf x}_{w}^{\widetilde{B}_{w}{\bf v}})=\max\{{\bf v}^{T}\widetilde{B}_{w}^{T}\Lambda_{w}{\bf g}_{u^{\prime}}^{w}\mid c_{\bf v}\neq 0\}$
$=\max\{{\bf v}^{T}(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}\mid c_{\bf v}\neq 0\}=F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}].$
Hence, we have
$\langle u,u^{\prime}\rangle_{\rm trop}=\beta_{u^{\prime}}(u)=\beta_{u^{\prime}}({\bf x}_{w}^{{\bf g}_{u}^{w}})+\beta_{u^{\prime}}(F_{u}^{w}(\widehat{\bf y}_{w}))=({\bf g}_{u}^{w})^{T}\Lambda_{w}{\bf g}_{u^{\prime}}^{w}+F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}].$
(ii) By (i) and the definition of $F$-invariant, we have
$(u\mid\mid u^{\prime})_{F}$ $=\langle u,u^{\prime}\rangle_{\rm trop}+\langle u^{\prime},u\rangle_{\rm trop}$
$=({\bf g}_{u}^{w})^{T}\Lambda_{w}{\bf g}_{u^{\prime}}^{w}+F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}]+({\bf g}_{u^{\prime}}^{w})^{T}\Lambda_{w}{\bf g}_{u}^{w}+F_{u^{\prime}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}].$
Since $\Lambda_{w}$ is skew-symmetric, we know that $({\bf g}_{u}^{w})^{T}\Lambda_{w}{\bf g}_{u^{\prime}}^{w}+({\bf g}_{u^{\prime}}^{w})^{T}\Lambda_{w}{\bf g}_{u}^{w}=0$. Thus
$(u\mid\mid u^{\prime})_{F}=F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}]+F_{u^{\prime}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}].$
Since both $F_{u}^{w}$ and $F_{u^{\prime}}^{w}$ are polynomials with constant term $1$, we have
$F_{u}^{w}[{\bf r}],\;F_{u^{\prime}}^{w}[{\bf r}]\;\in\mathbb{Z}_{\geq 0}$
for any ${\bf r}\in\mathbb{Z}^{n}$. So $(u\mid\mid u^{\prime})_{F}=F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}]+F_{u^{\prime}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]\;\in\mathbb{Z}_{\geq 0}$.
(iii) Say $u$ is a cluster monomial in seed $({{\bf x}_{w},\widetilde{B}_{w}})$. Then $F_{u}^{w}=1$. Then by (ii), we see that $(u\mid\mid u)_{F}=0$. ∎
Since the $F$-invariant $(u\mid\mid u^{\prime})_{F}$ can be written as
$(u\mid\mid u^{\prime})_{F}=F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}]+F_{u^{\prime}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}],$
we call the non-negative integer $F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}]$ the partial $F$-invariant of $(u,u^{\prime})$ at vertex $w\in\mathbb{T}_{n}$.
Remark 4.11.
Notice that $(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}=S({\bf g}_{u^{\prime}}^{w})^{\circ}$, where $({\bf g}_{u^{\prime}}^{w})^{\circ}\in\mathbb{Z}^{n}$ is the principal part of ${\bf g}_{u}^{w}\in\mathbb{Z}^{m}$. This suggests that the notion of partial $F$-invariant at vertex $w\in\mathbb{T}_{n}$ and the $F$-invariant in some sense only depend on the principal part of the mutation matrices (i.e., the exchange matrices) of $\mathcal{A}$ and a choice of the skew-symmetrizer $S$.
Remark 4.12.
Recently, Fei introduced and studied the tropical $F$-polynomials in [fei_2019a, fei_2019b] from the viewpoint of representation theory of finite-dimensional algebras and [fei_2019b, Theorem 3.22] suggests that the partial $F$-invariant coincides with the partial $E$-invariant for Jacobi-finite quivers with potentials.
Proposition 4.13.
Let $x_{i;t}$ be a cluster variable of $\mathcal{U}$, where $i\in[1,m]$ and $t\in\mathbb{T}_{n}$. Then for any good element $u\in\mathcal{U}$, we have
$(x_{i;t}\mid\mid u)_{F}=\begin{cases}s_{i}f_{i;u}^{t},&i\in[1,n];\\ 0,&i\in[n+1,m],\end{cases}$
where $s_{i}$ is the $(i,i)$-entry of the diagonal matrix $S=diag(s_{1},\ldots,s_{n})$ and $f_{i;u}^{t}$ is the $i$th component of the $f$-vector of $u$ with respect vertex $t$.
Proof.
We know that $F_{x_{i;t}}^{t}=1$ and ${\bf g}_{x_{i;t}}^{t}={\bf e}_{i}$, where ${\bf e}_{i}$ is the $i$th column of $I_{m}$. Now we use the vertex $t$ to calculate $(x_{i;t}\mid\mid u)_{F}$. By using the formula in Theorem 4.10 (ii), we have
$(x_{i;t}\mid\mid u)_{F}=0+F_{u}^{t}[(S\mid{\bf 0}){\bf e}_{i}].$
Then the result follows. ∎
Remark 4.14.
(i) The $F$-invariant $(-\mid\mid-)_{F}\colon\mathcal{U}^{\operatorname{\mathsf{good}}}\times\mathcal{U}^{\operatorname{\mathsf{good}}}\rightarrow\mathbb{Z}_{\geq 0}$ is a great improvement of the $f$-compatibility degree introduced by Fu and Gyoda in [Fu-Gyoda] and thus the compatibility degree introduced by Fomin and Zelevinsky in [fz-2003y]. The $f$-compatibility degree, defined using the components of $f$-vectors of cluster variables, is a function
$(-\mid\mid-)_{f}\colon\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{Z}_{\geq 0},$
where $\mathcal{X}$ is the set of unfrozen cluster variables in $\mathcal{U}$. By Proposition 4.13, when restricted to $\mathcal{X}$, the $F$-invariant and the $f$-compatibility degree are related as follows:
$(x_{i;t}\mid\mid x_{j;w})_{F}=s_{i}(x_{i;t}\mid\mid x_{j;w})_{f}.$
(ii) By construction, $F$-invariant is symmetric, i.e., $(u\mid\mid u^{\prime})_{F}=(u^{\prime}\mid\mid u)_{F}$ for any two good elements in $\mathcal{U}$. Applying to the case $u=x_{i;t}$ and $u=x_{j;w}$ with $i,j\in[1,n]$, we get
$s_{i}(x_{i;t}\mid\mid x_{j;w})_{f}=(x_{i;t}\mid\mid x_{j;w})_{F}=(x_{j;w}\mid\mid x_{i;t})_{F}=s_{j}(x_{j;w}\mid\mid x_{i;t})_{f}.$
This recovers the result in [Fu-Gyoda, Proposition 4.14 (2)].
Corollary 4.15.
Let ${\bf x}_{t}$ be a cluster of $\mathcal{U}$ and $u\in\mathcal{U}^{\operatorname{\mathsf{good}}}$. The following statements hold.
  • (i)
    If $(x_{k;t}\mid\mid u)_{F}=0$ for some $k\in[1,n]$, then the $k$th component $g_{k}^{t}$ of ${\bf g}_{u}^{t}=\operatorname{\mathsf{deg}}^{t}(u)$ is non-negative.
  • (ii)
    If $(x_{i;t}\mid\mid u)_{F}=0$ holds for any $i\in[1,n]$, then $u$ is a cluster monomial in ${\bf x}_{t}$.
  • (iii)
[…]
Proof.
The results follow from Proposition 4.13 and Corollary 3.14. ∎
Proposition 4.16.
Let $u$ be a $\bf g$}]-bigood element in $\mathcal{U}$, and let ${\bf x}_{t}^{\bf h}=\prod_{j=1}^{m}x_{j;t}^{h_{j}}$ be a cluster monomial of $\mathcal{U}$, i.e., ${\bf h}^{\circ}\coloneqq(h_{1},\ldots,h_{n})^{T}\in\mathbb{Z}_{\geq 0}^{n}$. Then
  • (i)
    $({\bf x}_{t}^{\bf h}\mid\mid u)_{F}=\sum_{j=1}^{n}h_{j}\cdot(x_{j;t}\mid\mid u)_{F}$.
  • (ii)
    $F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{{\bf x}_{t}^{\bf h}}^{w}]=\sum_{j=1}^{n}h_{j}\cdot F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{x_{j;t}}^{w}]$ for any vertex $w\in\mathbb{T}_{n}$.
Proof.
(i) Consider the canonical expression of $u$ with respect to the vertex $t$,
$u={\bf x}_{t}^{{\bf g}_{u}^{t}}F_{u}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t}),$
where ${\bf g}_{u}^{t}$ is the value of $\bf g$}] at vertex $t$. Since $u$ is a $\bf g$}]-bigood element, $F_{u}^{t}$ has the form:
$F_{u}^{t}=1+\text{middle terms }+y_{1}^{f_{1}^{t}}\cdots y_{n}^{f_{n}^{t}},$
where ${\bf f}_{u}^{t}=(f_{1}^{t},\ldots,f_{n}^{t})^{T}$ is the $f$-vector of $u$ with respect to vertex $t$. Since ${\bf x}_{t}^{\bf h}$ is a cluster momomial, we know that $F_{{\bf x}_{t}^{\bf h}}^{t}=1$ and $(h_{1},\ldots,h_{n})^{T}\in\mathbb{Z}_{\geq 0}^{n}$. By using the cluster ${\bf x}_{t}$ to calculate $({\bf x}_{t}^{\bf h}\mid\mid u)_{F}$, we have
$({\bf x}_{t}^{\bf h}\mid\mid u)_{F}$ $=0+F_{u}^{t}[(S\mid{\bf 0}){\bf h}]=F_{u}^{t}[(s_{1}h_{1},\ldots,s_{n}h_{n})^{T}]$
$=f_{1}^{t}s_{1}h_{1}+\ldots+f_{n}^{t}s_{n}h_{n}=\sum_{j=1}^{n}h_{j}\cdot(x_{j;t}\mid\mid u)_{F}.$
(ii) Let $w$ be a vertex in $\mathbb{T}_{n}$. We have
[…]
$=$ $\sum_{j=1}^{n}h_{j}\cdot\left(F_{x_{j;t}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]+F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{x_{j;t}}^{w}]\right)$
$=$ $\sum_{j=1}^{n}h_{j}\cdot F_{x_{j;t}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]+\sum_{j=1}^{n}h_{j}\cdot F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{x_{j;t}}^{w}]$
$=$ $F_{{\bf x}_{t}^{\bf h}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]+\sum_{j=1}^{n}h_{j}\cdot F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{x_{j;t}}^{w}].$
The last equality comes from the facts that $F_{{\bf x}_{t}^{\bf h}}^{w}=\prod_{j=1}^{m}(F_{x_{j;t}}^{w})^{h_{j}}=\prod_{j=1}^{n}(F_{x_{j;t}}^{w})^{h_{j}}$ and that for any two polynomials $F_{1},F_{2}\in\mathbb{Z}_{\geq 0}[y_{1},\ldots,y_{n}]$ and any vector ${\bf r}\in\mathbb{Z}^{n}$, we have
$(F_{1}F_{2})[{\bf r}]=F_{1}[{\bf r}]+F_{2}[{\bf r}],$
by Proposition 4.9 (ii).
On the other hand, we know that
$({\bf x}_{t}^{\bf h}\mid\mid u)_{F}=F_{{\bf x}_{t}^{\bf h}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]+F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{{\bf x}_{t}^{\bf h}}^{w}].$
By comparing the two equalities, we obtain $F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{{\bf x}_{t}^{\bf h}}^{w}]=\sum_{j=1}^{n}h_{j}\cdot F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{x_{j;t}}^{w}]$. ∎
Lemma 4.17.
Let $x$ and $z$ be two unfrozen cluster variables of $\mathcal{U}$. If $(x\mid\mid z)_{F}=0$, then they are contained in the same cluster.
Proof.
Since $(x\mid\mid z)_{F}=0$ and by Remark 4.14, the $f$-compatibility degree $(x\mid\mid z)_{f}$ is zero. Then the result follows from [Fu-Gyoda, Theorem 3.3].
Let us give another proof here. Consider the tropical point
${\bf g}_{z}]\coloneqq\{{\bf g}_{z}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}\in\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$
corresponding to the cluster variable $z$. Since $(x\mid\mid z)_{F}=0$ and by Corollary 4.15 (i), we know that the $k$th component $g_{k;z}^{t}$ of ${\bf g}_{z}^{t}$ satisfies $g_{k;z}^{t}\geq 0$ whenever $x=x_{k;t}$ for a vertex $t$.
On the other hand, by [cao-li-2020, Theorem 8], we know that there exists a seed $({\bf x}_{w},\widetilde{B}_{w})$ containing $x$ (say $x=x_{k;w}$) such that $g_{i;z}^{w}\geq 0$ for any $i\in[1,n]$ with $x_{i;t}\neq x$, i.e., for any $i\in[1,n]\setminus\{k\}$. Thus for the vertex $w$, we have $g_{j;z}^{w}\geq 0$ for $j=1,\ldots,n$, i.e., the principal part $({\bf g}_{z}^{t})^{\circ}$ of ${\bf g}_{z}^{t}$ belongs to $\mathbb{Z}_{\geq 0}^{n}$. Thus ${\bf x}_{w}^{{\bf g}_{z}^{w}}$ is a cluster monomial of $\mathcal{U}$. Hence, the two cluster monomials ${\bf x}_{w}^{{\bf g}_{z}^{w}}$ and $z$ have the same extended $g$-vector ${\bf g}_{z}^{w}$ with respect to vertex $w$. It is well-known that different cluster monomials have different extended $g$-vectors, cf. [GHKK18]. So we must have $z={\bf x}_{w}^{{\bf g}_{z}^{w}}$. This implies that $z$ is a cluster variable in ${\bf x}_{w}$. In particular, $z$ and $x=x_{k;w}$ are contained in the same cluster ${\bf x}_{w}$. ∎
Lemma 4.18.
Let $\{z_{1},\ldots,z_{p}\}$ be a subset of unfrozen cluster variables of $\mathcal{U}$. If $(z_{i}\mid\mid z_{j})_{F}=0$ for any $i,j\in[1,p]$, then $\{z_{1},\ldots,z_{p}\}$ is a subset of some cluster of $\mathcal{U}$.
Proof.
By Lemma 4.17, we know that any two cluster variables in $\{z_{1},\ldots,z_{p}\}$ are contained in the same cluster. Then the result follows from [cao-li-2020, Theorem 13]. ∎
The following result says that $F$-invariant is a very concrete and easily computable numerical invariant to distinguish whether the product of two cluster monomials is still a cluster monomial.
Theorem 4.19.
Let $u$ and $u^{\prime}$ be two cluster monomials in $\mathcal{U}$. Then the product $uu^{\prime}$ is still a cluster monomial in $\mathcal{U}$ if and only if $(u\mid\mid u^{\prime})_{F}=0$, i.e.,
$F_{u}^{t}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{t}]+F_{u^{\prime}}^{t}[(S\mid{\bf 0}){\bf g}_{u}^{t}]=0,$
for some (equivalently, any) vertex $t\in\mathbb{T}_{n}$.
Proof.
$\Longrightarrow$”: Suppose that the product $uu^{\prime}$ is a cluster monomial in $\mathcal{U}$. Say, $uu^{\prime}$ is a cluster monomial in cluster ${\bf x}_{t}$. Then both $u$ and $u^{\prime}$ can be viewed as cluster monomials in ${\bf x}_{t}$. Thus $F_{u}^{t}=1=F_{u^{\prime}}^{t}$. So we have $(u\mid\mid u^{\prime})_{F}=F_{u}^{t}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{t}]+F_{u^{\prime}}^{t}[(S\mid{\bf 0}){\bf g}_{u}^{t}]=0$.
[…]
By Proposition 3.7 and Proposition 4.16, we have
$(u\mid\mid u^{\prime})_{F}=\sum_{i=1}^{n}\sum_{j=1}^{n}h_{i}r_{j}(x_{i;s}\mid\mid x_{j;s^{\prime}})_{F}.$
Since all the numbers $h_{i},r_{j}$ and $(x_{i;s}\mid\mid x_{j;s^{\prime}})_{F}$ are non-negative for $i,j\in[1,n]$, the assumption $(u\mid\mid u^{\prime})_{F}=0$ implies $h_{i}r_{j}(x_{i;s}\mid\mid x_{j;s^{\prime}})_{F}=0$ for any $i,j\in[1,n]$. Hence, $(x_{i;s}\mid\mid x_{j;s^{\prime}})_{F}=0$ for any $i,j\in[1,n]$ with $h_{i}r_{j}\neq 0$.
Let $V=\{x_{i;s}\mid i\in[1,n],h_{i}\neq 0\}$ and $V^{\prime}=\{x_{j;s^{\prime}}\mid j\in[1,n],r_{j}\neq 0\}$. We have $(z\mid\mid z^{\prime})_{F}=0$ for any $z\in V$ and $z^{\prime}\in V^{\prime}$. Since $V$ and $V^{\prime}$ are subsets of clusters of $\mathcal{U}$, we have
$(z_{1}\mid\mid z_{2})_{F}=0\;\;\;\text{and}\;\;\;(z_{1}^{\prime}\mid\mid z_{2}^{\prime})_{F}=0$
for any $z_{1},z_{2}\in V$ and $z_{1}^{\prime},z_{2}^{\prime}\in V^{\prime}$.
Hence, $(z\mid\mid z^{\prime})_{F}=0$ holds for any two cluster variables $z,z^{\prime}$ in $V\cup V^{\prime}$. Then by Lemma 4.18, we know that $V\cup V^{\prime}$ is a subset of some cluster ${\bf x}_{t}$ of $\mathcal{U}$. Hence, the product $uu^{\prime}$ is a cluster monomial in ${\bf x}_{t}$. ∎
Example 4.20.
[…]
$x_{3}={x_{1}^{-1}x_{2}\cdot(1+\widehat{y}_{1}),}\;\;\;x_{4}={x_{1}^{-1}\cdot(1+\widehat{y}_{1}+\widehat{y}_{1}\widehat{y}_{2}),}\;\;\;x_{5}={x_{2}^{-1}\cdot(1+\widehat{y}_{2})}.$
Thus we have
$(x_{3}\mid\mid x_{4})_{F}$ $=F_{x_{3}}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{x_{4}}^{t_{0}}]+F_{x_{4}}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{x_{3}}^{t_{0}}]$
$=(1+y_{1})\begin{bmatrix}-1\\ 0\end{bmatrix}+(1+y_{1}+y_{1}y_{2})\begin{bmatrix}-1\\ 1\end{bmatrix}$
$=\max\{0,\;-1\}+\max\{0,\;-1,\;0\}$
$=0,$
$(x_{3}\mid\mid x_{5})_{F}$ $=(1+y_{1})\begin{bmatrix}0\\ -1\end{bmatrix}+(1+y_{2})\begin{bmatrix}-1\\ 1\end{bmatrix}$
$=\max\{0,\;0\}+\max\{0,\;1\}$
$=1.$
Then by Theorem 4.19, we know that $x_{3}x_{4}$ is a cluster monomial, while $x_{3}x_{5}$ is not.
Recall that the Newton polytope of a polynomial $F=\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}{\bf y}^{{\bf v}}\in\mathbb{Z}[y_{1},\ldots,y_{n}]$ is the convex hull of $\{{\bf v}\in\mathbb{N}^{n}\mid c_{\bf v}\neq 0\}$ in $\mathbb{R}^{n}$.
Corollary 4.21.
Let $u$ and $u^{\prime}$ be two cluster monomials in $\mathcal{U}$ and $t$ a vertex of $\mathbb{T}_{n}$. If the two $F$-polynomials $F_{u}^{t}$ and $F_{u^{\prime}}^{t}$ have the same Newton polytope, then the product $uu^{\prime}$ is a cluster monomial of $\mathcal{U}$.
Proof.
Since $u$ and $u^{\prime}$ are cluster monomials and by Theorem 4.10 (iii), we have
$(u\mid\mid u)_{F}=2F_{u}^{t}[(S\mid{\bf 0}){\bf g}_{u}^{t}]=0\;\;\;\text{and}\;\;\;(u^{\prime}\mid\mid u^{\prime})_{F}=2F_{u^{\prime}}^{t}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{t}]=0.$
We get $F_{u}^{t}[(S\mid{\bf 0}){\bf g}_{u}^{t}]=0=F_{u^{\prime}}^{t}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{t}]$. Since $F_{u}^{t}$ and $F_{u^{\prime}}^{t}$ have the same Newton polytope, we have $F_{u}^{t}[{\bf h}]=F_{u^{\prime}}^{t}[{\bf h}]$ for any ${\bf h}\in\mathbb{Z}^{n}$. Hence, we have
$(u\mid\mid u^{\prime})_{F}$ $=F_{u}^{t}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{t}]+F_{u^{\prime}}^{t}[(S\mid{\bf 0}){\bf g}_{u}^{t}]$
$=F_{u^{\prime}}^{t}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{t}]+F_{u}^{t}[(S\mid{\bf 0}){\bf g}_{u}^{t}]=0.$
Then by Theorem 4.19, we know that the product $uu^{\prime}$ is a cluster monomial. ∎
Definition 4.22.
Two tropical points ${\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ and ${\bf g}^{\prime}]=\{({\bf g}^{\prime})^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ are said to be sign-coherent, if for any vertex $t\in\mathbb{T}_{n}$ and any $k\in[1,n]$, we have $g_{k}^{t}\cdot(g_{k}^{\prime})^{t}\geq 0$, where $g_{k}^{t}$ and $(g_{k}^{\prime})^{t}$ are the $k$th components of ${\bf g}^{t}$ and $({\bf g}^{\prime})^{t}$.
Lemma 4.23.
Let ${\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ and ${\bf g}^{\prime}]=\{({\bf g}^{\prime})^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ be two tropical points in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$. If $\bf g$}] and $\bf g$$^{\prime}$] are sign-coherent, then the collection $\{{\bf g}^{t}+({\bf g}^{\prime})^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ is still a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$. We denote it by ${\bf g}\uplus{\bf g}^{\prime}]$.
Proof.
It suffices to show that the collection $\{{\bf g}^{t}+({\bf g}^{\prime})^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ satisfies the relations in (3.1). Thanks to the sign-coherence assumption, this is clear. ∎
Theorem 4.24.
Let $u$ be a $\bf g$}]-good element in $\mathcal{U}$ and $u^{\prime}$ a $\bf g$$^{\prime}$]-good element in $\mathcal{U}$. Suppose that $(u\mid\mid u^{\prime})_{F}=0$. Then the following statements hold.
  • (i)
    The two tropical points $\bf g$}] and $\bf g$$^{\prime}$] are sign-coherent. In particular, the tropical point ${\bf g}\uplus{\bf g}^{\prime}]$ can be defined.
  • (ii)
    The product $uu^{\prime}$ is a ${\bf g}\uplus{\bf g}^{\prime}]-$good element in $\mathcal{U}$.
Proof.
(i) We write ${\bf g}]=\{{\bf g}_{u}^{t}=(g_{1;u}^{t},\ldots,g_{m;u}^{t})^{T}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ and ${\bf g}^{\prime}]=\{{\bf g}_{u^{\prime}}^{t}=(g_{1;u^{\prime}}^{t},\ldots,g_{m;u^{\prime}}^{t})^{T}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$. Assume the contrary that the two tropical points $\bf g$}] and $\bf g$$^{\prime}$] are not sign-coherent to deduce a contradiction. Then there exist some vertex $s\in\mathbb{T}_{n}$ and some $k\in[1,n]$ such that $g_{k;u}^{s}\cdot g_{k;u^{\prime}}^{s}<0$. Without loss of generality, we can assume $g_{k;u}^{s}<0$ and $g_{k;u^{\prime}}^{s}>0$.
By Corollary 3.13, we know that the monomial ${\bf y}^{[-g_{k;u}^{s}]_{+}{\bf e}_{k}}=y_{k}^{[-g_{k;u}^{s}]_{+}}$ appears in the $F$-polynomial $F_{u}^{s}\in\mathbb{Z}_{\geq 0}[y_{1},\ldots,y_{n}]$ with coefficient $1$, where ${\bf e}_{k}$ is the $k$th column of $I_{n}$. Thus we have
$F_{u}^{s}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{s}]\geq([-g_{k;u}^{s}]_{+}{\bf e}_{k})^{T}(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{s}=[-g_{k;u}^{s}]_{+}\cdot s_{k}g_{k;u^{\prime}}^{s}.$
Then by $g_{k;u}^{s}<0$, $g_{k;u^{\prime}}^{s}>0$ and $s_{k}>0$, we get $F_{u}^{s}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{s}]>0$.
By using the cluster ${\bf x}_{s}$ to calculate the $F$-invariant $(u\mid\mid u^{\prime})_{F}$, we have
$(u\mid\mid u^{\prime})_{F}=F_{u}^{s}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{s}]+F_{u^{\prime}}^{s}[(S\mid{\bf 0}){\bf g}_{u}^{s}]\geq F_{u}^{s}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{s}]>0,$
which contradicts $(u\mid\mid u^{\prime})_{F}=0$. Hence, the two tropical points $\bf g$}] and $\bf g$$^{\prime}$] must be sign-coherent.
(ii) For any vertex $t$, consider the Laurent expansions of $u$ and $u^{\prime}$ with respect to vertex $t$:
$u={\bf x}_{t}^{{\bf g}_{u}^{t}}F_{u}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t}),\;\;\;u^{\prime}={\bf x}_{t}^{{\bf g}_{u^{\prime}}^{t}}F_{u^{\prime}}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t}),$
where $F_{u}^{t}$ and $F_{u^{\prime}}^{t}$ are polynomials in $\mathbb{Z}_{\geq 0}[y_{1},\ldots,y_{n}]$ with constant term $1$. Hence, we have
$uu^{\prime}={\bf x}_{t}^{{\bf g}_{u}^{t}+{\bf g}_{u^{\prime}}^{t}}H_{uu^{\prime}}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t}),$
where $H_{uu^{\prime}}^{t}\coloneqq F_{u}^{t}F_{u^{\prime}}^{t}$ is a polynomial in $\mathbb{Z}_{\geq 0}[y_{1},\ldots,y_{n}]$ with constant term $1$.
By (i) and Lemma 4.23, we know that ${\bf g}\uplus{\bf g}^{\prime}]\coloneqq\{{\bf g}_{u}^{t}+{\bf g}_{u^{\prime}}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ forms a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$. Then by the definition of good elements, we know that the product $uu^{\prime}$ is a ${\bf g}\uplus{\bf g}^{\prime}]-$good element in $\mathcal{U}$. ∎
4.3. Log-canonical cluster variables are compatible
In this subsection, we show that if two cluster variables are log-canonical in a $\mathsf{\Lambda}$-upper cluster algebra, then they are contained in the same cluster.
As before, we fix a Langlands-Poisson triple $(\mathcal{S}_{X},\mathcal{S}_{Y},\mathsf{\Lambda})$, and let $\mathcal{U}$ be the corresponding $\mathsf{\Lambda}$-upper cluster algebra. Denote by $S=diag(s_{1},\ldots,s_{n})$ the type of the compatible pair $(\widetilde{B}_{t_{0}},\Lambda_{t_{0}})$, namely, $\widetilde{B}_{t_{0}}^{T}\Lambda_{t_{0}}=(S\mid{\bf 0})$, where $\bf 0$} is the $n\times(m-n)$ zero matrix.
Recall that we can define a Poisson bracket $\{-,-\}$ on the ambient field $\mathbb{F}=\mathbb{Q}(x_{1;t_{0}},\ldots,x_{m;t_{0}})$ using the $m\times m$ skew-symmetric matrix $\Lambda_{t_{0}}=(\lambda_{ij;t_{0}})$ at the rooted vertex $t_{0}$:
$\{x_{i;t_{0}},x_{j;t_{0}}\}\coloneqq\lambda_{ij;t_{0}}\cdot x_{i;t_{0}}x_{j;t_{0}}.$
This Poisson bracket is compatible with the cluster pattern $\mathcal{S}_{X}$, that is, for any cluster ${\bf x}_{t}$ of $\mathcal{S}_{X}$, we have
$\{x_{i;t},x_{j;t}\}=\lambda_{ij;t}\cdot x_{i;t}x_{j;t},$
where $\lambda_{ij;t}$ is the $(i,j)$-entry of $\Lambda_{t}$, cf. [gsv-2003], [bz-2005, Remark 4.6].
Two elements $f_{1},f_{2}$ in $\mathbb{F}$ are said to be log-canonical with respect to the Poisson bracket $\{-,-\}$, if
$\{f_{1},f_{2}\}=cf_{1}f_{2}$
for some $c\in\mathbb{Q}$.
Clearly, any two cluster variables in the same cluster are log-canonical. In Theorem 4.26 (ii), we show the converse statement is also true.
Lemma 4.25.
Let $w$ be a vertex of $\mathbb{T}_{n}$ and $u$ an element in $\mathbb{F}$ with the form
$u={\bf x}_{w}^{\bf g}F(\hat{y}_{1;w},\ldots,\hat{y}_{n;w}),$
where $F=\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}{\bf y}^{{\bf v}}$ is a polynomial in $\mathbb{Z}[y_{1},\ldots,y_{n}]$ with constant term $1$. If $u$ and $x_{k;w}$ are log-canonical for some $k\in[1,n]$, then $f_{k}=0$, where $f_{k}$ is the maximal degree of $y_{k}$ in $F$.
Proof.
We know that $\Lambda_{w}=(\lambda_{ij;w})$ is the coefficient matrix of $\{-,-\}$ with respect to the cluster ${\bf x}_{w}$. Since $(\widetilde{B}_{w},\Lambda_{w})$ is a compatible pair of type $S$, we have $\widetilde{B}_{w}^{T}\Lambda_{w}=(S\mid{\bf 0})$. For $i\in[1,n]$ and $j\in[1,m]$, we denote by ${\bf e}_{i}$ the $i$th column of $I_{n}$ and $\widetilde{\bf e}_{j}$ the $j$th column of $I_{m}$. We have
$\{\hat{y}_{i;w},x_{j;w}\}$ $=\{{\bf x}_{w}^{\widetilde{B}_{w}{\bf e}_{i}},{\bf x}_{w}^{\widetilde{\bf e}_{j}}\}=\left((\widetilde{B}_{w}{\bf e}_{i})^{T}\Lambda_{w}\widetilde{\bf e}_{j}\right)\hat{y}_{i;w}x_{j;w}$
$=\left({\bf e}_{i}^{T}(\widetilde{B}_{w})^{T}\Lambda_{w}\widetilde{\bf e}_{j}\right)\hat{y}_{i;w}x_{j;w}=\left({\bf e}_{i}^{T}(S\mid{\bf 0})\widetilde{\bf e}_{j}\right)\hat{y}_{i;w}x_{j;w}$
$=\begin{cases}s_{j}\cdot\hat{y}_{j;w}x_{j;w},&i=j;\\ 0,&i\neq j.\end{cases}$
Now we consider the integer $k\in[1,n]$ such that $u$ and $x_{k;w}$ are log-canonical. For ${\bf v}=(v_{1},\ldots,v_{n})^{T}\in\mathbb{N}^{n}$, we have
$\{\widehat{\bf y}_{w}^{\bf v},x_{k;w}\}=s_{k}(v_{k}\widehat{\bf y}_{w}^{\bf v})x_{k;w}.$
For simplicity, we write $F(\widehat{\bf y}_{w})=F(\hat{y}_{1;w},\ldots,\hat{y}_{n;w})$. We have
$\{F(\widehat{\bf y}_{w}),x_{k;w}\}=\{\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}\widehat{\bf y}_{w}^{{\bf v}},x_{k;w}\}=s_{k}(\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}v_{k}\widehat{\bf y}_{w}^{\bf v})x_{k;w}.$
Now let us calculate $\{u,x_{k;w}\}=\{{\bf x}_{w}^{\bf g}F(\widehat{\bf y}_{w}),x_{k;w}\}$.
$\{{\bf x}_{w}^{\bf g}F(\widehat{\bf y}_{w}),x_{k;w}\}$ $=\{{\bf x}_{w}^{\bf g},x_{k;w}\}F(\widehat{\bf y}_{w})+{\bf x}_{w}^{\bf g}\{F(\widehat{\bf y}_{w}),x_{k;w}\}$
$=({\bf g}^{T}\Lambda_{w}\widetilde{\bf e}_{k})\cdot{\bf x}_{w}^{\bf g}\cdot x_{k;w}\cdot F(\widehat{\bf y}_{w})+{\bf x}_{w}^{\bf g}\cdot s_{k}(\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}v_{k}\widehat{\bf y}_{w}^{\bf v})x_{k;w}.$
Since $u$ and $x_{k;w}$ are log-canonical, there exists some $c\in\mathbb{Q}$ such that
$\{u,x_{k;w}\}=c\cdot ux_{k;w}=c\cdot{\bf x}_{w}^{\bf g}\cdot F(\widehat{\bf y}_{w})\cdot x_{k;w}.$
Thus we have the following equality in $\mathbb{F}$.
$({\bf g}^{T}\Lambda_{w}\widetilde{\bf e}_{k})\cdot{\bf x}_{w}^{\bf g}\cdot x_{k;w}\cdot F(\widehat{\bf y}_{w})+{\bf x}_{w}^{\bf g}\cdot s_{k}(\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}v_{k}\widehat{\bf y}_{w}^{\bf v})x_{k;w}=c\cdot{\bf x}_{w}^{\bf g}\cdot F(\widehat{\bf y}_{w})\cdot x_{k;w}.$
We get
(4.1) $s_{k}(\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}v_{k}\widehat{\bf y}_{w}^{\bf v})=(c-{\bf g}^{T}\Lambda_{w}\widetilde{\bf e}_{k})\cdot F(\widehat{\bf y}_{w}).$
Assume by contradiction that the maximal degree $f_{k}$ of $y_{k}$ in $F=\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}{\bf y}^{{\bf v}}$ is non-zero. Then the left side of (4.1) is non-zero, which implies $c-{\bf g}^{T}\Lambda_{w}\widetilde{\bf e}_{k}\neq 0$. Now we compare the constant term on both sides of (4.1). The constant term on the left side is 0 while the constant term on the right side is $(c-{\bf g}^{T}\Lambda_{w}\widetilde{\bf e}_{k})\cdot 1=c-{\bf g}^{T}\Lambda_{w}\widetilde{\bf e}_{k}\neq 0$. This is a contradiction. Hence, $f_{k}=0$. ∎
Theorem 4.26.
Let $u\in\mathcal{U}^{\operatorname{\mathsf{good}}}$ and $w\in\mathbb{T}_{n}$. The following statements hold.
  • (i)
    If $u$ and $x_{k;w}$ are log-canonical for some $k\in[1,n]$, then $(x_{k;w}\mid\mid u)_{F}=0$.
  • (ii)
    If $u$ is a cluster monomial and it is log-canonical with $x_{k;w}$, then the product $u\cdot x_{k;w}$ is a cluster monomial of $\mathcal{U}$. In particular, if two cluster variables are log-canonical, then they are contained in the same cluster.
  • (iii)
    If $u$ and $x_{k;w}$ are log-canonical for any $k\in[1,n]$, then $u$ is a cluster monomial in ${\bf x}_{w}$.
  • (iv)
[…]
Proof.
(i) This follows from Lemma 4.25 and Proposition 4.13.
(ii) By (i), we have $(x_{k;u}\mid\mid u)_{F}=0$. Since $u$ is a cluster monomial and by Theorem 4.19, we obtain that the product $u\cdot x_{k;w}$ is a cluster monomial of $\mathcal{U}$.
Both (iii) and (iv) follow from (i) and Corollary 4.15. ∎
$4.$ $F$-invariant for cluster algebras with trivial coefficients
In this subsection we show that the notion of $F$-invariant for a pair of cluster monomials can be defined for any upper cluster algebra with trivial coefficients, regardless of whether it is a $\mathsf{\Lambda}$-upper cluster algebra.
Recall that in the process of our discovery and definition of $F$-invariant, we work on $\mathsf{\Lambda}$-(upper) cluster algebras and thus (upper) cluster algebras of full rank, because the $F$-invariant is defined as the symmetrized sum of tropical invariant
$(u\mid\mid u^{\prime})_{F}=\beta_{u^{\prime}}(u)+\beta_{u}(u^{\prime})=\langle u,u^{\prime}\rangle_{\rm trop}+\langle u^{\prime},u\rangle_{\rm trop}$
and the tropical invariant heavily depends on the Poisson coefficient matrices $\mathsf{\Lambda}=\{\Lambda_{t}\mid t\in\mathbb{T}_{n}\}$. However, if we look at Theorem 4.10 (ii), we know that $(u\mid\mid u^{\prime})_{F}$ is given as follows:
$(u\mid\mid u^{\prime})_{F}=F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}]+F_{u^{\prime}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}],$
where $w$ is any vertex in $\mathbb{T}_{n}$. This formula eliminates the dependence of $(u\mid\mid u^{\prime})_{F}$ on the Poisson coefficient matrices $\mathsf{\Lambda}=\{\Lambda_{t}\mid t\in\mathbb{T}_{n}\}$. Thus we can define $F$-invariant for any (upper) cluster algebra with trivial coefficients whenever we have well-defined $g$-vectors and $F$-polynomials.
The philosophy in this subsection is that for each (upper) cluster algebra with trivial coefficients there is a natural $\mathsf{\Lambda}$-(upper) cluster algebra behind it.
Setting 4.27.
(i) Let $B$ be an $n\times n$ skew-symmetrizable matrix with a fixed skew-symmetrizer $S=diag(s_{1},\ldots,s_{n})$. Let ${\bf z}=(z_{1},\ldots,z_{n})$ and ${\bf x}=(x_{1},\ldots,x_{m})$ be the ordered sets of indeterminates, where $m=2n$.
(ii) Set $\widetilde{B}=\begin{bmatrix}B\\ I_{n}\end{bmatrix}$ and $\Lambda=\begin{bmatrix}{\bf 0}&-S\\ S&-SB\end{bmatrix}$. Since $\widetilde{B}^{T}\Lambda=(S\mid{\bf 0})$, the pair $(\widetilde{B},\Lambda)$ forms a compatible pair of type $S$.
(iii) Let $\mathcal{U}$ be the $\mathsf{\Lambda}$-upper cluster algebra with initial $\mathsf{\Lambda}$-seed $({\bf x},\widetilde{B},\Lambda)$ and let $\mathcal{U}^{\circ}$ be the upper cluster algebra with trivial coefficients whose initial seed is $({\bf z},B)$. Denote by $({\bf x}_{t},\widetilde{B}_{t},\Lambda_{t})$ the $\mathsf{\Lambda}$-seed of $\mathcal{U}$ at vertex $t\in\mathbb{T}_{n}$ and by $({\bf z}_{t},B_{t})$ the seed of $\mathcal{U}^{\circ}$ at vertex $t\in\mathbb{T}_{n}$, where ${\bf x}_{t}=(x_{1;t},\ldots,x_{m;t})$ and ${\bf z}_{t}=(z_{1;t},\ldots,z_{n;t})$.
(iv) For a vector ${\bf h}=(h_{1},\ldots,h_{m})^{T}\in\mathbb{Z}^{m}$, denote by ${\bf h}^{\circ}\coloneqq(h_{1},\ldots,h_{n})^{T}\in\mathbb{Z}^{n}$ the principal part of $\bf h$}.
Theorem 4.28 (Separation formula, [fomin_zelevinsky_2007]).
Keep the above setting. Let $u^{\circ}=\prod_{i=1}^{n}z_{i;t}^{a_{i}}$ be a cluster monomial in $\mathcal{U}^{\circ}$ and $u=\prod_{i=1}^{n}x_{i;t}^{a_{i}}$ the corresponding cluster monomial in $\mathcal{U}$. Denote by ${\bf g}_{u}^{w}\in\mathbb{Z}^{m}$ and $F_{u}^{w}\in\mathbb{Z}[y_{1},\ldots,y_{n}]$ the extended $g$-vector and $F$-polynomial of $u$ with respect to vertex $w\in\mathbb{T}_{n}$. Then the Laurent expansion of $u^{\circ}$ with respect to seed $({\bf z}_{w},B_{w})$ has a cannonical expression in terms of ${\bf g}_{u}^{w}$ and $F_{u}^{w}$:
$u^{\circ}={\bf z}_{w}^{({\bf g}_{u}^{w})^{\circ}}F_{u}^{w}({\bf z}_{w}^{B_{w}{\bf e}_{1}},\ldots,{\bf z}_{w}^{B_{w}{\bf e}_{n}}),$
where $({\bf g}_{u}^{w})^{\circ}\in\mathbb{Z}^{n}$ is the principal part of ${\bf g}_{u}^{w}\in\mathbb{Z}^{m}$ and ${\bf e}_{k}$ is the $k$th column of $I_{n}$.
Definition 4.29.
Keep the notation in the above theorem. We call ${\bf g}_{u^{\circ}}^{w}\coloneqq({\bf g}_{u}^{w})^{\circ}\in\mathbb{Z}^{n}$ and $F_{u^{\circ}}^{w}\coloneqq F_{u}^{w}\in\mathbb{Z}[y_{1},\ldots,y_{n}]$ the $g$-vector and $F$-polynomial of the cluster monomial $u^{\circ}\in\mathcal{U}^{\circ}$ with respect to vertex $w\in\mathbb{T}_{n}$.
In the following proposition, we prove that $F$-invariant can be naturally defined for upper cluster algebras with trivial coefficients.
Proposition 4.30.
Let $B$ be an $n\times n$ skew-symmetrizable matrix with a fixed skew-symmetrizer $S$. Let $\mathcal{U}^{\circ}$ be the upper cluster algebra with trivial coefficients whose initial seed is $({\bf z},B)$. Let $u^{\circ}$ and $v^{\circ}$ be two cluster monomials in $\mathcal{U}^{\circ}$. Then the integer
$(u^{\circ}\mid\mid v^{\circ})_{F}\coloneqq F_{u^{\circ}}^{w}[S{\bf g}_{v^{\circ}}^{w}]+F_{v^{\circ}}^{w}[S{\bf g}_{u^{\circ}}^{w}]$
only depends on $u^{\circ}$ and $v^{\circ}$, not on the choice of the vertex $w\in\mathbb{T}_{n}$.
Proof.
Let $\mathcal{U}$ be the $\mathsf{\Lambda}$-upper cluster algebra as in the Setting 4.27 and let $u,v$ be the cluster monomials of $\mathcal{U}$ corresponding to the cluster monomials $u^{\circ},v^{\circ}$ of $\mathcal{U}^{\circ}$. By Theorem 4.10 (ii), we have
$(u\mid\mid v)_{F}=F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{v}^{w}]+F_{v}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}],$
where $w$ is any vertex in $\mathbb{T}_{n}$. Clearly, we have $(S\mid{\bf 0}){\bf g}_{u}^{w}=S({\bf g}_{u}^{w})^{\circ}$ and $(S\mid{\bf 0}){\bf g}_{v}^{w}=S({\bf g}_{v}^{w})^{\circ}$, where $({\bf g}_{u}^{w})^{\circ},\;({\bf g}_{v}^{w})^{\circ}\in\mathbb{Z}^{n}$ are the principal parts of ${\bf g}_{u}^{w},\;{\bf g}_{v}^{w}\in\mathbb{Z}^{m}$. Thus
$(u\mid\mid v)_{F}=F_{u}^{w}[S({\bf g}_{v}^{w})^{\circ}]+F_{v}^{w}[S({\bf g}_{u}^{w})^{\circ}].$
By definition, we know that ${\bf g}_{u^{\circ}}^{w}=({\bf g}_{u}^{w})^{\circ},\;{\bf g}_{v^{\circ}}^{w}=({\bf g}_{v}^{w})^{\circ}$ and $F_{u^{\circ}}^{w}=F_{u}^{w},\;F_{v^{\circ}}^{w}=F_{v}^{w}$. So we have
$(u\mid\mid v)_{F}=F_{u^{\circ}}^{w}[S{\bf g}_{v^{\circ}}^{w}]+F_{v^{\circ}}^{w}[S{\bf g}_{u^{\circ}}^{w}]=(u^{\circ}\mid\mid v^{\circ})_{F},$
which is independent of the choice of the vertex $w\in\mathbb{T}_{n}$. ∎
Remark 4.31.
The key ingredients to define $F$-invariant is the (extended) $g$-vectors and $F$-polynomials. The $g$-vectors and $F$-polynomials are natural concepts in the setting of (upper) cluster algebras of full rank or $\mathsf{\Lambda}$-(upper) cluster algebras. In this setting, we can even define $g$-vectors and $F$-polynomials for many elements beyond cluster monomials.
In contrast, when we work on (upper) cluster algebras with trivial coefficents, we need to restrict to cluster monomials in order to have a natural definition of $g$-vectors and $F$-polynomials. On the other hand, we see that during the process of defining $g$-vectors and $F$-polynomials for an upper cluster algebra with trivial coefficients, we lift to an upper cluster algebra of full rank.
5. Comparison with $\Lambda$$\mathfrak{d}$-invariant in monoidal cluster categorification
5.1. Monoidal categorification of cluster algebras
Let $\mathbf{k}$ be a base field and let $\mathcal{C}$ be a $\mathbf{k}$-linear monoidal abelian category such that any object of $\mathcal{C}$ has finite length. We denote by $[M]$ the isomorphism class of an object $M\in\mathcal{C}$ and by $\operatorname{\mathsf{Irr}}(\mathcal{C})$ the set of isomorphism classes of simple objects in $\mathcal{C}$. Let $\mathcal{K}_{0}(\mathcal{C})$ be the Grothendieck group of $\mathcal{C}$. We know that
  • $\mathcal{K}_{0}(\mathcal{C})$ is a free $\mathbb{Z}$-module with a basis given by $\operatorname{\mathsf{Irr}}(\mathcal{C})$, i.e., $\mathcal{K}_{0}(\mathcal{C})=\bigoplus_{[S]\in\operatorname{\mathsf{Irr}}(\mathcal{C})}\mathbb{Z}[S]$.
  • $\mathcal{K}_{0}(\mathcal{C})$ is a ring with multiplicity induced by $M][N]\coloneqq[M\otimes N]$, where $M,N\in\mathcal{C}$.
Definition 5.1.
Let $M$ and $N$ be two simple objects in $\mathcal{C}$.
  • (i)
    We say that $M$ and $N$ commute, if $M\otimes N\cong N\otimes M$.
  • (ii)
    We say that $M$ and $N$ strongly commute, if $M$ and $N$ commute and $M\otimes N$ is simple.
  • (iii)
    We say that $M$ is real, if $M\otimes M$ is simple.
Definition 5.2.
Let $\mathcal{M}=(M_{1},\ldots,M_{r})$ be an ordered set of simple objects in $\mathcal{C}$.
  • (i)
    We say that $\mathcal{M}$ a commuting set, if $M_{i}$ and $M_{j}$ commute for any $i,j\in[1,r]$.
  • (ii)
    We say that $\mathcal{M}$ is a strongly commuting set, if $M_{i}$ and $M_{j}$ strongly commute for any $i,j\in[1,r]$. In particular, each $M_{i}$ is a real simple object.
Given a commuting set $\mathcal{M}=(M_{1},\ldots,M_{r})$ and a vector ${\bf a}=(a_{1},\ldots,a_{r})^{T}\in\mathbb{Z}_{\geq 0}^{r}$, denote
$\mathcal{M}({\bf a})\coloneqq M_{1}^{\otimes a_{1}}\otimes M_{2}^{\otimes a_{2}}\otimes\ldots\otimes M_{r}^{\otimes a_{r}}.$
Clearly, for any ${\bf a},{\bf c}\in\mathbb{Z}_{\geq 0}^{r}$, we have $\mathcal{M}(\mathbf{a})\otimes\mathcal{M}(\mathbf{c})\cong\mathcal{M}(\mathbf{a}+\mathbf{c})$.
Definition 5.3 (Monoidal categorification).
A finite length, $\mathbf{k}$-linear, monoidal abelian category $\mathcal{C}$ is a monoidal categorification of a cluster algebra $\mathcal{A}^{+}$ (with non invertible frozen variables), if there exists a $\mathbb{Z}$-algebra isomorphism
$\varphi\colon\mathcal{K}_{0}(\mathcal{C})\rightarrow\mathcal{A}^{+}$
such that the cluster monomials of $\mathcal{A}^{+}$ correspond to a subset of simple objects (up to isomorphism) in $\mathcal{C}$. We denote by $(\mathcal{A}^{+},\mathcal{C},\varphi)$ for this monoidal categorification.
The notion of monoidal categorification of cluster algebras is originally introduced by Hernandez and Leclerc [HL_2010]. Examples of cluster monoidal categorification include various monoidal subcategories of finite-dimensional modules over quantum affine algebras and quiver Hecke algebras [HL16, kkko-2018, kkop-2024], and the category of perverse coherent sheaves on the affine Grassmannian of $GL_{n}$ [CW19].
From now on, we fix a monoidal categorification $(\mathcal{A}^{+},\mathcal{C},\varphi)$ and denote
$\varphi(M)\coloneqq\varphi([M])\in\mathcal{A}^{+}$
for any object $M\in\mathcal{C}$. Let $({\bf x}_{t},\widetilde{B}_{t})$ a seed of $\mathcal{A}^{+}$ and let $M_{i;t}\in\mathcal{C}$ be a simple object such that $\varphi(M_{i;t})=x_{i;t}$, which is unique up to isomorphism. We call $\mathcal{M}_{t}\coloneqq(M_{1;t},\ldots,M_{m;t})$ a monoidal cluster and $(\mathcal{M}_{t},\widetilde{B}_{t})$ a monoidal seed. Following the terminology in the additive categorification, a simple object $M\in\mathcal{C}$ is said to be reachable if $\varphi(M)$ is a cluster monomial.
The following facts can be checked easily.
  • Each monoidal cluster $\mathcal{M}_{t}=(M_{1;t},\ldots,M_{m;t})$ is a strongly commuting set.
  • Let ${\bf a}\in\mathbb{Z}_{\geq 0}^{m}$, then $\varphi(\mathcal{M}_{t}({\bf a}))={\bf x}_{t}^{\bf a}$. In particular, $\mathcal{M}_{t}({\bf a})$ is a reachable simple object.
  • Each reachable simple object of $\mathcal{C}$ is real.
Proposition 5.4 ([HL_2010, Proposition 2.2]).
Let $(\mathcal{A}^{+},\mathcal{C},\varphi)$ be a monoidal categorification and $M$ an object in $\mathcal{C}$. Then $\varphi(M)$ is universally positive in $\mathcal{A}^{+}$, that is, for any cluster ${\bf x}=(x_{1},\ldots,x_{m})$ of $\mathcal{A}^{+}$, we have $\varphi(M)\in\mathbb{Z}_{\geq 0}[x_{1}^{\pm 1},\ldots,x_{m}^{\pm 1}]$.
5.2. $\mathsf{\Lambda}$-structure and $\mathsf{\Lambda}$-monoidal categorification
In this subsection, we fix a finite length, $\mathbf{k}$-linear, monoidal abelian category $\mathcal{C}$. For an object $M\in\mathcal{C}$, we denote by $\operatorname{\mathsf{hd}}(M)$ the head of $M$. For any two objects $M,N$ in $\mathcal{C}$, we denote by
$M\nabla N\coloneqq\operatorname{\mathsf{hd}}(M\otimes N)$
the head of the tensor product $M\otimes N$.
The $\Lambda$-invariant and $\mathfrak{d}$-invariant are introduced by Kang, Kashiwara, Kim and Oh [kkko-2018] in the representation theory of quiver Hecke algebras and by Kashiwara, Kim, Oh and Park [kkop-2020] in the representation theory of quantum affine algebras in the study of monoidal categorification of cluster algebras. The definitions of $\Lambda$-invariant (and hence $\mathfrak{d}$-invariant) in both settings are constructive. For the purpose of this paper, we will rely only on the properties of the $\Lambda$-invariant, rather than on the specific constructions. For this reason, we introduce the notion of $\mathsf{\Lambda}$-structure on a monoidal (abelian) category, which allows us to unify the definitions of $\Lambda$-invariant and $\mathfrak{d}$-invariant in different settings and unify the proofs for our results.
Definition 5.5 ($\mathsf{\Lambda}$-structure and $\Lambda/\mathfrak{d}$-invariant).
A $\mathsf{\Lambda}$-structure on a monoidal category $\mathcal{C}$ is a pair $\mathsf{\Lambda}=(\Omega,\Lambda)$ such that (i) and (ii) hold.
  • (i)
    $\Omega$ is a class of non-zero objects of $\mathcal{C}$ containing all the simple objects and closed under isomorphism, taking tensor products and non-zero subquotients.
  • (ii)
    ${\Lambda}\colon\Omega\times\Omega\rightarrow\mathbb{Z}$ is a map satisfying the following conditions.
    • (a1)
      The number $\Lambda(M,N)$ depends only on the isomorphism classes of $M$ and $N$.
    • (a2)
      For $M,N\in\Omega$, we set
      $\mathfrak{d}(M,N)\coloneqq\frac{\Lambda(M,N)+\Lambda(N,M)}{2}.$
      If $M$ and $N$ are real simple objects and they strongly commute, then $\mathfrak{d}(M,N)=0$.
    • (a3)
      Let $M,N,L\in\Omega$. If $L$ is a simple object, then
      $\Lambda(M\otimes N,L)$ $=\Lambda(M,L)+\Lambda(N,L),$
      $\Lambda(L,M\otimes N)$ $=\Lambda(L,M)+\Lambda(L,N).$
    • (a4)
      Let $M,N,L\in\Omega$ be simple objects. If $L$ is real and it strongly commutes with $N$, then
      $\Lambda(M\nabla N,L)=\Lambda(M,L)+\Lambda(N,L).$
      Dually, if $L$ is real and it strongly commutes with $M$, then
      $\Lambda(L,M\nabla N)=\Lambda(L,M)+\Lambda(L,N).$
    • (a5)
      Let $M,N\in\Omega$. Then for any subquotient $0\neq M^{\prime}$ of $M$ and subquotient $0\neq N^{\prime}$ of $N$, we have $\Lambda(M^{\prime},N^{\prime})\leq\Lambda(M,N)$.
We call the numbers $\Lambda(M,N)$ and $\mathfrak{d}(M,N)$ the $\Lambda$-invariant and $\mathfrak{d}$-invariant of the pair $(M,N)$. If $\mathsf{\Lambda}=(\Omega,\Lambda)$ is a $\mathsf{\Lambda}$-structure on a monoidal category $\mathcal{C}$, we simply say that $\Lambda\colon\Omega\times\Omega\rightarrow\mathbb{Z}$ is a $\mathsf{\Lambda}$-structure on $\mathcal{C}$.
Remark 5.6.
The above axioms of $\mathsf{\Lambda}$-structure extract some of key properties of the $\Lambda$-invariant appearing in the representation theory of quiver Hecke algebras [kkko-2018, Lem. 3.1.5, Lem. 3.2.3, Prop. 3.2.8, Lem. 3.2.13 ] and that of quantum affine algebras [kkop-2020, Prop. 3.9, Lem. 3.10, Cor. 3.17, Lem. 4.3]. More generally, these properties are consequences of the much more natural assumption that the monoidal abelian category $\mathcal{C}$ admits a system of renormalized $r$-matrices and has separable triple products in the sense of [CW19, Section 4].
A system of renormalized $r$-matrices is an assignment of a non-zero homomorphism
$\mathbf{r}_{M,N}\colon M\otimes N\to N\otimes M$
together with an integer $\Lambda(M,N)$ for each ordered pair $(M,N)$ in $\Omega$ satisfying certain compatibility conditions between $\mathbf{r}_{M,N}$ and $\Lambda(M,N)$, which we can regard as a weaker version of the structure of a braided monoidal category. In general, the homomorphisms $\mathbf{r}_{M,N}$ need neither to be isomorphisms nor to be natural in $(M,N)$. The $\Lambda$-invariant somehow controls how close these homomorphisms are to being natural. (Although [CW19, Definition 4.1] requires $\mathbf{r}_{M,N}$ and $\Lambda(M,N)$ to be defined for all $M,N\in\mathcal{C}$, there is no harm to weaken this requirement so that they are defined only for $M,N\in\Omega$ to obtain all the necessary consequences.) Having separable triple products is a weaker version of the rigidity in the monoidal category.
Many interesting non-commutative monoidal abelian categories admitting generic braidings satisfy these conditions. Examples include the category of finite-dimensional modules over symmetric quiver Hecke algebras and the quantum affine algebras [KKKO-2015, kkko-2018, kkop-2020], the category of finite-length representations of general linear groups over $p$-adic fields [LM18, Dros], and the category of equivariant Koszul-perverse coherent sheaves on the BFN spaces of triples (including the equivariant perverse coherent sheaves on affine Grassmannians as special cases) categorifying the $K$-theoretic Coulomb branches [CW19, CW23]. In all these examples, the renormalized $r$-matrices ${\bf r}_{M,N}$ are obtained from a canonical family of braiding operators $R_{M,N}(z)$, meromorphically depending on an additional continuous parameter $z$ by a suitable renormalization and a specialization as ${\bf r}_{M,N}=\lim_{z\to 1}(z-1)^{s}R_{M,N}(z)$ for some $s\in\mathbb{Z}$. Then, roughly speaking, the invariant $\Lambda(M,N)$ encodes the pole/zero order $s$ of $R_{M,N}(z)$ at $z=1$. In all the known cases, the $\mathsf{\Lambda}$-structure arises in this way.
Lemma 5.7.
Let $\Lambda\colon\Omega\times\Omega\rightarrow\mathbb{Z}$ be a $\mathsf{\Lambda}$-structure on $\mathcal{C}$. Let $\mathcal{M}_{w}=(M_{1},\ldots,M_{m})$ be a strongly commuting set of real simple objects and set $\lambda_{ij}\coloneqq\Lambda(M_{i},M_{j})$ for $i,j\in[1,m]$. Then the following statements hold.
  • (i)
    The matrix $\Lambda_{w}\coloneqq(\lambda_{ij})$ is an $m\times m$ skew-symmetric integer matrix.
  • (ii)
    For any ${\bf a},{\bf b}\in\mathbb{Z}_{\geq 0}^{m}$, we have
    $\Lambda(\mathcal{M}_{w}({\bf a}),\mathcal{M}_{w}({\bf b}))={\bf a}^{T}\Lambda_{w}{\bf b}.$
Proof.
(i) Since $\mathcal{M}_{w}$ is a strongly commuting set of real simple objects and by the condition (a2) in the definition of $\mathsf{\Lambda}$}-structure, we know that $\Lambda(M_{i},M_{j})+\Lambda(M_{j},M_{i})=0$ for $i,j\in[1,m]$. Thus the matrix $\Lambda_{w}$ is skew-symmetric.
(ii) Suppose that ${\bf b}={\bf b}_{1}+{\bf b}_{2}$ with ${\bf b}_{1},{\bf b}_{2}\in\mathbb{Z}_{\geq 0}^{m}$. Then $\mathcal{M}_{w}({\bf b}_{1})\otimes\mathcal{M}_{w}({\bf b}_{2})\cong\mathcal{M}_{w}({\bf b})$. Since $\mathcal{M}_{w}({\bf a})$ is simple and by the condition (a3), we have
$\Lambda(\mathcal{M}_{w}({\bf a}),\mathcal{M}_{w}({\bf b}))=\Lambda(\mathcal{M}_{w}({\bf a}),\mathcal{M}_{w}({\bf b}_{1}))+\Lambda(\mathcal{M}_{w}({\bf a}),\mathcal{M}_{w}({\bf b}_{2})).$
The similar reduction works for ${\bf a}\in\mathbb{Z}_{\geq 0}^{m}$ as well. Repeated reductions lead to the result. ∎
The following lemma plays a key role in the proof of the main theorem (Theorem 5.16) in this section.
Lemma 5.8.
Let $\Lambda\colon\Omega\times\Omega\rightarrow\mathbb{Z}$ be a $\mathsf{\Lambda}$-structure on $\mathcal{C}$. Let $M,N,L\in\Omega$ be simple objects, and let $S_{1},\ldots,S_{l}$ be a list of composition factors of $M\otimes N$. Suppose that the head $M\nabla N$ of $M\otimes N$ is simple. Then the following statements hold.
  • (i)
    If $L$ is real and it strongly commutes with $N$, then
    $\Lambda(M,L)+\Lambda(N,L)=\Lambda(M\nabla N,L)=\max\{\Lambda(S_{i},L)\mid i\in[1,l]\}.$
  • (ii)
    If $L$ is real and it strongly commutes with $M$, then
    $\Lambda(L,M)+\Lambda(L,N)=\Lambda(L,M\nabla N)=\max\{\Lambda(L,S_{i})\mid i\in[1,l]\}.$
Proof.
(i) Since the head $M\nabla N$ is simple, we know that $M\nabla N\cong S_{k}$ for some $k\in[1,l]$. As $L$ is real and it strongly commutes with $N$, the conditions (a3) and (a4) for the $\mathsf{\Lambda}$-structure $\Lambda\colon\Omega\times\Omega\rightarrow\mathbb{Z}$ imply
$\Lambda(S_{k},L)=\Lambda(M\nabla N,L)=\Lambda(M,L)+\Lambda(N,L)=\Lambda(M\otimes N,L).$
On the other hand, the condition (a5) implies
$\Lambda(S_{i},L)\leq\Lambda(M\otimes N,L)=\Lambda(M,L)+\Lambda(N,L)$
for any $i\in[1,l]$. Therefore, we have
$\Lambda(M,L)+\Lambda(N,L)=\Lambda(M\nabla N,L)=\max\{\Lambda(S_{i},L)\mid i\in[1,l]\}.$
(ii) The proof is similar to (i). ∎
We can see that the simple head condition in the above lemma plays an important role in the proof. In order to maintain this condition, we make the following assumption on the monoidal category $\mathcal{C}$, which is automatically satisfied once $\mathcal{C}$ admits a system of renormalized $r$-matrices and has the separable triple products in the sense of [CW19, Section 4]. In particular, this is true for all known examples of $\mathsf{\Lambda}$-structures.
Assumption 5.9.
Let $M,N\in\mathcal{C}$ be simple objects such that at least one of them is real. Then the head $M\nabla N$ of $M\otimes N$ is simple.
Definition 5.10 ($\mathsf{\Lambda}$-monoidal categorification).
Let $(\mathcal{A}^{+},\mathcal{C},\varphi)$ be a monoidal categorification and $\Lambda\colon\Omega\times\Omega\rightarrow\mathbb{Z}$ a $\mathsf{\Lambda}$-structure on $\mathcal{C}$. We say that $(\mathcal{A}^{+},\mathcal{C},\varphi,\Omega,\Lambda)$ is a $\mathsf{\Lambda}$}-monoidal categorification if the following conditions hold.
  • (c1)
    The monoidal category $\mathcal{C}$ satisfies Assumption 5.9.
  • (c2)
    The $\mathsf{\Lambda}$-structure and cluster structure on $\mathcal{C}$ are compatible in the sense that there exists a vertex $w\in\mathbb{T}_{n}$ such that $(\widetilde{B}_{w},\Lambda_{w})$ forms a compatible pair, i.e., $\widetilde{B}_{w}^{T}\Lambda_{w}=(S\mid{\bf 0})$ for some diagonal matrix $S=diag(s_{1},\ldots,s_{n})$ with $s_{j}\in\mathbb{Z}_{\geq 1}$. Here $\Lambda_{w}$ is the $m\times m$ skew-symmetric matrix given as in Lemma 5.7.
The teminology of $\mathsf{\Lambda}$-monoidal categorification is borrowed from [kkop-2020, Definition 6.7].
Remark 5.11.
Thanks to Proposition 5.13 below, for a $\mathsf{\Lambda}$-monoidal categorification, we know that $(\widetilde{B}_{t})^{T}\Lambda_{t}=(S\mid{\bf 0})$ holds for any vertex $t\in\mathbb{T}_{n}$.
Remark 5.12.
The examples of $\mathsf{\Lambda}$-monoidal categorifications contain those monoidal categorifications constructed by Kang, Kashiwara, Kim, and Oh [kkko-2018] using monoidal subcategories of representations of (symmetric) quiver Hecke algebras and by Kashiwara, Kim, Oh, and Park [kkop-2024] using monoidal subcategories of representations of quantum affine algebras. Notice that in [kkko-2018, kkop-2024], the authors always take $S=2I_{n}$ in condition (c2) of Definition 5.10.
Proposition 5.13 (cf. [kkko-2018, Proposition 7.1.2(e)], [kkop-2020, Proposition 6.4]).
Let $(\mathcal{A}^{+},\mathcal{C},\varphi,\Omega,\Lambda)$ be a $\mathsf{\Lambda}$-monoidal categorification. Then the following statements hold.
  1. (i)
    For any vertex $t\in\mathbb{T}_{n}$, $(\widetilde{B}_{t},\Lambda_{t})$ forms a compatible pair.
  2. (ii)
    For any edge ${t}$${t^{\prime}}$$\scriptstyle{k}$ in $\mathbb{T}_{n}$, $(\widetilde{B}_{t^{\prime}},\Lambda_{t^{\prime}})=\mu_{k}(\widetilde{B}_{t},\Lambda_{t})$ as mutation of compatible pairs.
In particular, $\mathcal{A}^{+}$ is a $\mathsf{\Lambda}$-cluster algebra.
Proof.
For the convenience of the reader, we display the proof here. Since $(\mathcal{A}^{+},\mathcal{C},\varphi,\Omega,\Lambda)$ is a $\mathsf{\Lambda}$-monoidal categorification, there exists a vertex $w\in\mathbb{T}_{n}$ such that $(\widetilde{B}_{w},\Lambda_{w})$ forms a compatible pair. Denote by $(\widetilde{B}^{\prime},\Lambda^{\prime})=\mu_{k}(\widetilde{B}_{w},\Lambda_{w})$ the compatible pair obtained from $(\widetilde{B}_{w},\Lambda_{w})$ by mutation in direction $k\in[1,n]$, where $\Lambda^{\prime}$ is given by the equation (2.5).
Let $w^{\prime}=\mu_{k}(w)$ be the vertex in $\mathbb{T}_{n}$ such that $w$ and $w^{\prime}$ are joined by the edge labeled by $k\in[1,n]$. We know that $\widetilde{B}^{\prime}=\widetilde{B}_{w^{\prime}}$. Denote by $\Lambda_{t}=(\lambda_{ij;t})$ for $t\in\mathbb{T}_{n}$ and $\Lambda^{\prime}=(\lambda_{ij}^{\prime})$. All we need to do is to prove $\Lambda_{w^{\prime}}=\Lambda^{\prime}$, i.e., $\lambda_{ij;w}=\lambda_{ij}^{\prime}$ for any $i,j\in[1,m]$.
We first consider the case $i\neq k$ and $j\neq k$. In this case,
$\lambda_{ij}^{\prime}=\lambda_{ij;w}=\Lambda(M_{i;w},M_{j;w})=\Lambda(M_{i;w^{\prime}},M_{j;w^{\prime}})=\lambda_{ij;w^{\prime}}.$
Now we consider the case $i\neq k$ and $j=k$. Denote by ${\bf b}_{k}=(b_{1k},\ldots,b_{mk})^{T}\in\mathbb{Z}^{m}$ the $k$th column of $\widetilde{B}_{w}$ and by
${\bf b}_{k}]_{+}=([b_{1k}]_{+},\ldots,[b_{mk}]_{+})^{T},\quad[-{\bf b}_{k}]_{+}=([-b_{1k}]_{+},\ldots,[-b_{mk}]_{+})^{T}.$
Since $(\widetilde{B}_{w},\Lambda_{w})$ is a compatible pair and $i\neq k$, the $(i,k)$-entry of $\Lambda_{w}\widetilde{B}_{w}$ is zero and thus
$\Lambda(M_{i;w},\mathcal{M}_{w}([-{\bf b}_{k}]_{+}))=\sum_{l=1}^{m}[-b_{lk}]_{+}\lambda_{il;w}=\sum_{l=1}^{m}[b_{lk}]_{+}\lambda_{il;w}=\Lambda(M_{i;w},\mathcal{M}_{w}([{\bf b}_{k}]_{+})).$
In the Grothendieck ring $\mathcal{K}_{0}(\mathcal{C})$, we have the relation:
$M_{k;w^{\prime}}][M_{k;w}]=[M_{k;w}\otimes M_{k;w^{\prime}}]=[\mathcal{M}_{w}([{\bf b}_{k}]_{+})]+[\mathcal{M}_{w}([-{\bf b}_{k}]_{+})].$
So either $M_{k;w}\nabla M_{k;w^{\prime}}\cong\mathcal{M}_{w}([-{\bf b}_{k}]_{+})$ or $M_{k;w}\nabla M_{k;w^{\prime}}\cong\mathcal{M}_{w}([{\bf b}_{k}]_{+})$. In either case, we have
$\Lambda(M_{i;w},M_{k;w}\nabla M_{k;w^{\prime}})=\Lambda(M_{i;w},\mathcal{M}_{w}([-{\bf b}_{k}]_{+}))=\Lambda(M_{i;w},\mathcal{M}_{w}([{\bf b}_{k}]_{+})).$
Since $M_{i;w}$ strongly commutes with $M_{k;w}$ and by the condition (a4) in Definition 5.5, we have
$\Lambda(M_{i;w},\mathcal{M}_{w}([-{\bf b}_{k}]_{+}))=\Lambda(M_{i;w},M_{k;w}\nabla M_{k;w^{\prime}})=\Lambda(M_{i;w},M_{k;w})+\Lambda(M_{i;w},M_{k;w^{\prime}}).$
Thus
$\lambda_{ik;w^{\prime}}=\Lambda(M_{i;w},M_{k;w^{\prime}})$ $=-\Lambda(M_{i;w},M_{k;w})+\Lambda(M_{i;w},\mathcal{M}_{w}([-{\bf b}_{k}]_{+}))$
$=-\lambda_{ik;w}+\sum_{l=1}^{m}[-b_{lk}]_{+}\lambda_{il;w}.$
By comparing with equation (2.5), we see that the $(i,k)$-entries of $\Lambda^{\prime}$ and $\Lambda_{w^{\prime}}$ are the same. Since both $\Lambda^{\prime}$ and $\Lambda_{w^{\prime}}$ is skew-symmetric, we know that the $(k,i)$-entries of $\Lambda^{\prime}$ and $\Lambda_{w^{\prime}}$ are also the same. Thus $\Lambda^{\prime}=\Lambda_{w^{\prime}}$. So $(\widetilde{B}_{w^{\prime}},\Lambda_{w^{\prime}})=\mu_{k}(\widetilde{B}_{w},\Lambda_{w})$ as mutation of compatible pairs. ∎
5.3. Comparison with $\Lambda$}-invariant and $\mathfrak{d}$-invariant
In this subsection, we compare the tropical invariant and $F$-invariant with the $\Lambda$}-invariant and $\mathfrak{d}$-invariant.
Lemma 5.14.
Let $(\mathcal{A}^{+},\mathcal{C},\varphi,\Omega,\Lambda)$ be a $\mathsf{\Lambda}$-monoidal categorification and $M$ a simple object in $\mathcal{C}$. Then for any monoidal cluster $\mathcal{M}_{w}$, there exist ${\bf d,h,h^{\prime}}\in\mathbb{Z}_{\geq 0}^{m}$ satisfying
$\mathcal{M}_{w}({\bf d})\nabla M\cong\mathcal{M}_{w}({\bf h})\quad\text{and}\quad M\nabla\mathcal{M}_{w}({\bf d})\cong\mathcal{M}_{w}({\bf h^{\prime}}).$
Proof.
For the vertex $w\in\mathbb{T}_{n}$, by the Laurent phenomenon, there exist a vector ${\bf d}\in\mathbb{Z}_{\geq 0}^{m}$ and a finite subset $\mathbb{B}\subseteq\mathbb{Z}_{\geq 0}^{m}$ such that
(5.1) $\varphi(M)\cdot\varphi(\mathcal{M}_{w}({\bf d}))=\varphi(M\otimes\mathcal{M}_{w}({\bf d}))=\varphi(\mathcal{M}_{w}({\bf d})\otimes M)=\sum_{{\bf b}\in\mathbb{B}}c_{{\bf b}}\varphi(\mathcal{M}_{w}({\bf b})),$
where $\{\mathcal{M}_{w}({\bf b})\mid{\bf b}\in\mathbb{B}\}$ is the set of composition factors of $M\otimes\mathcal{M}_{w}({\bf d})$ (and also those of $\mathcal{M}_{w}({\bf d})\otimes M$) and $c_{\bf b}\in\mathbb{Z}_{\geq 1}$ is the multiplicity of $\mathcal{M}_{w}({\bf b})$. Then, by the condition (c1) in Definition 5.10, we know that the heads $\mathcal{M}_{w}({\bf d})\nabla M$ and $M\nabla\mathcal{M}_{w}({\bf d})$ are simple. Hence, $\mathcal{M}_{w}({\bf d})\nabla M\cong\mathcal{M}_{w}({\bf h})$ and $M\nabla\mathcal{M}_{w}({\bf d})\cong\mathcal{M}_{w}({\bf h^{\prime}})$ for some ${\bf h,h^{\prime}}\in\mathbb{B}$. ∎
From now on, we fix a $\mathsf{\Lambda}$-monoidal categorification $(\mathcal{A}^{+},\mathcal{C},\varphi,\Omega,\Lambda)$. We know that $\mathcal{A}^{+}$ is a $\mathsf{\Lambda}$-cluster algebra.
Notations: Denote by $(\mathcal{S}_{X},\mathcal{S}_{Y},\mathsf{\Lambda})$ the Langlands-Poisson triple corresponding to the $\mathsf{\Lambda}$-cluster algebra $\mathcal{A}^{+}$. Let $M\in\mathcal{C}$ be an object such that $\varphi(M)$ is a good element in $\mathcal{A}^{+}$. For any vertex $t\in\mathbb{T}_{n}$, we denote by
${\bf g}_{M}^{t}\coloneqq{\bf g}_{\varphi(M)}^{t}\in\mathbb{Z}^{m}\;\;\text{and}\;\;F_{M}^{t}\coloneqq F_{\varphi(M)}^{t}\in\mathbb{Z}[y_{1},\ldots,y_{n}]$
the extended $g$-vector and $F$-polynomial of $\varphi(M)$ with respect to vertex $t$ and denote by
$\beta_{M}\coloneqq\beta_{\varphi(M)}\colon\mathbb{F}_{>0}=\mathbb{Q}_{\rm sf}(x_{1;t_{0}},\ldots,x_{m;t_{0}})\rightarrow\mathbb{Z}^{\rm max}$
the semifield homomorphism given by the good element $\varphi(M)$ as in Proposition 4.4.
Let $N$ be a reachable simple object, say $N\cong\mathcal{M}_{w}({\bf a})$. We know that $\varphi(N)={\bf x}_{w}^{\bf a}$ is a cluster monomial in $\mathcal{A}^{+}$ and thus it is a good element. We denote by
${\bf g}_{N,\pm}]=\{{\bf g}_{N,\pm}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$
the unique tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ determined by the condition ${\bf g}_{N,\pm}^{w}=\pm{\bf a}$. By Proposition 4.3, we know that
$\{\Lambda_{t}{\bf g}_{N,\pm}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$
are tropical points in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$. Then by Corollary 3.5, there exist unique semifield homomorphisms
$\beta_{N,\pm}\colon\mathbb{F}_{>0}=\mathbb{Q}_{\rm sf}(x_{1;t_{0}},\ldots,x_{m;t_{0}})\rightarrow\mathbb{Z}^{\rm max}$
such that
(5.2) $\beta_{N,\pm}({\bf x}_{t})=(\Lambda_{t}{\bf g}_{N,\pm}^{t})^{T}$
where $t$ is any vertex in $\mathbb{T}_{n}$. In particular, $\beta_{N,\pm}({\bf x}_{w})=\pm(\Lambda_{w}{\bf a})^{T}$.
Remark 5.15.
Since ${\bf g}_{N,+}^{w}={\bf a}$ is the extended $g$-vector of the cluster monomial $\varphi(N)={\bf x}_{w}^{\bf a}$ with respect to vertex $w$, we know that ${\bf g}_{N,+}^{t}={\bf g}_{N}^{t}$ for any vertex $t\in\mathbb{T}_{n}$ and thus
$\beta_{N,+}=\beta_{N}=\beta_{\varphi(N)}.$
For each object $M\in\mathcal{C}$, we know that $\varphi(M)$ is an element in $\mathbb{F}_{>0}=\mathbb{Q}_{\rm sf}(x_{1;t_{0}},\ldots,x_{m;t_{0}})$, by Proposition 5.4. Thus $\beta_{N,\pm}(\varphi(M))$ can be defined. For simplicity, we denote by
$\beta_{N,\pm}(M)\coloneqq\beta_{N,\pm}(\varphi(M))$
for any object $M\in\mathcal{C}$. Hence, if $\varphi(M)$ is a good element, we have
(5.3) $\beta_{N,+}(M)=\beta_{\varphi(N)}(\varphi(M))=\langle\varphi(M),\varphi(N)\rangle_{\rm trop}.$
Theorem 5.16.
Let $(\mathcal{A}^{+},\mathcal{C},\varphi,\Omega,\Lambda)$ be a $\mathsf{\Lambda}$-monoidal categorification. Then for any reachable simple object $N=\mathcal{M}_{w}({\bf a})$, the following statements hold.
  • (i)
    If $M$ is a simple object of $\mathcal{C}$, we have
    $\Lambda(M,N)=\beta_{N,+}(M),\quad\Lambda(N,M)=\beta_{N,-}(M).$
  • (ii)
    If $M$ is a simple object of $\mathcal{C}$ such that $\varphi(M)$ is a good element, we have
    $\Lambda(M,N)=\langle\varphi(M),\varphi(N)\rangle_{\rm trop},\quad\Lambda(N,M)=\langle\varphi(N),\varphi(M)\rangle_{\rm trop}$
    and hence
    $2\mathfrak{d}(M,N)=(\varphi(M)\mid\mid\varphi(N))_{F}.$
Proof.
(i) For the vertex $w\in\mathbb{T}_{n}$, as in the proof of Lemma 5.14, there are a vector ${\bf d}\in\mathbb{Z}_{\geq 0}^{m}$ and a finite subset $\mathbb{B}\subseteq\mathbb{Z}_{\geq 0}^{m}$ such that the equation (5.1) holds. Then $\{\mathcal{M}_{w}({\bf b})\mid{\bf b}\in\mathbb{B}\}$ is the set of composition factors of $\mathcal{M}_{w}({\bf d})\otimes M$.
On the one hand, the head $\mathcal{M}_{w}({\bf d})\nabla M$ is simple, by the condition (c1) in Definition 5.10. On the other hand, we know that $N=\mathcal{M}_{w}({\bf a})$ is real and it strongly commutes with $\mathcal{M}_{w}({\bf d})$. Thus Lemma 5.8 (ii) can be applied and we have
$\Lambda(N,\mathcal{M}_{w}({\bf d}))+\Lambda(N,M)=\Lambda(N,\mathcal{M}_{w}({\bf d})\nabla M)=\max\{\Lambda(N,\mathcal{M}_{w}({\bf b}))\mid{\bf b}\in\mathbb{B}\}.$
Thus
$\Lambda(N,M)$ $=$ $\max\{\Lambda(N,\mathcal{M}_{w}({\bf b}))\mid{\bf b}\in\mathbb{B}\}-\Lambda(N,\mathcal{M}_{w}({\bf d}))$
[…]
Now we calculate $\beta_{N,-}(M)$. By applying the semifield homomorphism $\beta_{N,-}\colon\mathbb{F}_{>0}\rightarrow\mathbb{Z}^{\rm max}$ to (5.1), we have $\beta_{N,-}(M)+\beta_{N,-}(\mathcal{M}_{w}({\bf d}))=\max\{\beta_{N,-}(\mathcal{M}_{w}({\bf b}))\mid{\bf b}\in\mathbb{B}\}$, i.e.,
$\beta_{N,-}(M)+\beta_{N,-}({\bf x}_{w}^{\bf d})=\max\{\beta_{N,-}({\bf x}_{w}^{\bf b})\mid{\bf b}\in\mathbb{B}\}.$
Since $N=\mathcal{M}_{w}({\bf a})$, we have $\beta_{N,-}({\bf x}_{w})=(\Lambda_{w}{\bf g}_{N,-}^{w})^{T}=-(\Lambda_{w}{\bf a})^{T}={\bf a}^{T}\Lambda_{w}$ and hence
$\beta_{N,-}(M)$ $=\max\{\beta_{N,-}({\bf x}_{w}^{\bf b})\mid{\bf b}\in\mathbb{B}\}-\beta_{N,-}({\bf x}_{w}^{\bf d})$
$=\max\{{\bf a}^{T}\Lambda_{w}{\bf b}\mid{\bf b}\in\mathbb{B}\}-{\bf a}^{T}\Lambda_{w}{\bf d}.$
Thus we obtain $\Lambda(N,M)=\beta_{N,-}(M)$. The proof of $\Lambda(M,N)=\beta_{N,+}(M)$ is similar.
(ii) The equality $\Lambda(M,N)=\langle\varphi(M),\varphi(N)\rangle_{\rm trop}$ is immediate from the first equality in (i) and (5.3). Let us prove the other one, i.e., $\Lambda(N,M)=\langle\varphi(N),\varphi(M)\rangle_{\rm trop}$. As $\varphi(M)$ is a good element and by (5.1), we know that the finite set $\mathbb{B}\subseteq\mathbb{Z}_{\geq 0}^{m}$ contains ${\bf b}_{0}\coloneqq{\bf g}_{M}^{w}+{\bf d}$ satisfying ${\bf b}-{\bf b}_{0}\in\widetilde{B}_{w}(\mathbb{Z}_{\geq 0}^{n})$ for all ${\bf b}\in\mathbb{B}$. Since
${\bf a}^{T}\Lambda_{w}({\bf b}-{\bf b}_{0})\in{\bf a}^{T}\Lambda_{w}\widetilde{B}_{w}(\mathbb{Z}_{\geq 0}^{n})=-{\bf a}^{T}(S\mid{\bf 0})^{T}(\mathbb{Z}_{\geq 0}^{n})\subseteq\mathbb{Z}_{\leq 0}$
for any ${\bf b}\in\mathbb{B}$, we deduce
$\max\{{\bf a}^{T}\Lambda_{w}{\bf b}\mid{\bf b}\in\mathbb{B}\}={\bf a}^{T}\Lambda_{w}{\bf b}_{0}={\bf a}^{T}\Lambda_{w}({\bf g}_{M}^{w}+{\bf d}).$
Combined with (5.3), it yields
(5.5) $\Lambda(N,M)={\bf a}^{T}\Lambda_{w}({\bf g}_{M}^{w}+{\bf d})-{\bf a}^{T}\Lambda_{w}{\bf d}={\bf a}^{T}\Lambda_{w}{\bf g}_{M}^{w}.$
On the other hand, as $N=\mathcal{M}_{w}({\bf a})$, we have ${\bf g}_{N}^{w}={\bf a}$, $F_{N}^{w}=F_{{\bf x}_{w}^{\bf a}}^{w}=1$ and thus
$\langle\varphi(N),\varphi(M)\rangle_{\rm trop}=({\bf g}_{N}^{w})^{T}\Lambda_{w}{\bf g}_{M}^{w}+F_{N}^{w}[(S\mid{\bf 0}){\bf g}_{M}^{w}]={\bf a}^{T}\Lambda_{w}{\bf g}_{M}^{w}.$
Thus, we get the desired equality $\Lambda(N,M)=\langle\varphi(N),\varphi(M)\rangle_{\rm trop}$. ∎
Since the reachable simple objects correspond to cluster monomials, which are always good elements in $\mathcal{A}^{+}$, we have the following consequence.
Corollary 5.17.
Let $(\mathcal{A}^{+},\mathcal{C},\varphi,\Omega,\Lambda)$ be a $\mathsf{\Lambda}$-monoidal categorification. If $M$ and $N$ are reachable simple objects, then we have
$\Lambda(M,N)=\langle\varphi(M),\varphi(N)\rangle_{\rm trop}\;\;\;\text{and}\;\;\;2\mathfrak{d}(M,N)=(\varphi(M)\mid\mid\varphi(N))_{F}.$
6. Comparison with $E$-invariant in additive cluster categorification
6.1. $E$-invariant in the theory of quivers with potentials
Throughout, let $Q=(Q_{0},Q_{1})$ be a quiver with vertex set $Q_{0}=\{1,\ldots,n\}$. We will assume that $Q$ has no cycles of length $\leq 2$. Such a quiver $Q$ corresponds to a skew-symmetric matrix $B_{Q}=(b_{ij})$, where
(6.1) $b_{ij}=\#\{j\rightarrow i\text{ in }Q_{1}\}-\#\{i\rightarrow j\text{ in }Q_{1}\}.$
For example, if $Q=1\leftarrow 2$, then $B_{Q}=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$.
For $m\in\mathbb{Z}_{\geq 0}$, let $\mathbb{C}Q_{m}$ be a $\mathbb{C}$-vector space with a $\mathbb{C}$-basis labeled by the set $Q_{m}$ of paths of length $m$ in $Q$. Clearly, $\mathbb{C}Q_{m}$ is finite-dimensional.
The completed path algebra of $Q$ is denoted by $\widehat{\mathbb{C}Q}\coloneqq\prod_{m\geq 0}\mathbb{C}Q_{m}$. A potential $W$ of $Q$ is an element $W=\sum_{m\geq 1}w_{m}$ of $\widehat{\mathbb{C}Q}$, where $w_{m}$ is a $\mathbb{C}$-linear combination of $m$-cycles of $Q$.
Let $(Q,W)$ be a quiver with potential, and $J=J(Q,W)$ the associated Jacobian algebra [DWZ08, Section 3], which is a quotient of $\widehat{\mathbb{C}Q}$ by an ideal $I_{W}$ (known as Jacobian ideal) determined by $W$. The quiver with potential $(Q,W)$ is said to be Jacobi-finite, if the corresponding Jacobian algebra $J=J(Q,W)=\widehat{\mathbb{C}Q}/I_{W}$ is finite-dimensional.
Example 6.1.
Consider the quiver $Q$ as follows:
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\beta_{3}}$${3\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\beta_{2}}$${2\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\beta_{1}}$
Then $Q_{2}=\{\beta_{2}\beta_{1},\;\beta_{3}\beta_{2},\;\beta_{1}\beta_{3}\}$. Notice that we read paths from right to left. If we take $W=\beta_{3}\beta_{2}\beta_{1}$, the corresponding Jacobian algebra $J=J(Q,W)$ is given by the quiver $Q$ with relations $\beta_{2}\beta_{1}=0,\;\beta_{3}\beta_{2}=0,\;\beta_{1}\beta_{3}=0$. In this case, $(Q,W)$ is Jacobi-finite.
A (finite-dimensional) representation of $Q=(Q_{0},Q_{1})$ is a tuple $M=(M_{i},M_{\alpha})_{i\in Q_{0},\alpha\in Q_{1}}$, where each $M_{i}$ is a finite-dimensional $\mathbb{C}$-vector space and $M_{\alpha}\colon M_{i}\rightarrow M_{j}$ is a linear map for each arrow $\alpha\colon i\rightarrow j$. Denote by $\underline{\dim}M=(\dim M_{1},\ldots,\dim M_{n})^{T}\in\mathbb{N}^{n}$ the dimension vector of $M$. Notice that in this paper, we always view vectors as column vectors.
A representation of $(Q,W)$ is a representation of $Q$ which is annihilated by the Jacobian ideal $I_{W}$. A decorated representation of $(Q,W)$ is a pair $\mathcal{M}=(M,V)$, where $M$ is a representation of $(Q,W)$ and $V=(V_{i})_{i\in Q_{0}}=(V_{1},\ldots,V_{n})$ is a tuple of finite-dimensional $\mathbb{C}$-vector spaces. We identify the representations of $(Q,W)$ with the finite-dimensional left modules of $J=J(Q,W)$.
A decorated representation $\mathcal{M}=(M,V)$ of $(Q,W)$ is said to be negative, if $M=0$; and it is said to be negative simple, if $M=0$ and there exists $i\in Q_{0}$ such that $V_{i}=\mathbb{C}$ and $V_{j}=0$ for any $j\neq i$.
The $g$-vectors, $F$-polynomials and $E$-invariants of decorated representations are introduced by Derksen, Weyman and Zelevinsky [DWZ10], which are reformulated nicely in [CLS-2015]. Now let us recall these notions following [CLS-2015, Section 3.1].
Definition 6.2 ($g$-vector, $F$-polynomial and $E$-invariant).
Let $\mathcal{M}=(M,V)$ and $\mathcal{M}^{\prime}=(M^{\prime},V^{\prime})$ be two decorated representations of $J=J(Q,W)$.
  • (i)
    The $g$-vector of $\mathcal{M}$ is a vector ${\bf g}({\mathcal{M}})\in\mathbb{Z}^{n}$ whose $i$th component is defined by
    $g_{i}(\mathcal{M})\coloneqq-\dim\operatorname{\mathsf{Hom}}_{J}(S_{i},M)+\dim\operatorname{\mathsf{Ext}}_{J}^{1}(S_{i},M)+\dim V_{i}.$
  • (ii)
    The $F$-polynomial of $\mathcal{M}=(M,V)$ is defined as follows:
    $F_{\mathcal{M}}(y_{1},\ldots,y_{n})\coloneqq\sum_{{\bf v}\in\mathbb{N}^{n}}\chi(\operatorname{\mathsf{Gr}}\nolimits_{\bf v}(M)){\bf y}^{\bf v}\;\;\in\mathbb{Z}[y_{1},\ldots,y_{n}],$
    where $\operatorname{\mathsf{Gr}}\nolimits_{\bf v}(M)$ is the submodule Grassmannian of $M$ with dimension vector $\bf v$} and $\chi$ is the Euler–Poincaré characteristic.
  • (iii)
    Let $\langle-,$-$\rangle$ be the standard inner product on $\mathbb{R}^{n}$. We define two integers:
    (6.2) $E^{\rm inj}(\mathcal{M},\mathcal{M}^{\prime})$ $\coloneqq$ $\dim\operatorname{\mathsf{Hom}}_{J}(M,M^{\prime})+\langle\underline{\dim}M,{\bf g}(\mathcal{M}^{\prime})\rangle,$
    (6.3) $E^{\rm sym}(\mathcal{M},\mathcal{M}^{\prime})$ $\coloneqq$ $E^{\rm inj}(\mathcal{M},\mathcal{M}^{\prime})+E^{\rm inj}(\mathcal{M}^{\prime},\mathcal{M}),$
    which are called the partial $E$-invariant and $E$-invariant of the pair $(\mathcal{M},\mathcal{M}^{\prime})$.
Remark 6.3.
It is easy to see that the $F$-polynomial $F_{\mathcal{M}}\in\mathbb{Z}[y_{1},\ldots,y_{n}]$ has a unique term of maximal degree ${\bf y}^{\underline{\dim}M}$ with coefficient $1$. Every other monomial in $F_{\mathcal{M}}$ is a factor of ${\bf y}^{\underline{\dim}M}$.
Corollary 6.4.
Let $\mathcal{M}=(M,V)$ and $\mathcal{M}^{\prime}=(M^{\prime},V^{\prime})$ be two decorated representations of $J=J(Q,W)$. If $\mathcal{M}=(M,V)$ is negative, i.e., $M=0$ and $V=(\mathbb{C}^{a_{1}},\ldots,\mathbb{C}^{a_{n}})$ for some ${\bf a}=(a_{1},\ldots,a_{n})^{T}\in\mathbb{Z}_{\geq 0}^{n}$, then we have
$E^{\rm inj}(\mathcal{M},\mathcal{M}^{\prime})=0,\;\;E^{\rm inj}(\mathcal{M}^{\prime},\mathcal{M})=\langle\underline{\dim}M^{\prime},{\bf a}\rangle=E^{\rm sym}(\mathcal{M},\mathcal{M}^{\prime}).$
Proof.
By definition, we have ${\bf g}(\mathcal{M})={\bf a}$. Then the results from (6.2) and (6.3). ∎
Proposition 6.5 ([DWZ10, Cor. 10.9], [CLS-2015, Lem. 3.4]).
Let $\mathcal{M}=(M,V)$ and $\mathcal{M}^{\prime}=(M^{\prime},V^{\prime})$ be two decorated representations of $J=J(Q,W)$. Let $d_{i}(V)=\dim V_{i}$. Suppose that $(Q,W)$ is Jacobi-finite. Then the following statements hold.
  • (i)
    We have
    $E^{\rm inj}(\mathcal{M},\mathcal{M}^{\prime})=\dim\operatorname{\mathsf{Hom}}_{J}(\tau^{-1}M^{\prime},M)+\langle\underline{\dim}M,\underline{\dim}V^{\prime}\rangle.$
    In particular, if $V^{\prime}=0=V$, we have
    $E^{\rm inj}(\mathcal{M},\mathcal{M}^{\prime})=\dim\operatorname{\mathsf{Hom}}_{J}(\tau^{-1}M^{\prime},M),\;\;E^{\rm inj}(\mathcal{M}^{\prime},\mathcal{M})=\dim\operatorname{\mathsf{Hom}}_{J}(\tau^{-1}M,M^{\prime})$
    and hence
    $E^{\rm sym}(\mathcal{M},\mathcal{M}^{\prime})=\dim\operatorname{\mathsf{Hom}}_{J}(\tau^{-1}M^{\prime},M)+\dim\operatorname{\mathsf{Hom}}_{J}(\tau^{-1}M,M^{\prime}).$
  • (ii)
    Let
    $0\rightarrow M\rightarrow\oplus_{i=1}^{n}I_{i}^{a_{i}}\rightarrow\oplus_{i=1}^{n}I_{i}^{b_{i}}$
    be a minimal injective presentation of $M$. Then the $i$th component of the $g$-vector ${\bf g}(\mathcal{M})$ of $\mathcal{M}$ is given by
    $g_{i}(\mathcal{M})=-a_{i}+b_{i}+d_{i}(V).$
Remark 6.6.
We remark that the above results can be extended to the Jacobian-infinite setting by considering the truncated Jacobian algebras, which are finite-dimensional, cf. [CLS-2015, Prop. 3.5, Lem. 3.4]. In particular, for any two decorated representations $\mathcal{M}$ and $\mathcal{M}^{\prime}$ of $(Q,W)$, we have
$E^{\rm inj}(\mathcal{M},\mathcal{M}^{\prime})\geq 0\;\;\;\text{and}\;\;\;\;E^{\rm sym}(\mathcal{M},\mathcal{M}^{\prime})\geq 0.$
6.2. Mutation-invariance of $E$-invariant
We fix a quiver with potential $(Q,W)$ and assign it to the rooted vertex $t_{0}$ of the $n$-regular tree $\mathbb{T}_{n}$. We will assume that $W$ is a non-degenerate potential [DWZ08, Section 7] so that we can obtain a family of quivers with potentials
$Q,W]\coloneqq\{(Q_{t},W_{t})\mid t\in\mathbb{T}_{n}\}\;\;\;\text{with}\;\;\;(Q_{t_{0}},W_{t_{0}})=(Q,W)$
by mutations of quivers with potentials [DWZ08, Section 5]. Denote by $J_{t}=J(Q_{t},W_{t})$ the Jacobian algebra of $(Q_{t},W_{t})$.
In [DWZ08, Section 10] Derkesn, Weyman and Zelevinsky introduced mutations of decorated representations, which are generalizations of BGP-reflections. Let $\mathcal{M}=(M,V)$ be a decorated representation of $(Q,W)=(Q_{t_{0}},W_{t_{0}})$. We can produce a family of decorated representations
$\mathcal{M}]\coloneqq\{\mathcal{M}^{t}=(M^{t},V^{t})\mid t\in\mathbb{T}_{n}\}\;\;\;\text{with}\;\;\;\;\mathcal{M}^{t_{0}}=\mathcal{M}$
by mutations of decorated representations, where $\mathcal{M}^{t}$ is a decorated representation of the quiver with potential $(Q_{t},W_{t})$ at vertex $t\in\mathbb{T}_{n}$. The decorated representation $\mathcal{M}$ of $(Q,W)$ is said to be negative-reachable, if there exists a vertex $w\in\mathbb{T}_{n}$ such that $\mathcal{M}^{w}=(M^{w},V^{w})$ is negative, i.e., $M^{w}=0$.
Remark 6.7.
The negative-reachable decorated representations of $(Q,W)$ play the role of cluster monomials in the additive categorification of cluster algebras and the negative simple ones play the role of initial cluster variables. See Theorem 6.10 below for a precise statement.
Theorem 6.8 ([DWZ10, Thm. 7.1]).
Let $\mathcal{M}=(M,V)$ and $\mathcal{N}=(N,U)$ be two decorated representations of $(Q,W)$, and let $\{\mathcal{M}^{t}\mid t\in\mathbb{T}_{n}\}$ and $\{\mathcal{N}^{t}\mid t\in\mathbb{T}_{n}\}$ be the families of decorated representations corresponding to $\mathcal{M}$ and $\mathcal{N}$ respectively. Then for any two vertices $t,t^{\prime}\in\mathbb{T}_{n}$, we have
$E^{\rm sym}(\mathcal{M}^{t},\mathcal{N}^{t})=E^{\rm sym}(\mathcal{M}^{t^{\prime}},\mathcal{N}^{t^{\prime}}).$
In particular, the $E$-invariant is a mutation invariant.
6.3. Comparison with $E$-invariant
In this subsection, we show that $E$-invariant coincides with $F$-invariant for skew-symmetric cluster algebras.
Setting 6.9.
(i) Let $(Q,W)=(Q_{t_{0}},W_{t_{0}})$ be a quiver with non-degenerate potential and $\mathcal{M}=\mathcal{M}^{t_{0}}=(M^{t_{0}},V^{t_{0}})$ a decorated representation of $(Q_{t_{0}},W_{t_{0}})$. By applying mutations, we have a family of quivers with potentials and a family of decorated representations
$[(Q,W)]\coloneqq\{(Q_{t},W_{t})\mid t\in\mathbb{T}_{n}\},\;\;\;\;[\mathcal{M}]\coloneqq\{\mathcal{M}^{t}\mid t\in\mathbb{T}_{n}\},$
where $\mathcal{M}^{t}=(M^{t},V^{t})$ is a decorated representation of $(Q_{t},W_{t})$.
(ii) Let $B_{Q}$ be the skew-symmetric matrix corresponding to $Q$ (see 6.1), and $\mathcal{A}_{Q}$ the cluster algebra with trivial coefficients such that its exchange matrix at vertex $t_{0}$ is $B_{Q}$.
(iii) For a cluster monomial $u={\bf x}_{w}^{\bf a}=\prod_{i=1}^{n}x_{i;w}^{a_{i}}$ of $\mathcal{A}_{Q}$, let $\{\mathcal{M}_{u}^{t}\mid t\in\mathbb{T}_{n}\}$ be the family of decorated representations determined by the following conditions:
  • For the vertex $w\in\mathbb{T}_{n}$, set $\mathcal{M}_{u}^{w}=(M_{u}^{w},V_{u}^{w})$ to be the negative decorated representation of $(Q_{w},W_{w})$ given by $M_{u}^{w}=0$ and $V_{u}^{w}=(\mathbb{C}^{a_{1}},\ldots,\mathbb{C}^{a_{n}})$.
  • For any edge ${t}$${t^{\prime}}$$\scriptstyle{k}$ in ${\mathbb{T}}_{n}$, $\mathcal{M}_{u}^{t}$ and $\mathcal{M}_{u}^{t^{\prime}}$ are related by mutation in direction $k$.
The following theorem summarizes some results on categorification of cluster algebras using decorated representations of quivers with potentials by Derksen, Weyman and Zelevinsky [DWZ10].
Theorem 6.10 ([DWZ10]).
Keep the above setting. The following statements hold.
  • (i)
    The family $\{\mathcal{M}_{u}^{t}=(M_{u}^{t},V_{u}^{t})\mid t\in\mathbb{T}_{n}\}$ of decorated representations only depends on $u$, not on the choice of $w\in\mathbb{T}_{n}$ and ${\bf a}\in\mathbb{Z}_{\geq 0}^{n}$ such that $u={\bf x}_{w}^{\bf a}$.
  • (ii)
    The correspondence $u\mapsto\{\mathcal{M}_{u}^{t}\mid t\in\mathbb{T}_{n}\}\mapsto\mathcal{M}_{u}:=\mathcal{M}_{u}^{t_{0}}$ induces a bijection from cluster monomials of $\mathcal{A}_{Q}$ to the isomorphism classes of negative-reachable decorated representations of $(Q,W)=(Q_{t_{0}},W_{t_{0}})$.
  • (iii)
    For any vertex $t\in\mathbb{T}_{n}$, we have
    ${\bf g}_{u}^{t}={\bf g}(\mathcal{M}_{u}^{t})\;\;\;\text{and}\;\;\;F_{u}^{t}(y_{1},\ldots,y_{n})=F_{\mathcal{M}_{u}^{t}}(y_{1},\ldots,y_{n}).$
    In particular, ${\bf f}_{u}^{t}=\underline{\dim}(M_{u}^{t})$, where ${\bf f}_{u}^{t}$ is the $f$-vector of $u$ with respect to vertex $t$ and the dimension vector here is viewed as column vector.
We take $S=I_{n}$ to be the fixed skew-symmetrizer for the exchange matrices of $\mathcal{A}_{Q}$, when we consider the $F$-invariant for cluster monomials in $\mathcal{A}_{Q}$.
Theorem 6.11.
Let $(Q,W)$ be a quiver with non-degenerate potential and $\mathcal{A}_{Q}$ the corresponding cluster algebra with trivial coefficients. Let $u,u^{\prime}$ be two cluster monomials of $\mathcal{A}_{Q}$, and let $\{\mathcal{M}_{u}^{t}\mid t\in\mathbb{T}_{n}\}$ and $\{\mathcal{M}_{u^{\prime}}^{t}\mid t\in\mathbb{T}_{n}\}$ be the families of decorated representations corresponding to $u$ and $u^{\prime}$. Then for any vertex $t\in\mathbb{T}_{n}$, we have
$(u\mid\mid u^{\prime})_{F}=E^{\rm sym}(\mathcal{M}_{u}^{t},\mathcal{M}_{u^{\prime}}^{t}).$
In particular, we have $(u\mid\mid u^{\prime})_{F}=E^{\rm sym}(\mathcal{M}_{u},\mathcal{M}_{u^{\prime}})$, where $\mathcal{M}_{u}:=\mathcal{M}_{u}^{t_{0}}$ and $\mathcal{M}_{u^{\prime}}:=\mathcal{M}_{u^{\prime}}^{t_{0}}$ are the decorated representations of $(Q,W)$ corresponding to $u$ and $u^{\prime}$.
Proof.
We can assume that $u$ is a cluster monomial in ${\bf x}_{w}$ for some $w\in\mathbb{T}_{n}$, say $u=\prod_{i=1}^{n}x_{i;w}^{a_{i}}$, where ${\bf a}=(a_{1},\ldots,a_{n})^{T}\in\mathbb{Z}_{\geq 0}^{n}$. In this case, $\mathcal{M}_{u}^{w}=(M_{u}^{w},V_{u}^{w})$ is given by $M_{u}^{w}=0$, $V_{u}^{w}=(\mathbb{C}^{a_{1}},\ldots,\mathbb{C}^{a_{n}})$. Thus we have
$F_{u}^{w}=1=F_{\mathcal{M}_{u}^{w}},\;\;\;\text{and}\;\;\;{\bf g}_{u}^{w}={\bf a}={\bf g}({\mathcal{M}_{u}^{w}}).$
By using the vertex $w$ to calculate $(u\mid\mid u^{\prime})_{F}$, we have
$(u\mid\mid u^{\prime})_{F}=0+F_{u^{\prime}}^{w}[{\bf g}_{u}^{w}]=F_{u^{\prime}}^{w}[{\bf a}]=({\bf f}_{u^{\prime}}^{w})^{T}{\bf a}=\langle{\bf f}_{u^{\prime}}^{w},{\bf a}\rangle,$
where ${\bf f}_{u^{\prime}}^{w}$ is the $f$-vector of $u^{\prime}$ with respect to the vertex $w$ and $\langle-,$-$\rangle$ is the standard inner product on $\mathbb{R}^{n}$. On the other hand, since $\mathcal{M}_{u}^{w}$ is negative and by Corollary 6.4, we have
$E^{\rm sym}(\mathcal{M}_{u}^{w},\mathcal{M}_{u^{\prime}}^{w})=\langle\underline{\dim}M_{u^{\prime}}^{w},{\bf a}\rangle.$
By Theorem 6.10 (iii), we have ${\bf f}_{u^{\prime}}^{w}=\underline{\dim}(M_{u^{\prime}}^{w})$ and thus $(u\mid\mid u^{\prime})_{F}=E^{\rm sym}(\mathcal{M}_{u}^{w},\mathcal{M}_{u^{\prime}}^{w})$. Then by the mutation-invariance of the $E$-invariant, we have
$(u\mid\mid u^{\prime})_{F}=E^{\rm sym}(\mathcal{M}_{u}^{t},\mathcal{M}_{u^{\prime}}^{t})$
for any vertex $t\in\mathbb{T}_{n}$. ∎
7. Dominant sets and oriented exchange graphs
In this section, we introduce the dominant sets for seeds of cluster algebras as a replacement of torsion classes for $\tau$-tilting pairs in $\tau$-tilting theory. Then we give applications to the exchange graphs of cluster algebras.
Throughout this section, we will focus on cluster variables and cluster monomials. For this reason, we simply work on cluster algebra with trivial coefficients, rather than $\mathsf{\Lambda}$-cluster algebras. By Proposition 4.30, the $F$-invariant between cluster monomials is well-defined in this case.
7.1. Oriented exchange graphs are acyclic
Let $B$ be an $n\times n$ skew-symmetrizable matrix with a fixed skew-symmetrizer $S=diag(s_{1},\ldots,s_{n})$. Let $\mathcal{A}$ be the cluster algebra with trivial coefficients whose initial seed is $({\bf x},B)$. The seed of $\mathcal{A}$ at vertex $t\in\mathbb{T}_{n}$ is denoted by $({\bf x}_{t},B_{t})$. In particular, we have $({\bf x}_{t_{0}},B_{t_{0}})=({\bf x},B)$.
Two seeds $({\bf x}_{t},B_{t})$ and $({\bf x}_{w},B_{w})$ of $\mathcal{A}$ are said to be equivalent if they are the same up to relabeling. We denote by ${\bf x}_{t},B_{t}]$ the equivalent class of $({\bf x}_{t},B_{t})$.
The exchange graph $\mathcal{H}=\mathcal{H}(\mathcal{A})$ of $\mathcal{A}$ is the graph whose vertices correspond to the seeds (up to seed equivalence) of $\mathcal{A}$ and whose edges correspond to seed mutations.
For a seed $({\bf x}_{t},B_{t})$ of $\mathcal{A}$, we denote by $C_{t}=C_{t}^{t_{0}}$ the $C$-matrix of the vertex $t$ with respect to the rooted vertex $t_{0}$.
Definition 7.1 (Green mutation and oriented exchange graph).
(i) A seed mutation $\mu_{k}({\bf x}_{t},B_{t})$ in $\mathcal{A}$ is called a green mutation, if the $k$th column of the $C$-matrix $C_{t}=C_{t}^{t_{0}}$ is a non-negative vector. Otherwise, it is called a red mutation.
(ii) The oriented exchange graph of $\mathcal{A}$ with initial seed $({\bf x}_{t_{0}},B_{t_{0}})$ is the quiver $\overrightarrow{\mathcal{H}}^{t_{0}}=\overrightarrow{\mathcal{H}}^{t_{0}}(\mathcal{A})$ whose vertices correspond to the seeds (up to seed equivalence) of $\mathcal{A}$ and whose arrows correspond to green mutations.
We remark that if two seeds $({\bf x}_{t},B_{t})$ and $({\bf x}_{w},B_{w})$ are related by once mutation, then either the mutation from $({\bf x}_{t},B_{t})$ to $({\bf x}_{w},B_{w})$ is a green mutation or the mutation from $({\bf x}_{w},B_{w})$ to $({\bf x}_{t},B_{t})$ is a green mutation.
Denote by $\mathcal{X}$ the set of cluster variables of $\mathcal{A}$. Given a cluster monomial $u={\bf x}_{t}^{\bf h}$ in ${\bf x}_{t}$, that is, ${\bf h}=(h_{1},\ldots,h_{n})^{T}\in\mathbb{Z}_{\geq 0}^{n}$, we define two subsets of $\mathcal{X}$:
  • The subset $\operatorname{\mathsf{supp}}(u)=\{x_{i;t}\mid i\in[1,n],\;h_{i}\neq 0\}\subseteq\mathcal{X}$ is called the support of $u={\bf x}_{t}^{\bf h}$. We remark that this set only depends on $u$, not on the choice of $t\in\mathbb{T}_{n}$ and ${\bf h}\in\mathbb{Z}^{n}$ such that $u={\bf x}_{t}^{\bf h}$.
  • For a vertex $w\in\mathbb{T}_{n}$ and a cluster variable $z$, recall from Proposition 4.30 that
    $(z\mid\mid u)_{F}=F_{z}^{w}[S{\bf g}_{u}^{w}]+F_{u}^{w}[S{\bf g}_{z}^{w}].$
    The subset $\operatorname{\mathsf{dom}}^{w}(u)=\{z\in\mathcal{X}\mid F_{z}^{w}[S{\bf g}_{u}^{w}]=0\}\subseteq\mathcal{X}$ is called the dominant set of $u={\bf x}_{t}^{\bf h}$ with respect to vertex $w\in\mathbb{T}_{n}$.
Remark 7.2.
Notice that $F_{z}^{w}[S{\bf g}_{u}^{w}]$ is the partial $F$-invariant of the pair $(z,u)$ at vertex $w\in\mathbb{T}_{n}$. That is, the dominant set is defined using the vanishing condition of the partial $F$-invariant.
Definition 7.3 (Dominant set of a seed).
Let $({\bf x}_{t},B_{t})$ be a seed of $\mathcal{A}$. The dominant set $\operatorname{\mathsf{dom}}^{w}[t]$ of $({\bf x}_{t},B_{t})$ with respect to a vertex $w\in\mathbb{T}_{n}$ is defined to be the dominant set of the multiplicity free cluster monomial $u_{t}=\prod_{i=1}^{n}x_{i;t}$ with full support, i.e.,
$\operatorname{\mathsf{dom}}^{w}[t]\coloneqq\operatorname{\mathsf{dom}}^{w}(u_{t})=\{z\in\mathcal{X}\mid F_{z}^{w}[S{\bf g}_{u_{t}}^{w}]=0\}.$
Clearly, if ${\bf x}_{t^{\prime}},B_{t^{\prime}}]=[{\bf x}_{t},B_{t}]$, i.e., the two seeds are equivalent, then $\operatorname{\mathsf{dom}}^{w}[t^{\prime}]=\operatorname{\mathsf{dom}}^{w}[t]$. If we want to view a cluster ${\bf x}_{t}$ as a set, we will write
${\bf x}_{t}]\coloneqq\{x_{1;t},\ldots,x_{n;t}\}.$
Proposition 7.4.
Let $u={\bf x}_{t}^{\bf h}$ be a cluster monomial of $\mathcal{A}$ in seed $({\bf x}_{t},B_{t})$. Then the following statements hold for any vertex $w\in\mathbb{T}_{n}$.
  • (i)
    ${\bf x}_{w}]\subseteq\operatorname{\mathsf{dom}}^{w}(u)$ and in particular, ${\bf x}_{w}]\subseteq\operatorname{\mathsf{dom}}^{w}[t]$.
  • (ii)
    ${\bf x}_{t}]\subseteq\operatorname{\mathsf{dom}}^{w}(u)$ and in particular, ${\bf x}_{t}]\subseteq\operatorname{\mathsf{dom}}^{w}[t]$.
  • (iii)
    $\operatorname{\mathsf{dom}}^{w}(u)=\bigcap_{x\in\operatorname{\mathsf{supp}}(u)}\operatorname{\mathsf{dom}}^{w}(x)$ and in particular, $\operatorname{\mathsf{dom}}^{w}[t]=\bigcap_{i=1}^{n}\operatorname{\mathsf{dom}}^{w}(x_{i;t})$.
Proof.
(i) Since $F_{x_{i;w}}^{w}=1$ for $i\in[1,n]$, we have $F_{x_{i;w}}^{w}[S{\bf g}_{u}^{w}]=0$. Thus $x_{i;w}\in\operatorname{\mathsf{dom}}^{w}(u)$ for $i\in[1,n]$, that is, ${\bf x}_{w}]\subseteq\operatorname{\mathsf{dom}}^{w}(u)$. In particular, if we take $u=u_{t}=\prod_{i=1}^{n}x_{i;t}$, we obtain
${\bf x}_{w}]\subseteq\operatorname{\mathsf{dom}}^{w}(u_{t})=\operatorname{\mathsf{dom}}^{w}[t].$
(ii) Fix $i\in[1,n]$. By using the vertex $t$ to calculate $(x_{i;t}\mid\mid u)_{F}=(x_{i;t}\mid\mid{\bf x}_{t}^{\bf h})_{F}$, we see $(x_{i;t}\mid\mid u)_{F}=0$. On the other hand, we know that
$(x_{i;t}\mid\mid u)_{F}=F_{x_{i;t}}^{w}[S{\bf g}_{u}^{w}]+F_{u}^{w}[S{\bf g}_{x_{i;t}}^{w}]$
and each term in this sum is non-negative. Thus we obtain $F_{x_{i;t}}^{w}[S{\bf g}_{u}^{w}]=0$, i.e., $x_{i;t}\in\operatorname{\mathsf{dom}}^{w}(u)$. Since $i\in[1,n]$ is arbitrary, we get ${\bf x}_{t}]\subseteq\operatorname{\mathsf{dom}}^{w}(u)$. In particular, if we take $u=u_{t}=\prod_{i=1}^{n}x_{i;t}$, we obtain ${\bf x}_{t}]\subseteq\operatorname{\mathsf{dom}}^{w}(u_{t})=\operatorname{\mathsf{dom}}^{w}[t]$.
(iii) By definition, a cluster variable $z$ belongs to $\operatorname{\mathsf{dom}}^{w}(u)$ if and only if
$F_{z}^{w}[S{\bf g}_{u}^{w}]=0.$
By Proposition 4.16 (ii), we have
$F_{z}^{w}[S{\bf g}_{u}^{w}]=F_{z}^{w}[S{\bf g}_{{\bf x}_{t}^{\bf h}}^{w}]=\sum_{j=1}^{n}h_{j}\cdot F_{z}^{w}[S{\bf g}_{x_{j;t}}^{w}].$
Since $h_{j}\geq 0$ and $F_{z}^{w}[S{\bf g}_{x_{j;t}}^{w}]\geq 0$ for any $j\in[1,n]$, we know that $F_{z}^{w}[S{\bf g}_{u}^{w}]=0$ if and only if $F_{z}^{w}[S{\bf g}_{x_{j;t}}^{w}]=0$ for any $j\in[1,n]$ with $h_{j}\neq 0$, i.e., for any $x_{j;t}\in\operatorname{\mathsf{supp}}(u)$. This is equivalent to say that $z$ belongs to $\bigcap_{x\in\operatorname{\mathsf{supp}}(u)}\operatorname{\mathsf{dom}}^{w}(x)$. Hence, $\operatorname{\mathsf{dom}}^{w}(u)=\bigcap_{x\in\operatorname{\mathsf{supp}}(u)}\operatorname{\mathsf{dom}}^{w}(x)$. In particular, if we take $u=u_{t}=\prod_{i=1}^{n}x_{i;t}$, we obtain $\operatorname{\mathsf{dom}}^{w}[t]=\operatorname{\mathsf{dom}}^{w}(u_{t})=\bigcap_{i=1}^{n}\operatorname{\mathsf{dom}}^{w}(x_{i;t})$. ∎
Corollary 7.5.
Let $u$ and $u^{\prime}$ be two cluster monomials of $\mathcal{A}$ such that $\operatorname{\mathsf{supp}}(u)=\operatorname{\mathsf{supp}}(u^{\prime})$. Then $\operatorname{\mathsf{dom}}^{w}(u)=\operatorname{\mathsf{dom}}^{w}(u^{\prime})$ for any vertex $w\in\mathbb{T}_{n}$. In particular, $\operatorname{\mathsf{dom}}^{w}[t]=\operatorname{\mathsf{dom}}^{w}(u)$ for any cluster monomial $u$ with $\operatorname{\mathsf{supp}}(u)=[{\bf x}_{t}]$.
Proof.
This follows from Proposition 7.4 (iii). ∎
For a seed $({\bf x}_{t},B_{t})$ of $\mathcal{A}$, we denote by $G_{t}=G_{t}^{t_{0}}$ the $G$-matrix of the vertex $t$ with respect to the rooted vertex $t_{0}$ and by
$\mathbb{R}_{\geq 0}G_{t}\coloneqq\mathbb{R}_{\geq 0}{\bf g}_{1;t}+\cdots+\mathbb{R}_{\geq 0}{\bf g}_{n;t}\subseteq\mathbb{R}^{n}$
the cone in $\mathbb{R}^{n}$ spanned by the columns of $G_{t}=({\bf g}_{1;t},\ldots,{\bf g}_{n;t})$. We also denote by
$\mathbb{Q}_{>0}G_{t}$ $\coloneqq\mathbb{Q}_{>0}{\bf g}_{1;t}+\cdots+\mathbb{Q}_{>0}{\bf g}_{n;t}\subseteq\mathbb{Q}^{n},$
$\mathbb{Z}_{\geq 1}G_{t}$ $\coloneqq\mathbb{Z}_{\geq 1}{\bf g}_{1;t}+\cdots+\mathbb{Z}_{\geq 1}{\bf g}_{n;t}\subseteq\mathbb{Z}^{n}.$
Lemma 7.6.
Let $({\bf x}_{t^{\prime}},B_{t^{\prime}})=\mu_{k}({\bf x}_{t},B_{t})$ be a green mutation in $\mathcal{A}$. Then there exist two vectors ${\bf v},{\bf v}^{\prime}\in\mathbb{Z}_{\geq 1}^{n}$ such that $G_{t}{\bf v}-G_{t^{\prime}}{\bf v}^{\prime}\in\mathbb{Z}_{\geq 0}^{n}$.
Proof.
Without loss of generality, we can assume that $k=1$. We write $G_{t}=({\bf g}_{1},{\bf g}_{2},\ldots,{\bf g}_{n})$, where ${\bf g}_{i}\in\mathbb{Z}^{n}$. Since $({\bf x}_{t^{\prime}},B_{t^{\prime}})=\mu_{k}({\bf x}_{t},B_{t})=\mu_{1}({\bf x}_{t},B_{t})$, the $G$-matrix $G_{t^{\prime}}$ has the form $G_{t^{\prime}}=({\bf g}_{1}^{\prime},{\bf g}_{2},\ldots,{\bf g}_{n})$.
By Theorem 3.10 (i), we know that $\mathbb{R}_{\geq 0}G_{t}$ and $\mathbb{R}_{\geq 0}G_{t^{\prime}}$ are two cones of dimension $n$ in $\mathbb{R}^{n}$ and they have a common facet $\mathcal{C}$ (cone of dimension n1n-1) given by $\mathcal{C}=\mathbb{R}_{\geq 0}{\bf g}_{2}+\ldots+\mathbb{R}_{\geq 0}{\bf g}_{n}$.
Let $S=diag(s_{1},\ldots,s_{n})$ be the fixed skew-symmetrizer for the exchange matrices of $\mathcal{A}$ and ${\bf c}_{1;t}$ the first column of the $C$-matrix $C_{t}=C_{t}^{t_{0}}$. By Theorem 3.10 (iii), we have $SC_{t}S^{-1}(G_{t})^{\rm T}=I_{n}$, equivalently, $(G_{t})^{\rm T}(SC_{t}S^{-1})=I_{n}$. So the first column vector $s_{1}^{-1}(S{\bf c}_{1;t})$ of $SC_{t}S^{-1}$ is an inner normal vector of the facet $\mathcal{C}$ of the cone $\mathbb{R}_{\geq 0}G_{t}$. So when we go along the direction $s_{1}^{-1}(S{\bf c}_{1;t})$, we can cross the common facet $\mathcal{C}$ from the cone $\mathbb{R}_{\geq 0}G_{t^{\prime}}$ to the cone $\mathbb{R}_{\geq 0}G_{t}$.
Now we take a vector ${\bf h}^{\prime}\in\mathbb{Z}_{\geq 1}G_{t^{\prime}}\subseteq\mathbb{R}_{\geq 0}G_{t^{\prime}}$. Then there exists $\lambda\in\mathbb{Q}_{>0}$ such that
${\bf h}\coloneqq{\bf h}^{\prime}+\lambda(s_{1}^{-1}(S{\bf c}_{1;t}))\in\mathbb{Q}_{>0}G_{t}\subseteq\mathbb{R}_{\geq 0}G_{t}.$
We can take a positive integer $p$ large enough such that $p{\bf h}\in\mathbb{Z}_{\geq 1}G_{t}$. Since $p{\bf h}\in\mathbb{Z}_{\geq 1}G_{t}\subseteq\mathbb{Z}^{n}$ and $p{\bf h}^{\prime}\in\mathbb{Z}_{\geq 1}G_{t^{\prime}}\subseteq\mathbb{Z}^{n}$, we have $p\lambda s_{1}^{-1}(S{\bf c}_{1;t})=p{\bf h}-p{\bf h}^{\prime}\in\mathbb{Z}^{n}$. Since $({\bf x}_{t^{\prime}},B_{t^{\prime}})=\mu_{k}({\bf x}_{t},B_{t})$ is a green mutation, we have ${\bf c}_{1;t}\in\mathbb{Z}_{\geq 0}^{n}$. So we have $p{\bf h}-p{\bf h}^{\prime}=p\lambda s_{1}^{-1}(S{\bf c}_{1;t})\in\mathbb{Z}_{\geq 0}^{n}$.
Since $p{\bf h}\in\mathbb{Z}_{\geq 1}G_{t}$ and $p{\bf h}^{\prime}\in\mathbb{Z}_{\geq 1}G_{t^{\prime}}$, there exist ${\bf v},{\bf v}^{\prime}\in\mathbb{Z}_{\geq 1}^{n}$ such that
$p{\bf h}=G_{t}{\bf v}\;\;\;\text{and}\;\;\;p{\bf h}^{\prime}=G_{t^{\prime}}{\bf v}^{\prime}.$
Thus $G_{t}{\bf v}-G_{t^{\prime}}{\bf v}^{\prime}=p{\bf h}-p{\bf h}^{\prime}\in\mathbb{Z}_{\geq 0}^{n}$. ∎
Lemma 7.7.
Let $F$ be a polynomial in $\mathbb{Z}[y_{1},\ldots,y_{n}]$ with constant term $1$. If $({\bf r}-{\bf r}^{\prime})\in\mathbb{Z}_{\geq 0}^{n}$ for some ${\bf r},{\bf r}^{\prime}\in\mathbb{Z}^{n}$ and $F[{\bf r}]=0$, then $F[{\bf r}^{\prime}]=0$.
Proof.
Let ${\bf y}^{\bf u}$ be a monomial appearing in $F$. We know that ${\bf u}\in\mathbb{Z}_{\geq 0}^{n}$. Since $({\bf r}-{\bf r}^{\prime})\in\mathbb{Z}_{\geq 0}^{n}$, we have ${\bf u}^{T}({\bf r}^{\prime}-{\bf r})=-{\bf u}^{T}({\bf r}-{\bf r}^{\prime})\leq 0$. Since $F[{\bf r}]=0$, we have ${\bf u}^{T}{\bf r}\leq 0$. Thus
${\bf u}^{T}{\bf r}^{\prime}={\bf u}^{T}({\bf r}^{\prime}-{\bf r}+{\bf r})={\bf u}^{T}({\bf r}^{\prime}-{\bf r})+{\bf u}^{T}{\bf r}\leq 0.$
Since $F$ has constant term $1$ and by the definition of $F[{\bf r}^{\prime}]$, we obtain $F[{\bf r}^{\prime}]=0$. ∎
Theorem 7.8.
If $({\bf x}_{t^{\prime}},B_{t^{\prime}})=\mu_{k}({\bf x}_{t},B_{t})$ is a green mutation, then $\operatorname{\mathsf{dom}}^{t_{0}}[t]\subsetneq\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$. In particular, the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}=\overrightarrow{\mathcal{H}}^{t_{0}}(\mathcal{A})$ of $\mathcal{A}$ is acyclic.
Proof.
Since $({\bf x}_{t^{\prime}},B_{t^{\prime}})=\mu_{k}({\bf x}_{t},B_{t})$ is a green mutation and by Lemma 7.6, there exist
${\bf v}=(v_{1},\ldots,v_{n})^{T},\;{\bf v}^{\prime}=(v_{1}^{\prime},\ldots,v_{n}^{\prime})^{T}\in\mathbb{Z}_{\geq 1}^{n}$
such that $G_{t}{\bf v}-G_{t^{\prime}}{\bf v}^{\prime}\in\mathbb{Z}_{\geq 0}^{n}$. Set $u=\prod_{i=1}^{n}x_{i;t}^{v_{i}}$ and $u^{\prime}=\prod_{i=1}^{n}x_{i;t^{\prime}}^{v_{i}^{\prime}}$. We know that
${\bf g}_{u}^{t_{0}}=G_{t}{\bf v},\;\;{\bf g}_{u^{\prime}}^{t_{0}}=G_{t^{\prime}}{\bf v}^{\prime}.$
Since $\operatorname{\mathsf{supp}}(u)=[{\bf x}_{t}],\;\operatorname{\mathsf{supp}}(u^{\prime})=[{\bf x}_{t^{\prime}}]$ and by Corollary 7.5, we have that
$\operatorname{\mathsf{dom}}^{t_{0}}[t]$ $=\operatorname{\mathsf{dom}}^{t_{0}}(u)=\{z\in\mathcal{X}\mid F_{z}^{t_{0}}[S{\bf g}_{u}^{t_{0}}]=0\}=\{z\in\mathcal{X}\mid F_{z}^{t_{0}}[SG_{t}{\bf v}]=0\},$
$\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$ $=\operatorname{\mathsf{dom}}^{t_{0}}(u^{\prime})=\{z\in\mathcal{X}\mid F_{z}^{t_{0}}[S{\bf g}_{u^{\prime}}^{t_{0}}]=0\}=\{z\in\mathcal{X}\mid F_{z}^{t_{0}}[SG_{t^{\prime}}{\bf v}^{\prime}]=0\}.$
Let $z\in\operatorname{\mathsf{dom}}^{t_{0}}[t]$, that is, we have $F_{z}^{t_{0}}[SG_{t}{\bf v}]=0$. Since
$SG_{t}{\bf v}-SG_{t^{\prime}}{\bf v}^{\prime}=S(G_{t}{\bf v}-G_{t^{\prime}}{\bf v}^{\prime})\in\mathbb{Z}_{\geq 0}^{n}$
and by Lemma 7.7, we get $F_{z}^{t_{0}}[SG_{t^{\prime}}{\bf v}^{\prime}]=0$. This implies that $z\in\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$. So we have $\operatorname{\mathsf{dom}}^{t_{0}}[t]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$.
In the rest part of the proof, we will show $x_{k;t^{\prime}}\in\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$ but $x_{k;t^{\prime}}\notin\operatorname{\mathsf{dom}}^{t_{0}}[t]$. Thus the inclusion $\operatorname{\mathsf{dom}}^{t_{0}}[t]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$ is strict.
By Proposition 7.4 (ii), we have ${\bf x}_{t}]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t]$ and ${\bf x}_{t^{\prime}}]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$. In particular,
$x_{k;t}\in\operatorname{\mathsf{dom}}^{t_{0}}[t]\;\;\;\text{and}\;\;\;x_{k;t^{\prime}}\in\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}].$
Since $\operatorname{\mathsf{dom}}^{t_{0}}[t]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$, we get $x_{k;t}\in\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$. Then by Proposition 7.4 (iii), we have $x_{k;t}\in\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}(x_{k;t^{\prime}})$. This implies
$F_{x_{k;t}}^{t_{0}}[S{\bf g}_{x_{k;t^{\prime}}}^{t_{0}}]=0.$
By using the vertex $t_{0}$ to calculate the $F$-invariant $(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}$, we have
$(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}=F_{x_{k;t}}^{t_{0}}[S{\bf g}_{x_{k;t}^{\prime}}^{t_{0}}]+F_{x_{k;t^{\prime}}}^{t_{0}}[S{\bf g}_{x_{k;t}}^{t_{0}}]=F_{x_{k;t^{\prime}}}^{t_{0}}[S{\bf g}_{x_{k;t}}^{t_{0}}].$
On the other hand, if we use the vertex $t$ to calculate $(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}$, we see that
$(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}=0+F_{x_{k;t^{\prime}}}^{t}[S{\bf g}_{x_{k;t}}^{t}]=(1+y_{k})[S{\bf e}_{k}]=s_{k},$
where ${\bf e}_{k}$ is the $k$th column of $I_{n}$ and $s_{k}$ is the $(k,k)$-entry of $S=diag(s_{1},\ldots,s_{n})$. Thus
$F_{x_{k;t^{\prime}}}^{t_{0}}[S{\bf g}_{x_{k;t}}^{t_{0}}]=(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}=s_{k}>0,$
which means $x_{k;t^{\prime}}\notin\operatorname{\mathsf{dom}}^{t_{0}}(x_{k;t})$. By Proposition 7.4 (iii), we know that $\operatorname{\mathsf{dom}}^{t_{0}}[t]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}(x_{k;t})$. Thus $x_{k;t^{\prime}}\notin\operatorname{\mathsf{dom}}^{t_{0}}(x_{k;t})$ implies that $x_{k;t^{\prime}}\notin\operatorname{\mathsf{dom}}^{t_{0}}[t]$.
Since $x_{k;t^{\prime}}\in\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$ but $x_{k;t^{\prime}}\notin\operatorname{\mathsf{dom}}^{t_{0}}[t]$, we know that the inclusion $\operatorname{\mathsf{dom}}^{t_{0}}[t]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$ is strict, i.e., we have $\operatorname{\mathsf{dom}}^{t_{0}}[t]\subsetneq\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$. ∎
Example 7.9.
Take $B=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$, ${\bf x}=(x_{1},x_{2})$ and $S=I_{2}$. The oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}$ of $\mathcal{A}=\mathcal{A}({\bf x},B)$ is as follows:
${t_{0}=\{x_{1},x_{2}\}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{x_{1}}}$$\scriptstyle{\mu_{x_{2}}}$${t_{1}=\{{x_{3}},x_{2}\}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{x_{2}}}$${t_{4}=\{x_{5},x_{1}\}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{x_{1}}}$${t_{2}=\{x_{3},{x_{4}}\}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{x_{3}}}$${t_{3}=\{{x_{5}},x_{4}\}}$
where $x_{3}=x_{1}^{-1}x_{2}\cdot(1+\widehat{y}_{1}),x_{4}=x_{1}^{-1}\cdot(1+\widehat{y}_{1}+\widehat{y}_{1}\widehat{y}_{2}),x_{5}=x_{2}^{-1}\cdot(1+\widehat{y}_{2})$. It is easy to check that
$\operatorname{\mathsf{dom}}^{t_{0}}(x_{1})=\{x_{1},x_{2},x_{5}\},\;\;\;\operatorname{\mathsf{dom}}^{t_{0}}(x_{2})=\{x_{1},x_{2},x_{3}\},\;\;\;\operatorname{\mathsf{dom}}^{t_{0}}(x_{3})=\{x_{1},x_{2},x_{3},x_{4}\},$
$\operatorname{\mathsf{dom}}^{t_{0}}(x_{4})=\{x_{1},x_{2},x_{3},x_{4},x_{5}\}=\operatorname{\mathsf{dom}}^{t_{0}}(x_{5}).$
Then by Proposition 7.4 (iii), we can get that
$\operatorname{\mathsf{dom}}^{t_{0}}[t_{0}]=\{x_{1},x_{2}\},\;\;\operatorname{\mathsf{dom}}^{t_{0}}[t_{1}]=\{x_{1},x_{2},x_{3}\},\;\;\operatorname{\mathsf{dom}}^{t_{0}}[t_{2}]=\{x_{1},x_{2},x_{3},x_{4}\},$
$\operatorname{\mathsf{dom}}^{t_{0}}[t_{3}]=\{x_{1},x_{2},x_{3},x_{4},x_{5}\},\;\;\operatorname{\mathsf{dom}}^{t_{0}}[t_{4}]=\{x_{1},x_{2},x_{5}\}.$
We can see that the initial cluster variables $x_{1},x_{2}$ are contained in any dominant set with respect to the vertex $t_{0}$, this is because the $F$-polynomials of initial cluster variables are $1$.
Let $\mathcal{H}_{0}$ be the set of seeds (up to seed equivalence) of $\mathcal{A}$, equivalently, $\mathcal{H}_{0}$ is the vertex set of the exchange graph of $\mathcal{A}$. For the rooted vertex $t_{0}\in\mathbb{T}_{n}$, we define an relation $\leq_{t_{0}}$ on $\mathcal{H}_{0}$ as follows: ${\bf x}_{w},B_{w}]\leq_{t_{0}}[{\bf x}_{t},B_{t}]$ if and only if there exists a sequence $\overleftarrow{\mu}$ of green mutations such that
$\overleftarrow{\mu}({\bf x}_{w},B_{w})]=[{\bf x}_{t},B_{t}],$
i.e., there exists a path from the vertex ${\bf x}_{w},B_{w}]$ to the vertex ${\bf x}_{t},B_{t}]$ in the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}=\overrightarrow{\mathcal{H}}^{t_{0}}(\mathcal{A})$ of $\mathcal{A}$.
If ${\bf x}_{w},B_{w}]\leq_{t_{0}}[{\bf x}_{t},B_{t}]$ and ${\bf x}_{t},B_{t}]\neq[{\bf x}_{w},B_{w}]$, we write ${\bf x}_{w},B_{w}]<_{t_{0}}[{\bf x}_{t},B_{t}]$.
Corollary 7.10.
The relation $\leq_{t_{0}}$ gives a partial order on $\mathcal{H}_{0}$, i.e., $(\mathcal{H}_{0},\leq_{t_{0}})$ is a poset.
Proof.
By Theorem 7.8, the oriented exchange graph of $\mathcal{A}$ is acyclic. The result follows. ∎
[…]
7.2. Oriented exchange graphs are Hasse quivers
[…]
Proposition 7.11.
Let $({\bf x}_{t},B_{t})$ be a seed of $\mathcal{A}$. Denote by
$\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])$ $=\{x\in\operatorname{\mathsf{dom}}^{t_{0}}[t]\mid F_{z}^{t_{0}}[S{\bf g}_{x}^{t_{0}}]=0,\;\forall z\in\operatorname{\mathsf{dom}}^{t_{0}}[t]\}$
$=\{x\in\operatorname{\mathsf{dom}}^{t_{0}}[t]\mid\operatorname{\mathsf{dom}}^{t_{0}}[t]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}(x)\}.$
Then $\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])=[{\bf x}_{t}]$ and in particular, ${\bf x}_{t}]$ is uniquely determined by the dominant set $\operatorname{\mathsf{dom}}^{t_{0}}[t]$.
Proof.
Fix $k\in[1,n]$. By Proposition 7.4 (ii), (iii), we know that
$x_{k;t}\in\operatorname{\mathsf{dom}}^{t_{0}}[t]\;\;\;\text{and}\;\;\;\operatorname{\mathsf{dom}}^{t_{0}}[t]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}(x_{k;t}).$
So we have $x_{k;t}\in\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])$. Since $k$ is an arbitrary integer $[1,n]$, we get ${\bf x}_{t}]\subseteq\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])$.
Now let us take $x\in\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])$. We know that $F_{z}^{t_{0}}[S{\bf g}_{x}^{t_{0}}]=0$ for any $z\in\operatorname{\mathsf{dom}}^{t_{0}}[t]$. Since $x_{k;t}$ is in $\operatorname{\mathsf{dom}}^{t_{0}}[t]$ for $k\in[1,n]$, we have $F_{x_{k;t}}^{t_{0}}[S{\bf g}_{x}^{t_{0}}]=0$ for $k\in[1,n]$.
Since $x\in\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}(x_{k;t})$, we have $F_{x}^{t_{0}}[S{\bf g}_{x_{k;t}}^{t_{0}}]=0$. Thus
$(x_{k;t}\mid\mid x)_{F}=F_{x_{k;t}}^{t_{0}}[S{\bf g}_{x}^{t_{0}}]+F_{x}^{t_{0}}[S{\bf g}_{x_{k;t}}^{t_{0}}]=0.$
Thus for any two cluster variables $z_{1},z_{2}$ in ${\bf x}_{t}]\cup\{x\}$, we have $(z_{1}\mid\mid z_{2})_{F}=0$. Then by Lemma 4.18, we know that ${\bf x}_{t}]\cup\{x\}$ is a subset of some cluster of $\mathcal{A}$. We must have $x=x_{i;t}$ for some $i\in[1,n]$. Thus we get
$\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])\subseteq[{\bf x}_{t}].$
Hence, $\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])=[{\bf x}_{t}]$. ∎
The following result is analogous to [air_2014, Proposition 2.26] in $\tau$-tilting theory.
Lemma 7.12.
Let $({\bf x}_{t},B_{t})$, $({\bf x}_{s},B_{s})$ and $({\bf x}_{w},B_{w})$ be three seeds of $\mathcal{A}$. If
$\operatorname{\mathsf{dom}}^{t_{0}}[w]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[s]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t],$
then ${\bf x}_{t}]\cap[{\bf x}_{w}]\subseteq[{\bf x}_{s}]$.
Proof.
By the definition of $\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[s])$, we can see $\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])\cap\operatorname{\mathsf{dom}}^{t_{0}}[s]\subseteq\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[s])$. By Proposition 7.11, we know that ${\bf x}_{t}]=\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])$ and ${\bf x}_{s}]=\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[s])$. Thus we have ${\bf x}_{t}]\cap\operatorname{\mathsf{dom}}^{t_{0}}[s]\subseteq[{\bf x}_{s}]$.
By Proposition 7.4 (ii), we know that ${\bf x}_{w}]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[w]$. Then by $\operatorname{\mathsf{dom}}^{t_{0}}[w]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[s]$, we have ${\bf x}_{w}]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[s]$. So we know that ${\bf x}_{t}]\cap[{\bf x}_{w}]\subseteq[{\bf x}_{t}]\cap\operatorname{\mathsf{dom}}^{t_{0}}[s]\subseteq[{\bf x}_{s}]$. ∎
Theorem 7.13.
The following statements are equivalent:
  • (a)
    There exists an arrow ${\bf x}_{w},B_{w}]\rightarrow[{\bf x}_{t},B_{t}]$ in the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}$ of $\mathcal{A}$.
  • (b)
    ${\bf x}_{w},B_{w}]<_{t_{0}}[{\bf x}_{t},B_{t}]$ and there exists no seed $({\bf x}_{s},B_{s})$ such that
    ${\bf x}_{w},B_{w}]<_{t_{0}}[{\bf x}_{s},B_{s}]<_{t_{0}}[{\bf x}_{t},B_{t}].$
Proof.
$(a)\Rightarrow(b)$": Suppose that ${\bf x}_{w},B_{w}]\rightarrow[{\bf x}_{t},B_{t}]$ is an arrow in the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}$ of $\mathcal{A}$. Then we know that ${\bf x}_{w},B_{w}]<_{t_{0}}[{\bf x}_{t},B_{t}]$.
Since the two seeds $({\bf x}_{t},B_{t})$ and $({\bf x}_{w},B_{w})$ differ by once mutation, for simplicity, we can assume $({\bf x}_{t},B_{t})=\mu_{1}({\bf x}_{w},B_{w})$. If we write ${\bf x}_{w}=(z_{1},z_{2},\ldots,z_{n})$, then ${\bf x}_{t}$ has the form ${\bf x}_{t}=(z_{1}^{\prime},z_{2},\ldots,z_{n})$.
If there exists a seed $({\bf x}_{s},B_{s})$ such that ${\bf x}_{w},B_{w}]<_{t_{0}}[{\bf x}_{s},B_{s}]\leq_{t_{0}}[{\bf x}_{t},B_{t}]$. Then we have $\operatorname{\mathsf{dom}}^{t_{0}}[w]\subsetneq\operatorname{\mathsf{dom}}^{t_{0}}[s]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t]$. By Lemma 7.12, we have $\{z_{2},\ldots,z_{m}\}=[{\bf x}_{t}]\cap[{\bf x}_{w}]\subseteq[{\bf x}_{s}]$. Since ${\bf x}_{w},B_{w}]<_{t_{0}}[{\bf x}_{s},B_{s}]$, we know that ${\bf x}_{s}]\neq[{\bf x}_{w}]$. Thus the two clusters ${\bf x}_{s}$ and ${\bf x}_{w}$ have exactly n1n-1 common cluster variables $z_{2},\ldots,z_{n}$. Then by [cao-li-2020, Corollary 3], we get ${\bf x}_{s},B_{s}]=[\mu_{1}({\bf x}_{w},B_{w})]=[{\bf x}_{t},B_{t}]$. So there exists no seed $({\bf x}_{s},B_{s})$ such that ${\bf x}_{w},B_{w}]<_{t_{0}}[{\bf x}_{s},B_{s}]<_{t_{0}}[{\bf x}_{t},B_{t}]$.
$(b)\Rightarrow(a)$": Since ${\bf x}_{w},B_{w}]<_{t_{0}}[{\bf x}_{t},B_{t}]$, there exists a non-empty sequence $\overleftarrow{\mu}$ of green mutations from ${\bf x}_{w},B_{w}]$ to ${\bf x}_{t},B_{t}]$. Since there exists no seed $({\bf x}_{s},B_{s})$ such that
${\bf x}_{w},B_{w}]<_{t_{0}}[{\bf x}_{s},B_{s}]<_{t_{0}}[{\bf x}_{t},B_{t}],$
the length of $\overleftarrow{\mu}$ must be $1$, i.e., ${\bf x}_{t},B_{t}]$ is obtained from ${\bf x}_{w},B_{w}]$ by once green mutation. So there exists an arrow $({\bf x}_{w},B_{w})]\rightarrow[({\bf x}_{t},B_{t})]$ in the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}$ of $\mathcal{A}$. ∎
The above theorem rules out the existence of the subquiver $\widetilde{A}_{1,d}$ ($d\geq 2$) of the form
[…]
in the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}$ of $\mathcal{A}$.
The following result is a direct consequence of the above theorem and it is analogous to the corresponding result [air_2014, Theorem 0.6] in the $\tau$-tilting theory.
Corollary 7.14.
[…]
7.3. Dominant sets and torsion classes
In this subsection we briefly explain why the dominant sets can be viewed as a replacement of torsion classes in $\tau$-tilting theory.
We fix a finite dimensional basic algebra $A$ over a field $\mathbf{k}$. Denote by $\operatorname{\mathsf{mod}}\nolimits A$ the category of finitely generated left $A$-modules, and by $\tau$ the Auslander-Reiten translation in $\operatorname{\mathsf{mod}}\nolimits A$. The isomorphism classes of indecomposable projective modules in $\operatorname{\mathsf{mod}}\nolimits A$ are denoted by $P_{1},\ldots,P_{n}$.
Given a module $M\in\operatorname{\mathsf{mod}}\nolimits A$, we denote by
  • $\operatorname{\mathsf{add}}\nolimits M$ the subcategory of $\operatorname{\mathsf{mod}}\nolimits A$ consisting of modules which are direct summands of $M^{\oplus p}$ for some $p>0$.
  • $\operatorname{\mathsf{Fac}}M$ the factor modules of the modules in $\operatorname{\mathsf{add}}\nolimits M$.
  • $\prescript{\bot}{}{M}\coloneqq\{X\in\operatorname{\mathsf{mod}}\nolimits A\mid\operatorname{\mathsf{Hom}}_{A}(X,M)=0\}$.
  • $M^{\bot}\coloneqq\{Y\in\operatorname{\mathsf{mod}}\nolimits A\mid\operatorname{\mathsf{Hom}}_{A}(M,Y)=0\}$.
  • $|M|$ the number of non-isomorphic indecomposable direct summands of $M$.
Definition 7.15.
Let $M$ be a module in $\operatorname{\mathsf{mod}}\nolimits A$ and $P$ a projective module in $\operatorname{\mathsf{mod}}\nolimits A$.
  • (i)
    $M$ is called $\tau$-rigid if $\operatorname{\mathsf{Hom}}_{A}(M,\tau M)=0$.
  • (ii)
    The pair $(M,P)$ is called $\tau$-rigid if $M$ is $\tau$-rigid and $\operatorname{\mathsf{Hom}}_{A}(P,M)=0$.
  • (iii)
    The pair $(M,P)$ is called $\tau$-tilting if $(M,P)$ is a $\tau$-rigid pair and $|M|+|P|=|A|=n$.
We will always consider modules and $\tau$-rigid pairs up to isomorphisms. A $\tau$-rigid pair $(M,P)$ is indecomposable if $M\oplus P$ is indecomposable in $\operatorname{\mathsf{mod}}\nolimits A$ and it is basic if both $M$ and $P$ are basic in $\operatorname{\mathsf{mod}}\nolimits A$. In a (basic) $\tau$-tilting pair $(M,P)$, it is known [air_2014, Proposition 2.3] that $P$ is uniquely determined by $M$.
A subcategory $\mathcal{T}$ of $\operatorname{\mathsf{mod}}\nolimits A$ is called a torsion class, if $\mathcal{T}$ is closed under quotients and extensions. A torsion class $\mathcal{T}$ is said to be functorially finite, if there exists a module $M\in\operatorname{\mathsf{mod}}\nolimits A$ such that $\mathcal{T}=\operatorname{\mathsf{Fac}}M$.
Let $\mathcal{C}$ be a subcategory of $\operatorname{\mathsf{mod}}\nolimits A$. A module $U\in\mathcal{C}$ is said to be Ext-projective in $\mathcal{C}$, if $\operatorname{\mathsf{Ext}}_{A}^{1}(U,\mathcal{C})=0$. We denote by $\mathcal{P}(\mathcal{C})$ the direct sum of one copy of each of the indecomposable Ext-projective objects in $\mathcal{C}$ up to isomorphisms.
Theorem 7.16 ([air_2014, Prop. 1.2 (b), Thm. 2.7]).
The following statements hold.
  • (i)
    There is a well-defined map $\Psi$ from $\tau$-rigid pairs to functorially finite torsion classes in $\operatorname{\mathsf{mod}}\nolimits A$ given by $(M,P)\mapsto\operatorname{\mathsf{Fac}}M$.
  • (ii)
    The above map $\Psi$ is a bijection if we restrict it to basic $\tau$-tilting pairs, which induces a natural order on the set of basic $\tau$-tilting pairs
    $(M,P)\leq(M^{\prime},P^{\prime})\overset{{\rm def.}}{\Longleftrightarrow}\operatorname{\mathsf{Fac}}M\subseteq\operatorname{\mathsf{Fac}}M^{\prime}.$
  • (iii)
    Let $\mathcal{T}$ be a functorially finite torsion class and denote by $(M,P)$ the basic $\tau$-tilting pair such that $\operatorname{\mathsf{Fac}}M=\mathcal{T}$. Then $M=\mathcal{P}(\mathcal{T})$.
One can refer to [air_2014, Theorem 2.19] for the definition of mutations of $\tau$-tilting pairs. Adachi, Iyama and Reiten [air_2014, Theorem 0.6] proved that mutations of $\tau$-tilting pairs correspond to the cover relations in the poset of $\tau$-tilting pairs (or functorially finite torsion classes). Some correspondences between cluster algebras and $\tau$-tilting theory are given in Table 1.
[…]
Table 1. Cluster algebras vs. $\tau$-tilting theory
Consider $M\in\operatorname{\mathsf{mod}}\nolimits A$ and let
$\bigoplus_{i=1}^{n}P_{i}^{b_{i}}\rightarrow\bigoplus_{i=1}^{n}P_{i}^{a_{i}}\rightarrow M\rightarrow 0$
be the minimal projective presentation of $M$ in $\operatorname{\mathsf{mod}}\nolimits A$. The vector
$\delta_{M}\coloneqq(a_{1}-b_{1},\ldots,a_{n}-b_{n})^{T}\in\mathbb{Z}^{n}$
is called the $\delta$-vector of $M$ and the vector ${\bf g}_{M}\coloneqq-\delta_{M}$ is called the $g$-vector of $M$.
For a $\tau$-rigid pair $(M,P)$, we define its $\delta$-vector and $g$-vector as follows:
$\delta_{(M,P)}\coloneqq\delta_{M}-\delta_{P},\;\;{\bf g}_{(M,P)}\coloneqq-\delta_{(M,P)}.$
With this definition, we can see that the $g$-vector ${\bf g}_{(0,P_{k})}$ of $(0,P_{k})$ is the $k$th column of $I_{n}$.
Remark 7.17.
Notice that the $\delta$-vectors defined here coincide with the $g$-vectors used in [air_2014]. For the considerations on the cluster algebra side, $g$-vectors defined here are the negative of the $\delta$-vectors.
Definition 7.18.
The $F$-polynomial $F_{M}$ of $M\in\operatorname{\mathsf{mod}}\nolimits A$ is defined to be
$F_{M}=\sum_{{\bf v}\in\mathbb{N}^{n}}\chi(\operatorname{\mathsf{Gr}}\nolimits_{\bf v}(M)){\bf y}^{\bf v}\in\mathbb{Z}[y_{1},\ldots,y_{n}],$
where $\operatorname{\mathsf{Gr}}\nolimits_{\bf v}(M)$ is the quotient module Grassmannian of $M$ with dimension vector $\bf v$} and $\chi$ is the Euler–Poincaré characteristic.
Since the zero module is a quotient module of $M$, the polynomial $F_{M}$ has constant term $1$. So we have
$F_{M}[{\bf r}]=\max\{{\bf v}^{T}{\bf r}\mid\chi(\operatorname{\mathsf{Gr}}\nolimits_{\bf v}(M))\neq 0\}\geq 0,\quad\forall{\bf r}\in\mathbb{Z}^{n}.$
For a $\tau$-rigid pair $(M,P)$, we define its $F$-polynomial as follows:
$F_{(M,P)}\coloneqq F_{M}\in\mathbb{Z}[y_{1},\ldots,y_{n}].$
In particular, we have $F_{(0,P_{k})}=1$ for $k=1,\ldots,n$.
Remark 7.19.
Notice that in Section 6 we define $F$-polynomial using submodules (see Definition 6.2), while in this section we define $F$-polynomial using quotient modules. The reason is that in this section we want to work with torsion classes (corresponding to $\tau$-tilting theory) rather than torsion-free classes (corresponding to $\tau^{-1}$-tilting theory). So the algebras/quivers in Section 6 correspond to the opposite algebras/quivers in this section. For example, in Section 6 we use the representations of the quiver $1\leftarrow 2$ to model the cluster algebra $\mathcal{A}$ with initial mutation matrix $B=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$, while in this section we would use the representations of the quiver $1\rightarrow 2$ to model $\mathcal{A}$.
Denote by $\mathcal{X}_{+}$ the set of indecomposable $\tau$-rigid modules in $\operatorname{\mathsf{mod}}\nolimits A$ and $\mathcal{X}$ the set of indecomposable $\tau$-rigid pairs in $\operatorname{\mathsf{mod}}\nolimits A$. We have
$\mathcal{X}=\mathcal{X}_{+}\sqcup\{(0,P_{1}),\ldots,(0,P_{n})\}.$
When we write $U\in\mathcal{X}$, we mean that $U$ is an indecomposable $\tau$-rigid pair, that is, either $U\in\mathcal{X}_{+}$ or $U=(0,P_{k})$ for some $k\in[1,n]$.
Following the definition of the dominant sets in cluster algebras, we define the dominant sets in $\tau$-tilting theory.
Definition 7.20.
Let $(M,P)$ be a $\tau$-tilting pair in $\operatorname{\mathsf{mod}}\nolimits A$. The dominant set $\operatorname{\mathsf{dom}}[M,P]$ and the reduced dominant set $\operatorname{\mathsf{dom}}_{+}[M,P]$ of $(M,P)$ are defined as follows:
$\operatorname{\mathsf{dom}}[M,P]$ $\coloneqq\{U\in\mathcal{X}\mid F_{U}[{\bf g}_{(M,P)}]=0\},$
$\operatorname{\mathsf{dom}}_{+}[M,P]$ $\coloneqq\{U\in\mathcal{X}_{+}\mid F_{U}[{\bf g}_{(M,P)}]=0\}.$
Since $F_{(0,P_{k})}=1$ for $k=1,\ldots,n$, the set $\{(0,P_{1}),\ldots,(0,P_{n})\}$ is contained in the dominant set $\operatorname{\mathsf{dom}}[M,P]$ for any $\tau$-tilting pair $(M,P)$. Thus we have
(7.1) $\operatorname{\mathsf{dom}}[M,P]=\operatorname{\mathsf{dom}}_{+}[M,P]\sqcup\{(0,P_{1}),\ldots,(0,P_{n})\}.$
For a module $X\in\operatorname{\mathsf{mod}}\nolimits A$ and a vector $\delta\in\mathbb{R}^{n}$, we write $\delta(X)=\langle\delta,\underline{\dim}X\rangle\in\mathbb{R}$ for the inner product of $\delta$ and the dimension vector of $X$.
Lemma 7.21 ([BKT-2014]*Prop. 3.1, [asai-2021]*Prop. 3.11).
Let $\delta\in\mathbb{R}^{n}$ and set
$\overline{\mathcal{T}}_{\delta}:=\{N\in\operatorname{\mathsf{mod}}\nolimits A\mid\text{for any quotient module X of }N,\;\;\delta(X)\geq 0\}.$
The following statements hold.
  • (i)
    The subcategory $\overline{\mathcal{T}}_{\delta}$ is a torsion class in $\operatorname{\mathsf{mod}}\nolimits A$.
  • (ii)
    Let $(M,P)$ be a basic $\tau$-tilting pair in $\operatorname{\mathsf{mod}}\nolimits A$ and $\delta=\delta_{(M,P)}$ the $\delta$-vector of $(M,P)$. Then $\operatorname{\mathsf{Fac}}M=\overline{\mathcal{T}}_{\delta}$.
Torsion classes of the form $\overline{\mathcal{T}}_{\delta}$ are called semistable torsion classes. By Lemma 7.21 (ii), all functorially finite torsion classes are semistable torsion classes.
Lemma 7.22 ([fei_2019a, Theorem 1.4]).
Let $U\in\operatorname{\mathsf{mod}}\nolimits A$ and let
$F_{U}=\sum_{{\bf v}\in\mathbb{N}^{n}}\chi(\operatorname{\mathsf{Gr}}\nolimits_{\bf v}(U)){\bf y}^{\bf v}\in\mathbb{Z}[y_{1},\ldots,y_{n}]$
be the $F$-polynomial of $U$. Then the convex hull $\mathcal{N}(F_{U})$ of $\{{\bf v}\in\mathbb{N}^{n}\mid\chi(\operatorname{\mathsf{Gr}}\nolimits_{\bf v}(U))\neq 0\}$ and the convex hull $\mathcal{N}(U)$ of the dimension vectors of the quotient modules of $U$ are the same.
Proof.
The corresponding result for submodules and $F$-polynomials defined using submodules is proved by Fei in [fei_2019a, Theorem 1.4]. The required result follows from Fei’s result by considering the modules over the opposite algebras. ∎
Proposition 7.23.
Let $(M,P)$ be a $\tau$-tilting pair in $\operatorname{\mathsf{mod}}\nolimits A$. Then we have
  • (i)
    $\operatorname{\mathsf{dom}}_{+}[M,P]=\mathcal{X}_{+}\cap\operatorname{\mathsf{Fac}}M=\{\text{indecomposable }\tau\text{-rigid modules in }\operatorname{\mathsf{Fac}}M\}$.
  • (ii)
    $\operatorname{\mathsf{dom}}[M,P]=\{\text{indecomposable }\tau\text{-rigid modules in }\operatorname{\mathsf{Fac}}M\}\sqcup\{(0,P_{1}),\ldots,(0,P_{n})\}$.
Proof.
(i) Let $\delta=\delta_{(M,P)}$ be the $\delta$-vector of $(M,P)$ and ${\bf g}={\bf g}_{(M,P)}=-\delta$ the $g$-vector of $(M,P)$. Let $U$ be an indecomposable $\tau$-rigid module in $\operatorname{\mathsf{mod}}\nolimits A$ and
$F_{U}=\sum_{{\bf v}\in\mathbb{N}^{n}}\chi(\operatorname{\mathsf{Gr}}\nolimits_{\bf v}(U)){\bf y}^{\bf v}\in\mathbb{Z}[y_{1},\ldots,y_{n}]$
the $F$-polynomial of $U$. We know that $U\in\operatorname{\mathsf{dom}}_{+}[M,P]$ if and only if
$F_{U}[{\bf g}]=\max\{{\bf v}^{T}{\bf g}\mid\chi(\operatorname{\mathsf{Gr}}\nolimits_{\bf v}(U))\neq 0\}=0.$
Thanks to Lemma 7.22, this is equivalent to
$\max\{\langle{\bf g},\underline{\dim}X\rangle\mid X\text{ is a quotient module of }U\}=0,$
equivalently, $\langle{\bf g},\underline{\dim}X\rangle\leq 0$ for any quotient module $X$ of $U$. This is also equivalent to $\delta(X)=-\langle{\bf g},\underline{\dim}X\rangle\geq 0$ for any quotient module $X$ of $U$. Then by Lemma 7.21, this is equivalent to $U\in\operatorname{\mathsf{Fac}}M$. Hence, we have
$\operatorname{\mathsf{dom}}_{+}[M,P]=\mathcal{X}_{+}\cap\operatorname{\mathsf{Fac}}M=\{\text{indecomposable $\tau$-rigid modules in $\operatorname{\mathsf{Fac}}M$}\}.$
(ii) This follows from (i) and the equality (7.1). ∎
Since $\operatorname{\mathsf{Fac}}M$ is the smallest torsion class containing all the indecomposable $\tau$-rigid modules in $\operatorname{\mathsf{Fac}}M$, we know that $\operatorname{\mathsf{Fac}}M$ is uniquely determined by the dominant set $\operatorname{\mathsf{dom}}[M,P]$. Thus we can view the dominant sets as a kind of replacement of torsion classes in $\tau$-tilting theory.
Example 7.24.
Let $A$ be the path algebra of the quiver $1\rightarrow 2$. We have the exact sequence
$0\rightarrow P_{2}\rightarrow P_{1}\rightarrow S_{1}\rightarrow 0$
in $\operatorname{\mathsf{mod}}\nolimits A=\operatorname{\mathsf{add}}\nolimits(P_{1}\oplus P_{2}\oplus S_{1})$. The Hasse quiver of the poset of functorially torsion classes in $\operatorname{\mathsf{mod}}\nolimits A$ is as follows:
${\mathcal{T}_{0}=\{0\}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{T}_{1}=\operatorname{\mathsf{add}}\nolimits(S_{1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{T}_{4}=\operatorname{\mathsf{add}}\nolimits(P_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{T}_{2}=\operatorname{\mathsf{add}}\nolimits(S_{1}\oplus P_{1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{T}_{3}=\operatorname{\mathsf{mod}}\nolimits A}$
The oriented exchange graph of $\tau$-tilting pairs is as follows:
${(0,P_{1}\oplus P_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{(0,P_{1})}}$$\scriptstyle{\mu_{(0,P_{2})}}$${(S_{1},P_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{(0,P_{2})}}$${(P_{2},P_{1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{(0,P_{1})}}$${(S_{1}\oplus P_{1},0)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{(S_{1},0)}}$${\mathcal{(}P_{1}\oplus P_{2},0)}$
The corresponding dominant sets are as follows:
$\operatorname{\mathsf{dom}}[0,P_{1}\oplus P_{2}]=\{(0,P_{1}),\;(0,P_{2})\},$
$\operatorname{\mathsf{dom}}[S_{1},P_{2}]=\{(0,P_{1}),\;(0,P_{2}),\;S_{1}\},$
$\operatorname{\mathsf{dom}}[S_{1}\oplus P_{1},0]=\{(0,P_{1}),\;(0,P_{2}),\;S_{1},\;P_{1}\},$
$\operatorname{\mathsf{dom}}[P_{1}\oplus P_{2},0]=\{(0,P_{1}),\;(0,P_{2}),\;S_{1},\;P_{1},\;P_{2}\},$
$\operatorname{\mathsf{dom}}[P_{2},P_{1}]=\{(0,P_{1}),\;(0,P_{2}),\;P_{2}\}.$
It is easy to check that
$F_{S_{1}}=1+y_{1},\;\;F_{P_{1}}=1+y_{1}+y_{1}y_{2},\;\;F_{P_{2}}=1+y_{2},$
${\bf g}_{S_{1}}=\begin{bmatrix}-1\\ 1\end{bmatrix},\;\;{\bf g}_{P_{1}}=\begin{bmatrix}-1\\ 0\end{bmatrix},\;\;{\bf g}_{P_{2}}=\begin{bmatrix}0\\ -1\end{bmatrix}$.
From the viewpoint of categorification of cluster algebras, the indecomposable $\tau$-rigid modules $S_{1},P_{1},P_{2}$ correspond to the non-initial cluster variables
$x_{3}={x_{1}^{-1}x_{2}\cdot(1+\widehat{y}_{1}),}\;\;\;x_{4}={x_{1}^{-1}\cdot(1+\widehat{y}_{1}+\widehat{y}_{1}\widehat{y}_{2}),}\;\;\;x_{5}={x_{2}^{-1}\cdot(1+\widehat{y}_{2})}$
in Example 7.9.
Acknowledgements
I would like to express my gratitude to Antoine De Saint Germain, Jiang-Hua Lu, and Jie Pan for their helpful discussions during my postdoctoral stay at The University of Hong Kong. I am also thankful to Bernhard Keller for his email exchanges and to Jiarui Fei for his valuable feedback on the first version of this manuscript. I am particularly indebted to Ryo Fujita for his helpful discussions, comments and remarks during my writing the part on monoidal cluster categorification.
This project, which began in 2023, has been partially supported by grants from the National Key R&D Program of China (2024YFA1013801), the Research Grants Council of the Hong Kong SAR, China (GRF 17307718, GRF 17306621, GRF 17303420), the National Natural Science Foundation of China (Grant No. 12071422), and the Guangdong Basic and Applied Basic Research Foundation (Grant No. 2021A1515012035).